A Complete Bibliography of Publications in Technometrics for the decade 2000–2009

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
      WWW URL: http://www.math.utah.edu/~beebe/

17 October 2019 Version 1.02

Title word cross-reference

1 [OH05]. 12 [MS01]. 2^{k−p} [BH00, LM02, MP00]. 2^{m−p} [KT03]. $69.95 [Zie01-52]. 80^{th} [Gho07]. C_{P} [Mal00]. D [GV03, Mit00]. E(s^{2}) [NC08]. F [Ano00e, SBD98]. g [FG06]. h [FG06]. K [SC04, CFP08, LD04]. m [BG06]. μ [BGD05]. n [CFP08]. p [BG06, MW05]. r [BG06]. T^{2} [CJFR05].

[?]and-Field:2006:MD. -circulant [LD04]. -Control [BGD05]. -Means [SC04]. -Optimal [GV03, Mit00, NC08]. [?]out-of-Carreras:2008:RSD. -Run [MS01].

0-444-50552-0 [Zie05b]. 0-471-23633-0 [Mai05]. 0-471-38491-7 [Zie01-52].

1900 [Zie00o].

2000 [Ano05s, Mye01, Zie02b]. 21st [Zie01-46, Zie02-45]. 2nd [Ano09h, Ano09i, Ano09a, Ano09m, Ano09d, Ano09e, Hey09, Hly09c, Ke09b, Rut09].

3 [Zie00m]. 3rd [Ano09g, Ano09b, Ano09f, Boo09a, Lip09].

4 [Zie00-44]. 43 [IM02]. 48 [Bri06, Buj06, Efr06, Hub06, Lan06]. 4th [Bar09, Hor09c].

5 [Das00].

6 [Zie02-30].

95 [Hin06].

Abbott [Pla02, Yat02]. Abdel [Ano05-41, Zie02y]. Abdelhak [Bab05]. Abdelmonem [Zie04f]. Abdi [Cut00]. Abdol [Lu04b]. Aberration [DT02, XD05, KT03]. Abrams [Ano05-56]. Accelerated [Din05, Dog01, MEH09, PM03, Pas06, SEM09, ZM06, Ano05-52]. Acceptance [Hem02, Kla01, Van02, Ano09a]. Accounting [CS02]. Accuracy [QW08, MR01]. Acharya [Mul04]. Achieving [TTL07, TJ09, Bre01]. Achim [Mye09a]. across [WB05]. Action [LWS03]. Active [AB03]. Actuarial [Kat09, Nku07, Zie04t, Car07a]. Adams [Ano07l]. Adaptive [Ahm04, CM03, FW06, GL09b, NK03, SHY04]. Addendum [Ano03a]. Adding [LSB06]. Additive [Chi07, Glu06]. Adi [Zie01-55, Zie04z]. Adjusted [LT09]. Adjustment [LBHB04, LS00, Neu03, Zie01-35, LCdC06]. Adjustments [Cla04a]. Adler [McC00]. Administrative [Bar08b]. Adrian [Ano05-49, Ano06-31, Zie02-45]. Adrienne [Ano06-34]. Advanced [Ano08t, Mar06, Xu06, Ano03g, Bar07a, Zie02k]. Advances [Ano03p, Ano05u, Ano09k, Zie00-32, Ano07e, Boo09b, Liu09b, Ord07a, Ord07b]. Affected [LWS03]. Afifi [Zie04f]. after [Bri06, Buj06, Efr06, Hub06, Lan06, Mal06]. against [DGS07]. Age [Boo05b, CLL09, Tho01, Yu07a]. Ageel [Mat01]. Ageing [Hor07]. Agreement [Nag04, Von06]. Agresti [Ano03i]. Ahmet [Neu08a]. Aided [Zie04f]. AIDS [Ano07e]. Air [Mye09c]. Aitken [Soy09]. Aitkin [Ke06]. Ajoy [Ano06j]. Ákos [Ano03-34]. Alan [Ano03i, Ano03-44, Liu04, Ng09, Oli03]. Albert [Ahm08c, Ano06-38, Fle09, HG02, Zie02x]. Albright [Zie04g]. Alec [Hey02]. Alex [Che08c, Chi06]. Alexander [Ano06-31, Ano08n, Von06]. Alfred [Ano05-42, Neu01, Zie00-44]. Algebra [Ahm08e]. Algorithm [BSC09, Li05, McC06a, Mit00, Sin06, Xu02, ZC06]. Algorithmic [GS07, Xu09]. Algorithms [Ano05-33, MWJ06, Zie02d, Ano03m, Ano05-37, Bag05, Mur07]. Alho [Ano06-42]. Ali [Ano07n, Moo01, Neu08a]. Alicia [Ye03, Zie00d]. Allan [Cab00, Cha01c, HI00, HG02]. Allen [Sin01, Sau07, Won04]. Allenby [Kib07]. Allison [Ano06q, Zie00-31]. Allocation [RS04]. Alloy [RWZF04]. alternatives [DGS07, Bor01]. Altman [Mai05]. Aluminum [RWZF04]. Alvin [Ano03x, Bur01a, Rut09]. Alwan [Bel01]. Alwin [Bar08a]. Alyson [Ano08l, Hor09a]. Aman [Zie00t]. Ambroise [Ano05-53]. America [So01]. Amilcar [Zie01t]. Amplification [JW04a, LM04, ME04, Miy04, NW04, Sud04]. Amsden [Sta00, Sta00]. Amsterdam [Zie05b]. Ana [MS07, Ng09]. Analyses [Ano09f, Fri01, Zie02i, Zie02j, Zie04l, Ano07p]. Analysis [Ahm08a, Alt01, Alt02, Ami00, Ami04, Ano03a, Ano03-45, Ano03-49, Ano03-41, Ano03n, Ano03-38, Ano04-29, Ano04z, Ano04-31, Ano05v, Ano05-47, Ano05-54, Ano05-53, Ano05-38, Ano05h, Ano06-33, Ano06e, Ano06-39, Ano06l, Ano06-41, Ano06-32, Ano06-40, Ano07n, Ano07e, Ano08j, Ano08g, Ano08r, Ano08f, Ano08p, AS01, BBJK04, Bas02, Bay02, Boo04a, Boo05b, BN01, Bre04, Bre07, Buk09, Bur02, Bur08a, Cah06, Cam02, Car07d, Car00, CHCFC04, CL09, Cha01c, CE02, Che08a, Con00, Coo00b, CLL09, DWG09, DLM^{+}07, Den04, Din05, DP02, Dri06, Dup05, Esp06b, Geo03, Gho03, GCS06, GS07, GP01, GL09b, Gri00, Gri07, Gun00, Gun04, Hap00, Hec01, Hec07, Hec08, Hey09, Hin05, HCSC07, Hol01, HRV05, HJM09, Huz00, Huz01, Jos06a, JD07, Kar07d, KFV^{+}07, Ke09a, Ken03]. Analysis [Key01, Kok01, Kus03, Kus06, Laz03b, LS05, Lin02, LEH03, LS00, Liu04, Lou01, Lu07, MW04, Mai07, Mar06, Mat01, Maz06a, McC05b, MV05, MBC00, Mee03, MSM03, Mil05, Min02, MBB05, Moo01, Mor05a, MMM08, Mye05b, Nas02, ND02, Nem02b, O’N01, Oak09, PYC07, PK05, Pfa04, Qiu02, RCGD06, Ram00a, Ray03b, RWH^{+}04, RA00, RACH09, Rut05, San08b, Sar05, Slu03, Sta04, SMB01, Sym05, Utl05, VD08, Was05, Weh04, Wlu03, WW00, WKT^{+}03, Xu06, YW05, Zel00, Zie00-35, Zie00g, Zie00w, Zie00-46, Zie00-32, Zie00j, Zie00-42, Zie01j, Zie01n, Zie01p, Zie04-27, Zie04r, Zie04g, Zie05a, ZTL08, ZMW08, Ano05-52, Ano06-39, Alt03, AW01, Ano03i, Ano03-40, Ano03x, Ano03-31, Ano04e, Ano05f, Ano05-52, Ano05-34, Ano05z, Ano05u, Ano05e, Ano06-41, Ano06-30, Ano06o, Ano07j]. Analysis [Ano07e, Ano07u, Ars05, Bel01, Boo03, Bre05a, Bri05, Bur08b, Cam01a, Cam01b, Cha02b, Cha00, Che01c, Cob02, Cob04, Dan01, Esm09a, Esm09b, Gat06, Gho01, Gra02, Gre04, Hec05, Kat01, Kat08b, Lip01, Lit01, Liu07a, Lor06, McC06c, Mic05, Nem00, Nem03, O’N05, Oli03, Oli05, Olw01a, Owe00, Por02, Qiu00, Qu07b, Rav04, Ray03a, Rum06, Sar06, Sau02b, Sca05, Sen06a, Sym01, Zie00n, Zie00-29, Zie00-30, Zie00-39, Zie01d, Zie01-36, Zie01-52, Zie01-31, Zie01a, Zie01c, Zie01f, Zie02o, Zie02-37, Zie02s, Zie04f, vT00, Ano03h, Ano06f, Boo00, Bur02, Chi06, Fin02, Ken02, MM00, McC06c, Mic05, Ruk02, Sin01, Zie01b, Zie02m, Zie04d]. Analytic [Bur01b, HPP01, SB04]. Analytical [Ahm07c, Ano05-46, Ano05-40, Zie01l, Zie04v]. Analytics [Nas08, Sen07]. Analyzing [Ano04g, Ker02, YJL07, Ano05-53, Ano06g, Das04, Kat04a, Von06, Zie02n]. Anand [Gol04]. Anantoly [Ruk02]. Anastasios [Fok07]. Anders [Bar08b, Lip05]. Andersen [Ano07u, Bre07, Dan01]. Anderson [Ano05-41, Car00, Qu07b, So01, Ano03y, Ano04e, MR00]. Andre [Ano06p, Ano03g, Ano07o, Ye00]. Andrea [Kat04b, Was05, Zie01-36]. Andreas [Ano03-46, Ano06-44, Zie01-51, Zie02q]. Andres [Che08a]. Andrew [Ano05-56, Ano06-27, Ano07x, Baj00, Boo04b, Che01c, Dav03, Gre04, Lew04, Rom06, Ros04, Ye03, Ye05, Zie00d, Zie01-42]. Andrzej [Bar08c, Hec09]. Angel [Kus02, Kus06]. Angela [Ami00]. Angles [JP01]. Anh [Ano05-53]. Animal [Zie00j]. Anita [Aba01a]. Ann [Ano05-28, Ano06-37, Ano08j, Zel05]. Annette [Zie02l]. Announcement [Kaf00a, Not03a, Ste09a]. ANOVA [Fre03, Gun03, RSB09]. Answers [Fam03]. Anthology [Ano06-38]. Anthony [Ano06-33, Gho07, Kat04b, Por02]. Antis [Ano05p]. Antonio [Neu08b]. Antony [KS07a, Esp06a]. anyone [Bri01]. Aparna [Bre06]. AppDev [Ano08o]. Application [Ahm06, BF05, CHCFC04, Che06, FLWZ00, FL07, GB01, Glu06, GV07, Li09, RSB09, Sán06, Zie01l, Kus00]. Applications [Ahm08g, Ahm08e, Ano03a, Ano03w, Ano03g, Ano05-58, Ano06-33, Ano07e, Ano08p, Ano08t, Boo05b, CGHL01, Cha01b, DP02, Gho01, Hly07, Jos04, KT05, Liu09b, Lu08, MW04, Mai01, McC06c, MX08, Mic05, Nel00b, Oli02, Oli09d, Ord07a, Ord07b, PM03, PS00, Qia01, Ruk01, San08a, Sen05, Tor01, Was04, Wei03, MW05, MN06, Ano05-55, Ahm07a, Ano06r, Ano07r, Bar07a, Cab04, Cab09, Cha01a, Che02b, Che02a, Den00, Edg04, Gri04, Kar07e, Ker03, Laz05, Lou01, Maz02, McC03b, Mee03, Mye08, Neu08c, Oli08, Pic06, Qia00, Qia02, Sau06, Sha04, Taa02, Zha05, Zie00g, Zie00z, Zie01t, Zie01-34, Zie01e, Zie04q, Zie04x, Zie04h]. Applied [Ano03-28, Ano05x, Cab03, Fot09, Glo05, Hem02, Ke09b, Mam09, Oli04, SC00, Zie00-33, Zie00t, Zie01-27, Zie04c, Ano06-42, Ano06-39, Ahm04, Alt02, Alt03, Ami04, Ano04f, Ano05f, Ano05-52, Ano05g, Ano05e, Ano06-39, Ano06h, Ano06-40, Ano09b, Bar09, Boo03, Bri04, Cah06, Con02a, Dob02, Esm09a, Esm09b, Gra00, Hec05, Hey09, Key01, Lew04, Mal05, Mic01, Mor05b, Mye09a, Nem00, Rag01, Ran03, San05, Sar06, Sen06a, Sho03, Sho04, Ye05, Zie00a, Zie00b, Zie01c, Zie02o, Zie02q, Zie02p, Zim01, Zim04, vT00]. Applying [Goo02, Nem04, Zie02r]. Approach [Ano06-37, AL03a, Che03, Den06, FKQS01, Gun03, GPR05, HPP01, HRV05, JW04a, Jos06a, JS07, LM04, LH02, Lin02, Liu09a, ME04, Miy04, NW04, Pol01, QD04, Sin05, Sud04, SMB01, Weh04, YJL07, ZSC06, ZTL08, vN02, Ahm06, AW05, Ano08m, Cha08b, Che07, Glo05, Gru02, Gun04, HG02, Liu02, Liu06, Oli04, Per04, Qia05, Ruk01, Rut00, Sar05, Sar00b, Sen06b]. Approaches [Ano05-56, Mar06, Tom07, ZCC08]. Appropriate [Qia05]. Approximate [JM00, BG06]. Approximation [Ano07t, GS08, HME08, MY08]. Approximations [CG09]. Arabie [Lip07c]. Arce [Liu06]. Architecture [Glu06]. Area [Ahm07c, Ano04v]. Areas [Ano09k]. Arithmetic [BG04]. ARMA [JTW00]. Armijo [Ano08l]. Armitage [Ano06m, Zie00l]. Armstrong [Ano05s, Mye00]. Army [MW05]. Arnaud [Sar03]. Arnholt [Ng09]. Arnljot [Zie04x]. Arnold [Boo09b, Rao01, Sea09]. Arrays [Den00, KJ09, MMM08, Xu02, ZZJ07]. Art [Ano07v, Kal03, Lip09, Maz00]. Arthur [Oli09d, Ale02]. Artificial [Ano05-33, Zel05]. Ascent [MX08]. Asit [Zie01-41]. Aspects [DLM^{+}07, DS00, Zie00-44, Zie01-39, Ano05h]. ASQ [Zie02h]. Assembly [MNS04]. Assessing [ACBJ09, LPP00, PHSL02]. Assessment [Ano05-43, Ano08s, BH08, Bar08c, BCBS08a, BCBS08b, BMS09, Gol08, HB08, Jon08, KL08, She00, SHY04, VDLB08, Zie00-46]. Assimilation [Wik05]. Assisted [Zie00h, Zie00g]. Assurance [Ano05-40]. Asymmetric [Mar06]. Asymptotic [Oli09a]. Asymptotics [WW00, Zie00c]. Atkinson [Hor09b, Kat04b, Por02, Zie00-32]. Atmospheric [Zie01-49, Wik05]. Attila [Zie01n]. Attoh [Sen06a]. Attoh-Okine [Sen06a]. Attributes [Kla01]. Audit [Ano05i]. Auget [Ano07e]. Augmented [FM04]. Aurelio [Neu09b]. Autocorrelated [AL03b, CM08, GP01]. Automatic [RF05]. Aven [Hua00]. Average [AL03b, CM03, CE02, HMT07, JCR01, RS05, RS06, TTL07, TJ09, ZTW07]. Averages [Rob00a]. Averaging [RdC05, Oli09b]. Avoiding [HME08]. Axel [Lor05]. Aydinli [Ano07v]. Ayyub [Ano03-32, Ke08, San08b, Bre04, Sen06a].

B [Ahm08a, Alt03, Ann06, Ano03-28, Ano03-47, Ano05-52, Ano05-49, Ano05-40, Ano05-56, Ano06-35, Ano06-39, Ano06y, Ano06-41, Ano07w, vB02, Cha05, Che01c, Con05, Fot09, Gre04, Hec09, HL06, Kar07a, LaL09, Law01, Laz03b, Liu04, Mee03, Mye05a, Per04, Pla01, Ray00, Ros04, Rut00, Sea09, Sin08, Soy04, Taa02, Zel05, Zie00u, Zie00m, Zie01-27, Zie01d, Zie01-42, Zie02n, Zie00u]. Baba [Ars05]. Babin [Qu07b]. Back [Ano00a, Ano00b, Ano00c, Ano00d, Ano01a, Ano01b, Ano01c, Ano01d, Ano02a, Ano02b, Ano02c, Ano02d, Ano03b, Ano03c, Ano03d, Ano03e, Ano04a, Ano04b, Ano04c, Ano04d, Ano05a, Ano05b, Ano05c, Ano05d, Ano06a, Ano06b, Ano06c, Ano06d, Ano07a, Ano07b, Ano07c, Ano07d, Ano08a, Ano08b, Ano08c, Ano08d]. Baclawski [Pic09]. Baddeley [Ano05-49]. Baecher [Mye05a]. Bai [Bar07a, Zha05]. Bailer [Ano06g]. Baker [Zie00h]. Bakir [Ano08n]. Balakrishnan [Ano03p, Ano05u, Ano07e, Ano09d, Boo09b, Che02a, Kar07e, Sea09, Zie00y, Zie00z, Zie01-34]. Balanced [HCSC07, KM04, LLI05, Mic05, PKV07]. Baldi [Owe04]. Balding [Ano03o, Mul08]. Balgobin [Zim01]. Ball [Zie02x]. Bands [JM01, Lom05]. Bandwidth [GG04, GLV02]. Bani [Zie01s]. Bank [LH02]. Banks [Zie00m, Zie00v]. Bärbel [Ano04i]. Barbour [Che08b]. Barker [HL06]. Barnard [Zie04k]. Barnett [Ano05-41, Zie00-44, Ano04s, Zie00f, Zie04h]. Barrentine [Rut00]. Barry [Boo09b, Jen08, Qu07b, Rao01, Zie01b]. Bartholomew [Ano03-45, Lip01]. Barton [Bin01]. Based [BE03, BGMS03, CG09, DJ07, FG01, Gun04, JLH^{+}06, KM04, KMM08, Mar05, MS00, NM04, NK03, ÖBMT02, Pfa04, QH01, Rob00a, Sar00b, Sen02, SMB01, Whi04, Zha06, ZDCS03, Zie02e, ZTW08, Alt06, Bar08b, Sca07, Sin09]. Basford [Hec01]. Basic [Ano06-37, Van04, Zie01-50, Zie02-51, Ano05j, Bri05, McC02, Mye00, Zie01d, Zie01e]. Basics [Ano03-46, Ano06-44, Boo05a, Zie01-51]. Basis [Qia02]. Basu [Zie01-41].

Batagelj [Pic07]. Batch [FK00, KML04]. Bates [Zie01z]. Bayes [Zie01f, Sin05, Gru08, Zie01f]. Bayesian [Ahm08d, Ano05k, Ano08i, Cha08b, Cla04a, Con01, Con05, Gan05, Gho08b, Gru02, HG02, Kat01, Ker03, Sho04, Ye03, Ye05, Zie00d, Ano05-56, BE03, BGD05, BFL02, CG09, Dri06, GP06, GLM07, GP01, GPR05, HJMW04, Jos06a, KJ09, Kun08, LLBN08, Li05, MSM03, QW08, RdC05, RSB09, RSS02, RACH09, TGLG09, Tom07, TH05, WCT09, ZM06, Ahm08b, Ano06i, Ano09c, Bur04, Cas05, Das07a, Gan02, Gho09, Gre04, Hor09a, Kib07, Mar03, Sea01, Sen02, Soy04, Soy09, Zel05, Zie02s]. Bazaraa [Mur07]. Be [RS05]. Beam [HTMW08]. Bean [Bre02]. Beata [Zie01-54]. Becki [Zie01x]. Bedford [Bas02]. Beginner [Fin05]. Behavior [Nel01]. Beirlant [Laz05]. Bell [Ano09c]. Belle [Ano05l, Pry03]. Belt [Mor05b]. Ben [Ano08n]. Benaroya [Esm06]. Benchmarking [Kar07a]. Bendat [Cha01c]. Benedetti [Zie04-27]. Bengt [Bar04]. Benichou [Zie02z]. Benjamin [Qu08, Ray03a]. Bennett [Ano06-38, Bai04, Zie00-50, Zie02x]. Benoit [Zie01-35]. Benteftifa [MM00]. Beran [Zie04w]. Berenson [Zie01-45]. Berger [Ano06h, Wei03, Zie02-41]. Bergman [Bar04]. Berliner [Zie01-49]. Bernard [Ano07g, Ano08n, Ke09b]. Bernd [Ano06t]. Berry [Ano06k, Zie01y, Zie01o, Ano08m, Liu02]. Berthold [Ano05v, Zie00-29]. Berthouex [Ano03-51]. Bertram [Moo01]. Best [Che03]. Beta [Oli08]. Beth [HI00]. Bethel [Gun03]. Bethlehem [Zie00h, Zie02b]. Bets [Cha02a]. Better [LM02]. Bettini [Ano05s]. Betty [Cut00]. between [GLV02]. Beyond [Ano03-35, Bzi04, Hec08, Mor04, Zie04o, Zie04p, Sen05, Zie02-51, Kaf03, Zie04k, Oli08, Vuk00]. Bhagat [Ano06x]. Bhat [Cab03]. Bhavani [Zie00i]. Bhisham [Mor05b]. Bhote [Zie01-55, Zie04z, Ano03-48, Zie01-55, Zie02-49, Zie04z]. Bianca [Gho09]. Bianchi [Lip07b]. Bias [JS03, JC06, Kva08]. Bias-Correction [JS03]. Biased [BF05, HK00]. Biedermann [Soy09]. Biemer [Ano03r]. Bilal [Ano03-32, Bre04, Ke08, San08b, Sen06a]. Bill [Jar08]. Billo [Lar08]. Bimal [Ye00, Zha05]. Binary [Ano04n, KLM08, VD08, FVB02, MN06]. Binning [MNS04]. Binomial [WT09, Jen02]. Bioenvironmental [Zie01g]. Bioinformatics [Bur09b, Mul04, Zie02g]. Biological [Ano05-39]. Biology [Was04]. Biomedicai [Ano09l]. Biomedical [Ano03-41]. Biometrika [Zie02t]. Biopharmaceutical [Zie01q]. Biostatistical [Zie00-33, Zie02v, Zie02u, Zie02a]. Biostatistics [Ano03u, Ano06m, Gol06b, Ano03t, Zie00l, Ano05k, Ano05l, Zie02w]. Birthday [Gho07]. Bisgaard [O’N03, Zie01h]. Bishop [Ano03o, Nea07]. Bit [HTMW08]. Bivariate [Ano09d]. Bjorn [Bre07, Dan01]. Black [Cab08, Qu07b, Car07b]. Black-Scholes [Cab08]. Blæsild [Was04]. Blanke [Bar07b]. Blind [AL03a, SA08]. Blischke [Ano03-50, Olw01b]. Block [CLM04, Khu06, NC08, VB09]. Blocked [CLY04, Ke05, MB04]. Blocking [CW02]. Bloomfield [Cha01c]. Blundell [Liu09b, Ord07a, Ord07b]. Bo [Bar04]. Bock [Pic07, Sch04, Sho05]. Boeck [Sen06b]. Bohnenkamp [Ano08o]. Böhning [Zie00g]. Boland [Kat09]. Bolch [Hly07]. Bolstad [Ano08i, Gan05]. Book [Aba01b, Aba01a, Ahm04, Ahm06, Ahm07b, Ahm07a, Ahm07c, Ahm08a, Ahm08d, Ahm08b, Ahm08g, Ahm08c, Ahm08f, Ahm08e, Ahm08h, Ahm09, Ale01, Ale02, All01, Alt01, Alt02, Alt03, Alt06, Ami00, Ami04, AC02, AW01, AW05, And08, Ann01, Ann03, Ann06, Ano03i, Ano03m, Ano03r, Ano03-27, Ano03-48, Ano03-36, Ano03-33, Ano03-35, Ano03w, Ano03p, Ano03q, Ano03-40, Ano03-34, Ano03z, Ano03-45, Ano03t, Ano03u, Ano03g, Ano03-49, Ano03x, Ano03v, Ano03s, Ano03-28, Ano03-39, Ano03-41, Ano03-42, Ano03o, Ano03f, Ano03l, Ano03-44, Ano03n, Ano03h, Ano03y, Ano03k, Ano03-29, Ano03-30, Ano03-37, Ano03j, Ano03-46, Ano03-43, Ano03-47, Ano03-38, Ano03-31, Ano03-32, Ano04l, Ano04-30, Ano04-29, Ano04e, Ano04n, Ano04v, Ano04s, Ano04t, Ano04i, Ano04j, Ano04w, Ano04z]. Book [Ano04m, Ano04r, Ano04h, Ano04q, Ano04-27, Ano04g, Ano04k, Ano04y, Ano04f, Ano04u, Ano04p, Ano04o, Ano04-28, Ano04x, Ano04-31, Ano05v, Ano05-46, Ano05i, Ano05f, Ano05-32, Ano05-42, Ano05-52, Ano05-27, Ano05-34, Ano05q, Ano05o, Ano05-36, Ano05k, Ano05-43, Ano05t, Ano05-35, Ano05y, Ano05g, Ano05j, Ano05-44, Ano05-47, Ano05r, Ano05s, Ano05m, Ano05x, Ano05-55, Ano05-49, Ano05n, Ano05-30, Ano05z, Ano05p, Ano05-40, Ano05w, Ano05-33, Ano05-39, Ano05-41, Ano05u, Ano05-57, Ano05-45, Ano05-56, Ano05-54, Ano05-53, Ano05-31, Ano05-28, Ano05-37, Ano05-48, Ano05l, Ano05e, Ano05-51, Ano05-38, Ano05-29, Ano05-50, Ano05h, Ano06-33, Ano06w, Ano06t, Ano06-42, Ano06-34, Ano06g, Ano06j, Ano06x, Ano06-43, Ano06r, Ano06p, Ano06z, Ano06i, Ano06e]. Book [Ano06-27, Ano06-35, Ano06n, Ano06-31, Ano06q, Ano06-39, Ano06l, Ano06k, Ano06y, Ano06m, Ano06-38, Ano06-41, Ano06u, Ano06-37, Ano06s, Ano06h, Ano06-30, Ano06-28, Ano06-36, Ano06f, Ano06-29, Ano06o, Ano06v, Ano06-32, Ano06-40, Ano07k, Ano07w, Ano07m, Ano07r, Ano07v, Ano07s, Ano07q, Ano07n, Ano07t, Ano07x, Ano07j, Ano07g, Ano07e, Ano07h, Ano07u, Ano07p, Ano07i, Ano07f, Ano07l, Ano07o, Ano08j, Ano08k, Ano08m, Ano08g, Ano08o, Ano08q, Ano08h, Ano08l, Ano08n, Ano08r, Ano08f, Ano08p, Ano08e, Ano08i, Ano09h, Ano09c, Ano09g, Ano09l, Ano09k, Ano09i, Ano09j, Ano09b, Ano09a, Ano09m, Ano09d, Ano09e, Ano09f, Ars05, Avi01, Bab02b, Bab02a, Bab05, Bag05, Bai04, Bai05, Bai07b, Bai07a, Baj00, Bak00, Bak01, Bar02]. Book [Bar05, Bar08a, Bar08b, Bar01, Bar04, Bar07c, Bar07b, Bar07a, Bar08c, Bar09, Bas00, Bas02, Bay02, Bel01, Bin01, Bla01, Bod02b, Bod02a, Boo00, Boo03, Boo04a, Boo04b, Boo05a, Boo05b, Boo07a, Boo07b, Boo08, Boo09c, Boo09b, Boo09a, Bor00, Bor02, Bor01, vB02, Bre02, Bre05a, Bre06, Bre05b, Bre01, Bre04, Bre07, Bri03, Bri04, Bri05, Bru00, Buk09, Bur01a, Bur02, Bur00, Bur01b, Bur03, Bur04, Bur06, Bur07, Bur08a, Bur08b, Bur09a, Bur09b, Bzi04, Cab00, Cab03, Cab04, Cab06, Cab08, Cab09, Cah06, Cam01a, Cam01b, Cam02, Car07d, Car07a, Car07b, Car07c, Car00, Cas05, Cha01a, Cha02a, Cha01b, Cha03, Cha01c, Cha02b, Cha04, Cha05, Cha08b, Cha09, Cha00, Che01a, Che03, Che01b, Che01c, Che02b, Che02a, Che07, Che08c, Che08b, Che08a, Chi06, Chi07]. Book [Cho06, Cla04a, Cob00, Cob02, Cob04, Con00, Con01, Con02a, Con02b, CL03, Con04, Con05, Cut00, Dan01, Das00, Das04, Das07b, Das07a, Dav03, Den00, Den04, Den06, Din05, Dob02, Dog01, Dog04, Dok08, Dok09, Dra04, Eas07, Edg04, Esm05, Esm06, Esm09a, Esm09b, Esp06a, Esp06b, Esp07, Eva01, Fah07, Fam03, Fin02, Fin05, Fle09, Fok07, Fot01, Fot04, Fot07a, Fot07b, Fot07c, Fot09, Fre00, Fre03, Fri00, Fri01, Fri05, Gan02, Gan05, Gan06, Gar00, Gat02, Gat06, Geo00, Geo03, Gho01, Gho02, Gho03, Gho05, Gho07, Gho08b, Gho08a, Gho09, Gib01, Glo01, Glo05, Glo06, Gol04, Gol06a, Gol06b, Gol07, Goo02, Gra00, Gra02, Gre04, Gri00, Gri04, Gri07, Gro00, Gru02, Gru05, Gru08, Gun03, Gun04, Gun06, Gup00, Hap00, Hec01]. Book [Hec05, Hec07, Hec08, Hec09, Hem02, Hen01, HL06, Hes02, Hes04, Hey01, Hey02, Hey06, Hey09, Hil00, Hin05, Hin06, Hly07, Hly09a, Hly09b, Hly09c, Hof00, HR05, Hol01, Hol08, Hor06, Hor07, Hor09b, Hor09a, Hor09c, Hua00, Hus08, Hut04, HI00, Huz01, HG02, Huz03, Iye02, Jar04, Jar08, Jar09, Jen02, Jen08, Jen09c, Jen09a, Jen09b, Jia03, Joh00, Joh03, Jos04, Kaf03, Kal03, Kal06, Kap03, Kar07b, Kar07d, Kar07c, Kar07e, Kar07f, Kar07a, Kat03b, Kat03a, Kat04a, Kat04b, Kat08a, Kat09, Kat01, Kat08b, Ke05, Ke06, Ke07, Ke08, Ke09a, Ke09b, Ken02, Ken03, Ker02, Ker03, Ket00, Key01, Kia00, Kia01, Kib07, Kim00, Kim04, Kim05, Kok01, KS07b, KS07a, Kuh03, Kuh05b, Kuh05a, Kuo00, Kus00, Kus02, Kus03]. Book [Kus06, LaL01, LaL04a, LaL04b, LaL09, Lar08, Law01, Laz03a, Laz03b, Laz05, Lew04, Li06, LH06, LZ02, Lin02, Lip01, Lip05, Lip06a, Lip06b, Lip07c, Lip07b, Lip07a, Lip09, Lit01, Liu02, Liu04, Liu06, Liu07a, Liu07b, Liu09a, Liu09b, Loe07, Lor01, Lor05, Lor06, Lou01, Lu04a, Lu04b, Lu07, Lu08, Lum01, Mag09, Mag00, MM00, Mai05, Mai07, Mai08, Mal05, Mam08, Mam09, Mar03, Mar06, Mat01, Maz02, Maz06b, Maz06a, Maz00, McC00, McC03a, McC07, McC08, McC03b, McC06a, McC02, McC06c, McC03c, McC04, McC05a, McC05b, McG02, McG07, Mee03, MS07, Mic01, Mic03, Mic05, Min02, Mit02, Moo01, Moo03, Mor04, Mor05a, Mor05b, MR00, Mul01, Mul04, Mul08, Mul02, Mur07, Mye00, Mye01, Mye05a, Mye05b, Mye08, Mye09b, Mye09c, Mye09a, Nag04]. Book [Nag06, Nas01, Nas02, Nas04, Nas08, Nea07, Nel00a, Nel01, Nem04, Nem08, Nem00, Nem02b, Nem02a, Nem03, Neu01, Neu03, Neu05, Neu08c, Neu08a, Neu08b, Neu09b, Neu09a, Ng05, Ng06, Ng08, Ng09, Nku07, O’N01, O’N03, O’N05, Ode00b, Ode00a, Oli03, Oli04, Oli05, Oli07, Oli08, Oli09a, Oli09c, Oli09b, Olw01a, Olw01b, Ord07a, Ord07b, Owe00, Owe03, Owe04, Owe06, Pan00, Ped01, Per04, Pet03a, Pfa04, Pic06, Pic07, Pic09, Pig06, Pla01, Pla02, Por02, Pry03, Qia00, Qia01, Qia02, Qia05, Qia07, Qiu00, Qiu02, Qiu05, Qiu08, Qu07b, Qu07a, Qu08, Rag01, Ram00a, Ram00b, Ran03, Rao00, Rao01, Rav04, Ray03a, Ray00, Ray03b, Rob00b, Rob02, Rob05, Rob06, Rom03, Rom06, Ros04, Ros05, Ruk01, Ruk02, Rum00, Rum04, Rum06, Rut02, Rut05, Rut09, Rut00]. Book [San03b, San03a, San05, San07, San08b, San08a, Sar00a, Sar02, Sar03, Sar05, Sar06, Sar00b, Sau02b, Sau02a, Sau05, Sau06, Sau07, Sca02, Sca04, Sca05, Sca07, Sch04, Sea01, Sea02, Sea05, Sea09, Sen00, Sen05, Sen07, Sen02, Sen06a, Sen06b, Sha02, Sha04, She00, She07, Sho03, Sho04, Sho05, Sin01, Sin06, Sin07, Sin08, Sin09, Slu03, Smi01, Smi02, Smi06, So01, Soy04, Soy09, Spi07, Sta00, Sta04, Ste00b, Ste04, Str01, Sul01, Sym01, Sym05, Taa02, Tob00, Tob03, Tor01, Umb01, Utl05, Van04, Vol05, Von03, Von06, Vuk00, Wal03, Wal05, Was01, Was02, Was04, Was05, Weh02, Weh04, Wei03, Wik03, Wik05, Wil02, Wil05, Wlu03, Wlu06, Won04, Woo00, Xu06, Yat01, Yat02, Ye00, Ye03, Ye05, Yeh05, Yeh07, Yu07b, Zel00, Zel01, Zel05]. Book [Zha05, Zie00u, Zie00-35, Zie00c, Zie00p, Zie00m, Zie00-40, Zie00e, Zie00-45, Zie00n, Zie00i, Zie00o, Zie00d, Zie00-51, Zie00-29, Zie00g, Zie00f, Zie00a, Zie00-48, Zie00-44, Zie00-49, Zie00r, Zie00q, Zie00s, Zie00x, Zie00y, Zie00z, Zie00w, Zie00v, Zie00-36, Zie00-46, Zie00-30, Zie00-47, Zie00-27, Zie00-33, Zie00-41, Zie00-32, Zie00-37, Zie00l, Zie00j, Zie00t, Zie00-39, Zie00h, Zie00k, Zie00-50, Zie00-38, Zie00-42, Zie00-34, Zie00-31, Zie00b, Zie00-28, Zie00-43, Zie01-43, Zie01-28, Zie01j, Zie01-51, Zie01-27, Zie01-38, Zie01-41, Zie01d, Zie01i, Zie01-50, Zie01-36, Zie01s, Zie01x, Zie01-44, Zie01-47, Zie01-49, Zie01-39, Zie01r, Zie01-52, Zie01y, Zie01v, Zie01-33, Zie01-31, Zie01-53, Zie01u, Zie01-42, Zie01-35, Zie01-40, Zie01-29, Zie01a, Zie01m, Zie01-55, Zie01-46, Zie01z, Zie01c, Zie01-30, Zie01o, Zie01t, Zie01n, Zie01k, Zie01q]. Book [Zie01g, Zie01l, Zie01h, Zie01p, Zie01w, Zie01-37, Zie01f, Zie01-54, Zie01-34, Zie01e, Zie01b, Zie01-32, Zie01-48, Zie01-45, Zie02-42, Zie02-31, Zie02-36, Zie02-35, Zie02-28, Zie02-48, Zie02-51, Zie02i, Zie02j, Zie02-46, Zie02v, Zie02-45, Zie02-30, Zie02t, Zie02l, Zie02o, Zie02-38, Zie02-41, Zie02-50, Zie02m, Zie02y, Zie02u, Zie02-33, Zie02-34, Zie02q, Zie02-49, Zie02h, Zie02p, Zie02f, Zie02e, Zie02c, Zie02d, Zie02-37, Zie02a, Zie02g, Zie02x, Zie02b, Zie02-27, Zie02-29, Zie02-39, Zie02k, Zie02-47, Zie02-40, Zie02-32, Zie02r, Zie02-44, Zie02-43, Zie02s, Zie02n, Zie02w, Zie02z, Zie03b, Zie03c, Zie03a, Zie04a, Zie04-27, Zie04u, Zie04i, Zie04s, Zie04q, Zie04n, Zie04c, Zie04b, Zie04k, Zie04o, Zie04y, Zie04e, Zie04l, Zie04w, Zie04j, Zie04f, Zie04x, Zie04p, Zie04z, Zie04m, Zie04d, Zie04r, Zie04v, Zie04h, Zie04g, Zie04t]. Book [Zie05b, Zie05a, Zie06, Zim00, Zim01, Zim02, Zim04, vN02, vT00]. Booker [Ker02]. Bootstrap [Ano05-35, CLM04, FG01, JS03, JC06, Weh02, dCGM05, Ano08e, Bab05, Kus00, Was01]. Borgelt [Slu03]. Boris [Sen00]. Borror [Maz06a]. Bos [Mam08]. Bosq [Ode00b]. Both [CR06, GHB^{+}07]. Boundaries [Ano07x]. Bounded [Oli08]. Bounds [Ano09c, FW00]. Bouvier [Ano04-27]. Bowerman [Zie01i]. Box [Ano07p, Pig06, O’N03, Zie01h]. Boyadjieva [Rob02]. Bozdogan [Ars05, Ano04w]. Braatz [Zie02-29]. Brad [Zie00-28]. Bradley [Ahm09, Sca07, Zie00d, Zie01f]. Brady [Hec09]. Brain [McC07]. Branching [HJM09]. Brandimarte [Boo09c]. Brandt [Ram00a]. Brani [Ket00, Lu08, Sen02]. Braun [Hly09a, Gun04]. Breakdown [CCD09]. Breakthrough [Ale01, Zie04k, Ano05m]. Breeding [Hec01]. Brémaud [Ode00a]. Brenda [Hec09]. Brereton [Boo04a]. Brettner [Tob03]. Brewer [LaL04a]. Breyfogle [Zie01x, Ano03q, Gar00]. Brian [Ano05j, Ano06e, Ano06-40, Ano09f, Den04, Fri01, Hey01, Ke06, Lip09, Neu09a, Zie01k, Zie02i, Zie02j, Zie02o]. Briggs [Bai04]. Brill [Hly09b]. Brillinger [Ano06-33]. Britt [Bar08b]. Britz [Law01]. Bro [Boo05b]. Broad [Ahm06]. Brocklebank [Ano04t]. Brockwell [Ano03s]. Brown [Zie00a, Ano03-51, Rag01]. Browne [Ano05-56]. Bruce [Ano04z, Ano04y, Ano06-42, Ano06-31, Din05, Kal06, Rut09, Zie01i, Zie02-44]. Brunner [Gho03]. Bruno [Nas02]. Brunsdon [Zie06]. Bryan [Ano05-32, Zie02-43]. BSE [Zie01-39]. Buck [Ano07x]. Buehler [KL02]. Buncher [KS07b]. Buntinas [Neu05]. Burdick [Maz06a]. Burlew [Zie00-41]. Burman [KB05, MS01]. Burnham [Ano03y, MR00]. Burr [Maz06b]. Burrill [Bre01, Hof00]. Burt [Utl05]. Bury [Ann01]. Burzykowski [Ano06-43]. Bushman [Zie00-28]. Business [Alt02, Ano03-36, Ano05-36, Ano05-43, Ano05m, Jen08, Kal03, Sar05, She00, Zie04y, Ano06-29, Car07b, Zie01i]. Butler [Sta00]. Buydens [Zie00u]. Buyse [Ano06-43]. Byron [Ano03-39, Ano04-31, Dob02].

C [Alt01, Ano03-51, Ano03p, Ano03-34, Ano03x, Ano03o, Ano04t, Ano04j, Ano04h, Ano04k, Ano05-46, Ano05-27, Ano05-47, Ano05s, Ano05p, Ano05-41, Ano05u, Ano06i, Ano06l, Ano06-41, Ano06s, Ano07q, Ano07j, Ano07h, Ano09e, Bar09, Bel01, Boo09b, Bor01, Bri03, Bri04, Bri06, Buj06, Bur01a, Cha05, Che03, Che01b, Das00, Edg04, Efr06, Gho07, Gol06b, Gra02, Gri00, Hes04, Hil00, HR05, Hol01, Hor09b, Hor09a, Hub06, Kap03, Kat04b, KS07b, Kus00, Lan06, Law01, Loe07, Maz06a, Mor05b, Mur07, O’N03, Pla01, Pla02, Qu07b, Ran03, Rao01, Rob06, Rut09, Sea09, Smi01, Soy04, Soy09, Spi07, Vol05, Yat02, Yeh07, Zie00u, Zie00-45, Zie00-48, Zie00x, Zie00y, Zie00z, Zie00w, Zie00-32, Zie00-28, Zie00-43, Zie01-52, Zie01v, Zie01-33, Zie01a, Zie01-46]. C [Zie01z, Zie01g, Zie01h, Zie01p, Zie01-32, Zie02-31, Zie02-36, Zie02v, Zie02p, Zie02-27, Zie04d]. Cache [GV07]. Caitlin [Ano07x]. Calafiore [San07]. Calculated [Cha02a]. Calculating [CCD09, Tom07]. Calculus [Ano03g, Ke09b, Zie00o]. Calibration [BG05, FJ08, For08, GB01, HSR09, MW04, TG00, Zie04a, Gat02]. Call [BJL^{+}07]. Cambridge [Ano03-47, Ano07w]. Campbell [Zie00m]. Campolongo [Che08a, Was05]. Can [Gel06]. Cancer [Ano07e, Boo07a]. Cannings [Ano03o]. Cannot [Gel06]. Cantor [Liu04]. Cao [Lu04b]. Capability [Ahm08h, Boy01, Pol01, Vuk00, Yat02, Zim00]. Caponetto [Zie02-40]. Cappé [Nag06]. Carbon [Kva08]. Care [Ano05-56]. Cariboni [Che08a]. Carl [Ahm09, Kar07c, Sca07]. Carl-Erik [Kar07c]. Carleton [Aba01a]. Carlin [Gre04, Zie00d, Zie01f]. Carlo [Ahm08d, Ano03-34, Ano05-31, Bor00, Con01, Den06, Hes02, Kim00, McC04, Ode00a, Sar03, BF08, Jon04, LCdC06, MR01, Kuo00]. Carmen [Ano05-56]. Carmichael [Ano05i]. Carmona [Nas02, Zie05a]. Carol [Zie02w, Ano09l]. Carr [Ano05o]. Carrera [Ano06r]. Carriquiry [Ye03, Zie00d]. Carroll [Kat04a, Wal05]. Carter [Tob00]. Case [BF05, Mye09c, Ano03-50, vB02, Ye03, Zie00d]. Casebook [Sym01]. Casella [Zie02-41, Ano05-31, Kim00]. Cassandras [Ano09h]. Castillo [Rao01, Gho09, Neu03]. Caswell [Ano06-42]. Cate [Ano03-37]. Categorical [Ano08r, Das04, Das07a, GP01, Hap00, JW04a, LM04, ME04, Miy04, NW04, Sud04, Zie01-53, vT00, Ano03i, Zie01j]. Categorization [GLM07].

Catherine [Ano09l]. Causal [Hea03, Maz00, Ye05]. Causality [Zel01]. Causation [Bur03]. Cause [Bre07, Dan01]. Cautious [AK04, BF05]. Cavanagh [Zie02-48]. Cawse [Zie03a]. Celebrating [Gho07]. Censored [BSG00, HV07, JM00, JM01, KCM00, LS00, Nel00b, SNS07, Ste04, YW05]. Censoring [Kun08, MEM05, NCB04]. Census [So01]. Census-Taking [So01]. Centered [Zie01m]. Century [Zie01-46, Zie02-45]. Cepstral [Ano00e, SBD98]. Ceric [Lew04]. Cerioli [Kat04b]. Cerrito [Sin08]. Cesa [Lip07b]. Cesa-Bianchi [Lip07b]. Cetama [Bur00]. Chain [Ahm08d, Kuo00, Jon04, MR01]. Chains [Hly07, Ode00a, Ruk01]. Chakraborti [Ano04q]. Chakravarty [Sen00]. Challenges [GS07, Yu07a]. Chambers [Zie04d]. Champigny [Ano05s]. Chan [CL03, Qia07, Zie01-36, Zie02-50]. Chance [Fri05, HI00, Ano07f, Kok01, Neu09a, Zie01k]. Chances [Ano05n]. Chandra [Mit02]. Change [COR08, Wik03, ZH06, ZTL08]. Change-Point [COR08, ZH06]. Changepoint [GG04, HZ05]. Changes [CE02, TH05]. Chaos [CL03, Sca05]. Chap [Ano03t, vT00]. Chapman [Ano03-27]. Characteristics [WR02]. Characterization [KML04]. Charles [Ano05m, Ano06-38, Ano06-41, Gol06b, Ken03, Rag01, Von03, Zie00p, Zie01j]. Charlton [Zie06]. Chart [CM03, JTW00, LPP00, Rob00a, SJ02, CJFR05, Nel01]. Charting [CA06, HMT07, LT09, Umb01]. Charts [AL03b, BGD05, CM08, JWT^{+}02, JCR01, LPP00, RS04, RS05, RS06, Umb01, WM05]. Chatfield [Ano04-29, Bor02]. Chatterjee [Ano07n, Moo01]. Chattopadhyay [Sar06]. Chatwin [Ano05-41]. Chau [Boo05a]. Checking [AB04, Bed00]. Checks [McC05a]. Chemical [Atk05, Boo04a, Boo05b, Nem02b, Zie02h, FVB02, Neu08a]. Chemistry [Hor06, Ano03n, Ano05-46, Ano05-40, Zie01l, Zie01-54, Zie04v]. Chemoinformatics [Zie04e]. Chemometrics [Ano05-33, Zie00u, Zie04v, Boo04a, Boo05a, Zie01l]. Chen [Con01, Zha05, Zie02s, Ano03l, Che08b, Glo06, Zim01]. Cheng [Gup00, Zie02d]. Cherkassky [Ano08k, Lor01]. Chernick [Ano03u, Ano08e, Was01]. Cheryl [O’N05]. Chevret [Nem08]. Chi [Gup00]. Chi-Lun [Gup00]. Chiang [Zie02-29, IM02]. Chih [Cab00]. Chih-Ling [Cab00]. Childers [Esm05]. Chilès [Zie00r]. Chilko [Bre05b]. Chin [Ano09d, Hor07]. Chin-Diew [Ano09d, Hor07]. Cho [Ano04y, Oli05]. Choice [Ano07f, CW02]. Cholette [Kar07a]. Chong [Fre03]. Choosing [LSW09, Qia05]. Chosen [HN03]. Chow [Ano04-28, Ano05-54, Zie00j, Zie01q]. Chowdhury [Bar07c]. Chris [Ano04-29, Ano06-41, Bor02, Hap00, Sau02b, Zie06]. Christensen [Fam03, Zie02k]. Christian [Ano05-31, Ano09k, Kim00, Mye09a, Slu03, Zie04g, Zie04t, Mye05a]. Christl [Zie01-39]. Christophe [Ano05-53]. Christopher [Gan06, Kok01, Nea07, Zie04g]. Christos [Ano09h]. Christy [Che08c, Chi06]. Chromatography [Zie01n]. Chronologies [Ano07x].

Chuang [Che08c, Chi06]. Chuang-Stein [Che08c, Chi06]. Chung [Ano04-28, Ano05-54, Zie00j, Zie01q, Ano07f]. Cihon [O’N05]. Cinar [Neu08a]. circulant [LD04]. Circulation [RMGR09]. Civil [Baj00]. Claeskens [Oli09b]. Claims [CLL09, FL07]. Clark [Bur03, Mye01, Zie00h, Zie04f, Ami04]. Clarke [O’N01, Ano03-41]. Clarkson [Ano06-41]. Class [ALKP02, Mor00a, SP07, Zie01-55, Zie04z]. Classical [Gun00, RG07, Con02b]. Classics [Law00, Mor00b, Var00]. Classification [Ano05-51, DT02, Epi08, GCS06, JL07, SR04, Soy04, Bar02, Pic07, Zie04a, Sul01]. Classified [Ano08r]. Classifiers [Bag05]. Claude [Bre01, Hof00]. Clayton [Ano05r]. Clemen [Zie02-32]. Clifford [Lou01]. Clinical [Ano03-33, Ano03f, Ano03j, Ano04-31, Ano05-56, Ano05-54, Che07, Chi06, Hus08, Sin07, Zie02w, Zie04-27, Zie04l, Zie04-27]. Clinton [Sca05]. Clive [Ano05-41, Pan00]. Clock [JS03, JC06]. Closed [VD07, dC02]. Closed-Loop [VD07, dC02]. Cluster [Ros04]. Clustering [Mai01, RG07]. Cochran [Ano06-38, Ano05m]. Code [MBC00, Kuh05a]. Coded [ASW^{+}06]. Codes [RBM08, RSB09, RBM11]. Coding [Hec07]. Cody [Ano06-40, Hol08, San03b]. coefficient [Bri01]. Coefficients [BK04]. Cognition [Fri05, Zie00e]. Coleman [Zie03b]. Coles [Smi02]. Collaborators [Ano09n]. Collani [Jen02]. Collecting [Iye02]. Collection [Zie01d, Zie00h]. Collett [Ano04n, Ano04o]. Collision [RHOYL06]. Colored [HTMW08]. Colosimo [Gho09]. Colton [Ano06m, Zie00l, Zie02w]. Column [HWY02]. Combinations [MWJ06]. Combinatorial [Den06, Zie03a]. Combined [LaL04a]. Combining [HG06, SA08, WB05, Bag05]. Comment [BH08, Bat08, Cha08a, CW08, Gol08, Han08, HB08, Jon08, KL08, LI08, Old08]. Comments [IM02, Mal00]. Committee [Zie02h]. Common [Ano03j]. Communication [Rom06, AC02]. Companion [Ano05w, Zie04c]. Company [Dok09]. Comparative [Rut05, Zie00f]. Comparing [Esp06b, FS08]. Comparison [BV01, FHR06, Lom05, MBC00, Tom07, Pet03a, McC03a]. Comparisons [HWY02, JM00, Mil05, RS06, Zie00-34]. Compatibility [JLSZ08]. Competing [Con02b, CL05, And08]. Competing-Risks [CL05]. Complete [Bai05, Ste04, YW05, Zie00-51, Kuh05a]. Complex [Ano05-38, Cha09, CG09, Fam03, Oak09, RBM08, ZZ07, RBM11]. Compliance [Van02]. Component [Ano03-31, Ano06-27, HRV05, Mai07, Ray03b, Jon04]. Components [ALKP02, Che02b, Chi01, HN03, Mar05, IM02, MW05, Mic03]. Composite [Mee01, MBDD09]. Compound [ZZJ07]. compounds [FVB02]. Comprehensive [Ahm08h, Goo02]. COMPSTAT [Zie02b]. Computation [GJS09, Ano06i, Con01]. Computational [Ano05t, GS07, LH06, Mar06, Nas04, Owe03, Ram00a, Zie02b, Ano06j, Kus02, Ng05, Owe06]. Computationally [RSS02, ZSC06]. Computations [Ahm08e, MV05]. Computer [BBP^{+}07, BBS09, BCL^{+}09, Den04, Dri06, Glo06, Glu06, GV07, HSNB09, HSR09, Hly07, HJM09, LaL04b, LS05, LBH^{+}06, LSW09, McC03b, MBC00, MB06, MMM08, Oak09, QWW08, RBM08, RWH^{+}04, RSB09, SdHSV03, WCT09, Zie01l, RBM11, Alt06, Bod02a, Zie00g, Zie00h, Zie04f]. Computer-Aided [Zie04f]. Computer-Assisted [Zie00g]. Computer-Based [Alt06]. Computing [Ano08h, Bat08, Cha08a, CW08, Gra00, Gui08, Han08, Kus03, LI08, Mai05, Mul04, Old08, Wil08a, Wil08b, Hes02, Lip09, Zie02-40]. Concentrated [JP01]. Concentrations [Dup05]. Concepts [Ano03m, Ano08k, Lor01, Lou01, Zie00p, Kat03a]. Conceptual [Dra04]. Conditional [Rao01]. conditioning [PDDL02]. Condra [Zie02f]. Conduct [Ano03f, Sin07]. Confidence [Ahm04, BG05, Chi01, DM08a, DS00, GD03, HME08, JM00, JM01, KM04, Lom05, Oli09c, IM02, MW05]. Confusion [Vuk00]. Congdon [Das07a, Gan02, Sho04]. Congress [Liu09b, Ord07a, Ord07b]. Conjecture [Maz00]. Connie [Maz06a]. Conover [Gro00]. Consider [MS01]. Constantine [Ye03, Zie00d]. Constrained [GV07, RF05, SdHSV03]. Constraints [Smy00, VD07]. construct [KB05]. Constructed [NC08]. Constructing [Chi01, Xu02, IM02]. Construction [Ano07x, But04, HWY02, JL07, Mit00, PKV07, Rut05, Xu09]. Consultant [Rum00]. Consulting [AC02]. Contamination [MY08, Zie03b]. Contemporary [Car07b, Nem04, So01]. Content [FG01]. Content-Corrected [FG01]. Context [BGMS03, JC06]. Context-Based [BGMS03]. Contingency [Che03]. Continual [Bre01, Hof00]. Continuous [Ano05-42, Oli08, Ano09d, Ruk01]. Continuous-Time [Ruk01]. Contour [RBM08, RBM11]. Control [Ahm08h, Ano03-43, AL03b, AK04, Bar08c, BGD05, BGMS03, CM03, CM08, CA06, DW06, DM08b, Edg04, Fre00, Fri00, Gho09, HZ05, HMT07, Hof00, Hor06, Huz01, Jos03, Kaf03, KML04, LWS03, LC00, O’N03, Pla01, RS04, RS06, Rob00a, Ste00a, SJ02, Var00, Wlu06, WM05, ZH06, Zie00-47, Zie00-42, Zie02-30, Zim02, CJFR05, Ano05-27, Ano09a, Bar07b, Hes04, Maz06b, Mit02, Mor04, Neu03, Neu08b, Qiu08, Spi07, Yat02, Zie01v]. Controlled [YM00]. Controller [FFG06, TTL07, TJ09, dC02]. Convenience [GHB^{+}07]. Cook [Gra00, Sán06]. Cookbook [Zie02c]. Cooke [Bas02, Car07d]. Coordinate [Che06]. Copeland [Esp06b]. Coping [Gat02]. Copulas [Ray00]. Corbett [Zie05b]. Cornell [Zie02-28]. Corporations [Ale01]. Corrected [DS00, FG01]. Correction [Ano00e, Ano04-32, Ano05-58, Ano08t, Ano08s, JS03, JC06]. Correlated [DWG09, Zho01]. correlation [Bri01]. Correlations [ŽH09]. Correspondence [Hec07]. Corrosion [FHR06, RWZF04]. Counting [SM04, Liu07a]. Counts [So01]. Couper [Ano05-29, Ano05-50, Zie00h]. Course [Fot09, Hly09a, Kim05, Kok01, Sea09, Utl05, Vol05, Zie01a, LaL01]. Courtroom [Zie01-44, Zie02r]. Covariance [HMT08, Hub01, JS07, LS00, PP01a, PP01b, RW01, Bur02]. Covariates [RLR06, ST06]. Coverage [LWY02, RHOYL06, WT09]. Cowan [Owe00, Ram00a]. Cowell [Ped01]. Cox [Ano03-29, Ano07r, Ano07s, Che01a, Gho07, LZ02, Mic03, Ano03-29, Dok08, Zie02t]. Craig [Kal06]. Cramer [Bar05, JLSZ08]. Crawley [Kus03, Ng06]. Credit [Kla01]. Creighton [Ano05-28, Ano08j, Ano09g, Zie04n]. Creveling [Ano05p]. Cristianini [Bre05a]. Criteria [BV01, DT02, RACH09, SA08]. Criterion [Zho01, KT03]. Critical [Ano08s, BN01, Din05, Gui08, VDLB08]. Cross [Ano04-31, Ano08r, Boo08, Den06, HTMW08, Ano03k, Ano04-31]. Cross-Classified [Ano08r]. Cross-Cultural [Ano03k]. Cross-Entropy [Boo08, Den06]. Cross-Over [Ano04-31, Ano04-31]. Cross-Section [HTMW08]. Crossing [Ano07x, Hly09b]. Crossley [Yat01]. Crowder [Ano07l, Con02b]. Crowe [Hey02]. Crowley [Zie04-27]. Cultural [Ano03k]. Cumulative [CLM04, GS08, JM01, RS05, Umb01]. Cupello [Zie01x]. Current [Ano05-41]. Curvature [PHSL02]. Curve [JQ09, Kuh03, Qia00, Zie02x]. Curves [BG05, Qia05, She09]. Customer [Bar04, Sin01, Zie01m]. CUSUM [LC00, LPP00, QH01]. Cusums [UL07]. Cuthbert [Zie01r]. Cycle [Ng08]. Cynthia [Zie00h, Zim00].

D [Aba01b, Aba01a, Ahm08b, AW05, Ano03-35, Ano03-28, Ano03o, Ano03-50, Ano04j, Ano04m, Ano04g, Ano04u, Ano05p, Ano05w, Ano05-54, Ano05l, Ano06-42, Ano06i, Ano06q, Ano06-39, Ano06-29, Ano07m, Ano07r, Ano07s, Ano07l, Ano08p, Bab05, Bak01, Bzi04, Cab00, Cab06, Cab09, Cam02, Car00, Che01a, Che03, Che08b, Che08a, Con00, Dog04, Dok08, Geo03, Gho08b, Gol07, Hol01, Hor09b, Hor09c, Jen09c, Jen09b, Key01, Kia00, Kus00, Laz03b, Mam09, McC03a, Mic03, Mor04, Mor05a, Mul01, Mur07, Mye09c, Ode00b, Olw01b, Rom03, Sch04, Sho05, Soy04, Ste04, Sul01, Tob00, Wei03, Zie00u, Zie00-49, Zie00-34, Zie00-31, Zie00-43, Zie01-27, Zie02-51, Zie02t, Zie03b, Zie04o, Zie04p, Zim01, OH05]. Dabbene [San07]. D’Agostino [Che08c]. Dagum [Kar07a]. Dahaene [Car07a]. Daley [Sca04]. Dalgaard [Ano09i, San03a]. Dali [Bar07a]. Dallas [Bur02, Nem00]. Damien [Bar08c, Ke09b]. Dan [Ano06n]. Dana [Hey02]. Dani [Ahm08d, Kuo00]. Daniel [Ami00, Ano05q, Ano05-55, Boo07b, Bre05b, Con05, Mag09, Mye09b, O’N03, Sym01, Zie01-38, Zie01r, Zie01a, Zie01h]. Dankmar [Zie00g]. Darwinian [ZC06]. Daryl [Sca04]. DasGupta [Oli09a]. Data [Ahm07b, Ahm07c, Alt03, Ano03a, Ano03i, Ano03-40, Ano03-45, Ano04w, Ano04o, Ano05v, Ano05-42, Ano05-52, Ano05-47, Ano05-55, Ano06q, Ano06-39, Ano06-41, Ano06s, Ano06o, Ano06v, Ano06-40, Ano07h, Ano08j, Ano08k, Ano09m, BBJK04, BH08, Bar08b, Boo04a, BCBS08a, BCBS08b, Bri05, Bur02, Cah06, CM08, Cha01c, Cha09, Cha00, Cob04, Coo00b, DLM^{+}07, Epi08, Esm09a, Esm09b, Gat06, Gho01, Gho03, GS07, GP01, Gol07, Gol08, GM01, GHB^{+}07, Gre04, HG06, HPP01, HB08, Hec01, Hec07, HV07, Hol01, HC00, HI00, Huz00, HG02, HN03, Iye02, JM00, JM01, Jen08, JLH^{+}06, JL07, Jon08, Kat04b, KL08, KD05, Kok01, KCM00, Kus03, Kus06, LLBN08, LZ02, Lin02, Lit01, LS00, Lor01, Lu04b, Lu07, MY08, McC06c, Mee03, Mic05, Nel00b, Nem00]. Data [Nem03, Neu09b, O’N05, Owe00, Per04, Qu07b, RCGD06, RG07, RA00, Rum06, San03b, San05, Sar06, Sau02b, Sho05, Sin08, Slu03, SNS07, ST06, SB04, SP07, Sym01, Utl05, VD08, Wik05, WW00, Woo00, Ye05, YW05, Zie00n, Zie00-29, Zie00-46, Zie00-32, Zie01-28, Zie01j, Zie01d, Zie01-52, Zie01y, Zie01f, Zie02o, Zie02n, Zie05a, ZMW08, MN06, Ano03-49, Ano05-53, Ano06-39, Ano03-49, Ano03-41, Ano03h, Ano04n, Ano04g, Ano05-55, Ano06g, Ano07m, Ano08g, Ano08n, Ano08r, Avi01, Boo00, Bre06, Bur06, Car00, Das04, Das07a, Fin02, Fok07, Geo03, Hap00, Hut04, Jar04, Jen09c, Joh03, Kat04a, Ken02, Kia00, Laz03b, Min02, She07, Sin01, Ste04, Wlu06, Zie01r, Zie01-53, Zie04d, Zie04r, Ano03m, Ano03l, Ano05o, Ano05-53, Ano06k, Boo07b, Cab04]. Data [Dra04, Gun04, Jen09a, Lou01, Mag09, Mag00, Mar06, Mul04, Nas08, Pic07, Ram00a, Zie00i, Zie01o, Zie01n, Zie02c, Zie04g]. Data-Analytic [HPP01]. Data-Chosen [HN03]. Data-Mining [Ano05-55, KD05]. Database [Ano04z]. Databases [KFV^{+}07, WM02]. Datasets [Cha09, IM05, KS07a, Mai01, MZ02]. Datta [Bur09b, Sar06]. Dave [Ano07l]. David [Ano03-45, Ano03v, Ano03y, Ano04l, Ano04n, Ano04t, Ano04o, Ano05v, Ano05j, Ano05m, Ano05-56, Ano05-48, Ano06-33, Ano06-28, Ano06-32, Ano07r, Ano07s, Ano07q, Ano07h, Ano07i, Ano09j, Bar02, Bod02a, Cha04, Cha09, Con02a, Fah07, Gho07, Gol06b, Gru08, Hin06, Kat01, Lip01, Liu07a, MR00, Mul08, Mye05b, Nas01, Nku07, Ped01, Qu08, Ros04, Sca04, Sch04, Sea05, Sha02, Sho05, Umb01, Wal05, Wil02, Ye03, Zie00m, Zie00-29, Zie00-30, Zie00t, Zie00-38, Zie01-50, Zie01-52, Zie01c, Zie01w, Zie01-45, Zie02-37, Zie03b, Zie04i, Zie04y, Zie05b]. Davida [Sta00]. Davidson [Cab06]. Davies [Zie04a]. Davis [Zie01j, Ano03s, Ano05-47, Ken03, LaL09]. Davison [Gho07, Kus00, HR05]. Dawid [Ped01]. Dayanand [Zie00b, Zie01-28]. Dean [Ami00, Ano09a, Wlu06]. Debajyoti [Zie02s]. Debnath [Cha02b]. Deborah [Boo04b, Cha01a, Rum04]. DeCarlo [Kal06]. Decelerated [Sin05]. Decision [Ahm08b, Ano05-43, Bla01, Car07b, Mar03, Oak09, She00, Zie02-33, Jen08]. Decision-Theoretic [Oak09]. Decisions [KLM08, Ano06y, Ano06v, Bur04, Gri07, Zie02-32, Gri07]. Decomposition [HC00]. Decompositions [Cha09]. DeCoursey [Jos04]. Decreasing [Pro00]. Defective [DM08b]. Defects [DM08a]. Degradation [BK04, CLM04, SEM09]. Degree [GPR05]. Dekking [Yeh07]. Del [Gho09, Neu03]. Delaney [Ano04g]. Delay [Ano00e, SBD98]. Delay-Fired [Ano00e, SBD98]. Delfiner [Zie00r]. Delivery [Ano08o]. Delwiche [Hor09c]. Demanding [RSS02]. DeMaris [Ano05-42]. Deming [Kaf03]. Demography [Ano06-42, Ano06-42]. Demonstration [MEM05, MN06]. Denison [Soy04, Ros04]. Dennis [Gat06, Gra00, Hey01]. Densities [FS08]. Density [Cab09, GCS06, PC01]. Denuit [Car07a]. Dependence [Car07d, CS02, Hor07, ŽH09]. Dependency [McC06b]. Dependent [BGMS03, Car07a, Hut04, RLR06]. Derek [Sin01]. Derivative [JWT^{+}02, JQ09]. Derived [Won04]. Derong [Qiu08]. Derr [AC02]. Descriptive [Bur08a]. Descriptors [Epi08]. Deshpande [Ahm07b]. Design [Alt01, AW01, Ano03f, Ano03-38, Ano05-38, Ano06n, Ano06o, Ano07r, Ano09g, AL03b, Bin01, BC07, BV01, Bzi04, CM08, CL09, Che01a, CE02, CA06, DW06, DT02, Den04, DS06, FLWZ00, FFG06, Glu06, Gol06a, GL09b, HJM09, HWY02, Huz01, Iye02, JW02, Jos03, Jos06a, KJ09, Ke05, KD05, MB04, MS01, Mor00b, Neu08c, O’N01, O’N03, Oli09d, Pig06, QD04, RBM08, RHOYL06, Rob02, Rut00, Rya08, San07, Sar00b, Sin07, Wei03, ZZJ07, Zie00-35, Zie00p, Zie00n, Zie00-48, Zie00w, Zie00-39, Zie02f, Zie03a, Zie04o, RBM11, Ano04-31, Ano05-54, Zie04-27, Ano08f, Eas07, HL06, Loe07, Nel00a, Qia01, Sta04, Tob00, Xu06, Zie00-32, Zie00-37, Zie00-42, Ami00, AW01, Ano05p, Ano05-54, Ano06l, Ano08f, Ano09e, Din05, Esp06a, Gri00, Hey02, Ke09a, Maz06a, Mor05a, Nem02b]. Design [Won04, Xu06, Zie00j, Zie01p, vN02]. Designated [KL02]. Designed [AW05, GCLH01, Wlu03, YJL07]. Designing [For08, Hea03, Ano04g, Lip07a]. Designs [AB03, ACBJ09, ALKP02, Ano05-58, Atk05, BFL02, BS03, BV01, BH00, But04, CW02, CLY04, CG09, FM04, GV03, GJS09, Hor09b, KCV02, KT05, LWY02, LN00, LL03, LLY03, LP06, LST07, ML01, Mee01, MS05, Mit00, Mor00a, MBDD09, NC08, PKV07, Rut02, SdHSV03, TG01, VB09, Voe05, WLER06, XCW04, XD05, Xu09, Zho01, KT03, KB05, LM02, LD04, MP00, Ano06h, vB02]. Desirable [LST07]. Desk [Yat01]. Destination [Li05]. Destructive [SEM09]. Desu [Ste04]. Detect [TH05]. Detecting [Ano00e, SBD98, WR02]. Detection [ASW^{+}06, Bod02a, COR08, Coo00a, Hub01, JLH^{+}06, JQ09, PP01a, PP01b, RW01, WM02, XCW04, ZSC06, Zie02-29, Zie03b]. Determination [CHvdW06, HSR09, TTL07, TJ09]. Determine [DM08b]. Determining [Bla01]. Deterministic [HC00, RSB09]. Deterministic/Stochastic [HC00]. Deutsch [Ano05r]. Developing [Hin06]. Development [TG00, Zie00j, Zie03a]. Deviations [YM00]. Devices [LBHB04]. Devore [Ano05g, Mic01, Min02, Sau02b, Zie04m]. Dey [Ano06i, Zie01s]. DFSS [Alt06]. Dhillon [Ann06]. Diagnosability [Che06, ZDCS03]. Diagnosing [AS01]. Diagnosis [DM08b, Zie02-29, Bar07b]. Diagnostic [Ano03-42, Coo00b, Por02, RA00, McC05a]. Diagnostics [BO09, Boo07a, MBG07, SX07, Kuh03]. Diamond [Rut02]. Diane [Zie04-27]. Dickey [Ano04t]. Dickinson [Ano04q]. Dictionary [Ano03-47, Ano04-30, Ano07w]. Didier [Zie00-37]. Diebold [Sar00a]. Dielman [Alt02, Ano05f]. Dietrich [Neu01]. Diew [Ano09d, Hor07]. Differential [Bur09a]. Differentiation [RF05]. Diggle [Ano03h, Sin09, Sym05]. Digital [FR07]. Dijksterhuis [Cha05]. Dillman [Zie02-46]. Dimensional [Car07d, CG09, Dri06, LHB^{+}02, MZ02]. Dimensionality [Lin02]. Dimitris [Str01]. Dineen [Cab08]. Dioxide [dCGM05]. Dipak [Zie01s]. Direct [Ano07g]. Directional [Mai08]. Dirk [Boo08, Den06]. Disciplinary [Ano07x]. Discovering [Ano05-55]. Discovery [HI00, HG02, KBS08, LT09, O’N03, Ano04w, Pig06, Zie01h]. Discrete [Ano06-34, Ano07m, Ano08g, Ano09h, Che06, FFG06, Kia00, dC02, Ano05q, Fot01]. Discrete-Part [Che06]. Discrete-Time [FFG06, Fot01]. Discriminant [PYC07]. Discrimination [HPP01, Zie01-28]. Discussion [Bri06, Buj06, Coo00b, DS00, Efr06, Hub06, Hub01, Lan06, RW01, SBF^{+}08, WKT^{+}03, LM04, ME04, Miy04, NW04, Sud04]. Disease [Ano03a, Mee03, Ano05-56]. Dispersion [BN01, MZ02, ML01]. Display [Mul02, Utl05]. Distance [Ano08m, Liu02, Sán06]. Distances [Oli02]. Distributed [FHR06, KFV^{+}07]. Distribution [BSG00, BKG09, Boo09c, CLM04, CS01, FG06, GS08, Jen02, Kar07a, KMM08, Kun08, LPP00, NM04, NCB04, PS00, Weh02, ZS09, Bas00, Bru00, Huz03]. Distributions [Ahm07a, Ahm08c, Ann01, JM01, MEM05, Oli08, Sen05, SHY04, Weh02, WW00, Zie04t, Ano05q, Ano06-34, Ano09d]. Disturbance [AK04, dC02]. diverse [FVB02]. Division [Zie02h]. Dmitrienko [Che08c, Chi06]. DNA [Ano05-54, Ano05-53, Boo07a, Hly09c, Mul08, Das07b]. Do [Ano05-53, Gel06, Gol07]. Dobrivoje [Ano06j]. Dobson [Zie02l]. Dodge [Gho07, Ano04-30]. Dodson [Hey01]. Does [Bri01, Pla01]. Dolores [Ng09]. Domhof [Gho03]. dominance [DGS07]. Dominici [Mye09c]. Dominique [Cut00, Zie01-35]. Don [Ano05i, Zie02-46]. Donald [Bar01, Che01c, Esm05, Gre04, Jar04, Law01, Nel01, Pfa04, Rum06, Tob03, Vuk00, Wil05]. Donev [Hor09b]. Donna [Ano03-42, Ye03]. Donnelly [Zie01-39, Bai05]. D’Orazio [Kar07f]. Dorp [Oli08]. Dose [Nem08]. Dose-Finding [Nem08]. Double [TTL07, TJ09]. Doucet [Sar03]. Douglas [Ano03-37, Ano04h, Ano05i, Ano05-27, Ano06t, Ano06l, Ano06-41, Ano07j, Bri04, Gra02, Gri00, Kat04a, Maz06a, Ram00b, Spi07, Umb01, Zie00e, Zie01-49, Zie01v, Zie01z, Zie01p, Zie02-31, Zie02-36, Zie02p, Zie02-27, Zie04u]. Dowdy [Bre05b]. Dräger [Jen02]. Drähici [Ano05-53]. Draper [Ano07p]. Driven [Mag00]. Drivers [GM01]. Dror [Zie00-34]. Drug [Ano04-28]. Dual [QD04, CFP08]. Dual-Response [QD04]. Duane [Bar08a]. Duckworth [Zie04y]. Duda [Bar02]. Dudewicz [Weh02]. Dugard [vB02]. Dummies [Rum04]. Dummies(R) [Kal06]. Duncan [Hly09a]. Dunlop [Hol01]. Dunn [Ami04, Zie02o]. Dunne [Liu09a]. Dunson [Ano09j]. Durbin [Hin05]. Durrett [Hly09c]. Dynamic [Fot01, Fre00, NK03, RCGD06, RLR06, Sán06, YM00]. Dynamics [Lu04b].

each [BG06]. Eamonn [Hor06]. Earl [Zie01m]. Early [WM02]. Earth [Zie00s]. Easterling [Rya08]. Easy [Ale02, Zie00-41]. Eckes [Zie02-47]. Econometric [Ano04f, Cob04]. Econometrics [Liu09b, Mye09a, Ord07a, Ord07b, Ano05x, Ng05, Zie01-29]. Economic [Ano06y, Lip07a, Zie00t]. Economics [Bar05, Liu09b, Ord07a, Ord07b, Zie04t, Alt02, Bre04]. Ecotoxicology [Zie01-47]. ed [Ano09h, Ano09g, Ano09i, Ano09b, Ano09a, Ano09m, Ano09d, Ano09e, Ano09f, Bar09, Boo09a, Dok09, Hey09, Hly09c, Hor09c, Ke09b, Lip09, Rut09]. Eddy [Ano07g]. Edelman [Cut00]. Edgar [Gho03, Neu01]. edited [Zie01-52]. Editor [GD04, Ans03, GL09a, Kaf00b, Kaf01a, MLC09, Not02, Not03b, Not04, Pet03b, PCdC04, Rya08, Sit05, Sit06, SS07, Sit07, Ste08b, Ste09b, Ste09c, dCGRB08]. Editor-Elect [SS07]. Editorial [Ano04-32, Ano09n, Kaf00a, Not03a, Ste09a]. Edmund [Ano03n]. Edward [Ano03-27, Ano05-54, Ano06-39, Ano06-37, Ano06s, Ano09a, Weh02, Wlu06, Zie00q, Zie00-46, Zie01b]. Edwards [Wil02]. Effect [Ano09j, Ahm06]. Effective [AC02]. Effectiveness [JC06]. Effects [Ano06q, BG05, BN01, FK00, GB01, Lan01, ML01, MBDD09, ZDCS03, Gho08a, Zie01z]. Efficiencies [Voe05]. Efficient [CL07, RSS02, RS04, Xu09, YJL07, ZS09, ZSC06]. Efromovich [Qia00]. Efron [Ahm09]. Efstratios [Ano06n]. Eigenanalysis [Smy00]. Eight [Ano06m]. El-Shaarawi [Zie02y, Ano05-41]. Elart [Jen02]. Eleanor [Ano05-29]. Elect [SS07]. Electrodynamics [RMGR09]. Electronics [Tob03]. Elementary [Hol01, Sca04, Jen09b, Kar07b, Maz06b, Sau06]. Elements [Ano09m, Zie03c, Cab03, Gho02, Mam09, Owe03, Sar00a, Zie00k]. Elementwise [MY08]. Elephants [Car07c]. Eliciting [GCLH01, Ker02]. Eliminating [PW01]. Elimination [MWJ06]. Elisa [Ano03-40]. Elizabeth [Ano03-49, Ano05-29, Ano07j, Gra02]. Ellen [Ano05-37]. Elliptical [Sen05]. Ellis [Wlu06]. Elsevier [Zie05b]. Elston [Zie02u]. Eltinge [Zie02-46]. Embracing [Yu07a]. Emerling [Law01]. EMP [Jen09c]. Empirical [Dri06, GS08, Lin02, Zie01f]. Employing [Smy00]. Emulation [RMGR09]. Emulators [BO09]. Encounters [Kok01]. Encyclopedia [Ahm08h, Ano06m, Zie00m, Zie00l, Zie01q, Zie02y, Zie02z]. Enders [Ano06-33, Ano04f]. Endpoints [Ano06-43]. Enduring [Hun00]. Energy [Sán06]. Engel [Zie04e]. Engineering [All01, Alt06, Ano05-58, Bar07c, Bod02b, Bre04, CGHL01, Esm06, Gri07, Hey01, Hey06, Huz00, Joh00, Jos04, JW04b, KD05, KT05, Liu07b, Mam09, McC06a, Neu08c, Olw01a, San08b, Sau07, Sta04, Wei03, Zie01d, Zie02d, Zie04m, Cla04b, Aba01b, Ann01, Baj00, Lu08, Mye05a, Ng08, Rom03, Sen00, Ano04h, Ano06n, Zie02-27]. Engineers [Ano03-32, Ano05g, Ano09b, Esp06a, Ke08, Lar08, Mic01, Rut02, Sau05, Zie00s, Zie05b, Zim01, Ann06, Ano03-51, Ano07k, Bri04, Jen02, Kim04, Mam08, Ram00a, Yeh05, Zie02p]. English [Bri03]. Enn [Huz03]. Enrico [Ano05h]. Enrique [Gho09, Neu03, Rao01]. Enterprise [Jen08, Jen09a, LaL09, Sin08]. Entropy [Boo08, Den06]. Environment [Zie00-44, Ano04i]. Environmental [Ano03-51, Ano05-39, Ano05-41, Ano06g, Ano06r, Mye09c, San08a, Smi01, Zie00-46, Zie01u, Zie01t, Zie02-43, Zie03b, Zie04u, Lum01, Zie04h]. Environments [JLSZ08]. Environmetrics [Zie02y]. Epidemiologic [Zie02z]. Epidemiological [Ano05-45]. Epidemiology [Cla04a, Mye09c, Ano05-57, Ano05-57, Ano06-39, Zie01-42, Zie02v, Zie02u, Ano06o, Zie00n]. Equation [Ano06-31, Cha08b, Fin05, Lip05]. Equations [Bur09a, Zie01r, Cha03, Gan06]. Equipment [Ahm07c]. Equivalence [Laz03a]. Equivalent [PKV07]. Ergonomics [FR07]. Eric [Ano03z, Ano07g, Bur07, Fot07b, Gol06b, Nag06]. Erich [Dok09]. Erik [Kar07c, Rum00]. Eriksson [Nem03]. Erjavec [Bod02b]. Erkki [Ano05-38, LH06]. Errata [RBM11]. Erricos [Ano08h]. Error [Bar08a, Cla04a, HCSC07, Khu06, Sco01, Smi01, Gup00]. Errors [BG05, CR06, MJH00, YS05, Bur00]. Essence [Cho06]. Essential [Gho01]. Essentials [Alt06]. Estimate [CE02]. Estimated [JCR01, CJFR05].

Estimates [GCS06, MZ02, PC01]. Estimating [ALKP02, Cha03, CR06, Gan06, GHB^{+}07, HTMW08]. Estimation [Ahm07c, Ano05y, Ano05-45, Ano09c, Bur00, Bur04, CLM04, Che02b, DWG09, DM08b, GG04, Hes04, HK00, Hub01, JS03, KCV02, Li09, Mam08, Ode00b, PKV07, PP01a, PP01b, PW01, Qia00, RBM08, RG07, RW01, Sán06, Smy00, Whi04, ZS09, ŽH09, Jon04, RBM11, Ano04v, Zie04j, Kar07c]. Estimator [YW05]. Estimators [JC06, SNS07]. Eubank [Bak00]. Eugenia [Wik05]. Eun [Von06]. Eva [Ano05-49]. Evaluating [Ano05-29, HTMW08, RACH09]. Evaluation [Ano05-51, Ano06-43, Bar01, Hly07, Ke09a, LPP00, MW05, Ano05-56, Cob00, Neu08a, Ano06p]. Evans [Zie00-30, Ano05-36]. Event [Ano07u, Ano09h, Bre06, SP07]. Events [Ano03a, FL07, Mee03]. Everitt [Ano06e, Zie02o, Zie02n, Ahm08a, Ano03-47, Ano06-40, Ano07w, Ano09f, Fri01, Neu09a, Zie01k, Zie02i, Zie02j]. Everson [Ray03b]. Everyday [Bai04]. Evidence [Gol07, Mul08, Jar08]. Evolution [ZC06, Ano06-43, Hly09c]. Ewens [Zie02g]. Exact [BSC09, DM08a, Gho05, ND02, Ano08g, Lit01]. Example [Gun04, Hol08, Zie00-38, Ano07n, Moo01]. Example-Based [Gun04]. Examples [Ano08o, Sar06, Utl05, Ano03j, Ano09g, Bur09a]. Exceedance [ZCC08]. Excel [Ano03-43, Jos04, Zie00-47, Zie00-38, Zie01-45, Zie04g, Lar08]. Excellent [Zie02-51]. Excel(R) [Sho03]. Exceptional [Das07b]. Exchangeable [ST06]. Expansion [LLBN08]. Expected [FW06]. Experiment [GCLH01, LSW09, RBM08, Rut02, RBM11]. Experimental [ALKP02, Ann03, Ano08f, vB02, DS06, Eas07, HL06, Hor09b, Mit00, Mor00a, Mor00b, Nel00a, Oli09d, Rya08, Xu06, Zho01, Zie00-48, Zie00-32, Zie00-37, Zie00-42, Gol06a, Sar00b, Wei03, Zie03a]. Experimentation [Ano08s, FW06, VDLB08]. Experimenters [Pig06]. Experimenting [Gib01]. Experiments [AW01, AW05, Ano04g, Ano08f, Ano09g, BS03, BN01, CL09, DP02, Esp06a, GL09b, HSNB09, HSR09, Hea03, HTMW08, HJM09, Iye02, Jos06a, JD07, KBS08, KCV02, KCM00, Lan01, LS05, LBH^{+}06, LS00, MB04, MB06, Mil05, Neu08c, PHSL02, QW08, QWW08, RWH^{+}04, RACH09, Rut00, SdHSV03, Tom07, Won04, Xu06, YJL07, ZZJ07, Zie00-35, vN02, PDDL02, Ami00, Ano03-38, Ano06l, Bin01, Bzi04, Che01a, Den04, Gho03, Gri00, Ke05, Nem08, O’N01, Rob02, Rut05, Wlu03, Zie00p, Zie00w, Zie01p, Zie02f, Alt01, Zie02-28]. Expert [Ker02, Ped01, RMGR09]. Explained [Dog01, GP06, Won04]. Explanation [Pro00]. Explanatory [Sen06b]. Exploration [Min02]. Exploratory [Zie01-52, Kus06]. Exploring [Ano07h, Kat04b]. Exponential [SHY04]. Exponentially [AL03b, CM03, CE02, HMT07, HMT08, JCR01, RS05, RS06, TTL07, TJ09, ZTW07]. Exposition [Ahm08h]. Expression [Ano03-49, Ano05-53]. Extending [LZ02]. External [YM00]. Extrapolation [GV07]. Extreme [Ano08p, Cam02, FHR06, LWS03, MN06, Smi02, Bri01, Ano04i, MS07]. Extreme-Value [FHR06]. Extremes [Dup05, Laz05]. Extremum [Van02]. Eye [Von06, Ano06-31].

F [Ahm08d, Ahm08b, Alt01, Ano03-33, Ano03-41, Ano03-29, Ano03-38, Ano05-32, Ano05n, Ano05-30, Ano06h, Ano06-30, Ano08g, Bai04, Bar08a, Che08a, Das07b, Esp06b, Eva01, Gat02, Gol06a, Hor09a, Joh03, Kok01, Lew04, Loe07, Ng09, Oli03, Qu07b, Rom06, Sch04, Sea02, Sho05, Soy04, Soy09, Ste00b, Was02, Wil05, Yeh07, Zie00h, Zie02-27, Zie02-43, Zie04q]. Fabrizio [Kat01, San07]. Factor [AS01, FW06, GCLH01, Lip01, XCW04, KT03, Ano03n]. Factor-Analysis [AS01]. Factorial [Ano05-58, BS03, BH00, BN01, But04, CL09, CLY04, Gho03, KBS08, KT05, Lan01, LN00, LL03, Loe07, ML01, MB04, Mil05, Voe05, XD05, Xu09]. Factors [AB03, HJM09, KCM00, MBDD09, QWW08]. Fagerhaug [Bre07, Dan01]. Fahrmeir [Zie02e]. Fail [Boy01]. Failure [Fin02, Geo03, HTMW08, JW04a, KL02, LM04, McC06b, MEM05, ME04, Miy04, NW04, Nel00b, Pro00, Sud04]. Failures [NM02]. Fairclough [Zie04-27]. False [KBS08, LT09]. Families [Oli08, Ahm08c]. Family [SHY04]. Fan [Ano03-34, Lu04a]. Fanchi [Ano07k]. Fang [Kuh03]. Faraway [Rob05]. Farebrother [Kat03a, Zie00o]. Farnum [Ano05g, Mic01]. Fasser [Tob03]. Fast [CG09, GJS09, LC00, ŽH09]. Fatigue [Pas03]. Fatigue-Limit [Pas03]. Fault [Bar07b, JLH^{+}06, LT09, Zie02-29]. Fault-Tolerant [Bar07b]. Fearn [Zie04a]. Feed [Jos03]. Feed-Forward [Jos03]. Feedback [DW06]. Feedforward [Rao00]. Feinberg [Hey02]. Feldman [McC00]. Felinger [Zie01n]. Felsovályi [Ano03-34]. Feo [Zie04k]. Ferguson [Zie01-39]. Ferhan [Neu08a]. Feridun [Zie00-44]. Ferligoj [Pic07]. Fernandez [Cab04]. Ferraty [Lu07]. Ferreira [MS07]. Fetterman [Gun03]. Fiber [GLV02]. Field [LHB^{+}02, MEH09, SB04]. Fields [BLC01, Ode00a, Pic06, ŽH09, Ano08p, Bar05, Iye02]. Fienberg [Ano07q, Ano08r, So01]. Fifty [Ste08a]. Filia [Ano09l]. Filip [Ano08k, Lor01]. Filling [CL07]. Filter [LH02, Li09]. Filtering [Ano09c]. Filtering/Tracking [Ano09c]. Filters [CA06]. Finance [Ano04i, Ano05-48, Ano08p, Cab08, Ke09b]. Financial [Ano03z, Ano04u, Ano06f, Bur07, Fot07b, Lu04b, Qia07, Zie02m, Zie05a]. Finding [Nem08]. Findlay [Ano03-27]. Fine [Rao00]. Fingerprint [DJ07]. Fingerprint-Based [DJ07]. Finite [FL07, Sha02]. Finite-Horizon [FL07]. Finkelstein [Zie02-44]. Finkenstädt [Ano04i]. Finn [Mar03, Car00]. Fiona [Ano03-45]. Fionn [Hec07]. Fired [Ano00e, SBD98]. First [Hly09a, Kim05, Kok01, LaL01, Sea09, Vol05]. Fisher [Ano05l, Qia01, Qia01]. Fission [Ano06-30, Bur08b]. Fit [Bed00, BDNN04, CS01, NM04, PS00]. Fits [MS00]. Fitted [GD03]. Fitting [Qiu04, RF05, Weh02, Zie00o, Zie01r]. Fitzgerald [Sca02]. Fitzmaurice [Ano05e]. Fixed [Ano06q]. Flannery [Lip09]. Fleiss [Ano04y]. Fleming [Liu07a]. Flexible [CS02]. Flgueredo [Neu09b]. Flood [Cam01a]. Florac [Aba01a]. Flow [LHB^{+}02]. Flowgraph [Huz00, Bre06]. flows [Bri01]. Floyd [Ano05-50]. Fluid [LHB^{+}02]. Focuses [ŽH09]. Fokianos [Ray03a]. Folded [MS01, MS05]. Folded-Over [MS01, MS05]. Foldover [LL03, LLY03, LM02]. Fons [Ano03k]. Foo [Boo05a]. Foo-Tim [Boo05a]. Forecast [NK03, Sán06]. Forecast-Based [NK03]. Forecasting [Ano04t, LH02, Lu04b, She09, Ano03s, Ano05-36, Ano06-42, Ano06j, Bor02, Sar00a]. Forecasts [FL07]. Forensic [Fah07, Mul08, Soy09]. Forest [Mye09b]. Foreword [Hun00]. Formula [Cab08]. Formulae [Zie01-38]. Formulated [BSC09]. Formulation [HZ05, Gol06a]. Forrest [Ano03q, Gar00, Zie01x]. Fortran [Hin06]. Fortuna [Zie02-40]. Fortunato [Bab02a]. Forty [Kaf00c]. Forward [Jos03, Kat04b]. Fotheringham [Zie06]. Foundations [Ano07s, Bas02, Zie02-51]. Fourier [Cha01c]. Fowler [Ano05-50]. Fox [Hey09, Zie04c]. Fractal [Sca07]. Fractal-Based [Sca07]. Fractional [BS03, BH00, But04, Lan01, LL03, ML01, MB04, Voe05, XD05, Xu09]. Fractionated [Jos06a, LS00, VB09]. fractions [LM02]. Frailty [ST06]. Fraley [Ano06-41]. Framework [BBP^{+}07, KML04]. Francesca [Mye09c, Was05]. Francis [Ke06, Rut05, Sar00a]. Frank [Gho03, Qiu02, Wal03, Zie00-46]. Franklin [Qia01]. Fransconi [Owe04]. Frantisek [McC03c]. Fred [Cob04, Mor05b, Zie01r]. Frederic [Ano06-33, Lu07]. Frederick [Ano07q, Ano07h, Rob06, Zie01-52]. Freedman [Ano06-28]. Frees [Ano06-39]. Freitas [Sar03]. Frequency [Cam01a, Cha02b, Nas02, Smy00]. Freund [Zie01-33, Zie04s, Zie04n, Moo03]. Frieden [Per04]. Friedman [Ano09m, Zie03c]. Friedrich [Ano07g]. Friendly [Zie04a, Zie01-53]. Friis [Ano03u]. Froidevaux [Zie01t]. Front [Ano00f, Ano00g, Ano00h, Ano00i, Ano01e, Ano01f, Ano01g, Ano01h, Ano02e, Ano02f, Ano02g, Ano02h, Ano03-52, Ano03-53, Ano03-54, Ano03-55, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano05-59, Ano05-60, Ano05-61, Ano05-62, Ano06-45, Ano06-46, Ano06-47, Ano06-48, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano08u, Ano08v, Ano08w, Ano08x]. Frontiers [Zie02-30]. FRS [Ano07l]. Function [Ano08m, Liu02, Lom05, Qia02, ZCC08, BF08]. Functional [Alt03, Ano06-41, CGHL01, Epi08, HPP01, Lu07, MBG07, SX07, dCGM05, Ano06-41]. Functionally [JD07]. Functions [AB04, Bed00, Chi01, JS07, SC00, IM02, Glo01]. Fundamental [Zie00p]. Fundamentals [Bab02b, Nku07, Yeh05]. Funk [Neu05]. Funkenbusch [AW05]. Futing [Pet03a]. Future [Bat08, Cha08a, CW08, Han08, Kaf00c, LI08, NM02, Old08, SBF^{+}08, Wil08a, Wil08b]. Fuzzy [Ano08t, Bar07a, Qiu08, Qiu08].

G [Ahm07b, Ano03v, Ano04-31, Ano05k, Ano05x, Ano06-35, Ano06-29, Ano06-32, Ano07j, Ano09h, Ano09a, Bai07a, Bar02, Boo04a, Bur01b, Bur09b, Cha01c, Cob04, Con01, Din05, Esm05, Fin05, Gan06, Gib01, Gra02, Hor09a, Huz01, Iye02, Kib07, Oli09b, Ped01, Pig06, Ros04, Ruk01, Rut05, Rut09, Sen05, Soy04, Wlu06, Zie00u, Zie00e, Zie00x, Zie01-43, Zie01j, Zie01-37, Zie02-31, Zie02b, Zie02-40, Zie02s]. Gabor [Lip07b, Nas02, Zie01-46]. Gad [Zie00-48]. Gail [Zie02z]. Gained [CG09]. Galaxy [Huz03]. Galbraith [Ano03-45, Ano06-30]. Galbreith [Bur08b]. Galecki [Hec09]. Galen [Law01]. Gallhofer [Ke09a]. Galli [Ano05s]. Galwey [Hec08]. Gamerman [Ahm08d, Kuo00]. Games [Lip07b]. Gamma [ASW^{+}06, BKG09, KMM08, BG06]. Gamma-Ray [ASW^{+}06]. Gao [Boo05a]. Garbolino [Soy09]. Gardiner [Sar00b]. Gareth [Sar02, Zie00-37]. Garmt [Cha05]. Garrett [Ano03-49, Ano05e]. Garthwaite [Ano03-39]. Gary [Sta04, Zie01j]. Gases [Nem02a]. Gasteiger [Zie04e]. Gastwirth [Zie01-44]. Gatelli [Che08a]. Gatsonis [Ye03, Zie00d]. Gauge [Boy01, Maz06a]. Gaussian [Bas00, BO09, LS05, Li09, LBH^{+}06, NM04, QWW08, Pic06]. GCM [RMGR09]. Gebhardt [Esp07]. Geert [Ano06-43, Ano07e, Avi01, Chi06]. Geladi [Qiu00, Boo05b]. Gelman [Boo04b, Gre04, Ye03, Ye05, Zie00d]. Gen [Zie02d]. Gene [Ano03-49, Ano05-53, Qu08]. General [Ahm08g, CHCFC04, Chi01, LLI05, RMGR09, SP07, ZTW07, IM02]. Generalizations [Sha04]. Generalized [CM08, CS01, DT02, DP02, DS06, Hey09, KM04, MBG07, MW05, MJH00, Mye08, Sen06b, Weh02, WLER06, ZS09, Zie02l, Zie02e, Cha03, Chi07, Gho05, Gho08a, Glo05, Lip05, Qiu05, Von03, Zie01s, Zie02-31]. Generation [Ano05-55, Bor00, McC04]. Genetic [Ano05-33, Ano05-37, MWJ06, ZC06, Zie02u, Zie02d]. Genetics [Zie02u, Ano03o]. Genichi [Alt06, Bar07c, Jia03]. Genome [Ano07e]. GENSTAT(R) [Che01b]. Gentle [Neu09b, Ahm08e, Ano05t, Bor00, McC04, Owe03]. Genton [Sen05]. geoENV [Ano06r, Zie01t]. Geof [Owe06]. Geoff [Ano06-40, Ano09f, Zie02j]. Geoffrey [Ano05-53, Ano07j, Gra02, Mor05a, O’N01, Sha02, Zie01-43, Zie02-31]. Geographically [Zie06]. Geohydrology [Rag01]. Geology [Ano05-47, Ano05h, Was04, Zie00q]. Geometric [Rob00a, Lin02]. Geophysical [Ano06-33]. Geophysicists [Mye05b]. George [Ano04l, Ano05k, Ano05-54, Ano05-31, Ano07p, Ano08q, Bri04, Bur02, Cab04, Kim00, Kok01, LaL01, O’N03, Oli03, Pig06, Ruk01, San08b, Sar00b, Zie01a, Zie01h, Zie01w, Zie02-41, Zie02p, Zie02-27, Zie02-47, Zie04y]. Geosciences [Ano05o]. Geoscientists [Zie05b]. Geostatistical [Ano05r, Smi01]. Geostatistics [Ano06r, Cha01b, Cob02, Mye01, Zie01t, Ano04p, Mye00, Sin09, Zie00-36, Ano05s, Cob00, Zie00r, Zie00q, Zie00s, Zie01u]. Geotechnical [Mye05a]. Gerald [Ano05l, Fri00, Neu05, Pry03, Sho03, Ste00b, Zim02]. Gerhard [Nas08, Zie02e]. Gertsbakh [Tor01]. Gettinby [Sar00b]. Geurts [Ano07g]. Ghiocel [Ano06n]. Ghosh [Zie00w, Zie01s, Zie00-35, Zie00c]. Gibbons [Ano04q, Zie03b]. Gibbs [Ode00a]. Giesbrecht [Rut05]. Gilbert [Ahm08f]. Gilchrist [Glo01]. Giles [Zie00t]. Gill [Mai05]. Gillespie [Hec09]. Ginevan [Zie04u]. Giovanni [Ano03-49, Zie02-33]. Giudici [San05]. Giulio [Boo09c]. Giuseppe [San07]. Give [LC00]. Given [GV03]. Givens [Owe06]. Glaz [Zie01-34, Zie02-39]. Glen [Con00, Goo02, Owe00, Ram00a]. Glenn [Ano03j, Ano05-43, Bla01, Maz00, She00]. Glidden [Gol06b]. GLIM4 [Ke06]. Global [BGW03, Che08a]. Glymour [Bur03]. Gnedenko [Sen00]. Go [Car07c]. Goegebeur [Laz05]. Goggin [Zie05b]. Gokhan [Ano07v, Ano08n]. Golberg [Oli05]. Goldstein [Gru08, Ano06w, Zie02-34]. Golyandina [Ruk02]. Gómez [Zie01t, Ano06r]. Gómez-Hernández [Zie01t, Ano06r]. Gonzalo [Liu06]. Good [Ano03f, Ano05-35, Ano06z, Ano06u, Cha00, Sin07, Zie01-30, Zie02-35, Zie02r]. Goodness [BDNN04, CS01, NM04]. Goodness-of-Fit [CS01, NM04]. Goos [Ke05]. Goovaerts [Car07a, Cob00]. Gordon [Ano06k, Ano06v, Che07, Zie01y, Zie01o, Ano03-36, Sar03, Sul01]. Gorman [Zie01r]. Goswami [Fot09]. Goupy [Ano09g]. Govindarajulu [Gho02, Kat08b]. Gower [Cha05]. Graham [Zie02o]. Grahn [Qiu00]. Grambsch [LZ02]. Grammar [Bai07b]. Granfeldt [Was04]. Grant [Zie02g]. Graphical [ACBJ09, Esp06b, PYC07, SNS07, Wil02, ZZ07, Bin01, Hec01, Slu03]. Graphics [Ano08o, Gun04, Bai07b, Gra00, Gun06, KS07a]. Graphs [Sea05, Ahm08b, Mar03]. Graubard [Zie01b]. Gray [Cab06, Ano05j]. Green [Mor05b, Ros05, Zie04-27, Kat04a]. Gregory [Cab03, Mye05a, Zie02g]. Greiner [Hly07]. Griffith [Sym01]. Grissom [Ahm06]. Groundwater [Gat02]. Group [Ano08s, Pet03a, VDLB08]. Groves [Ano05-50, Zie02-46]. Growth [Kuh03]. Gu [Ano06-41, Fre03]. Guang [Boo05a]. Guangbin [Ng08]. Gubbins [Mye05b]. Guide [AC02, AW05, Ano03f, Ano08j, Ano08e, Bai05, Cab08, Cha00, Fin05, Hec09, LSW09, McC00, Neu09a, Rum06, Sin07, Zie00-51, Zie04a, Ano06-41, Ano05-43, Che08c, Chi06, Goo02, Hol08, Ker02, LaL09, San03b, She00, Van04, Was01]. Guided [KD05, TGLG09]. Gumpertz [Rut05]. Gunst [Ano03-38]. Gunter [Hly07]. Gupta [Mor05b]. Gustafson [Cla04a]. Guy [Ano05i]. Gygi [Kal06].

H [Ano03-35, Ano03u, Ano03-39, Ano03-32, Ano04u, Ano05-41, Ano05e, Ano06-31, Ano06v, Ars05, Boo03, Bur04, Bzi04, Cab09, Cam01a, Cam01b, Cas05, Cha04, Che08b, Con00, Fle09, Hly09b, HG02, Kar07e, Ke08, Kim04, Lip09, Lu08, McC06a, Mor04, Mor05b, Mur07, Oli05, Owe06, Pic07, Sea09, Umb01, Wik03, Yeh07, Zie00-32, Zie00-34, Zie01-43, Zie01m, Zie02-31, Zie02-36, Zie02-51, Zie02-30, Zie02y, Zie02-34, Zie02z, Zie04o, Zie04p]. H.-H [Pic07]. H.-J [Zie02-30]. Haan [MS07]. Hadamard [DT02]. Hadi [Moo01, Ano07n]. Hadjicostas [Ano05-40]. Hafed [MM00]. Hair [Qu07b]. Hal [Ano06-42, Gre04]. Half [Gui08]. Half-Normal [Gui08]. HALT [Dog01]. Hamada [Hor09a, Alt01]. Hamed [Cam01a]. Hamilton [Das00]. Hamparsum [Ano04w]. Han [Esm06]. Hand [Ano07r, Ano07s, Ano07l, Ano05v, Zie00-29, Zie01i]. Handbook [Ahm08a, Ahm08h, Ano06e, Ano06-38, Fri01, Hem02, Jen02, Kus02, Liu07b, Zie02i, Zie02j, All01, Ano06n, Bar01, Bar07c, Hey01, Mor05a, Ahm07a, Ano03p, Ano03o, Ano04j, Ano04k, Ano05t, Ano05u, Ano06s, Ano07i, Ano08h, Ano09f, Nas01, Zie00u, Zie00x, Zie00y, Zie00z, Zie00w, Zie00v, Zie00t, Zie04i]. Hang [Qia07]. Hangi [Bar07a]. Hanneke [Ano03-37]. Hans [Ano04p, Qiu00, Zie00-36]. Harald [Bay02]. Hard [Zie02-32]. Hardeo [Mat01, McC06c, Mic05]. Hardin [Cha03]. Hardle [Ano07v, Hec05, Lor05]. Hare [Law01]. Harkness [Ano03k]. Harlow [Cho06]. Harman [Ahm08f]. Harney [Bur04]. Harold [Ano04g]. Harper [Mye01]. Harrell [Wal03]. Harrington [Con00, Liu07a]. Harry [Ale01, Ano06v, Ano09c, Dog01, Hor09a, Sen07]. Hart [Bar02]. Harvey [Ano06w, Ano06-27]. HASA [Dog01]. HASS [Dog01]. Hastie [Ano09m, Zie03c]. Hatcher [Ano06-37, Van04]. Haupt [Ano05-37, Ano05-37]. Havard [Pic06]. Having [HSNB09]. Hawkins [Umb01]. Hayes [Rom06]. Haykin [Qia02]. Hayler [Ano06-36]. Haym [Esm06]. Hazard [Nel00b]. Hazards [BCL^{+}09]. Head [LC00]. Heagerty [Ano03h, Ano05l]. Health [Ano03u, Ano05-56, Ano07e, Zie00-45, Zie00-44, Zie01g, Zie01b, Zie02-34, Ano04r, Mye09c]. Heavy [McC00]. Heckard [Ano05-30, Sea02]. Hedayat [Den00]. Hedibert [Ahm08d]. Heiberger [Utl05]. Heijden [Zie02b]. Heike [KS07a]. Held [Pic06]. Heldt [Bru00]. Helen [Zie00a]. Helge [Bor01, Wlu03]. Helsel [Gat06]. Henderson [Bai07a]. Henk [Vol05]. Henry [Kap03]. Herbert [Cas05, Eva01]. Hermann [Hly07]. Hernández [Zie01t, Ano06r]. Herrmann [Zie00e]. Hershberger [LaL01]. Hertzberg [Ano07r, Ano07s]. Hervé [Cut00]. Hesketh [Ahm08a, Fri01, Lip05, Zie02n]. Hess [Ano03-38]. Heterogeneous [Khu06]. Heteroscedastic [BG05, CR06]. Heumann [Ano09k, Ano09k]. Heuristics [Zie01-29]. Hicks [Zie00p]. Hidden [Nag06]. Hierarchical [GP06, Gel06, QW08, Sin05]. Hierarchically [BSC09]. Higdon [Ye03]. Higgins [Fot04]. High [Car07d, CG09, Dri06, LHB^{+}02, MZ02, QW08, RWZF04, Zie03a]. High-Accuracy [QW08]. High-Dimensional [CG09, Dri06, LHB^{+}02, MZ02]. High-Strength [RWZF04]. Highly [Ros05]. Hilbe [Cha03]. Hildebrand [Ano05j]. Hill [O’N03, Zie01h]. Hinde [Ke06]. Hinkelmann [Ano08f, Xu06]. Hinkley [Kus00]. Hirji [Ano08g]. Hiroshi [Mar06, Qiu05]. Hisashi [Ng05]. History [Ano07u, Boo09a, Zie00o]. Hjort [Ros05, Oli09b]. Hoaglin [Ano07h, Zie01-52, Ano07q]. Hoang [Liu07b]. Hoar [Zie01-49]. Hoboken [Mai05]. Hochberg [Zie00-34]. Hocking [Ano03w]. Hoehn [Gat02]. Hoerl [Law01, Zie02-42, Ano04m]. Hoeting [Owe06]. Hoffherr [Con00]. Hoffmann [Glo05]. Hofmann [Ano08n, KS07a]. Hogg [Ano09b]. Hohn [Zie00q]. Hoisington [Zie01m]. Holford [Ano05-57]. Holger [Ano04i, Ano05-34]. Holland [Ano05m, Utl05]. Hollander [Ram00b]. Holmes [Soy04]. Honour [Ano07l, Gho07]. Horizon [FL07]. Hosmer [Con02a, Zie01c]. Hou [Cab09]. Hougaard [Ken02]. Houlding [Cob02]. Houtkoop [Ano03-37]. Houtkoop-Steenstra [Ano03-37]. Howard [Sta00]. Howell [CL03]. Høyland [Zie04x]. Hsu [Sea01]. Hua [Ano03-42]. Huaguang [Qiu08]. Hubele [Zie02-27]. Huber [Ano09l]. Huber-Carol [Ano09l]. Hubert [Lip07c]. Huet [Ano04-27]. Hui [Con01, Zie02s]. Human [FR07]. Hundred [Zie02t]. Hunter [Pig06, Pig06]. Hurwicz [Lip07a]. Huzurbazar [Bre06]. Hwang [Nas02]. Hydrogeology [Cha01b]. Hydrologic [Wik03]. Hydrology [Ano08p]. Hypercube [BV01].

Hypercubes [CL07]. Hyperrectangle [MB06]. Hypotheses [Laz03a, Ano05-35]. Hypothesis [Ano05y].

Iacus [Bur09a]. Ian [Ano03-39, Gri07]. Ibrahim [Zie02s, Con01]. Icenogle [Ano03-43, Zie00-47]. Ideas [Ano05j]. Identification [LT09, SA08, VD07, dC02]. Identify [JP01]. Identifying [AB03, AL03a, BN01, GM01, Lan01]. Idiot [Bai05]. Idle [HWY02]. Ignored [MJH00]. Igor [Sen00]. II [McC06c, Zie01t, Ord07b]. III [Bur02, Jen09c, LM02, Liu09b]. ILL [PDDL02]. ILL-conditioning [PDDL02]. Illustrated [Hec01]. Illustration [BBJK04]. Ilya [Tor01]. Image [MM00, Qiu00, WLLZ07, Lor06]. Impact [Cla04a]. Imperfect [Jen09c]. Implementation [Nas02, Zie04p, Bar08c]. Implementing [Yat02, Ano03q, Gar00]. Importance [GCLH01, KL02, Pic05]. important [KT03]. Improved [Smy00, Ste09b, WT09]. Improvement [Ale02, Ann03, Bai07a, FW06, Rob02, Tob03, Zie02f, Bod02b, Bre01, Cah06, Fri00, Hof00, Umb01, Zie01-40, Zim02, vN02]. Improving [Law01]. Including [Gra00]. Incomplete [Boo00, NC08, Ye05]. Incomplete-Data [Ye05]. Incorporating [BC07, KCM00]. Independent [Mai07, Ray03b]. Index [ST02, FVB02]. Indices [Ahm08h, Zim00]. Individual [SJ02]. Induced [JD07]. Induction [Ahm08f]. Industrial [Ano08t, Ano08s, AK04, Bar07a, Edg04, GP01, Kaf01c, SBF^{+}08, VDLB08, Zie02-29]. Industries [Zie02h, Zie02h]. Industry [Ano05-43, Boo09b, Sar05, She00, Ano04k, Ano06x, KS07b, Pla01, Tob03, Zie02q]. Inertial [WM05]. Inference [Ano07s, Ano09m, Bur04, Bur09a, Gho05, JL07, Kun08, LLBN08, Li05, Mul01, RSS02, Sau02a, Sen02, Soy09, Ye05, Ahm08d, Ano03-39, Ano03y, Ano04q, Dok08, Kat08a, Kok01, LaL04a, MR00, Rob06, Zel01, Zie00f, Zie00x, Zie02-41, Nag06]. Inferring [ZZ07]. Influence [HBL03, Peñ05, BF08]. Influential [Coo00a]. Informal [Neu09a]. Information [Ano00j, Ano01i, Ano02i, Ano03-56, Ano04-37, Ano05-63, Ano06-49, Ano07-29, Ano08y, BC07, CG09, JW04a, LM04, ME04, Miy04, NW04, Nai07, SS07, Sud04, WB05, WM02, Yu07a, Zie00h, Mul02]. Ingram [Ahm08c]. Initial [LC00]. Innovation [Pig06]. Input [AK04, HSNB09, MBC00]. Inputs [DWG09, RCGD06]. Insight. [Zie00-27]. Inspection [Boy01, SC00]. Inspired [Ano05-33]. Institute [Zie04k]. Insua [Kat01]. Insurance [Ano08p]. Integral [JWT^{+}02]. Integrals [Yu07b]. Integrated [Gun03, RWH^{+}04, Sco01]. Integrating [QW08, Zie00-28]. Intelligence [Ano06j, Jen08, Zel05]. Intelligent [Ano05v, Zie00-29]. Intensity [BDNN04]. Intensive [She07]. Interaction [Glu06, PHSL02, XCW04]. Interactions [MS01, ZZ07, KT03]. Interdisciplinary [Hec01]. Interest [Zie02h]. Intergranular [RWZF04]. Interlaboratory [GB01]. Intermediate [Utl05, Hol01]. International [Hly09b]. Internet [Owe04, Ano05w]. Interobserver [Nag04]. Interplay [GLV02]. Interpolation [Woo00]. Interpolators [BGW03]. Interpretation [Ano03-45, Wlu06]. Interpreting [Goo02, McG02]. Interval [DM08b, JM00, LS00]. Interval-Censored [LS00]. Intervals [Chi01, DS00, HJMW04, HME08, KM04, KMM08, LLI05, NM02, WT09, BG06, IM02, MW05, Ahm04]. Intervention [MV05]. Intro [Sch04]. Introduction [Ano04e, Ano07i, Ano08f, BBS09, Cab06, Che08b, Chi07, Kus03, Law00, Ng06, Ray00, Rob00b, Rut00, Sar05, Sca04, Smi02, Var00, Yeh07, Zie02l, Ano05v, Ano05i, Ano05k, Ano05-48, Bay02, Geo00, Mai07, MS07, Neu09b, Ano03r, Ano03s, Ano04l, Ano05-27, Ano05y, Ano05x, Ano05z, Ano06t, Ano06u, Ano07j, Ano08i, Ano09h, Ano09g, Boo09c, Bur01b, Bur09b, Cha01b, Fah07, Fot04, Gan05, Gra02, Hec08, Hes04, Ke09b, Li06, O’N01, Oli04, Oli05, Pic09, Rob06, Sau02b, Sau05, Sau07, Sin08, Spi07, Wil02, Zie00-30, Zie01v, Zie01w, Zie04j]. Introductory [Smi06, Ano03t, Ano03u, Ano09i, Kat08a, McG07, San03a]. Intrusion [Bod02a]. Invariant [Weh04]. Inventories [Mye09b]. Inverse [Bas00, LN06, LH06, MBB05, Mye05b, NM04, SC04]. Investigations [Ano07r, Ano05-54, Jar09]. Investors [Ano03-27]. Ionescu [Zie00-43]. Irini [Ano03-45]. Irizzary [Ano03-49]. Irmtraud [Ke09a]. Irwin [Nel00a]. Isabella [Ye03, Zie00d]. Isaksson [Zie04a]. ISBN [Mai05, Zie01-52, Zie05b]. Iskander [Bab05]. ISO [McG02]. Isobel [Mye01]. Isotonic [Glu06]. Isotropic [ŽH09]. Issue [BBS09, Nai07]. Issues [Mai05, MNS04, Ano05-41]. Item [Sen06b]. Items [DM08b]. IV [Ano06r, Zie00d]. Ivan [Rob02]. Iverson [Ano08o].

J [Ahm06, Alt03, Ano03-45, Ano03s, Ano03o, Ano03h, Ano03k, Ano04v, Ano05v, Ano05-32, Ano05j, Ano05p, Ano05-56, Ano05-53, Ano05l, Ano05-50, Ano06k, Ano06-38, Ano06-41, Ano06-37, Ano06-29, Ano07k, Ano07r, Ano07s, Ano07t, Ano07g, Ano07i, Ano07l, Ano08m, Ano08q, Ano08n, Bar05, Bar01, Bod02a, Boo00, Boo09a, Bri05, Bru00, Car07a, Che03, Che01b, Che08a, Den00, Den04, Din05, Dob02, Eva01, Fin02, Fot04, Gri04, Gro00, Hap00, Hec01, Hin05, Hly09a, Hor09c, HI00, HG02, Huz03, Jar04, Jar09, Jen09c, Jos04, Kat04a, Kat09, Kia01, Kok01, Kus03, Law01, Laz03b, Lip01, Liu02, McC05b, Mic03, Mul08, Nas01, Nel01, Ng06, O’N03, Oli03, Oli04, Oli07, Oli08, Oli09d, Ped01, Pfa04, Pig06, Pla01, Qu07b, Rob00b, Rob05, Rob06]. J [Ros05, Rum06, Rut02, San08b, San08a, Sca02, Sea01, Sea05, Sin09, So01, Sym01, Sym05, Vuk00, Wal05, Weh02, Zie00u, Zie00e, Zie00-29, Zie00-49, Zie00-28, Zie01d, Zie01y, Zie01-33, Zie01-31, Zie01-46, Zie01o, Zie01h, Zie02-46, Zie02-30, Zie02l, Zie02g, Zie02-43, Zie04i, Zie04s, Zie04q, Zie04d, Zie05b]. Jaccard [FVB02]. Jaccard-Tanimoto [FVB02]. Jacek [Kaf03]. Jackie [Ano08o]. Jacqueline [Lip07c, Zie04-27]. Jacques [Ano09g, Zie02z]. Jaffe [Eva01]. Jaime [Ano06r, Zie01t]. James [Ahm08e, Ano03-27, Ano03-38, Ano05o, Ano05t, Ano05e, Ano05-50, Ano06-38, Bor00, Cah06, Cha03, Con00, Fot04, Gru02, Hes04, HG02, Kaf03, Kar07d, Kat04a, Ker03, LaL04b, LaL09, Mai07, McC04, McG02, Owe03, Pla02, Yat02, Zie00-30, Zie00h, Zie01x, Zie01e, Zie03a, Zie04v]. Jan [Ano06-27, Bar07b, Laz05, Zie04w]. Janacek [Sar02]. Jane [Ano03-45, Ker02, Zie02u, Zie04v]. Janet [Ano03k, Che07]. Janice [AC02, Law01]. Janis [Law01]. Janke [Rob06]. Jari [LH06]. Java [Ano08o, Hec07]. Jay [Ale02, Ano05g, Ano06-38, Mic01, Min02, Sau02b, Zie00-50, Zie02x, Zie04m]. Jayant [Ahm07b, Kuh05b]. Jean [Ami04, Ano04q, Ano05-29, Ano07e, Zie00r, Zie00h]. Jean-Louis [Ano07e]. Jean-Paul [Zie00r]. Jeff [Alt01, Loe07, Mai05]. Jeffrey [Ano06s, Ano06-40, Bai04, Das04, Smi01]. Jelke [Zie00h, Zie02b]. Jen [Ano05-54, Zie00j]. Jen-Pei [Ano05-54, Zie00j]. Jeng [DS00]. Jennifer [Ano05-29, Owe06]. Jensen [Zie05b, Ano05-49, Hua00, Mar03]. Jenson [Ahm08b]. Jerald [Joh03]. Jeremy [Car00]. Jerome [Ano09m, Zie03c]. Jerry [Zie00v, Zie05b]. Jesper [Nas04]. Jessica [Ano05-44, Ano05-30, Sea02]. Jesus [Ano06r]. Jewell [Ano05-57]. Jeya [Mit02]. Jia [KS07b]. Jia-Yeong [KS07b]. Jiahui [Ano03z, Bur07, Fot07b]. Jian [Kuh03]. Jian-Xin [Kuh03]. Jiang [Mye08]. Jianjun [Buk09]. Jianqing [Lu04a]. Jiju [Esp06a]. Jim [Ano06-38, Fle09, Zie02x]. Jiming [Mye08]. Jinfang [Gan06]. JMP [Ano09g, Ano08j, Zie04n, Ano05-28, Ano08j]. JMP(R) [Ano06-37, Jen09b]. Joachin [Qiu02]. Joan [Esp07, Weh04]. Jobe [Zie01d]. Jock [McC06a]. Joel [Hey06]. Joglekar [Gol04]. Johan [Laz05]. Johann [Zie04e]. Johannes [Ano03-37, Ano09b, Bre01, Hof00, Oli09d]. Johansson [Nem03]. John [Ahm06, Ano03-33, Ano03-40, Ano04t, Ano05-47, Ano05-39, Ano05-28, Ano06g, Ano06-39, Ano06-29, Ano07h, Ano07l, Ano08j, Ano08h, Bak01, Bod02b, vB02, Bre05a, Bru00, Cha05, Den00, Edg04, Geo03, Gho08a, Glo05, Gun04, Gup00, Hey09, Hly09a, Ke06, Maz06b, Mye05a, O’N05, Rav04, Smi06, Was02, Zie01r, Zie01-52, Zie01-32, Zie02-28, Zie02-46, Zie02a, Zie04-27, Zie04c, Kaf01b]. Johnson [Ano06-34, Ano06-29, Bar08c, Bur02, LaL04b, Nem00]. Jointly [MBDD09]. Jolivet [Ano04-27]. Jolliffe [Ano03-39, Ano03-31]. Jonathan [Ano05-56]. Jones [Ano04-31, Che01b, Ano03-39, Sca04]. Jong [Ano05-45, Zie00u]. Jordaan [Gri07]. Jørgen [Was04]. Jos [Ano06-39]. Jose [Boo09b, Rao01, Ano06r, Neu09b, Zie01z]. Jose-Maria [Boo09b]. Joseph [Ano03-36, Ano04y, Cha03, Con01, Hin06, Lar08, Qu07b, She07, Str01, Zie01-44, Zie01-34, Zie02-39, Zie02s, Zie04k, Zim01]. Journalists [Ano06-38]. Journel [Ano06p]. Journey [Kal03, Sen05]. Joy [Qia01]. Jozef [Ano05-55, Laz05]. Judah [McC02]. Judea [Zel01]. Judgment [Ker02]. Judith [Ano05-29, Ano06-39, Gru02, Zie00e]. Jugulum [Alt06, Jia03]. Juha [Ano06-42]. Julian [Cha01c, Rob05]. Julien [Sca05]. July [Kaf01b]. Jump [JQ09, Lor06, Qiu04]. Jun [Ano04-28, Hes02]. Junbin [Boo05a]. June [Kaf01b]. Jupp [Mai08]. Juran [Zie04k]. Jureä [Oli07].

Kaas [Car07a]. Kai [Ano07f, Kuh03, Kuh05b]. Kai-Tai [Kuh03]. Kaipio [LH06]. Kalbfleisch [Geo03]. Kales [Joh00]. Kalman [Li09]. Kalnay [Wik05]. Kamel [Neu08c]. Kaminskiy [Olw01a]. Kanefuji [Ars05]. Kannan [Din05]. Kantardzic [Ano05-55, Ano03m]. Kanti [Mai08]. Kantz [Ano05-34, Cam01b]. Karen [Esp06b, Zie01-36]. Karian [Weh02]. Karim [Ano08g]. Kariya [Qiu05]. Karl [Ann01]. Karlaftis [Cob04]. Kass [Ye03, Zie00d]. Katherine [Neu09b]. Kathleen [Hec09]. Katz [Boo09a]. Kay [Zie00-49]. Kayihan [Neu08a]. Kazunori [Sin06]. Kedem [Ray03a]. Kee [Ano06h]. Keenan [Ano03-34]. Keiding [Ano07u]. Keith [Ano05-56, Gun03, Qu07a]. Keki [Ano03-48, Zie01-55, Zie02-49, Zie04z]. Keller [Ano08l, Sho03]. Keller-McNulty [Ano08l]. Kemp [Ano06-34]. Kempson [O’N01]. Kempthorne [Ano08f, Xu06]. Ken [Car07b, LaL04a]. Kenneth [Ano03y, Ano05x, Ano08m, Gru05, Hem02, Liu02, MR00, Zel00, Zie00p, Zim04]. Kenward [Ano04-31]. Keppel [Mor05a]. Kept [Dok09]. Kernel [GCS06, Qiu04, Bre05a]. Kernels [SS01]. Kettaneh [Nem03]. Kettaneh-Wold [Nem03]. Kevin [Zel05]. Key [Tom07]. Keyfitz [Ano06-42]. Khaled [Cam01a]. Khattree [Ano04k, Zie00b, Zie01-28]. Khuri [Ye00, Ano03g, Ano07o]. Kim [Ahm08g, Ano05-53, Ahm06]. Kim-Anh [Ano05-53]. Kindle [Dok09]. Kinetic [Atk05]. Kinnaert [Bar07b]. Kinzelbach [Gat02]. Kirby [Lin02]. Kirk [Zie04j]. Kishor [Hly07, McC03b]. Kitanidis [Cha01b]. Kitchens [Bri05]. Kjell [Rum00]. Klaus [Ano08f, Jen02, Xu06]. Klefsjö [Bar04]. Kleiber [Zie04t]. Klein [Rav04, Ano03v, Ano06-32]. Kleinbaum [Ano03v, Ano06-32, Key01]. Klieber [Mye09a]. Klir [San08b]. Klugman [Ano06v]. Knott [Lip01]. Knottnerus [Con04]. know [Bri01]. Knowledge [Ano04w, Ano05-55, RMGR09, Ano03-37]. Knudsen [Boo07a]. Koch [Ano05-40, Zie01j]. Koenker [Lip06b]. Kokoska [Zie01-38]. Kolassa [Ano07t]. Koller [Ano05-43, Bla01, She00]. Kolmogorov [DGS07]. Kolmogorov-Smirnov-type [DGS07]. Koning [Fle09]. Konstantinos [Ray03a]. Kontoghiorghes [Ano08h]. Koopman [Ano06-27, Hin05]. Korb [Zel05]. Korn [Ano05-54, Zie01b]. Koronacki [Kaf03]. Kotz [Oli08, Sha04, Zie00m, Zim00, Ahm08h, Ano06-34, Zie04t]. Koutras [Che02a]. Kovar [Gat02]. Kraaikamp [Yeh07]. Kragh [Ano07u]. Krause [Ano03-46, Ano06-44, Zie01-51, Zie02q]. Kriging [Jos06b, LS05, Woo00]. Krishnamoorthy [Ahm07a]. Krishnan [Fot07c]. Kristine [Ano09c]. Krivtsov [Olw01a]. Kroese [Boo08, Den06]. Kroonenberg [Esm09a, Esm09b]. Kruse [Slu03]. Krzanowski [Jar09, Rob00b, Zie01-31]. Kuehl [AW01]. Kulkarni [Ahm08f, Huz01]. Kuncheva [Bag05]. Kung [Ano03h, Ano05-45, CL03, McC05a]. Kung-Jong [Ano05-45]. Kung-Sik [CL03]. Kung-Yee [Ano03h]. Kunio [Li06]. Kuo [Gri04]. Kupper [Key01]. Kurata [Qiu05]. Kurowicka [Car07d]. Kurz [MM00]. Kvam [Lu08]. Kyriakdis [Ano06p].

L [Ahm08h, Ano03-49, Ano03h, Ano03-43, Ano03-38, Ano04y, Ano05p, Ano05-56, Ano05-54, Ano05-37, Ano06-34, Ano09c, Bai04, Bak00, Boo00, Bur04, Cah06, Cha02b, Che08b, Cho06, Cob04, Edg04, Esm05, Geo03, Hin06, HI00, IM02, Kar07b, Key01, Kus02, Kus06, LaL01, LaL04b, McG02, Nem02b, Nem02a, Nem03, Oli09b, Ped01, Qu07b, Rao00, Rav04, Rut05, Sau02b, Sca02, She07, Woo00, Yat01, Yeh07, Yu07b, Zie00u, Zie00m, Zie00-30, Zie00-47, Zie00-27, Zie00-32, Zie00-39, Zie00h, Zie00k, Zie01i, Zie01-44, Zie01-49, Zie01b, Zie01-45, Zie02-46, Zie02-41, Zie02-29, Zie02-40, Zie04-27, Zie04y, Zie04m, Zie04d, Zie04g, Zie05b]. Laboratory [Boo04a, Hor06]. Labs [Cha01a]. Lachin [Ano03-33, Zie02a]. Lack [Bed00]. Ladiray [Zie01-35]. Lafortune [Ano09h]. LAGO [ZSC06]. Lahiri [Hut04]. Lai [Ano07f, Hor07, Ano09d]. Laird [Ano05e]. Lake [Zie05b]. Lamballais [Ano07g]. Lambda [Weh02]. Lamberton [Ke09b]. Lamprecht [Cah06, McG02]. Lan [Che07]. Lange [Gru05, Zel00, Zim04]. Langer [Gho03]. Language [Ano06-40]. Lapeyre [Ke09b]. Laplace [PS00]. Large [Ano05-58, Ano07g, GLM07, KS07a, KT05, Xu09, Zie00k]. Large-Eddy [Ano07g]. Large-Sample [Zie00k]. Large-Scale [GLM07]. Larose [Ano05-55, Boo07b, Mag09]. Larry [Ano06-37, Bri05, Fot07a, Kal03, Rut00, Sau06, Sym01, Van04, Zie04b, Zie05b]. Lars [Ano03r]. Larsen [Ano05w]. Latent [Ano09j, Lip05, Lip01]. Latin [CL07, MB06]. Latouche [Bur01b]. Lattin [Kat04a]. Laurens [MS07]. Lauritzen [Ped01]. Lavorgna [Zie02-40]. Lawless [Joh03]. Lawrence [Das00, Lip07c]. Lawson [Ano05-56, Bod02b, Ros04, Zie01-42]. Lawyers [Zie02-44]. Layne [Sym01]. Layout [FKQS01]. Layth [Bel01]. Leading [Ano04m]. Leaps [FW00]. Leardi [Ano05-33]. Learning [Ahm08f, Ano03v, Ano06-32, Ano09m, Bur09b, Lip07b, Den06, Nea07, Zie03c, Ano08k, Hol08, Lor01]. Least [Bor01, MJH00, Qiu05, BG06]. Ledolter [Ano09b, Hof00, Oli09d, Bre01]. Lee [Ano03-40, Ano05-28, Ano08j, Ano09g, Cab06, Gho08a, Zie04n, Cas05, Cha08b, Oli03]. Lefevre [Din05]. Lehman [Ano06-37, Ano05-28, Ano08j]. Lehmann [Dok09, Zie00k]. Lehtonen [Ano05-38]. Leland [Bai07b]. Lele [Weh04]. Lemeshow [Con02a, Zie00-40, Zie01c]. Lemmon [Hin06]. Lemuel [Hus08, Zie04l]. Length [Kva08, LPP00, RWZF04]. Lenz [Zie02-30]. Leonard [Sea01]. Leonhard [Pic06]. Leonid [Lip07a]. Léopold [Hec05]. Lepkowski [Ano05-50]. Lessler [Ano05-29]. Letter [Ans03, GD04, GL09a, MLC09, Pet03b, PCdC04, Rya08, dCGRB08]. Level [But04, CL09, CW02, CLY04, LL03, LLY03, MWJ06, Mor00a, XCW04, Hly09b]. Levels [Xu02, dCGM05, Oli09c]. Leveraged [BMS09]. Levin [Ano04y, Nel00a, Zie02-44]. Levine [Zie00-38, Zie01-45]. Levy [Zie00-40]. Lewi [Zie00u]. Lewis [Zie00-37]. Leyland [Zie02-34]. Li [Ye05, Hen01, McC05a]. Liang [Ano03h, Boo05a, Nas02]. Liangyue [Lu04b]. Liao [Pet03a]. Lid [Ros05]. Lidia [Rob02]. Life [Ano05-52, Kun08, McC06b, MEH09, PM03, Pas06, WW00, ZM06, Zie04-27, Bai04, Ahm07b, Ahm08c, Ng08]. Life-Data [WW00]. Lifetime [Joh03]. Likelihood [CM08, DS00, LS05, Sau02a, SR04, YW05, Zha06, Pan00, Sau02a]. Limit [BG04, Jos06b, Pas03]. Limited [Ano05-42]. Limits [FG01, KL02, KM04, RS05]. Limnios [Ano08l, Ano09l, Zie00-43]. Linda [Hil00]. Lindner [Ano09e]. Lindsey [Kia01, McC05b, Oli04]. Linear [Ahm08g, Ano03w, Ano07j, Ano07i, Ano09k, AB04, Bed00, CR06, CA06, CCD09, Coo00a, DP02, DS06, Fam03, Gho08a, Glo05, Gol06b, Gra02, Gru08, HBL03, Hey09, LBHB04, MW04, Mal05, MBG07, Moo03, Mye00, Mye08, Peñ05, Qiu04, Rob06, Sen06b, SX07, Von03, WR02, WLER06, Ye00, YW05, ZDCS03, Zie00o, Zie01s, Zie02-31, Zie02l, Zie02e, Zie02k, ZTW07, Avi01, Bor01, Bur01a, Hec09, Mye08, Oli03, Qu07a, Rob05, Rut09]. Linfield [Ano03-51]. Ling [Cab00]. Linn [Fre00]. Linoff [Ano06k, Zie01y, Zie01o]. Liquids [Nem02a]. Lisa [Ano05-54, Cho06]. Littell [Moo03, Zie04n, Zie01-33]. Little [Hor09c, Laz03b, Ann03, Zie02-46]. Liu [Ano05-54, Hes02, Qiu08, Zie00j]. Livingston [Ano06-38]. Lloyd [Ano05l, Zie02f, Hap00]. Loader [Pan00]. Local [CHCFC04, DWG09, HBL03, PW01, Qiu04, TGLG09, Pan00]. Localization [KLM08]. Location [ASW^{+}06, CHvdW06, MZ02, MEM05]. locations [BG06]. Log [MEM05]. Log-Location-Scale [MEM05]. Logic [Ano08t, Bar07a]. Logistic [Con02a, GLM07, Gol06b, Ano03v, Zie00-31]. Logistics [Boo09c]. Logit [Bar05]. Lognormal [BG04]. Lohr [Zie00-39]. Lomax [Fin05]. Lon [McC07]. Long [Whi04]. Long-Memory [Whi04]. Longford [Ahm07c]. Longitudinal [Ano03h, Ano05e, Ano06q, Ano06-39, Avi01, Bur06, Gho03, Lip05, She07, Ano06-39, San03b]. Longnecker [Kim05]. lonosphere [RMGR09]. Look [BGD05]. Loop [VD07, dC02]. Lopes [Ahm08d]. Lopuhaa [Yeh07]. Lora [Hor09c]. Loss [FFG06, ZCC08, Ano06v]. Lot [But04, Van02]. lots [LCdC06]. Louis [Ano07e, San08a, Zie01f]. Louise [Eva01]. Lovelace [Zim00]. Low [MY08, QW08]. Low-Accuracy [QW08]. Low-Rank [MY08]. Lowen [Sca07]. Lucy [Fah07]. Ludmila [Bag05]. Ludwig [Zie02e]. Ludwik [MM00]. Lugosi [Lip07b]. Lui [Ano05-45]. Lumelskii [Sha04]. Lumley [Ano05l]. Lun [Gup00]. Lundström [Kar07c]. Lunneborg [Lou01]. Lunze [Bar07b]. Luu [Zie00-37]. LuValle [Din05]. Lyberg [Ano03r]. Lyle [Zie02u]. Lyman [Ano05j, Kar07b, Kim05]. Lynne [Law01].

M [Ami04, Ano03-27, Ano03-36, Ano03-33, Ano03o, Ano03-43, Ano03-32, Ano04-27, Ano05i, Ano05-44, Ano05s, Ano05-55, Ano05-30, Ano05p, Ano05-40, Ano05-54, Ano05e, Ano05-29, Ano05-50, Ano06-42, Ano06n, Ano07r, Ano07s, Ano07l, Ano08k, Ano08p, Ano08i, Bab05, Bre04, Cab06, Cam02, Car07a, Cha03, Che01b, Che08a, Esm09a, Esm09b, Fri00, Gan05, Gho09, Gol04, Gru02, Kar07f, Kat03b, Kat04a, Ke08, Ker02, Kib07, LZ02, MM00, Maz06a, McC00, McG07, Mit02, Mur07, Nag04, Nea07, Neu05, Neu09b, O’N01, O’N03, Oli05, Oli09c, Pla01, San08b, Sau05, Sea02, Sen06a, Soy04, Sta00, Ste04, Umb01, Utl05, Wal05, Yeh07, Zie00u, Zie00e, Zie00-45, Zie00-49, Zie00-47, Zie00-41, Zie00h, Zie00-38, Zie01-39, Zie01z, Zie01h, Zie01-48, Zie01-45, Zie02-46, Zie02t, Zie02a, Zie02b, Zie02-40]. M [Zie04y, Zie04j, Zie04r, Zie05b, Zim02]. M.-A [Ano04-27]. Mac [Ano03-51]. Machine [Bur09b, Den06, MSM03, Nea07, Neu01]. Machines [Pla02]. MacKay [McC06a]. Macro [Zie00-41]. Maddala [Zie00x]. Made [Ale02, Zie00-41]. Magni [Bay02]. Mahalanobis [Jia03, WKT^{+}03].

Maindonald [Gun04]. Maintenance [Hey06, MSM03, Ano03-50, Tor01]. Makes [Esp07]. Making [Ano05-43, She00, Bla01, Car07b, Zie02-33, Jar04, Zie02-32]. Malinowski [Ano03n]. Mallick [Soy04, Zie01s]. Mallows [Bri06, Buj06, Efr06, Hub06, Lan06]. Malvin [Sca07]. Man [Con01]. Management [Ale01, Ano06-36, Hly09b, Smi01, Wei03, Ano06-35, Joh00, Kuh05b, Qia07, Zie02-43]. Manager [Ano03f, Sin07]. Managers [Zie01-45, Zie04g, Ano05j]. Managing [Zie01x]. Mandallaz [Mye09b]. Manly [Ano05-32, Zie02-43]. Mannering [Cob04]. MANOVA [Gho05]. Manufacture [Ano03-36]. Manufacturing [AS01, AL03a, BLC01, Buk09, Che06, McC06a, MNS04, SA08, ZZ07, ZDCS03, dC02, Gol04]. Many [LP06]. Mapping [Ano05-56, Qu08, Smi01]. Marc [Ano06-43, Sen05]. Marcel [Bar07b]. Marchette [Bod02a, Sea05]. Marcia [Rut05]. Marco [Kat04b, Por02, Was05]. Marcoulides [LaL01]. Marcus [Zie01d]. Mardia [Mai08]. Margaret [Ano05-51, Mye00, Zie01u]. Margins [Bar08a]. Margo [So01]. Mari [Ano04r]. Maria [Boo09b, Ng09, Rao01]. Marian [Zie01-36]. Marianna [Sha04]. Mario [Bai04, McC06c, Mic05]. Marjorie [Ano03-43, Zie00-47]. Mark [Ano06o, Olw01a, Sen06b, Yat01, Zie00n, Zie01-49, Zie01-45]. Marketing [Ano05-58, KT05, Oli09d, Ano04z, Kib07]. Markets [Ano03-27]. Markos [Che02a]. Markov [Hly07, Mag09, Nag06, Pic06, Ruk01, BLC01, Jon04, LHB^{+}02, MR01, Ahm08d, Kuo00, Ode00a]. Markovian [Pic05]. Marlene [Lor05]. Marshall [Ahm08c]. Martens [Bay02, Bay02]. Martijn [Ano06h]. Martin [Ano05-29, Ano07l, Con02b, KS07a, Lip01, Neu05, Zie00h, Zie02-45, Zie06, Ano06t]. Martinez [Huz03, Kus02, Kus06, Kus02, Kus06]. Martingale [BE03]. Martingales [Yu07b]. Martz [Hor09a]. Marvin [Cab09, Zie04x]. Mary [Ker02]. Mask [ASW^{+}06]. Masking [CL05]. Mason [Ano03-38, Edg04]. Massart [Zie00u]. Massive [JL07, Mai01]. Mastering [Zie01y]. Matching [Kar07f, Ano04x]. Materials [Qiu02, Ste09b, Zie03a]. Mateu [Zie04q]. Math [Ano07k]. Mathematical [Ano06-42, Ano08l, Boo09b, Cab08, Cha01a, Geo00]. Mathematics [Mam09, Zie02-51, Boo09a, Nku07]. Mathew [Ye00]. Mathieu [Zie00-37]. Matignon [Jen09a]. MATLAB [Lip07c]. MATLAB(R) [Kus02, Kus06]. Matrices [DT02, Li05, Lip07c, MY08]. Matrix [Bur01b, Cha09, HMT08, Hub01, PP01a, PP01b, RW01, Ahm08e, Kar07d]. Matroid [Che06]. Matter [Ano00a, Ano00b, Ano00c, Ano00d, Ano00f, Ano00g, Ano00h, Ano00i, Ano01a, Ano01b, Ano01c, Ano01d, Ano01e, Ano01f, Ano01g, Ano01h, Ano02a, Ano02b, Ano02c, Ano02d, Ano02e, Ano02f, Ano02g, Ano02h, Ano03b, Ano03c, Ano03d, Ano03e, Ano03-52, Ano03-53, Ano03-54, Ano03-55, Ano04a, Ano04b, Ano04c, Ano04d, Ano04-33, Ano04-34, Ano04-35, Ano04-36, Ano05a, Ano05b, Ano05c, Ano05d, Ano05-59, Ano05-60, Ano05-61, Ano05-62, Ano06a, Ano06b, Ano06c, Ano06d, Ano06-45, Ano06-46, Ano06-47, Ano06-48, Ano07a, Ano07b, Ano07c, Ano07d, Ano07y, Ano07z, Ano07-27, Ano07-28, Ano08a, Ano08b, Ano08c, Ano08d, Ano08u, Ano08v, Ano08w, Ano08x]. Matters [Dra04]. Matthew [Cob04]. Matthias [Zie01l]. Maura [Zie01j]. Maurer [Wei03]. Maximin [SdHSV03]. Maximization [JW04a, LM04, ME04, Miy04, NW04, Sud04]. Maximize [AB03]. Maximum [YW05]. Maxwell [Ano04g, Dra04]. May [Zie04f]. Maynard [Ano03-37]. McBean [Zie00-46]. McBride [Ano06-35]. McCabe [Zie04y, Ano04l, Zie01w]. McClish [Ano03-42]. McConnell [Kuh05a]. McConway [Che01b]. McCuen [Ano03-32, Ke08, Wik03]. McCulloch [Von03, Gol06b, Kib07]. McDonald [Mai05, Mam09]. McKnight [Neu09b]. McLachlan [Ano05-53, Sha02]. McLean [Dog01]. McNulty [Ano08l]. McPherson [Goo02]. McQuarrie [Cab00]. McShane [Ano05-54]. Meadows [Zie01x]. Mean [BG04, CE02, HZ05, ÖBMT02, RS06, TH05, UL07, Van02]. Means [Esp06b, MWJ06, ND02, SC04, Pfa04]. Measure [JW04b, Cla04b]. Measurement [BMS09, Bur00, Cha01c, CR06, Gup00, Ste00a, VD08, Zie02h, Zie04u, Ars05, Cla04a]. Measurements [Kva08, LWS03, Bar08a, Ken03, Kia01]. Measures [Car07a, GP06, Gho05, Gol06b, Ahm08h, Nag04, Yu07b]. Measuring [LBHB04, Aba01a]. Mechanical [Qia01, Ann03, Tob00]. mechanics [MR01]. Mechanisms [Lip07a]. Media [LHB^{+}02]. Medical [Ano04o, Ano05-51, Ano06-40, Gol07, Liu04, Zie02-33, Zie02n]. Medicine [Ano03-42, Zie00a]. Meeker [DS00]. Meer [Hly07]. Meester [Yeh07]. Megavariate [Nem03]. Mehmed [Ano03m, Ano05-55]. Meier [Ano05-46]. Melania [And08]. Melvin [Ano03-46, Ano06-44, Zie01-51]. Memorabilia [Ano07f]. memoriam [Kaf01b]. Memory [Whi04]. Mendenhall [Kar07b]. Meng [Ye05]. Mesbah [Ano07e]. Messy [Bur02]. Métais [Ano07g]. Metcalfe [Baj00]. Method [AS01, CHCFC04, Chi01, Den06, Esp06b, For08, HG06, HWY02, JTW00, JW04a, LM04, ME04, Miy04, NW04, Sud04, ZS09, ŽH09, IM02, LCdC06, Boo08, Che08b, Zie01-35]. Methodologies [Che02b, Maz02, Sar06]. Methodology [Ano07o, Kaf01c, Ano05-50, McG02, Rao00, Zie02-36]. Methods [Aba01b, Ahm04, Ano03m, Ano03-40, Ano03-41, Ano03-42, Ano03j, Ano04r, Ano04y, Ano05-46, Ano05-33, Ano05-41, Ano05-57, Ano05-38, Ano06-43, Ano06-35, Ano06q, Ano06u, Ano07r, Ano07t, Ano07e, Ano08m, Ano08l, Ano08e, Ano09l, BBJK04, Bag05, Bin01, Boo07b, Bre05a, BN01, Bur01b, Cha00, Che01c, COR08, Cob04, Con01, Gan06, Gho05, Glo06, Gol04, Gol06b, Hly09b, Hut04, HN03, Joh03, Kar07a, Kat09, Ken03, KD05, KCV02, KMM08, Kuh05b, Kus00, LaL01, Lit01, Liu02, McC03a, McC02, McC06c, MBC00, Mic05, Mye09c, Nem08, Nem00, Ng05, Oli07, PHSL02, Qia00, Ram00a, Rom06, San07, Sar03, Slu03, Smy00, Soy04, Ste04, Was01, Was02, Won04, ZM06, Zie00-45, Zie00-33, Zie01-41, Zie01-42, Zie01-40, Zie01f, Zie02v, Zie02g, Zie03b, Zie04x, Zie04h, ZTW08, Ahm07b]. Methods [Ano03k, Ano04s, Ano05-32, Ano05-31, Ano06z, Ano08k, Bab02b, Bar09, Bas02, Bor00, Gar00, Gho02, Gru08, Hey06, Hin05, Key01, Kim00, Kim05, Lar08, Lor01, Lu04a, Mag00, McC04, Nas04, Ram00b, Ran03, Sca04, Sea01, Weh02, Wik03, Wil05, Yat01, Zie00y, Zie02-35, Zie02a, Zie02z, Zie04s, Ano03w, Ano03x, Ano05-29, Ano07l, Kat03b]. Metrics [Mai01]. Meulman [Lip07c]. Meyer [Ker02]. Mi [Esm06]. Micah [Mai05]. Michael [Alt01, Ano03u, Ano03-29, Ano04-31, Ano05v, Ano05-36, Ano05w, Ano05-54, Ano06k, Ano06-36, Ano08e, Bar07b, Bar08c, Bre02, Din05, Gru08, Hor09a, Ke07, Kim05, Kus03, Lin02, Mai05, Ng06, Qia05, Sar05, Str01, Was01, Woo00, Zie00-29, Zie00q, Zie01i, Zie01y, Zie01-53, Zie01o, Zie01-37, Zie02-44, Zie04u]. Michalidis [Bur09b]. Michele [Zie00-41, Bur00]. Michiko [Sin06]. Mick [Ano05-29, Ano05-50, Zie00h]. Mickey [Ami04]. Microarray [Ano03-49, Ano05-54, Ano05-53, BH08, BCBS08a, BCBS08b, Gol08, HB08, Jon08, KL08]. Microarrays [Ano05-53, Boo07a]. Microsimulator [MBB05]. Microsoft [Sho03]. Microsoft(R) [Jos04, Zie00-38, Zie01-45, Zie04g]. Microstructures [Qiu02]. Mielke [Ano08m, Liu02]. Miguel [McC06c, Mic05]. Mike [Zie00d]. Mikel [Ale01]. Mikhail [Ano09l]. Militino [Ng09]. Millard [Lum01, Zie02q, Ano07x]. Miller [Esm05, Zie04v, Ano03-44, Cab03, Zie04v]. Milliken [Bur02]. Million [KS07a]. Min [Hor07]. Mind [Ano05-30, Sea02]. Minefields [WR02]. Miner [Sin08]. Mineral [Ano06p]. Miner^{T M} [Jen08, Jen09a]. Mines [WR02]. Ming [Con01, Gri04, Zie02s]. Ming-Hui [Con01, Zie02s]. Minghui [Zim01]. Minimal [RHOYL06]. Minimizers [PW01]. Minimum [DT02, RWZF04, Sco01, KT03]. Mining [Ano03m, Ano03l, Ano04w, Ano05-55, Ano06k, Ano06s, Ano09m, Boo07b, Cab04, Cha09, Jen08, Jen09a, JL07, KD05, Mag09, Mul04, Sin08, Zie01o, Zie02c, San05, Slu03, Zie00i, Zie01y]. MINITAB [Bai07a, HI00]. Minquez [Boo09b]. Misclassification [Cla04a, JL07]. Mises [JLSZ08]. Missing [Fok07, Laz03b, Ahm07c, Neu09b]. Mission [Din05]. Mission-Critical [Din05]. Misspecification [ACBJ09, Pas06]. Misused [Eva01]. Mitchel [Ano03v, Ano06-32]. Mitchell [Zie02z]. Mitra [Bur09b, Mul04]. Mitsuo [Zie02d]. Mixed [Avi01, Che02b, Gho05, Hec08, Hec09, Khu06, LLI05, Mye08, Qu07a, Von03, Xu02, Ye00, ZDCS03, Zie00a, Zie01z]. Mixed-Effects [ZDCS03, Zie01z]. Mixture [DP02, KCV02, PHSL02, PDDL02, Sha02]. Mixtures [Ano07p, SS01, Zie00g, Zie02-28]. MLwiN [Ano05-56]. MODA [Zie00-32]. Modarres [Olw01a]. Model [ACBJ09, Ano09j, AL03b, AK04, AB04, Atk05, BK04, BFL02, BC07, Cab00, CS02, CL05, FJ08, LN00, Oli09b, PM03, Pas03, Qu07a, RdC05, RMGR09, Sha04, SHY04, ST02, WLER06, ZH06, Zie00-32, Jon04, PDDL02, LZ02, Ano03y, MR00, Oli09b, Sin09]. Model-Based [Sin09]. Model-Checking [AB04]. Model-Oriented [Zie00-32]. Model-Robust [LN00, PM03, RdC05]. Modeling [Ano04z, Ano05-42, Ano06i, BBS09, Bla01, Boo09b, Buk09, Cha08b, Cob02, FM04, FR07, Gel06, GLV02, GM01, Gri04, Hly07, Huz00, IM05, Ket00, LH02, Lip05, Mor00a, Olw01b, QW08, Qiu08, RWZF04, San08b, Sau02a, Sco01, Sen07, Sen06a, She09, Smi02, Wal03, Wik05, WLLZ07, Ye05, Ano05r, Ano06-31, Bur01b, Fin05, Lew04, Zie02k, Ano03z, Bur06, Bur07, Fot07b, Huz01, LZ02, Owe04, Wik03, Zie02-33]. Modelling [Ano04j, Che01b, Gat02, Glo01, Hec08, Ke06, Oli04, Zie02-34, Zie02e, Car07d, Dob02, Gan02, Hil00, Rob00b, Ros04, Sho04, Wil02, Ano04n, Ano04o, Bur00, Lu04b]. Models [Ahm07b, Ahm08g, Ano03m, Ano05-52, Ano06y, Ano06-28, Ano06v, Ano07l, Ano09l, Ano09k, AB04, Avi01, Bar05, BE03, BBP^{+}07, BCL^{+}09, Bed00, BV01, Bor01, Bre06, Bur01a, CHCFC04, Che02b, Chi07, CCD09, Con05, CG09, DWG09, Das07b, Das07a, DS06, Esm06, GP06, Gho08a, GB01, Glo05, Glu06, GV07, Hec09, Hey06, Hly09b, Hly09c, Hua00, Huz00, Joh03, Kat03a, Ke07, Ker03, Khu06, KM04, Kuh03, LHB^{+}02, LS05, LLI05, LBHB04, LBH^{+}06, Lip01, MBG07, McC03a, McC02, McC06c, Mic05, MMM08, Mye08, Oak09, OH05, QWW08, RCGD06, Ray03a, RLR06, RSB09, RSS02, RF05, Rob05, Rob06, Rut09, Sán06, Sen06b, SX07, Slu03, ST06, SB04, SP07, SMB01, TG00, WCT09, WLER06, YS05, Zel01, ZZ07, ZDCS03, Zie00a, Zie00-43, Zie01z, Zie04x, Ano03w]. Models [Ano06w, Ano06-27, Boo07b, Car07a, Dog04, Fam03, Fot01, Fre03, Gho05, Gol06b, Hey09, HR05, Lip05, Lor05, McC03c, Moo03, Nag06, Qu07a, Rao01, Sen02, Sha02, Sho05, Sin06, Vol05, Von03, Ye00, Zie01s, Zie02-31, Zie02l, Zie02e, Ano03-27, Ano07m, Kia00, Kia01, She07]. Modern [Ahm07c, Ano06t, Bab02b, Fot04, Loe07, Rya08, Yeh07, Ano03-28, Ano08l, Bod02b, Eas07, Zie00-33, Zie01-27]. modification [FVB02]. Modified [MWJ06, YW05]. Modular [AW05]. Modulating [ASW^{+}06]. Moeschberger [Rav04]. Mogens [Bar07b]. Mohamed [Nag04]. Mohammad [Olw01a]. Mohammed [Mat01]. Mohler [Ano03k]. Molecular [Ano06-43]. Molecule [LWY02]. Molenberghs [Ano06-43, Chi06, Ano07e, Avi01]. Møller [Nas04]. Moment [NM04, XD05]. Moment-Based [NM04]. Monahan [Was02]. Money [Esp07]. Monitoring [BGMS03, Che07, Gho09, GP01, Hus08, JTW00, JWT^{+}02, LT09, NK03, RS06, Zie00-46, Zie01-43, ZTW07, ZTW08, Bod02a]. Monroe [Zie00e]. Monte [Ahm08d, Ano03-34, Bor00, Den06, Kuo00, McC04, Ode00a, Sar03, BF08, Jon04, LCdC06, MR01, Ano05-31, Con01, Hes02, Kim00]. Monte-Carlo [Den06]. Montes [Zie04q]. Montgomery [Ano07j, Bri04, Gra02, Zie02-31, Zie02p, Zie02-27, Ano04h, Ano05-27, Ano06l, Gri00, Maz06a, Spi07, Zie01v, Zie01p, Zie02-36]. Moore [Ano04l, Zie01w, Zie04y, Zie01-50]. Morgan [Zie00-28, Dob02]. Morris [Ahm09]. Morrison [Wil05]. Morton [Zie01-32]. Moses [Ano05z]. Moshe [Ano08q]. Mosteller [Ano07h, Zie01-52, Ano07q]. Motion [FR07]. Moulines [Nag06]. Mounir [Ano07e]. Moustaki [Ano03-45]. Moving [AL03b, CM03, CE02, HMT07, HMT08, JCR01, RS05, RS06, Rob00a, TTL07, TJ09, ZTW07]. Moyé [Hus08, Zie04l]. MR2248365 [Bri06, Buj06, Efr06, Hub06, Lan06]. MR2414514 [TJ09]. MR2655651 [RBM11]. Mücklich [Qiu02]. Mukerjee [Loe07]. Mukhopadhyay [Sar06, Kat08a]. Mulier [Ano08k, Lor01]. Muller [Gun03, Key01, Qu07a, Sen02, Lor05, McC03a, Iye02]. Mullins [Hor06]. Multi [Boo05b, Nem03]. Multi- [Nem03]. Multi-Way [Boo05b]. Multidimensional [ME05, Ano03-29, Bur08a]. Multilevel [GP06, Gel06, Lip05, Ano06w, Zie02-34]. Multimedia [Mul04]. Multimodel [Ano03y]. Multiple [HG06, HBL03, JW02, KBS08, ML01, MX08, YS05, Zie00-34, LCdC06, Zie00-34, Zie04l]. Multiple-Method [HG06]. Multiple-Response [MX08]. Multiple-Target [JW02]. Multiresponse [Hec01, Lan01]. Multiscale [GCS06]. Multistage [Buk09, But04, LT09, ZDCS03, ZTL08]. Multistate [Bre06]. Multistratum [TG01]. Multivariable [Key01]. Multivariate [Ahm08g, Ano03-45, Ano03x, Ano04e, Ano05h, Ano06-37, AS01, AL03a, Ars05, Atk05, BBJK04, Boo00, Boo03, Cho06, CS02, DS06, Dup05, FJ08, FG06, GV07, HMT07, HMT08, Hec05, Hub01, JP01, Kat03b, Kat04a, Kat04b, Ken02, LLBN08, Nem00, PP01a, PP01b, Pol01, QH01, Qu07a, QD04, Rag01, RF05, RW01, RMGR09, RVVA04, ST06, SJ02, TG00, TTL07, TJ09, ZH06, Zie00b, Zie01-31, Zie02o, Zie04a, Zie04f, ZTW07, BF08, Ano05-57, Ano04p, Ano05-32, Bab02a, Bay02, Con05, Edg04, Kuh05b, LaL01, Qiu00, Qu07b, Wil05, Zie00-35, Zie00-36, Zie01-28, Zie02e]. Multiway [Esm09a, Esm09b]. Mun [Von06]. MUNDA [Bur08a]. Murdoch [Hly09a]. Murray [Ke06]. Murrell [Gun06]. Murtagh [Hec07]. Murthy [Ano03-50, Dog04, Olw01b]. Musicology [Zie04w]. Muth [Zie01e]. Muzaffar [Neu08c]. MVT(R) [Ano05m]. Myers [Zie02-31, Zie02-36, Smi01]. Myerson [Ano06y]. Myles [Ram00b, Ano05-56]. Myunghee [Ano04y].

N [Ahm08g, Ahm09, Ano03p, Ano03-28, Ano03-50, Ano04v, Ano04j, Ano05s, Ano05u, Ano07e, Ano08n, Ano09d, Boo09b, Cha04, Che01a, Che02a, Chi07, Den00, Dog04, Hor09b, Hut04, Kar07e, Ke09a, Nem03, Oli09b, Olw01b, Ran03, Rob02, Sea09, Str01, Zie00y, Zie00z, Zie00b, Zie00-43, Zie01-28, Zie01-27, Zie01-34, Zie03a, vB02]. Nachlas [Hey06]. Næs [Zie04a]. Nagaraj [Lum01]. Nagaraja [Cha04, Sea09]. Naik [Zie00b, Zie01-28]. Nan [Ano05e]. Nancy [Ano03-42, McC08]. Nando [Sar03]. Nandram [Zim01]. Nanny [Gho07]. Nanotubes [Kva08]. Narayan [Cab03]. Nathan [Ano06-42]. Natural [Cob00]. Nature [Jar08, Ano05-33]. Nature-Inspired [Ano05-33]. Naumann [Zie01m]. Naus [Zie02-39]. Nearly [CL07, Xu02]. Nearly-Orthogonal [Xu02]. Needs [Bar04]. Neerchal [Lum01]. Neil [Ano06-27, Boo03, Kal06, Sar03, Zie00-42, Zie01-39]. Nekrutkin [Ruk02]. Nelder [Ano07l, Gho08a]. Nello [Bre05a]. Nelsen [Ray00]. Nelson [Ano05-52, Esp06b, Ano05-52, Mee03]. Ness [Gup00]. Nested [HJM09]. Network [Bod02a, IM05, MBB05, Rao00]. Networks [Ahm08b, CGHL01, DLM^{+}07, Hea03, Hly07, KLM08, Li05, Liu09a, Mar03, Ped01, Qia02, Soy09, Ano05-33, Cas05, Cut00, Zie00-49]. Neubauer [Ano09a, Wlu06]. Neuilly [Bur00]. Neuman [Zie02-48]. Neural [Ano05-33, Cas05, IM05, Liu09a, Rao00, Zie00-49, Cut00]. Neutron [HTMW08]. Newey [Liu09b, Ord07a, Ord07b]. Newman [Zie02v]. News [Ste07]. Newton [Rum00]. Next [Ano05-55]. Ng [Kar07e]. Ngai [Qia07, Zie02-50]. Niall [Ano07l]. Nicholas [Ahm07c, Ano05g, Ano05-57, Dra04, Hec08, Mic01]. Nicholls [Zie00h].

Nichols [Ano06-36]. Nicholson [Zel05]. Nickerson [Fri05]. Nicolo [Lip07b]. Niels [Ano07u]. Nielsen [Ano06-43, Ahm08b]. Nigel [Bar09]. Nikolaidis [Ano06n]. Nikolaos [Ano08l, Ano09l]. Nikulin [Ano09l]. Nil [Sen06a]. Nils [Ros05]. Nina [Ruk02]. Ninth [Liu09b, Ord07a, Ord07b]. Nishisato [Ars05, Bur08a]. Nitis [Kat08a, Sar06]. Nizam [Key01]. NJ [Mai05]. no [Bri06, Buj06, Efr06, Hub06, IM02, Lan06]. Noise [KCM00, LSB06, PK05]. Noisy [HC00]. Nolan [Cha01a, Boo04b, Hey01]. Non [Li09]. Non-Gaussian [Li09]. Noncentral [Mee01]. Nondetects [Gat06]. Nonhomogeneous [BDNN04]. Nonlinear [Ano04-27, Ano09c, BK04, Bur08a, For08, Gan06, GJS09, HV07, Lu04b, PW01, Sen06b, Soy04, SMB01, Ano05-34, Cam01b, Liu06, Lu04a, Mur07, Sca02]. Nonnormal [RACH09]. Nonorthogonal [HK00, MS05]. Nonparametric [Ahm08c, Ano07i, Bar09, CHvdW06, Che03, Fot04, Fot07a, GG04, Gro00, Li06, Lom05, Lu04a, Nas01, Pol01, Ran03, Zie04i, ZTW08, Ano04q, Bak00, Gho03, Lor05, Lu07, Lu08, Ode00b, Qia00, Ram00b, Ste04]. Nonparametrics [Cas05, Zie00c]. Nonquadratic [FFG06]. Nonregular [CLY04, LLY03, LST07, XD05]. Nonresponse [Kar07c, Zie02-46]. Nonstandard [RACH09]. Nonstationary [COR08, Sca02]. Nora [Ano03-37]. Norbert [Zie00e]. Norm [Ano06-37]. Norma [Zie02-27]. Normal [Gui08, HME08, KMM08, UL07]. Normal-Based [KMM08]. Normality [LP06, Kap03, Sen05, Nel01]. Normalized [ŽH09]. Norman [Ano06-34, Ano07p]. Northorn [Ano06e]. Note [CHvdW06, CO01, Gui08, Ste09b, TJ09, UL07]. Notz [Den04]. Novel [Kal03]. Nozer [Aba01b, Gho08b]. Null [GS08]. Number [Bor00, McC04, MEM05, NM02]. Numbers [Ano06-38, GV03]. Numerical [Lar08, Gan06, Lip09, Mai05, Was02, Zel00]. Nunnari [Zie02-40]. Nychka [Zie01-49].

O [Alt03, AW01, Ano06-41, Bai04, Bar02, Sen06a, Zie02-44]. Objective [Ker03]. Observation [Coo00a, ZMW08]. Observational [Ano03-30]. Observations [SJ02, Zho01, Zie00o, BG06]. Observed [Pro00, Oli09c]. Obuchowski [Ano03-42]. O’Connell [Zie01i]. O’Connor [Rom03]. Offen [Chi06]. Offset [JS03, JC06]. Ogden [Gho01]. O’Gorman [Ahm04]. Ohser [Qiu02]. Ojeda [McC06c, Mic05]. Okine [Sen06a]. Olea [Zie00s]. Oligonucleotide [BH08, BCBS08a, BCBS08b, Gol08, HB08, Jon08, KL08]. Olive [Ami04]. Oliver [Zie01u]. Olivia [Zie02c]. Olivier [Ano07g, Nag06]. Olkin [Ahm08c]. Olofsson [Lip06a]. Olsen [Sau02b]. Olson [Zie02u, Ano03-46, Ano06-44, Zie00-30, Zie01-51]. Olwell [Umb01]. Oncology [Zie04-27]. One [BG06, FKQS01, FW06, KM04, LLI05, Zie02t]. One- [LLI05]. One-Factor-at-a-Time [FW06]. One-Sided [KM04, BG06]. One-Way [FKQS01, KM04, LLI05]. Operating [Cla04b, JLSZ08, JW04b]. Operational [Sen07]. Operations [Hly09b, Oli09d]. Optical [GLV02]. Optics [Per04]. Optimal [Ano06h, BC07, CW02, CA06, DM08a, FM04, GD03, GV03, KJ09, Ke05, KD05, LL03, LLY03, MNS04, Mit00, NCB04, NC08, Pas03, RHOYL06, XCW04, ZZJ07, Gri04]. Optimality [BV01]. Optimization [CHCFC04, Den06, Hes04, JW04a, LM04, ME04, Miy04, NW04, RdC05, Sud04, TGLG09, Alt01, Ano06t, Gho09, Olw01b, Zie01-43, Zie02d, Gru05, Zie01-29]. Optimum [Atk05, DM08b, Iye02, Hor09b]. Order [BSC09, CA06, GD03, PKV07, RF05, Sea09, Zie00y, Zie00z, Cha04]. Ordered [GP01]. Orders [Car07a, Ano08q]. Ordinal [LLBN08]. Ordys [Bar08c]. O’Reilly [Zie00h]. Oriented [Zie00-32]. Origin [Li05]. Origin-Destination [Li05]. O’Rourke [Ano06-37]. Orthogonal [CL07, LLY03, Mar05, MMM08, Xu02, ZZJ07, Den00]. Oscar [Ano08f, Xu06]. Other [Ano03a, Ano08p, Bar05, CHCFC04, Mee03, Oli08, Won04]. Ott [Ano05j, Kim05, Wlu06, Kar07b]. Otto [Zie01l]. Outdoor [CLM04]. Outlier [Hub01, PP01a, PP01b, RW01]. Outliers [JP01]. Output [Ano08o, Dri06, MBC00]. ová [Oli07]. Oxford [Ano04-30]. Ozone [Dup05].

P [Aba01b, Ano03r, Ano03y, Ano04l, Ano05-57, Ano05-56, Ano05-31, Ano05-29, Ano05-50, Ano06h, Ano06-40, Ano07p, Boo08, Cha01b, Che01b, Cob04, Con00, Den06, Eas07, Glo05, Kar07a, Kib07, Kim00, LZ02, Lip09, Liu07a, Lum01, Mai05, Mic03, MR00, Nel00a, Ode00a, Pig06, Ran03, Rav04, Sar00b, Soy09, Str01, Wal05, Yeh07, Zie00u, Zie00-40, Zie00-32, Zie00l, Zie00h, Zie01-41, Zie01-40, Zie01g, Zie01w, Zie01f, Zie02-48, Zie02-30, Zie04y]. P.-T [Zie02-30]. Padhraic [Owe04]. Pahkinen [Ano05-38]. Paik [Ano06-33, Ano04y]. Pairs [FM04]. Palazoglu [Neu08a]. Palit [Ano06j]. Palmer [Zie02u]. Palta [Ano04r]. Pan [Kuh03]. Pande [Zie02-48]. Pandu [Lew04]. Panel [Ano06-39, SBF^{+}08]. Pang [Cab09]. Panjer [Ano06v, Sen07]. Paolo [Boo09c, Owe04, San05]. Paper [Mal06, Bri06, Buj06, Efr06, Hub06, Lan06]. Papers [Ano07r, Ano07s, Ano07q, Gho07, SS07, Ahm09]. Paradigm [Kaf03]. Parallel [Ano08h, ZC06]. Parameter [Alt01, Ano09c, BS03, Bur04, DW06, GJS09, JW02, Jos03, KJ09, LLBN08, PW01, QD04, ZZJ07, ŽH09, Mam08]. Parameters [AK04, BKG09, DM08b, HSR09, JCR01, LHB^{+}02, LP06, Qia05, CJFR05]. Parametric [Ahm08c, Ano05-35, Ano07i, AB04, Bed00, JM01, Lu04a, Nas01, Sco01, SP07, SMB01, Zie04i, DGS07]. Pareto [CS01, ZS09]. Park [Zie01-43]. Parmigiani [Ano03-49, Zie02-33]. Parr [Zie02c]. Part [Che06, Zie00u, BH00]. Partial [CL09, GM01, MS00]. Particular [DP02]. Pass [Boy01]. Pass-Fail [Boy01]. Past [Kaf00c]. Pat [vB02, Esp07]. Path [RWZF04]. Patricia [Nem02a, Sin08]. Patrick [Ano03h, Ano05l, Neu09b, Rom03, Zie05b]. Pattern [Bag05, BLC01, Bre05a, Dav03, Liu09a, Sea05, TGLG09, Ano06x, Bar02, Nea07]. Patterns [AL03a, LH02, SA08, Lin02, Sym05]. Paul [AW05, Ano03-51, Ano03r, Ano03-39, Ano03-30, Ano06q, Ano06-38, Ano08m, Boo05b, Cla04a, Con04, Gun06, Joh00, Kat04a, Liu02, Lu08, Qia02, Qiu00, Qu07a, Sch04, Sen06b, Sho05, Wei03, Zie00r, Zie00-31]. Pauler [Ye03].

Paulo [Sin09]. Pause [Zie00-45]. Pavlov [Sen00]. Pawitan [Gho08a, Sau02a]. Pearl [Zel01]. Pearn [Ahm08h]. Pearson [Fot01]. Peck [Ano07j, Gra02, Sau02b, Min02]. Peel [Sha02]. Pei [Ano05-54, Zie00j]. Peihua [Lor06]. Peña [O’N03, Zie01a, Zie01h]. Penalized [LS05, ME05]. Peng [Mye09c]. Pensky [Sha04]. Pepe [Ano05-51]. Percival [Che01c]. Percy [Hly09b]. Performance [Bar08c, CE02, Hly07, JCR01, Law01, LPP00, Neu08a, RACH09, Zie02-51]. Perkins [Bur09b]. Permutation [Bab02a, Ano05-35, Ano08m, Liu02, Zie01-30]. Perspective [And08, CL03, Gho08b]. Perspectives [Con04, Ye05]. Persson [Liu09b, Ord07a, Ord07b]. Perturbation [Ruk01]. Pesarin [Bab02a]. Peter [Ano03s, Ano03h, Ano03k, Ano05-46, Ano06m, Ano09i, Bar02, Bar09, Bur03, Cha01c, Das07a, Esp06b, Fin02, Gan02, Ke05, Lip06a, Mag00, Mai08, Ros05, San03a, Sca02, Sen02, Sho04, Sin09, Sym05, Zie00-34, Zie01-29, Zie02-48, Zie02b]. Petre [Ano05z]. Petrochemical [BBJK04]. Petroleum [Zie00q, Zie05b]. Petruccelli [Zim01]. Ph [Ano03k]. Phaedon [Ano06p]. Pham [Liu07b]. Phan [Zie00-37]. Phan-Tan-Luu [Zie00-37]. Pharmaceutical [KS07b, Zie00j, Zie01e, Zie02q, Che08c, Zie00-37]. Phase [MW04, ZTL08]. Philip [Ano05-41, Kat09, Ken02, Ped01]. Philippe [Lu07]. Phillip [Ano03f, Ano05-35, Ano06z, Ano06u, Cha00, Sin07, Zie01-30, Zie02-35, Zie02r]. Phipps [Lip07c]. Phiroz [Ano06x]. Physical [Ano09b, Pic05, RWH^{+}04]. Physics [GLV02]. Picek [Oli07]. PID [Neu08b]. Pieces [Bak01]. Piecewisely [Qiu04]. Piegorsch [Ano06g, Zie02y]. Pierre [Cob00, Owe04, Zie00r]. Piersol [Cha01c]. Pieter [Esm09a, Esm09b]. Pinheiro [Zie01z]. Pintilie [And08]. Pitting [FHR06]. Plackett [KB05, MS01]. Plackett-Burman [KB05]. Plain [Bri03]. Plan [Kun08]. Plane [Fam03]. Planning [Alt01, SEM09, ZM06, Rut05]. Plans [Ano05-52, BSG00, Hem02, LL03, LLY03, NCB04, PM03, Pas03, Pas06]. Plant [Hec01, Boo04a]. Plot [BS03, GV03, Gui08, Ke05, KCV02, MB04, PKV07, RACH09, KB05]. Plots [GV03, MS00, SNS07]. Plotting [Nel00b]. PLUS [Utl05, Zie04c, Zie00-42, Ano03-46, Ano06-44, Fot07b, Kus03, Lum01, Zie01-51, Zie01-27, Zie01z, Zie02i, Zie02n, Zie05a, Bur07, Zie00-33]. PLUS(R) [Ano03z, Ano06u]. Poduri [Che02b, Maz02]. Point [CCD09, COR08, Sca04, Sca07, Sym05, ZH06, ZTL08]. Poisson [Bru00, BDNN04, WT09]. Polansey [Oli09c]. Polhemus [Zie00-42]. Policy [Zie01-32]. Politics [So01]. Politis [Str01]. Pollution [Mye09c]. Polynomial [BGW03, BSC09, DWG09, GPR05]. Pooling [GP06]. Pools [RHOYL06]. Popovic [Ano06j]. Population [Ano04r, BG06]. Populations [Zie00-40]. Porous [LHB^{+}02]. Portfolio [Ano06t]. Possible [Mil05]. Post [VB09]. Post-Fractionated [VB09]. Poursat [Ano04-27]. Power [Ano03-48, MV05]. Powerful [Zha06]. pp [Zie01-52]. pp. [Mai05, Zie05b]. Prabhakar [Ano03-50, Dog04, Olw01b]. Practical [And08, Ano05-43, CM08, Cha00, Che08c, Chi06, DS00, Hec09, Ker02, LSW09, McC00, Sar00b, She00, Sta00, WW00, Yat02, AW05, Ano05-36, Ano05-39, Ano05-37, Ano05-38, Cob02, Gib01, Gro00, Mye01, Nas02, Neu08b, Rom03, Rut02, Sar02, Zie00-38]. Practice [Ano04l, Bak01, Gun00, Sar00b, Zie01-50, Zie01w, Zie04y, Ano06-28, Kar07f, Lu07, Ray03b, Sar03, Was05, Zie01i, Zim00]. Practitioner [Rum06, Was01]. Practitioners [Ano08e]. Preben [Was04]. Precedence [Kar07e]. Precedence-Types [Kar07e]. Predict [MEH09, ZCC08]. Predictability [Wik05]. Predicting [GPR05, dCGM05, Ano08n]. Prediction [BG04, Bur03, FL07, HJMW04, HSNB09, KMM08, LWS03, NM02, ÖBMT02, Olw01b, BG06, Ano05-51, Ano09m, Ode00b, Lip07b]. Predictions [McC03c]. Predictors [HBL03, MBG07]. Prentice [Geo03]. Preparation [Nas08]. Prescott [Zie00a]. Presence [Kun08]. Present [Kaf00c]. Preserving [KFV^{+}07]. Press [Gru02, Lip09, Ker03]. Presser [Ano05-29]. Pressure [FK00]. Preventive [Tor01]. Price [Moo01]. Primer [Ano05n, Nem02a, Che08a, Hor09c]. Principal [HRV05, HN03, Mar05, Ano03-31]. Principles [AW01, Ano04s, Gri04, Jar09, Ker03, Ray03b, WLLZ07, Hem02, Dok08, Zie01-31]. Prior [BC07]. Priors [JD07]. Privacy [KFV^{+}07]. Privacy-Preserving [KFV^{+}07]. Probabilistic [FJ08, Fri05, Gri07, Hey06, Kat09, Soy09, Zie00-43, Bas02, Ped01, San07]. Probabilities [HME08, WT09]. Probability [AB03, Ann03, Bri04, Gui08, Jos04, KL02, LPP00, Mam09, Pic09, Sau05, Sau06, SC00, SNS07, Yeh05, Yeh07, Zie01-38, Zie02p, Ano03-35, Ano05n, Oli09a, Zim04, Ano03-32, Ano06y, Bre02, Cab08, Esm05, Esm06, Fot07c, Hen01, Hly09c, Ke08, Kim04, LaL04b, Lip06a, McC03b, Ng09, Per04, Qia01, San08a, Taa02, Zie04m]. Probablity [Neu09a]. Problem [Per04, QD04, Sar05, Ste00b, Zie02-51, MR01]. Problems [CHCFC04, DP02, HK00, HME08, Kaf01c, WM02, Cha05, LH06]. Procedure [BGW03, BGMS03, QH01, Qiu04]. Procedures [Ano07i, Hem02, JM00, LSB06, Neu01, Zie00-46, Cha01c, Nas01, Zie04i]. Proceedings [Zie02b]. Process [Ahm08h, Ano06-36, AK04, BBJK04, Bar01, BO09, Bel01, BGMS03, Cah06, DM08b, Edg04, Fri00, Gho09, HZ05, JLH^{+}06, JWT^{+}02, Kaf03, KML04, KCV02, LT09, LEH03, LBH^{+}06, Mor04, Nel01, Neu01, Neu03, Neu08a, Pla01, Pol01, QWW08, RdC05, RS06, Ste00a, TH05, Var00, YM00, Yat02, ZH06, Zie01-43, Zie02h, Zim02, vN02, Aba01a, Zie04p, Bar08c, Tob03, Wlu06, Zim00]. Processes [Ano04j, AS01, AL03b, AL03a, BE03, BGMS03, BDNN04, But04, CE02, Che06, FFG06, Kim04, Liu07a, McC05b, Ode00b, Pic05, Sca04, Whi04, ZZ07, ZDCS03, ZTL08, ŽH09, dC02, Ano07s, Buk09, Cab03, Esm05, Fot07c, Fot09, Lip06a, McC06a, Sca07]. Processing [Liu06, Lor06, Zie01n, Bab05, Cab06, MM00, Sca02]. Procrustes [Cha05]. Producing [JS07]. Product [Ano03a, Ano05-58, KT05, Mee03, MEH09, vN02]. Production [BBJK04, YM00]. Productivity [GM01]. Products [CLL09]. Professionals [Ano04u]. Professor [Ano07l]. Profile [SMB01]. Profile-Based [SMB01].

Profiles [KML04, MW04, ZTW07, ZTW08, Mul08]. Programmer [Hol08, San03b]. Programming [Ano06-40, Fre00, Hly09a, Mur07, Zie00-41]. Progressive [Kun08, NCB04]. Progressively [BSG00]. Projection [BBJK04, LST07, XD05]. Projective [XCW04]. Promislow [Nku07]. Pronzato [Zie00-32]. Properties [CJFR05, HWY02, LST07, WM05]. Proportion [GHB^{+}07]. Proportional [JWT^{+}02]. Proportions [Ano04y, Esp06b]. Proposed [BN01]. Proschan [Che07]. Proximity [Lip07c]. Psychology [Fri05]. Public [Zie01g, Zie01-32]. Pugesek [Ano06-31]. Purohit [Ahm07b]. Purposes [Bar08b]. Putting [Bak01]. Pythagorean [Con04]. Pyzdek [All01, Zie00-51].

Qi [Con01]. Qi-Man [Con01]. Qing [Ruk01]. Qiu [Lor06]. Qiwei [Lu04a]. Quadratic [PYC07, YS05, PDDL02]. Qualification [Neu01]. Qualimetrics [Zie00u]. Qualitative [HSNB09, QWW08]. Quality [Ahm08h, Ann06, Ano03-36, Ano03-43, Ano05-27, Ano06-35, Ano09a, Bai07a, BH08, Bar07c, Bay02, Bod02b, Bre01, BCBS08a, BCBS08b, Fri00, GP01, Gol08, HB08, Hof00, Hor06, Jon08, KL08, Kuh05b, LC00, Maz06b, McC08, Mit02, Neu03, O’N03, Spi07, Umb01, Wlu06, WM05, Yat01, Zie00-47, Zie00-42, Zie01v, Zie01-40, Zie01h, Zie02-30, Zie04-27, Zie04u, Zim02, Ano03r, Sta00, Zie01-55, All01, Ano05-40, Bar04, Bru00, Esp07, HL06, Rob02, Ste00b]. Quality-Control [LC00]. Quantification [Zie03b]. Quantify [BCL^{+}09]. Quantifying [Smi01]. Quantile [Glo01, Lom05, ZCC08, Lip06b]. Quantitative [HSNB09, Mul02, QWW08, Ano04r, Ano05-41]. Quantum [Ale02]. Quenneville [Zie01-35]. Questionnaires [Ke09a, Ano05-29]. Questions [Fam03]. Queueing [Fre00, Hly07]. Queues [Ode00a]. Queuing [McC03b].

R [Ann03, Ano03-27, Ano03-48, Ano03w, Ano03p, Ano03u, Ano03-41, Ano03n, Ano03y, Ano03k, Ano03-30, Ano03-50, Ano04j, Ano04k, Ano05i, Ano05o, Ano05y, Ano05j, Ano05u, Ano05-57, Ano05-56, Ano06-33, Ano06i, Ano06-39, Ano06s, Ano07k, Ano07x, Ano07p, Ano08p, Ano08e, Bab02b, Bin01, Bur09a, Cab00, Cam02, Car07d, Car07a, Che01a, Che02b, Dok08, Esp06b, Gat06, Geo00, Gho01, Gra00, Gun04, Hin06, Hor09b, Jen09a, Kaf03, Kar07b, Kar07d, Kat03a, Kib07, Kim05, Kus02, Kus06, Laz03b, Liu06, Liu07a, Maz02, McC00, McC08, McC06a, Mic03, MR00, Mul02, Mye09c, Nem04, Oli08, Olw01b, Per04, Pla01, Qia01, Rum00, Sin01, Smi06, Utl05, Von03, Wal05, Was01, Weh04, Wlu06, Zie00p, Zie00x, Zie00y, Zie00z, Zie00w, Zie00-30, Zie01-55, Zie01-46, Zie01g]. R [Zie02-48, Zie02t, Zie02-49, Zie02g, Zie02-29, Zie02-40, Zie04c, Zie04z, Zie04d, Zim00, Ano06e, Ano06u, Ano09i, Chi07, Hec07, Hly09a, Mye09a, Ng06, Ng09, Oli07, Pic09, Rob05, San03a, Gun06]. R&D [Gol04]. R&R [Maz06a]. R.-D [Ano08p, Cam02]. R/S [Ano06u]. R/S-PLUS(R) [Ano06u]. Rabe [Ahm08a, Fri01, Lip05, Zie02n]. Rabe-Hesketh [Ahm08a, Fri01, Lip05, Zie02n]. Rabinowitz [Sau06]. Radhakrishna [Bor01]. Radial [Qia02]. Radmacher [Ano05-54]. Rae [Rum00]. Rafael [Ano03-49]. Raftery [Zie02-45]. Raghavarao [Ste04]. Rahul [Loe07]. Rainer [Ano07g]. Rajesh [Alt06, Jia03]. Ralph [Che08c, KS07b]. Ramachandra [Cam01a]. Ramaswami [Bur01b]. Ramon [Moo03, Zie01-33, Zie04n]. Ramsay [Alt03, Ano06-41]. Rand [Ano05y, Bab02b, Nem04]. Randall [Bak00, Fin05, Zie00-34]. Randolph [Ano05z]. Random [BK04, BLC01, BG04, BG05, CE02, Dra04, Esm05, FK00, Fot07c, Gho08a, GB01, GHB^{+}07, Hen01, Iye02, Kim04, KM04, LHB^{+}02, LLI05, McC06c, Mic05, Pas03, Pic06, SB04, ŽH09, Cha01c, Ano09j, Bor00, McC04, Sea05]. Random-Coefficients [BK04]. Random-Effects [BG05, GB01]. Randomization [Ano03-33]. Randomized [San07]. Randy [Ano05-37, Ste07]. Range [ŽH09, Pfa04]. Ranjit [vN02]. Rank [MY08, QH01]. Rank-Based [QH01]. Ranked [Zha05]. Rao [Ano06s, Bor01, Cam01a, Zie01-46, Ano03p, Ano04v, Ano04j, Ano04k, Ano05u, Ano06i, Che02b, Fot09, Maz02, Sin01, Zie00x, Zie00y, Zie00z, Zie00w, Zie01g]. Rasmus [Ano06-43, Boo05b]. Rässler [Ano04x]. Rate [KBS08, LT09, Pro00]. Rate-Adjusted [LT09]. Rater [Von06]. Rates [Ano04y, Esp06b, GV07]. Ratio [CM08, DS00, Zha06]. Ratner [Ano04z]. Ratto [Che08a, Was05]. Rausand [Zie04x]. Ravindra [Zie00b, Zie01-28]. Ray [ASW^{+}06, Ano05i]. Raymond [Fri05, Zie02-31, Zie02-36]. Rayner [Che03]. Reactor [LWS03]. Read [Zie00m]. Real [Ano03-27, Sar06]. Real-World [Sar06]. Really [Gol07]. Reasoning [Ahm08f, Bai04, Zel01, Ano03l, Fri05]. Rebecca [Zie04r]. Receptor [CS02]. Recipes [Lip09]. Recognition [Ano06x, BLC01, DJ07, Nea07, Dav03, Liu09a, Sea05]. Reconciliation [Kar07a]. Recovery [HC00]. Recurrence [ZMW08]. Recurrences [Ano03a, Mee03]. Recurrent [Ano03a, FL07, SP07, Mee03]. Recursive [Sán06]. Redmond [Zie02w]. Reduced [BV01]. Reducing [McC06a]. Reduction [JLH^{+}06, Lin02, Zie01-28]. Reese [Hor09a]. Reference [Tom07, Yat01]. Refresher [Ano07k]. Reginald [Zie00h]. Region [ZCC08]. Regions [BG05, GD03, BF08]. Register [Bar08b]. Register-Based [Bar08b]. Regression [Alt02, Ano03v, Ano05f, Ano06q, Ano07j, Bak00, BSC09, BF05, CHCFC04, Coo00a, Coo00b, CO01, GLM07, GB01, GG04, Gra00, Gra02, Gun00, Gup00, GPR05, HBL03, Hec08, HV07, Hey09, HK00, HN03, JQ09, Key01, LN06, Lor06, Mar05, ME05, McC03c, Oli02, Oli03, Oli05, OH05, Pan00, Peñ05, Por02, Qiu04, RA00, RVVA04, SC04, SB04, SMB01, YW05, Zie00-31, ZTW08, Ami04, Ano03-44, Ano04-27, Ano07i, Con02a, Li06, Lip06b, Mal05, Soy04, Wal05, Zie01-33, Zie01-37, Zie04c, Zie06, Ano05-42, Ano07n, Cab00, Gol06b, Gun03, Moo01, Ray03a, Wal03, Zie04n]. Regressions [FW00]. Regularized [Qia02]. Reid [Che01a, Con00]. Reilly [Zie02-32]. Reiss [Ano08p, Cam02]. Reiter [Lip07a]. Rejoinder [BCBS08b, Wil08b]. Rekab [Neu08c]. Related [Ano07o, Ano09k, HN03, Rag01, Ruk02, SC00, Sin06, Won04]. Relating [Nel00a]. Relationship [HJMW04, Pas06]. Relationships [CR06, Zie00o]. Reliability [Ann03, Ano03-32, Ano03-50, Ano06n, BSG00, Bar08a, BE03, Bru00, CLL09, Dog01, FK00, Gat02, Gri04, Ke07, Ke08, KMM08, Law00, McC03b, MEH09, MSM03, Ng08, RLR06, Rom03, Sen00, Tob00, WM02, Zie01-41, Zie04x, CFP08, MN06, Ano03p, Ano08l, Din05, Hey02, Hor07, Hor09a, Hua00, Zie00-43, Zie04z, Ann06, Gho08b, Hey01, Hey06, Hil00, Joh00, Ke07, Mye05a, Olw01a, Olw01b, Sta04, Tor01, Zie02f]. Reliable [Ahm08f]. Remacre [Ano05s]. Remaining [DM08a]. Reminiscences [Dok09]. Rencher [Rut09, Ano03x, Bur01a]. René [Nas02, Yu07b, Zie05a]. Renewal [LEH03]. Repairable [LEH03, Zie01-41]. Repairs [Ano03a, Mee03]. Repeated [Gho05, Gol06b, Ken03, Kia01]. Replication [CL09]. Report [Kaf00b, Kaf01a, Not02, Not03b, Not04, Sit05, Sit06, Sit07, Ste08b, Ste09c]. Representation [Cab09, Lip07c]. Requirements [MEM05, Ke07]. Resampling [Ano06u, Lou01, Ano06z, Cha00, Hut04, Zie02-35]. Research [Ahm06, AW01, Ano03j, Ano05-58, Hly09b, KT05, Sen06a, Ano04o, Ano04-28, Ano04-31, Bre05b, Ke09a, Liu04, Nem02b, Zie00e]. Researcher [Mor05a]. Researchers [Ano08e]. Reserves [Ano06p]. Reservoir [Ano05r]. Residual [MS00]. Residuals [YM00]. resolution [LM02]. Resources [Cob00, RS04]. Response [ACBJ09, ALKP02, Ano05-42, BV01, CO01, FM04, GD03, GD04, JW04a, JW04b, Khu06, LM04, LC00, MX08, MLC09, ME04, Miy04, Mor00a, MBDD09, NW04, PP01b, QD04, RF05, Rya08, Sen06b, Sud04, TG01, WKT^{+}03, YM00, YS05, Ano07p, Ano07o, Zie02-36]. Responses [Atk05, FHR06, RACH09, SX07]. Restrictions [DP02]. Results [Ano05m, MEH09, Zie02h, Zie00-28]. Reuven [Boo08, Den06]. Review [Aba01b, Aba01a, Ahm04, Ahm06, Ahm07b, Ahm07a, Ahm07c, Ahm08a, Ahm08d, Ahm08b, Ahm08g, Ahm08c, Ahm08f, Ahm08e, Ahm08h, Ahm09, Ale01, Ale02, All01, Alt01, Alt02, Alt03, Alt06, Ami00, Ami04, AC02, AW01, AW05, And08, Ann01, Ann03, Ann06, Ano03i, Ano03m, Ano03r, Ano03-27, Ano03-48, Ano03-36, Ano03-33, Ano03-35, Ano03w, Ano03p, Ano03q, Ano03-40, Ano03-34, Ano03z, Ano03-45, Ano03t, Ano03u, Ano03g, Ano03x, Ano03v, Ano03s, Ano03-28, Ano03-39, Ano03-41, Ano03-42, Ano03o, Ano03f, Ano03l, Ano03-44, Ano03n, Ano03h, Ano03y, Ano03k, Ano03-29, Ano03-30, Ano03-37, Ano03j, Ano03-46, Ano03-43, Ano03-47, Ano03-38, Ano03-31, Ano03-32, Ano04l, Ano04-30, Ano04-29, Ano04e, Ano04n, Ano04v, Ano04s, Ano04t, Ano04i, Ano04j, Ano04w, Ano04z, Ano04m]. Review [Ano04r, Ano04h, Ano04q, Ano04-27, Ano04g, Ano04k, Ano04y, Ano04f, Ano04u, Ano04p, Ano04o, Ano04-28, Ano04x, Ano05v, Ano05-46, Ano05i, Ano05f, Ano05-32, Ano05-42, Ano05-27, Ano05-34, Ano05q, Ano05o, Ano05-36, Ano05k, Ano05-43, Ano05t, Ano05-35, Ano05y, Ano05g, Ano05j, Ano05-44, Ano05-47, Ano05r, Ano05s, Ano05m, Ano05x, Ano05-49, Ano05n, Ano05-30, Ano05z, Ano05p, Ano05-40, Ano05w, Ano05-33, Ano05-39, Ano05-41, Ano05u, Ano05-45, Ano05-31, Ano05-28, Ano05-37, Ano05-48, Ano05l, Ano05e, Ano05-51, Ano05-38, Ano05-29, Ano05-50, Ano05h, Ano06-33, Ano06w, Ano06t, Ano06-34, Ano06g, Ano06j, Ano06x, Ano06r, Ano06p, Ano06z, Ano06i, Ano06e, Ano06-27, Ano06-35, Ano06n, Ano06-31, Ano06q, Ano06l, Ano06k, Ano06y, Ano06m, Ano06u]. Review [Ano06s, Ano06h, Ano06-30, Ano06-28, Ano06-36, Ano06f, Ano06-29, Ano06o, Ano06v, Ano06-32, Ano07k, Ano07w, Ano07m, Ano07r, Ano07v, Ano07s, Ano07q, Ano07n, Ano07t, Ano07x, Ano07j, Ano07g, Ano07e, Ano07h, Ano07u, Ano07p, Ano07i, Ano07f, Ano07l, Ano07o, Ano08j, Ano08k, Ano08m, Ano08g, Ano08o, Ano08q, Ano08h, Ano08l, Ano08n, Ano08r, Ano08f, Ano08p, Ano08e, Ano08i, Ano09h, Ano09c, Ano09g, Ano09l, Ano09k, Ano09i, Ano09j, Ano09b, Ano09a, Ano09m, Ano09d, Ano09e, Ano09f, Ars05, Avi01, Bab02b, Bab02a, Bab05, Bag05, Bai04, Bai05, Bai07b, Bai07a, Baj00, Bak00, Bak01, Bar02, Bar05, Bar08a, Bar08b, Bar01, Bar04, Bar07c, Bar07b, Bar07a, Bar08c, Bar09, Bas00, Bas02, Bay02, Bel01, Bin01, Bla01, Bod02b, Bod02a]. Review [Boo00, Boo03, Boo04a, Boo04b, Boo05a, Boo05b, Boo07a, Boo07b, Boo08, Boo09c, Boo09b, Boo09a, Bor00, Bor02, Bor01, vB02, Bre02, Bre05a, Bre06, Bre05b, Bre01, Bre04, Bre07, Bri03, Bri04, Bri05, Bru00, Buk09, Bur01a, Bur02, Bur00, Bur01b, Bur03, Bur04, Bur06, Bur07, Bur08a, Bur08b, Bur09a, Bur09b, Bzi04, Cab00, Cab03, Cab04, Cab06, Cab08, Cab09, Cah06, Cam01a, Cam01b, Cam02, Car07d, Car07a, Car07b, Car07c, Car00, Cas05, Cha01a, Cha02a, Cha01b, Cha03, Cha02b, Cha04, Cha05, Cha08b, Cha09, Cha00, Che01a, Che03, Che01b, Che01c, Che02b, Che02a, Che07, Che08c, Che08b, Che08a, Chi06, Chi07, Cho06, Cla04a, Cob00, Cob02, Cob04, Con00, Con01, Con02a, Con02b, CL03, Con04, Con05, Cut00, Dan01, Das00, Das04, Das07b, Das07a, Dav03]. Review [Den00, Den04, Den06, Din05, Dob02, Dog01, Dog04, Dok08, Dok09, Dra04, Eas07, Edg04, Esm05, Esm06, Esm09a, Esm09b, Esp06a, Esp06b, Esp07, Eva01, Fah07, Fam03, Fin02, Fin05, Fle09, Fok07, Fot01, Fot04, Fot07a, Fot07b, Fot07c, Fot09, Fre00, Fre03, Fri00, Fri01, Fri05, Gan02, Gan05, Gan06, Gar00, Gat02, Gat06, Geo00, Geo03, Gho01, Gho02, Gho03, Gho05, Gho07, Gho08b, Gho08a, Gho09, Gib01, Glo01, Glo05, Glo06, Gol04, Gol06a, Gol06b, Gol07, Goo02, Gra00, Gra02, Gre04, Gri00, Gri04, Gri07, Gro00, Gru02, Gru05, Gru08, Gun03, Gun04, Gun06, Gup00, Hap00, Hec01, Hec05, Hec07, Hec08, Hec09, Hem02, Hen01, HL06, Hes02, Hes04, Hey01, Hey02, Hey06, Hey09, Hil00, Hin05, Hin06, Hly07, Hly09a]. Review [Hly09b, Hly09c, Hof00, HR05, Hol01, Hol08, Hor06, Hor07, Hor09b, Hor09a, Hor09c, Hua00, Hus08, Hut04, HI00, Huz01, HG02, Huz03, Iye02, Jar04, Jar08, Jar09, Jen02, Jen08, Jen09c, Jen09a, Jen09b, Jia03, Joh00, Joh03, Jos04, Kaf03, Kal03, Kal06, Kap03, Kar07b, Kar07d, Kar07c, Kar07e, Kar07f, Kar07a, Kat03b, Kat03a, Kat04a, Kat04b, Kat08a, Kat09, Kat01, Kat08b, Ke05, Ke06, Ke07, Ke08, Ke09a, Ke09b, Ken02, Ken03, Ker02, Ker03, Ket00, Key01, Kia00, Kia01, Kib07, Kim00, Kim04, Kim05, Kok01, KS07b, KS07a, Kuh03, Kuh05b, Kuh05a, Kuo00, Kus00, Kus02, Kus03, Kus06, LaL01, LaL04a, LaL04b, LaL09, Lar08, Law01, Laz03a, Laz03b, Laz05, Lew04, Li06, LH06, LZ02, Lin02, Lip01, Lip05, Lip06a, Lip06b]. Review [Lip07c, Lip07b, Lip07a, Lip09, Lit01, Liu02, Liu04, Liu06, Liu07a, Liu07b, Liu09a, Liu09b, Loe07, Lor01, Lor05, Lor06, Lou01, Lu04a, Lu04b, Lu07, Lu08, Lum01, Mag09, Mag00, MM00, Mai05, Mai07, Mai08, Mal05, Mam08, Mam09, Mar03, Mar06, Mat01, Maz02, Maz06b, Maz06a, Maz00, McC00, McC03a, McC07, McC08, McC03b, McC06a, McC02, McC06c, McC03c, McC04, McC05a, McC05b, McG02, McG07, Mee03, MS07, Mic01, Mic03, Mic05, Min02, Mit02, Moo01, Moo03, Mor04, Mor05a, Mor05b, MR00, Mul01, Mul04, Mul08, Mul02, Mur07, Mye00, Mye01, Mye05a, Mye05b, Mye08, Mye09b, Mye09c, Mye09a, Nag04, Nag06, Nas01, Nas02, Nas04, Nas08, Nea07, Nel00a, Nel01, Nem04, Nem08, Nem00, Nem02b, Nem02a, Nem03, Neu01, Neu03, Neu05, Neu08c, Neu08a, Neu08b]. Review [Neu09b, Neu09a, Ng05, Ng06, Ng08, Ng09, Nku07, O’N01, O’N03, O’N05, Ode00b, Ode00a, Oli03, Oli04, Oli05, Oli07, Oli08, Oli09a, Oli09c, Oli09b, Olw01a, Olw01b, Ord07a, Ord07b, Owe00, Owe03, Owe04, Owe06, Pan00, Ped01, Per04, Pet03a, Pfa04, Pic06, Pic07, Pic09, Pig06, Pla01, Pla02, Por02, Pry03, Qia00, Qia01, Qia02, Qia05, Qia07, Qiu00, Qiu02, Qiu05, Qiu08, Qu07b, Qu07a, Qu08, Rag01, Ram00a, Ram00b, Ran03, Rao00, Rao01, Rav04, Ray03a, Ray00, Ray03b, Rob00b, Rob02, Rob05, Rob06, Rom03, Rom06, Ros04, Ros05, Ruk01, Ruk02, Rum00, Rum04, Rum06, Rut02, Rut05, Rut09, Rut00, Rya08, San03b, San03a, San05, San07, San08b, San08a, Sar00a, Sar02, Sar03, Sar05, Sar06, Sar00b, Sau02b, Sau02a, Sau05, Sau06, Sau07, Sca02, Sca04, Sca05]. Review [Sca07, Sch04, Sea01, Sea02, Sea05, Sea09, Sen00, Sen05, Sen07, Sen02, Sen06a, Sen06b, Sha02, Sha04, She00, She07, Sho03, Sho04, Sho05, Sin01, Sin06, Sin07, Sin08, Sin09, Slu03, Smi01, Smi02, Smi06, So01, Soy04, Soy09, Spi07, Sta00, Sta04, Ste00b, Ste04, Str01, Sul01, Sym01, Sym05, Taa02, Tob00, Tob03, Tor01, Umb01, Utl05, Van04, Vol05, Von03, Von06, Vuk00, Wal03, Wal05, Was01, Was02, Was04, Was05, Weh02, Weh04, Wei03, Wik03, Wik05, Wil02, Wil05, Wlu03, Wlu06, Won04, WKT^{+}03, Woo00, Xu06, Yat01, Yat02, Ye00, Ye03, Ye05, Yeh05, Yeh07, Yu07b, Zel00, Zel01, Zel05, Zha05, Zie00u, Zie00-35, Zie00c, Zie00p, Zie00m, Zie00-40, Zie00e, Zie00-45, Zie00n, Zie00i, Zie00o, Zie00d, Zie00-51, Zie00-29, Zie00g, Zie00f, Zie00a]. Review [Zie00-48, Zie00-44, Zie00-49, Zie00r, Zie00q, Zie00s, Zie00x, Zie00y, Zie00z, Zie00w, Zie00v, Zie00-36, Zie00-46, Zie00-30, Zie00-47, Zie00-27, Zie00-33, Zie00-41, Zie00-32, Zie00-37, Zie00l, Zie00j, Zie00t, Zie00-39, Zie00h, Zie00k, Zie00-50, Zie00-38, Zie00-42, Zie00-34, Zie00-31, Zie00b, Zie00-28, Zie00-43, Zie01-43, Zie01-28, Zie01j, Zie01-51, Zie01-27, Zie01-38, Zie01-41, Zie01d, Zie01i, Zie01-50, Zie01-36, Zie01s, Zie01x, Zie01-44, Zie01-47, Zie01-49, Zie01-39, Zie01r, Zie01-52, Zie01y, Zie01v, Zie01-33, Zie01-31, Zie01-53, Zie01u, Zie01-42, Zie01-35, Zie01-40, Zie01-29, Zie01a, Zie01m, Zie01-55, Zie01-46, Zie01z, Zie01c, Zie01-30, Zie01o, Zie01t, Zie01n, Zie01k, Zie01q, Zie01g, Zie01l, Zie01h, Zie01p, Zie01w, Zie01-37, Zie01f, Zie01-54, Zie01-34, Zie01e, Zie01b, Zie01-32, Zie01-48, Zie01-45, Zie02-42, Zie02-31, Zie02-36, Zie02-35, Zie02-28]. Review [Zie02-48, Zie02i, Zie02j, Zie02-46, Zie02v, Zie02-45, Zie02-30, Zie02t, Zie02l, Zie02o, Zie02-38, Zie02-41, Zie02-50, Zie02m, Zie02y, Zie02u, Zie02-33, Zie02-34, Zie02q, Zie02-49, Zie02h, Zie02p, Zie02f, Zie02e, Zie02c, Zie02d, Zie02-37, Zie02a, Zie02g, Zie02x, Zie02b, Zie02-27, Zie02-29, Zie02-39, Zie02k, Zie02-47, Zie02-40, Zie02-32, Zie02r, Zie02-44, Zie02-43, Zie02s, Zie02n, Zie02w, Zie02z, Zie03b, Zie03c, Zie03a, Zie04a, Zie04u, Zie04i, Zie04s, Zie04q, Zie04n, Zie04c, Zie04b, Zie04k, Zie04o, Zie04y, Zie04e, Zie04l, Zie04w, Zie04j, Zie04f, Zie04x, Zie04p, Zie04z, Zie04m, Zie04d, Zie04r, Zie04v, Zie04h, Zie04g, Zie04t, Zie05b, Zie05a, Zie06, Zim00, Zim01, Zim02, Zim04, vN02, vT00]. Reviews [Ano03-49, Ano04-31, Ano05-52, Ano05-55, Ano05-57, Ano05-56, Ano05-54, Ano05-53, Ano06-42, Ano06-43, Ano06-39, Ano06-38, Ano06-41, Ano06-37, Ano06-40, Cha01c, Zie02-51, Zie04-27]. Revolution [Zie02-47]. Revolutionized [Gun00]. Revolutionizing [Ale01]. Rex [Ano06-30, Bur08b]. Reyment [Ano05h]. Riani [Kat04b, Por02]. Ribeiro [Sin09]. Ricardo [Zie00s]. Riccardo [Ano05-33]. Richard [Ale01, Ano03s, Ano03-38, Ano03-32, Ano05-46, Ano05-54, Ano05h, Bar02, Boo04a, Bur03, Fin05, Hly09c, Kar07b, Ke08, Kim04, Liu09b, Maz06a, Ord07a, Ord07b, Ray03b, Sca02, Sch04, Sho05, Utl05, Wik03, Zie00o, Zie01i, Zie01u]. Richardson [Ros05]. Richtsmeier [Weh04]. Ridge [Ano07p, DP02, GD03, HK00, PK05]. Rigdon [Zie01-41]. Right [McC07, SNS07, YW05]. Right-Brain [McC07]. Right-Censored [SNS07, YW05]. Rinaldo [Taa02]. Rio [Ano05s]. Rios [Kat01]. Ripley [Ano03-28, Zie01-27]. Risk [Bas02, Gho08b, Ke07, Neu09a, Olw01a, Qia07, Sen07, Zie00-46, Zie00-30, Ano05-45, Ano05-43, Bla01, Bre04, She00, Zie02-37]. Risks [And08, Car07a, CL05, Con02b, McC03a]. Risto [Ano05-38]. river [Bri01]. Rizzotto [Zie02-40]. Robert [Ahm06, Ahm09, AW01, Ano03u, Ano03-41, Ano03-38, Ano05-30, Ano05-31, Ano05-50, Ano09b, Ano09m, Bab05, Bai05, Boo09b, Bur06, Cab06, Con00, Edg04, Kim00, Liu09a, O’N01, Ped01, Sea02, Sta00, Wei03, Ye03, Zie00d, Zie02-48, Zie02-46, Zie02u, Zie02-32, Zie03b, Zie03c]. Roberts [Ray03b, McC07]. Robin [Bai07a, Das07b, Zie00a]. Robinson [Ano06-33, Gib01]. ROBPCA [HRV05]. Robust [Alt06, Ano05y, Atk05, BS03, BF08, Coo00b, DW06, DS06, Dup05, FJ08, GJS09, Hub01, HRV05, JW02, Jos03, KJ09, KLM08, LN00, MZ02, Mar05, MY08, MS00, Oli02, PM03, Pas06, PP01a, PP01b, QD04, RdC05, RA00, RW01, RVVA04, TGLG09, TG00, Zho01, ZZJ07, Zie00x, Zie01-52, Kat01, Oli07, Por02]. Robustness [JW04b, Cla04b, O’N03]. Rodeiro [Ano05-56]. Roderick [Zie02-46]. Rodger [Ano09e]. Rodolphe [Das07b]. Roes [Pla01]. Roger [Ano04m, Ano06y, Bas02, Law01, Lip06b, Mye09c, Ray00, Zie00e, Zie00-37, Zie02-42, Zie02-41]. Roland [Zie01t, Zie02-48]. Rolph [Qu07b, Zie01-32]. Rom [Zie00-34]. Romano [Str01]. Ron [Hol08, San03b]. Ronald [Ano03w, Ano04m, Ano06-40, Fam03, Fot01, Qu07b, Zie02-42, Zie02k]. Rong [Hen01]. Root [Bre07, Dan01]. Roots [YS05]. Rootzén [Ano04i]. Rosenbaum [Ano03-30]. Rosenberger [Ano03-33]. Rosenblatt [McC02]. Ross [Geo03, McG07, Sau05]. Rossi [Kib07]. Rossman [HI00, HG02]. Rothgeb [Ano05-29]. Rovers [Zie00-46]. Rowe [Con05]. Rowland [Ano06-36]. Roxy [Min02, Sau02b]. Roy [Nem02b, vN02]. Rubin [Gre04, Laz03b]. Rubinstein [Boo08, Den06]. Rud [Zie02c]. Rudestam [Rum00]. Rudolf [Moo03, Slu03, Zie01-33, Zie04s, Zie04n]. Rue [Pic06]. Ruey [Ano06f, Zie01a, Zie02m]. Ruggeri [Kat01]. Rules [Neu09a, Pry03, Zie01k]. Rumsey [Rum04]. Run [LPP00, MS01, SC00, TH05, Xu09]. Run-Length [LPP00]. Run-Related [SC00]. Runger [Zie02-27, Bri04, Zie02p]. Running [Pla02]. Runs [Xu02, Che02a]. Runwei [Zie02d]. Ruppert [Wal05, Ano05-48]. Russell [Bin01, Zie00-34, Zie02-29]. Ruth [Ami04]. Ruud [Fle09]. Ryan [Eas07, Zie01-40]. Rydén [Nag06].

S [Ahm08a, Ann06, Ano03-49, Ano03-47, Ano04l, Ano04-27, Ano06k, Ano06f, Ano06-40, Ano07w, Ano07v, Ano07n, Ano08n, Ano09f, Ars05, Bar05, Bur09b, Cab09, Cha01c, Che02b, Che08a, Das04, Das07b, Den00, Esp06b, Fri05, Glo06, Gre04, Gru02, Hem02, Hes02, Hin05, Hor09a, Hut04, Kat03b, Ken03, Ker03, Lu04b, Maz02, McC03b, Moo01, Mur07, Nem03, Nku07, Oli08, Sea01, Sta04, Tob00, Utl05, Zie00u, Zie00-40, Zie00x, Zie00w, Zie01j, Zie01-51, Zie01-50, Zie01r, Zie01y, Zie01a, Zie01z, Zie01k, Zie01w, Zie02-48, Zie02i, Zie02j, Zie02o, Zie02n, Zie04c, Zie04y, Zie04g, Ano06-41, Ano03z, Ano03-28, Ano03-46, Ano06-44, Bur07, Fot07b, Kus03, Lum01, Nas02, Zie00-33, Zie01-51, Zie01-27, Zie01z, Zie02i, Zie02n, Zie05a]. S-PLUS [Utl05, Zie04c, Ano03-46, Ano06-44, Fot07b, Kus03, Lum01, Zie01-51, Zie01-27, Zie01z, Zie02i, Zie02n, Zie05a, Bur07, Zie00-33]. S-PLUS(R) [Ano03z, Ano06u]. Saar [Huz03]. Safety [Ann06]. Sahai [Mat01, McC06c, Mic05]. Saibal [Sar06]. Saisana [Che08a]. Saito [Mar06]. Sall [Ano05-28, Ano08j]. Sallie [Ano08l]. Sally [Zie01-32]. Saltelli [Che08a, Was05, Zie01-36]. Sam [Qia00, Zie00-27]. Samaradasa [Gho05, Lit01]. Sample [CHvdW06, CG09, LSW09, MEM05, TTL07, TJ09, Van02, WW00, Zha06, Zie00k, Ano04s, Con04]. Samples [BKG09, GHB^{+}07, VD07, dCGM05]. Sampling [Ano05i, Ano09a, BSG00, Bru00, Gho02, Hem02, Kla01, LaL04a, MB06, Nem02a, Pic05, RSS02, RS04, SC00, Smi01, Zha05, Kar07b, Zie00-35, Maz02, Mye09b, Zie00-40, Zie00-39, Zie02-38]. Samprit [Ano07n, Moo01]. Samuel [Ahm08h, Ano06-34, Sha04, Zie00m, Zie04t, Zim00]. Sanchez [Ano06r]. Sanchez-Vila [Ano06r]. Sandra [Jen09b]. Sanford [Gra00, Mal05]. Sanjeev [Ahm08f]. Santner [Den04]. Sarabia [Boo09b, Rao01]. Saris [Ke09a]. Sarjinder [Car07c]. Särndal [Kar07c]. SAS [Ano08o, Cab04, Gun03, Hol08, Hor09c, Sin08, Ano06-40, Ahm08g, Ano09f, Hor09b, Nas08, Utl05, Zie02j, Ano08o]. SAS(R) [Ano03j, Ano06-37, Ano06-40, Che08c, Chi06, Jen08, Jen09a, LaL09, San03b, Van04, Zie00-31, Zie00b, Zie01-28, Zie01j, Ano06q, Ano03-34, Ano04t, Liu04, Moo03, Zie00-41, Zie01-33]. Satisfaction [Sin01, Bar04]. Saul [Lip09]. Savage [Zie00-27]. Savazzi [Ano05h]. Scale [GLM07, MBG07, MEM05, SM04]. Scales [CLL09, Mar05, WB05]. Scaling [Ano03-29]. Scan [PC01, Zie01-34, Zie02-39]. Scans [Che02a]. Scanu [Kar07f]. Schaalje [Rut09]. Schaeffer [Ano03-37, Kar07b]. Schafer [Boo00, Hin06, She07]. Schbath [Das07b]. Schechter [Zie00e]. Scheines [Bur03]. Scheme [TH05]. Schemes [CW02, LT09, LC00, ZTW07]. Scherer [Ano06t]. Schilling [Ano09a, Wlu06, Yu07b]. Schimek [Zie01-37]. Schinazi [Taa02]. Schlotzhauer [Jen09b]. Schoenberg [Ano06-33]. Scholes [Cab08]. Scholkopf [Ano08n]. School [Car07c]. Schott [Kar07d]. Schreiber [Ano05-34, Cam01b]. Schroeder [Ale01]. Schulze [Neu01]. Schumacker [Fin05]. Schwarz [Zie00e]. Science [Ahm09, Bre02, Hly07, Hly09b, Kal03, Lu08, McC03b, Pic07, SS07, Zie01-44, Zie02-43, Esm06, Kat09, LaL04b, Mam09, Mul01, Qiu02, Rom06, San08a, Soy09]. Sciences [Ano07e, Zie00m, Ano03u, Boo05b, McC02, Neu05, Rag01, San08b, Wei03, Zie00-45, Zie01-49, Zie04m, Zie04t]. Scientific [Hes02, Jar09, Lip09]. Scientist [Ano07k, Mai05]. Scientists [Gru02, Jen02, Lar08, Mam08, Ram00a, Ano03-45, Ano03-32, Ano05g, Ano05-39, Ano09b, Esp06a, Fah07, Ke08, Mic01, Rut02, Sau05, Zie00s, Zie01u, Zim01]. Sclove [Zie04y]. Score [ST02]. Scott [Ano03-49, Ano03h, Ano04g, Esm05, LaL01, Zie01-36]. Screening [Ano08s, VDLB08, XCW04]. Sean [Ano03-34, Cab08, Esm06]. Search [Hes04, TGLG09, Bur03, Kat04b]. Searle [Von03]. Seasonal [LH02, Whi04, Zie01-35]. Sebastian [Gho03]. Seber [Oli03, Kok01]. Second [BSC09, CA06, GD03, PKV07, RF05]. Second-Order [BSC09, CA06, GD03, PKV07, RF05]. Section [HTMW08]. Secure [KFV^{+}07]. Security [Glo06]. Seeing [Ano05-44]. Segers [Laz05]. Seila [Lew04]. Seismic [Ano00e, SBD98]. SELC [MWJ06]. Selected [Ahm09, Ano07r, Ano07s, Ano07q]. Selecting [KT03, MBC00]. Selection [Ano03-44, Ano03y, BGW03, BFL02, BC07, CL05, DT02, FM04, GG04, LWY02, LBH^{+}06, LSB06, MR00, OH05, Oli09b, RSB09, SHY04, TVW05, YJL07, ZC06, FVB02, Ano09j, Cab00]. Selective [MNS04]. Self [Ano03v, Ano06-32, HMT07, SJ02, CFP08]. self-dual [CFP08]. Self-Learning [Ano03v, Ano06-32]. Self-Starting [HMT07, SJ02]. Selvin [Zie00-33]. Semiconductor [BLC01]. Semiconductors [HTMW08]. Semifolding [MP00]. Semiparametric [Ahm08c, FKQS01, Lor05, MSM03, YW05, Fok07, Wal05]. Semiparametrics [Ano07v]. Sen [Zie01g]. Senn [Ano04-31]. Sennott [Fre00]. Sense [Jar04]. Sensing [Che06]. Sensitivity [ACBJ09, CHCFC04, Che08a, DWG09, MMM08, Oak09, RCGD06, SMB01, Was05, Zie01-36]. Sensor [KLM08]. Separation [AL03a, Con05, SA08]. Sequence [Ano07e]. Sequences [Hly09c]. Sequential [BGD05, COR08, MWJ06, RBM08, RBM11, Sar06, LCdC06, Kat08b, Sar03]. Serially [Zho01]. Series [Ano03z, Ano03s, Ano05-34, Ano06-33, Ano06j, Ano07s, Bur07, Cab00, Cam01b, Che01c, COR08, Fot07b, Hin05, Hly09b, Lu04a, McC03c, Mye05b, Ray03a, Ruk02, Sca05, She09, Zie01a, Zie04r, Ano04-29, Ano04t, Ano04f, Ano06f, Cha01c, Kar07a, McC05a, Sar02, Zie00c, Zie02-50, Zie02m, Ano07t, Bor02]. Service [Oli09d]. Seshadri [Bas00]. Set [Zha05]. Setting [Ke07]. Settings [GJS09]. Setup [LCdC06]. Several [FS08, WLER06]. Shaarawi [Zie02y, Ano05-41]. Shade [Law01]. Shaefer [San08a]. Shafer [Maz00]. Shaikh [Neu08c]. Shaked [Ano08q]. Shalabh [Ano09k]. Shanbhag [Ano04j]. Shane [Hor09a]. Shanthikumar [Ano08q]. Shao [Con01, Ano04-28, Boo05a]. Shape [Epi08]. Shapes [Ano07h, Weh04]. Sharon [Zie00-39]. Shawe [Bre05a]. Shawe-Taylor [Bre05a]. Shayle [Von03]. Shayne [Zie00-48]. Shein [Ano04-28, Ano05-54, Zie00j, Zie01q]. Shein-Chung [Ano04-28, Ano05-54, Zie00j, Zie01q]. Sheldon [McG07, Sau05]. Shephard [Ano06-27]. Sherali [Mur07]. Sheskin [Ano07i, Nas01, Zie04i]. Shetty [Mur07]. Shewhart [RS05]. Shi [Buk09]. Shiboski [Gol06b]. Shifts [HZ05]. Shin [Alt06]. Shirley [Bre05b]. Shizuhiko [Bur08a]. Short [BH08, BCBS08a, BCBS08b, Gol08, HB08, Jon08, KL08, TH05, ŽH09]. Short-Range [ŽH09]. Short-Run [TH05]. shortfall [DGS07]. Shoukri [Zie00-45, Nag04]. Should [RS05]. Shuttle [FK00]. Sidani [Neu09b]. Sided [KM04, LLI05, UL07, BG06]. Siegmund [Qu08, Ram00a]. Siem [Ano06-27]. Sigma [Ale01, Ale02, Ano03-36, Ano03-35, Ano04u, Ano06-36, Bai07a, Bzi04, Gar00, Gol04, Kal03, Kal06, Mor04, Mor05b, Rum06, Won04, Zie02-48, Zie02-51, Zie02-47, Zie04k, Zie04o, Zie04p, Ano03q, Ano04m, Ano05p, Sau07, Zie00-51, Zie01x, Zie01m, Zie02-49, Zie04o]. Sigma^{T M} [Ano03-48]. Signal [Bab05, Cab06, Cha02b, Con05, HPP01, HC00, Liu06, ME05, Sca02, Zie01n]. Signals [Ano00e, BJL^{+}07, Hen01, SBD98]. Significance [Ahm04]. Significant [Lan01]. Sik [Cha08b, CL03]. Sik-Yum [Cha08b]. Silverman [Ano06-41, Alt03]. Simar [Hec05]. similarity [FVB02]. Simon [Aba01b, Ano05-54, Chi07, Cob02, Cob04, Qia02, Gol07]. Simonoff [Das04]. Simple [Chi01, IM02]. Simplified [Ale02, Bre07, Rut00, Sta00, Con00]. Simplifying [RF05]. Simulation [Ahm08d, Ano07g, Den06, Hes04, Lew04, Ode00a, Pic05, Zie00-30, Ano04j, Zie00v, Bur09a, Qia07]. Simulations [LHB^{+}02]. Simultaneous [HSR09, JM01, TVW05]. Sincich [Zie00-38]. Singer [Ano06-39, Ano05-29]. Singh [Car07c]. Singhal [Ano06n]. Single [HCSC07, KJ09, ST02, ZZJ07, vB02]. Single-Case [vB02]. Single-Error-Term [HCSC07]. Singpurwalla [Aba01b, Gho08b]. Singular [Ruk01]. Sinha [Ye00, Zha05, Zie02s]. Sir [Ano07r, Ano07s, Gho07]. Sirken [Zie00e]. Site [Ste09b]. Sitter [Ste07]. Sivo [Ano03-34]. Six [Ano03-48, Ano03q, Ano04m, Ano05p, Ano06-36, Gar00, Gol04, Kal03, Mor05b, Rum06, Sau07, Won04, Zie00-51, Zie01x, Zie01m, Zie02-48, Zie02-49, Zie02-47, Zie04k, Zie04o, Zie02-51, Ale01, Ale02, Ano03-36, Ano03-35, Ano04u, Bai07a, Bzi04, Kal06, Mor04, Zie02-51, Zie04o, Zie04p]. Sixten [Kar07c]. Size [CHvdW06, LSW09, MEM05, TTL07, TJ09, Zie04t]. Sizes [Ahm06, GV03, Xu09]. Skew [Sen05]. Skew-Elliptical [Sen05]. Skiena [Cha02a]. Skillicorn [Cha09]. Skinner [Zie04d]. Skrondal [Lip05]. Slaughter [Hor09c]. Sliced [LN06]. Sloane [Den00]. Slutsky [Ano05p]. Small [Ahm07c, BKG09, vB02, CG09, Gan06, HG06, IM05, VD07, WW00, Xu02, Ano04v]. Small-Area [Ahm07c]. Small-n [vB02]. Small-Sample [WW00]. Smarter [Gar00]. Smeeton [Bar09, Ran03]. Smeyers [Zie00u].

Smilde [Boo05b]. Smirnov [DGS07]. Smith [Ano06-40, Sca02, Fin02, Fri00, Gol06a, Nem02a, Soy04, Ste00b, Zim02]. Smola [Ano08n]. Smooth [Glu06, Pol01, She09]. Smoothing [HCSC07, Qiu04, RSB09, Bak00, Fre03, Zie01-37]. Smyth [Owe04]. Snee [Ano04m, Zie02-42]. Soares [Zie01t]. Social [Ano03-45, Ano05-42, Jar09, Mai05]. Soft [Mul04, Zie02-40]. Software [Aba01b, Aba01a, BE03, Hin06, JLSZ08, Mai01, RLR06, Sar00b, Gun03, Hec09, Zie00b, Zie01-28]. Software-Based [Sar00b]. Solids [Nem02a]. Solka [Ano06s]. Solomon [Mic03]. Solutions [Gar00]. Solving [Per04, Sar05, Zie02-51, Ste00b]. Some [CGHL01, Con04, Mal00, RS06, Woo00, KT03]. Somersalo [LH06]. Soofi [Lu04b]. Soong [Yeh05]. Sophia [Fri01, Lip05, Zie02n]. Søren [O’N03, Zie01h]. Sorin [Ano05-53]. Souraya [Neu09b]. Source [Con05, SA08]. Sources [ASW^{+}06]. Space [Ano06-27, CL07, FK00, Hin05, JS07]. Space-Filling [CL07]. Space-Shuttle [FK00]. Spaces [SR04]. Spall [Hes04]. Sparks [Zie01-47]. Sparse [CCD09, LN06]. Spatial [AL03a, Cob02, Iye02, PC01, Sym01, Sym05, WB05, Woo00, ZCC08, Zie01-42, Nas04, Ros04, Zie04q]. SPC [JTW00, Pla02, McC07, Sta00, Yat02]. Special [BBS09, Nai07]. Specification [Rao01, PDDL02]. Spectral [Ano05z, FS08, Zie04r]. Spectrum [HTMW08]. Speed [Li09, Ano03-49, Cha01a]. Spencer [Ano06-42]. Sperlich [Ano07v, Lor05]. Spheres [JS07]. Spiegelhalter [Ano05-56, Ped01]. Spirer [Eva01]. Spirtes [Bur03]. Spline [Bak00, Fre03, RSB09]. Split [BS03, But04, GV03, Ke05, KCV02, MB04, PKV07, RACH09, KB05]. Split-Lot [But04]. Split-Plot [BS03, GV03, Ke05, KCV02, MB04, PKV07, RACH09, KB05]. Splitstone [Zie04u]. Sports [Ano06-38, Fle09, Zie00-50]. Sprent [Bar09, Ran03, Mag00]. Springer [Liu07b]. Sprott [Mul01, Sca05]. Spurrier [Bak01]. Square [Sco01]. Squares [Bor01, MJH00, Qiu05]. Sr. [Che08c]. SriRaman [Din05]. Srivastava [Kat03b]. SSA [Ruk02]. Stability [TTL07, TJ09]. Stacey [San08a]. Stage [Ano08s, BFL02, BMS09, VDLB08]. Stamatis [Zie02-51, Ano03-35, Ano04u, Bzi04, Mor04, Zie02-51, Zie04o, Zie04p]. Standard [YS05, Zie01-38]. Standardization [Ano03-37]. Stanley [Ano05-29, Bre05b, Con02a, Lip07a, Zie01c, Zie04y]. Staroswiecki [Bar07b]. Start [Ano05-28, Ano08j, LC00]. Started [SB08]. Starting [HMT07, SJ02]. Stat [Cha01a]. STATA [Das00, Ahm08a, Fri01]. State [BGMS03, GS08, Hin05, Ano06-27]. State-Dependent [BGMS03]. STATGRAPHICS [Zie00-42]. Statistic [Ano00e, GS08, KL02, Peñ05, SBD98]. Statistical [Ahm04, Ahm07b, Ahm07a, Ahm08a, Ahm08f, AW01, Ano03o, Ano03j, Ano04-30, Ano04e, Ano04q, Ano05-32, Ano05-52, Ano05-27, Ano05j, Ano05-31, Ano05-51, Ano05h, Ano06w, Ano06e, Ano06-35, Ano07r, Ano07s, Ano07e, Ano07i, Ano08l, Ano09m, Ano09f, Bab02b, Bar08b, Bar09, Bat08, BCL^{+}09, BGMS03, Boo09b, BJL^{+}07, Cab06, Car00, Cha08a, CL03, CW08, DLM^{+}07, Dok08, Edg04, Fri01, Gan02, Gar00, Geo03, Gho01, Han08, HZ05, Hec05, Hec09, Hin06, Hly09a, Jar08, Kaf01c, KML04, Kat03a, Kat08a, Kim00, Kim05, Kuh03, Kuh05b, LaL01, LI08, Law01, Laz03a, LEH03, Lit01, Liu06, Liu09a, Mag00, MM00, Mai05, Maz06b, McC02, McG02, MNS04, MBB05, Mor04, Nas01, Nem04, Old08, Oli07, Per04, Ram00a, Ram00b, Ran03, Rao01, RHOYL06, Rob00b, Rob06, RWZF04, Rum00]. Statistical [Rut05, Sar05, Sar00b, Sau02a, Sco01, Sea05, Sin06, Smi02, Spi07, Ste00a, Ste04, Sym01, Taa02, Var00, Weh02, Weh04, Wik03, Wil08a, Wil08b, Wil05, WLLZ07, Yat01, Yat02, Yu07a, ZH06, ZSC06, Zie00m, Zie00f, Zie01v, Zie01e, Zie02i, Zie02j, Zie02-30, Zie02e, Zie03c, Zie04i, Zie04x, MR01, Ano03-49, Ano06-40, Aba01b, AC02, Ann01, Ano03-40, Ano03-39, Ano03-41, Ano03-42, Ano03-43, Ano03-38, Ano04w, Ano04z, Ano04-27, Ano04y, Ano04x, Ano05-46, Ano05-45, Ano06-42, Ano06-43, Ano06-28, Ano06-29, Ano08p, Ano09l, Bai04, Bel01, Cam02, Che01b, Che07, Cob04, Con00, Dav03, Fle09, Fri00, Glo01, Glo06, Gol04, Gol07, Gup00, Hap00, Hof00, HR05, Hus08, Jar09, Joh03, Kaf03, Kar07f, Kat09, Ke06, Ken03, Ket00, Kus03, Laz03b, LH06, McC06a, McC05b]. Statistical [Mit02, Mul01, Mye09c, Nem08, Neu01, Neu03, Neu08c, O’N05, Owe00, Pet03a, Pla01, Pry03, Qia05, Qiu02, Rom06, Sen00, Sym05, Utl05, Wlu03, Ye00, Zie00-45, Zie00-44, Zie00-46, Zie00-47, Zie00-42, Zie00-43, Zie01-43, Zie01-41, Zie01-44, Zie01-39, Zie01-42, Zie01-40, Zie02-42, Zie02-41, Zie02g, Zie03b, Zie04u, Zie04s, Zie04t, Zie05a, Zim02]. Statistically [Car07c]. Statistician [Dok09, Ahm07c]. Statisticians [Ano05-49, Zel00]. Statistics [Ami04, Ann03, Ano03-35, Ano03p, Ano03-28, Ano03-32, Ano04j, Ano04k, Ano05g, Ano05-39, Ano05u, Ano06-38, Ano06u, Ano06s, Ano06-40, Ano07s, Ano07l, Ano08j, Ano09i, Ano09b, Bak01, Bar08b, Bod02b, Boo09b, Bri04, Bri05, Car07b, CGHL01, Cha01a, Che08c, Cla04a, Con05, Das07b, Dra04, Fah07, FR07, Geo00, Gho07, GLV02, Goo02, Gru08, HI00, HG02, Jen09b, JL07, Ke08, Ker03, Kib07, Kim04, Kus02, LaL04b, Lip06a, Lu08, Lum01, McC03b, Mic01, Mor05b, Mye05a, Nai07, Nas04, Nel00a, Ng05, Ng09, Ode00b, Oli04, Oli09a, Qu08, Rag01, San03a, San08a, Sau02b, Sau05, Sau07, Sho03, SBF^{+}08, Van04, Ye03, Yeh05, Zie00d, Zie00x, Zie00y, Zie00z, Zie00w, Zie00-38, Zie00b, Zie01-27, Zie01-38, Zie01i, Zie01l, Zie01-34, Zie01e, Zie02q, Zie02p, Zie02b, Zie02r]. Statistics [Zie04q, Zie04m, Zie04h, Zim01, Ano06-37, Ahm08e, Ano03g, Ano03-47, Ano04l, Ano04h, Ano05t, Ano05-44, Ano05-30, Ano05w, Ano05-28, Ano06-37, Ano07w, Ano07t, Ano08h, Ano08i, Bai05, Boo04b, Bur01a, Cha04, Eva01, Fot04, Fot07a, Gan05, Gro00, Hec01, Hen01, Kar07d, Kat03b, Liu07b, Mai08, Mar06, McG07, Neu09a, Owe03, Owe06, Rut09, Sea02, Sea09, Smi06, Was02, Yeh07, Zie00t, Zie01-50, Zie01q, Zie01g, Zie01w, Zie01-32, Zie02-34, Zie02-27, Zie02-39, Zie04b, Zie04y, Ano03-51, Ano04-28, Ano05-47, Ano05-57, Ano05-48, Ano06-30, Baj00, Bre05b, Bri03, Bur08b, Das00, Hol01, Hor06, Huz03, Jos04, KS07b, LaL09, Laz05, Min02, Neu05, Ng06, Pig06, Rum04, Was04, Zie00-48, Zie00-49, Zie00-50, Zie01-47, Zie01-46, Zie01-48, Zie01-45, Zie02-45, Zie02-44, Zie02-43, Zie04w, Zie04v].

Statistics [Zie05b]. Stats [Sch04, Sho05]. Stauffer [Gat02]. Steady [GS08]. Steady-State [GS08]. Steele [Ano03-45]. Steen [Boo07a]. Steenstra [Ano03-37]. Steepest [MX08]. Stefan [Hly07, Laz03a, Lor05, McC06a]. Stefano [Bur09a, Was05]. Steffen [Ped01]. Stein [Che08c, Che08b, Chi06, Zie00-44, Woo00]. Steiner [McC06a]. Step [Ano06-37, CE02, Van04]. Step-by-Step [Ano06-37, Van04]. Stepanski [Ano06-37, Ano06-37]. Stephan [Zie00-38, Zie01-45]. Stephane [Ano09h]. Stephanie [Zie04-27]. Stephen [Ano03-34, Ano04-31, Ano07q, Ano08r, Gol06b, Gol07, O’N03, Ray03b, So01, Zie01-38, Zie01d, Zie01h, Zie01-48, Zie02v]. Stephens [Ano07l, Hem02]. Steps [Sta00, vN02]. Stereology [Ano05-49]. Stern [Gre04]. Steve [Kuh05a, Zie00-33]. Steven [Ano03-43, Cha02a, Lum01, Rob06, Sca07, Zie00-47, Zie01-41, Zie01m, Zie02-38, Zie02q]. Stewart [Zie06, Ano05x, Qu07a]. Stigler [O’N03, Zie01h, Zie01-48]. Stochastic [Ahm08d, Ano04j, Ano07s, Bur01b, Bur09a, Cab03, Dob02, FFG06, Fot09, Hes04, Hly09b, HC00, Huz01, Ke09b, Lip06a, McC03a, McC05b, Ode00b, Ros05, Vol05, Ano08q, Fre00, Hor07, Hua00]. Stoica [Ano05z]. Stokes [Zie01j]. Stone [Mai07]. Stopping [DM08a]. Stork [Bar02]. Stoyan [McC03a]. Strategies [BN01, Gib01, Hes02, Hof00, MNS04, Wal03]. Strategy [Ale01, BFL02, Jia03]. Stratified [GHB^{+}07]. Stream [Buk09]. Streams [GS07]. Strength [KMM08, RWZF04, Sha04]. Stress [KMM08, Pas06, Sha04]. Stress-Life [Pas06]. Stress-Strength [KMM08, Sha04]. strings [FVB02]. Strip [VB09]. Strip-Block [VB09]. Stroup [Moo03]. Structural [Fin05, Lip05, Lip07c, Ano06-31, Cha08b]. Structure [Ahm08c, Ruk02]. Structured [Ano08n, Ros05]. Structures [Mar06]. Stuart [Ano06v, Law01, Pig06, Smi02, Sar05]. Student [Van04]. Studies [Ano03-50, Ano07e, Gun00, HG06, Ye03, Zie00d, Zie00j, Zie04-27, Ano03-34, Ano03-30, Maz06a, Zie01-49]. Studio [Ano08o]. Study [Ano06o, Bar08a, Che06, Lin02, Mye09c, ZDCS03, Zie00n, BF08, Bri01]. Stufken [Den00]. Stulajter [McC03c]. Subhabrata [Ano04q]. Subhash [Weh04]. Subir [Bar07c, Zie00-35, Zie00c]. Subject [CE02, CR06]. Subjective [Ker03]. Subjectivity [Gru02]. Suboptimal [PW01]. Subsampling [Str01]. Subset [ÖBMT02, Ano03-44]. Subsets [BSC09]. Sudha [Ahm07b]. Sue [Ano05-37]. Sujay [Sar06]. Sujit [Zie01s]. Sulfur [dCGM05]. Sullivan [Ano05-51]. Sum [GS08, RS05, Umb01]. Sumantra [Sen00]. Sung [Zie01-43]. Supercomputer [GL09b]. Supersaturated [AB03, BFL02, NC08, LD04]. Supplemental [Ste09b]. Support [BJL^{+}07, Oli08]. Suren [Ano06n]. Surface [ACBJ09, ALKP02, Ano07o, BV01, FM04, Khu06, Mor00a, MBDD09, TG01, YS05, Zie02-36]. Surfaces [Ano07p, GD03, Qiu04, RF05]. Surrogate [Ano06-43]. Survey [Ahm07c, Ano03r, Ano03k, Ano04s, Ano05-29, Bar08a, Con04, Kar07b, Ke09a, LaL04a, Zie00-35, Zie00e, Zie00h, Zie04d, Ano05-50, Zie02-46]. Surveys [Kar07c, Ano05-38, Zie01b]. Survival [Ano03-40, Ano04o, Ano05u, Ano07e, Fin02, Gol06b, Ken02, LZ02, Liu04, Liu07a, ST06, Zie01c, Zie02s, Ano06-32, Ano07u, Rav04]. Susan [Ano05n, Hor09c, Zie00e]. Susanne [Ano04x, Zie04f]. Sushmita [Mul04]. Svolba [Nas08]. Swersey [Oli09d]. Sylvan [Zie02-39]. Sylvia [Ros05]. Sylvie [Nem08]. Symmetric [MBDD09]. Symmetry [Smy00]. System [BMS09, Glu06, KL02, KD05, VD08, VD07, WKT^{+}03, Zie01-33, Ano08o, Zie00-31, Zie01j, Zie04x]. Systems [Che06, Huz00, JW02, LEH03, Mam09, NK03, Ste00a, CFP08, Ano06-33, Ano09h, Ano09l, Fre00, Huz01, Ped01, Ros05, Zie01-41, Zie02-29]. Székely [Zie01-46].

T [Ahm07c, Ahm08b, Ano03-40, Ano03t, Ano03-39, Ano03-31, Ano04e, Ano05-29, Boo07b, Bur09b, Cam01b, Car00, Che01c, Che07, Che08a, Dob02, Hec09, Kim05, LZ02, Lip09, Mag09, Maz06b, Mye05a, Ng09, Rom03, Ros04, Sau07, Sca02, Soy04, Sta00, Weh04, Won04, Yeh05, Zie00l, Zie01i, Zie02-45, Zie02-30, Zie02-32, vT00]. Table [Che03, Zie01-48]. Tables [Ano07h, Zie01-38]. Tacit [Ano03-37]. Tadikamalla [Lew04]. Taguchi [Alt06, Bar07c, Jia03, vN02, Alt06, WKT^{+}03, Bar07c]. Tague [McC08]. Tai [Kuh03]. Tails [McC00]. Takayuki [Mar06]. Takeaki [Qiu05]. Takezawa [Li06]. Taking [So01]. Tamhane [Hol01]. Tan [Zie00-37]. Tanimoto [FVB02]. Tanizaki [Ng05]. Tanner [Zie02-45]. Tanniru [Sin01]. Tanur [Zie00e, Gru02]. Taqqu [McC00]. Tarantola [Was05, Che08a]. Tarconi [Soy09]. Target [JW02, KLM08]. Tarter [Qia05]. Taskar [Ano08n]. Tatham [Qu07b]. Taylor [Bre05a, O’N05, Che01b]. Teaching [Boo04b]. Technical [Ano09l, TJ09]. Technique [RSS02]. Techniques [AB04, Bab05, Jar09, JLH^{+}06, JS03, Liu04, Lu04b, Mye09b, O’N05, Qia07, Sar00b, Ano06k, Bre07, Dog01, Nem04, Ruk02, Zie01o]. Technologies [Ano08t, Bar07a]. Technology [Boo09b, Joh00, Nai07, SS07, Yu07a]. Technometrics [Bri06, Buj06, Efr06, Hub06, IM02, Lan06, Hun00, Kaf00c, Mor00b, Ste08a, SB08]. Teich [Sca07]. Telecommunications [Ano04i]. Tell [Gol07]. Temperatures [LWS03]. Temporal [Kar07a]. Terence [Zie02-32]. Terje [Hua00]. Term [HCSC07]. Terms [Ano04-30]. Terrell [Geo00]. Terrence [Rao00]. Terry [Alt02, Ano03-49, Ano05f, Cha01a, Zie00-38]. Test [Ano05-52, CHvdW06, PM03, Pas03, Pas06, SEM09, DGS07, MW05]. Testing [Ano05-52, Ano05-29, BK04, BKG09, KBS08, Kun08, LP06, McC06b, ML01, PM03, Per04, Sin05, Sta04, Van02, MN06, Ann03, Ano05y, Che03, Kap03, Laz03a, Oli09d]. Tests [Ahm04, Ano05-35, Ano05-51, BDNN04, CS01, Din05, JLSZ08, Kar07e, MEM05, MEH09, NM04, PS00, Rob00a, ST02, Ye00, ZM06, Zha06, Bab02a, Zie00-34, Zie01-30]. Teugels [Laz05]. Teukolsky [Lip09]. Text [GLM07, JL07, Ano03v, Ano06-32]. Their [HJMW04, Kus00, Mye08, Sen05]. Theodore [Ano05-57, Ano06m, Sau07, She07, Won04, Zie02w, San08a]. Theoretic [Oak09]. Theoretical [Ano07s, Pro00]. Theory [Ahm08g, Ahm08e, Ano06-28, Ano08k, Bar08c, Cab08, Car07a, Che01a, Che06, Con04, Den00, Fam03, FLWZ00, Fok07, Gho02, Gru08, Kar07f, Law00, Laz05, Liu09b, Loe07, Lor01, Lu07, McC06c, MS07, Mic05, Mur07, Mye05b, Nel00b, Oli09a, Ord07a, Ord07b, Pas03, Pic06, Qia00, Qia02, Qu07a, Sca04, Sha04, Tor01, Woo00, Zha05, Zie00y, Zie04x, Zim00, Ahm08f, Ano09e, Zie00k]. Thermosphere [RMGR09]. Thermosphere-lonosphere [RMGR09]. Therneau [LZ02]. Theus [KS07a]. Thinkers [McC07]. Thinking [Ano06i, Ano06-29, Fle09, Cho06, Law01, Zie02-42, Car07c]. Thode [Kap03]. Thomas [Ahm04, All01, Ano05-34, Ano05l, Ano08n, Den04, Eas07, HL06, Liu07a, Mor05a, Sea01, Ye00, Zie00-51, Zie01-40, Zie01f, Zie04e, Ano08p, Cam02]. Thompson [Ano03-27, Kaf03, Zie02-38]. Thomson [Jar08]. Three [CW02, MBC00, Mor00a, Mor00b, XCW04]. Three-Level [CW02, Mor00a, XCW04]. Throughput [Zie03a]. Thumb [Pry03]. Thuraisingham [Zie00i]. Tiao [O’N03, Zie01a, Zie01h]. Tibshirani [Ano09m, Zie03c, Ahm09]. TIE [RMGR09]. TIE-GCM [RMGR09]. Tijms [Vol05]. Tim [Bas02, Boo05a, Pet03a, Zie01-47]. Time [Ahm07b, Ano03z, Ano03s, Ano04-29, Ano04t, Ano04f, Ano05-34, Ano06j, Ano06f, Ano07s, BDNN04, Bre06, Bur07, Cab00, Cam01b, Cha01c, Cha02b, Che01c, COR08, FFG06, Fot07b, FW06, Geo03, JS07, Kar07a, Lu04a, McC03c, McC05a, Nas02, RCGD06, Ray03a, RLR06, Ruk02, Sar02, Sca05, She09, Zie00c, Zie01a, Zie02m, Zie04r, McC05b, Ano06-33, Bor02, Fot01, Hin05, Mye05b, Ruk01, Zie02-50]. Time-Dependent [RLR06]. Time-Frequency [Cha02b, Nas02]. Time-Series [Sca05, Zie04r, Bor02]. Time-to-Event [Bre06]. Time-Truncated [BDNN04]. Time-Varying [RCGD06]. Times [McC06b, Bri01]. Timm [Ahm08g, Boo03]. Timothy [Bri03, Zie01-49]. Tinku [Mul04]. Tinsley [Rob06]. Titterington [Zie02t, Zie00-49]. Tobias [Nag06, Zie00-34, Hor09b]. Todd [Gho01]. Todinov [Ke07]. Todman [vB02]. Together [Bak01]. Tolerance [FG01, HJMW04, KM04, KMM08, LLI05, WT09, BF08]. Tolerant [Bar07b]. Tom [Bre07, Dan01, Zie04a]. Tomas [Zie04a]. Tomasz [Ano06-43]. Tomer [Ano06-31]. Tomography [Mai01]. Tong [CL03]. Tony [Kar07e, Zie04a]. Toolbox [McC08]. Tools [ACBJ09, Ano04-27, Ano05-53, Bre07, Hof00, MSM03, PYC07, SB04, Zie04u, Ano07x]. Top [Ale01]. Topics [Ano07o]. Tormod [Zie04a]. Torresani [Nas02]. Torsten [Liu09b, Ord07a, Ord07b]. Torster [Ano06e]. Tourangeau [Zie00e]. Toutenburg [Bor01, Wlu03]. Townend [Ano05-39]. Townsend [Esp07]. Toxicologists [Zie00-48]. Track [Ano06-30, Bur08b]. Tracking [JL07, Ano09c]. Tradition [Hun00]. Tragic [Ste07]. Transform [Boo05a]. Transformation [Atk05]. Transformations [Coo00b, CO01, RA00]. Transforms [Cha02b, Nas02]. Transient [BJL^{+}07]. Transport [Li05]. Transportation [Cob04]. Tranter [Nem02b]. Treatments [FHR06]. Trees [Ano09c, Jen08]. Trend [LEH03]. Trend-Renewal [LEH03]. Trends [Ano07h]. Trevor [Ano03-29, Ano09m, Zie03c]. Trewn [Kuh05b]. Trial [Hec01]. Trials [Ano04-31, Ano05-56, Che07, Chi06, Gol07, Zie04-27, Ano04-31, Zie04-27, Ano03-33, Ano03f, Ano05-54, Hus08, Sin07, Zie02w, Zie04l]. Triola [Bai04].

Trip [Pla01]. Trivedi [Hly07, McC03b]. Troubleshooting [Wlu06]. Troutt [Cab09]. Truncated [BDNN04]. Trusting [Zie02h]. Tsai [Cab00]. Tsay [Ano06f, KS07b, Zie01a, Zie02m]. Tsiatis [Fok07]. Tufte [Mul02]. Tukey [Bri06, Buj06, Efr06, Hub06, Lan06, Zie01-52, Ano07h, Hec01, Kaf01b, Mal06, Tho01]. Tuning [HSR09, LSB06, dC02]. Turkman [Zie00-44]. Turner [Zie00p]. Tutorial [Mai07]. Tutz [Zie02e]. Twisk [Ano06-39]. Two [Ano08s, BBJK04, BFL02, BMS09, But04, CL09, CW02, CLY04, JLSZ08, Law00, LL03, LLY03, LLI05, Ste00a, UL07, Var00, VDLB08, Zha06, Jon04, KT03]. two-component [Jon04]. two-factor [KT03]. Two-Level [But04, CL09, CW02, CLY04, LL03, LLY03]. Two-Sample [Zha06]. Two-Sided [LLI05, UL07]. Two-Stage [Ano08s, BFL02, BMS09, VDLB08]. type [DGS07, JM00]. Types [Kar07e].

U [Cab03]. Uduehi [Bar08c]. Ugarte [Ng09]. Ullah [Zie00t]. Ultimate [Ano03-48, Zie02-49]. Unbalanced [KM04, LLI05, McC06c]. Unbounded [GS08]. Uncertain [Ano03l, AK04]. Uncertainties [RCGD06]. Uncertainty [AL03b, GJS09, Gri07, Sen06a, Smi01, WLER06, Bre02, Gat02, San07, Car07d, San08b]. Unconditional [BDNN04]. Understanding [SB04, Yat02, Cha09, Zie01-52]. Unequal [ND02]. Unified [AW05, Che07, Den06, Geo00]. Uniform [FLWZ00, LWY02]. Units [GM01]. Univariate [Ano06-37, For08, Qu07a, Ahm08g, Ano06-34]. Universes [ZC06]. Unknown [GPR05]. Unmixing [Con05]. Unobserved [Ano06-27]. Unreplicated [BN01, ML01, Mil05]. Unwin [KS07a]. Update [Zie00m]. Upper [BG04, KL02]. Urdan [Bri03]. Usage [CLL09]. Use [BBJK04, KB05]. Used [RS05, WW00]. useful [Bri01]. User [Ano06-41, Zie04a]. User-Friendly [Zie04a]. Uses [Bru00]. Ushakov [Sen00]. Using [Ahm08a, Ano03-43, Ano06e, Ano06q, Ano06-37, Ano06-40, Ano08j, Ano08o, Ano09f, BCL^{+}09, Cab04, Che01b, Che08c, Chi06, CG09, DT02, Fri01, Gar00, GCS06, GM01, GHB^{+}07, Gun03, Gun04, HZ05, Hec09, HTMW08, Huz00, Jen08, Jen09c, Jen09a, Jen09b, JS03, JP01, Kus03, LaL09, LLBN08, LS05, Li05, LS00, LWS03, McC03c, MEH09, MS01, Mil05, MS05, MMM08, Nas08, Ng06, RdC05, RLR06, RHOYL06, RF05, Sán06, Sau02a, Sin08, Ste00a, Van04, Van02, WM02, ZZ07, Zie00-47, Zie00-33, Zie00-38, Zie00-42, Zie00-31, Zie01j, Zie01-45, Zie02i, Zie02j, Zie02n, Zie04n, ZTW07, vN02, FVB02, KT03, LCdC06, Ano06-35, Smi06]. Utts [Ano05-30, Sea02, Ano05-44]. Uwe [Hua00].

V [Ahm07b, Ahm08b, Ano05r, Ano08n, Ano09b, Ano09a, Baj00, Bas00, Bre06, Bur01b, Che02a, Fot09, Gol06b, Huz01, Kus00, Mai07, Mai08, Mar03, Mye01, Pic07, Qia02, Ruk02, Wlu06, Zie00p]. Valentin [Cut00]. Validation [BBP^{+}07, WCT09]. Value [Bla01, FHR06, FW06, MS07, Tom07]. Values [Ano04i, Ano08p, Gui08, MBC00, MW05, Cam02, Smi02]. Vandeginste [Zie00u]. Vardeman [Zie01d]. Variability [AS01, AK04]. Variable [Ano09j, BG04, LBH^{+}06, Lip01, Lip05, LSB06, OH05, RSB09, TVW05, YJL07, ZC06].

Variables [CR06, HSNB09, KCV02, MBC00, MJH00, PK05, WT09, WLER06, Ano05-42]. Variance [Ami04, ALKP02, AB04, Bed00, Chi01, GP06, HZ05, HCSC07, MM00, McC06c, Mic05, RS06, Zie04j, IM02, MW05, Hec08, Mat01, Mic03, Che02b]. Variances [Khu06, ND02, RCGD06]. Variation [AL03a, Buk09, KML04, McC06a, SA08]. Variogram [RG07]. Varying [RCGD06]. Vasiliy [Olw01a]. vCJD [Zie01-39]. Veaux [Sch04, Sho05]. Vedel [Ano05-49]. Vehicular [Li09]. Velleman [Sch04, Sho05]. Venables [Ano03-28, Zie01-27]. Venkatarama [Fot07c]. Verbeke [Avi01, Zie00u]. Verdinelli [Zie00d, Ye03]. Vere [Sca04]. Vere-Jones [Sca04]. Verification [Sta04]. Vertical [Cab09]. Verzani [Smi06]. Vessels [FK00]. VI [Ano07g, Ye03]. via [Cas05]. Vic [Ano04s, Ano05-41, Zie00f, Zie00-44, Zie04h]. Victor [Boo09a]. Vidakovic [Ket00, Lu08, Sen02]. Vidal [Ano05-56]. Vieu [Lu07]. View [BJL^{+}07]. Vijver [Ano03k]. Vila [Ano06r]. Ville [Jen08]. Vincent [Huz03]. Vining [Ano07j, Gra02, Zie01-43, Zie02-31]. Virginia [Ami04, Zie04f]. Vishwanathan [Ano08n]. Visioli [Neu08b]. Visual [Mul02]. Visualization [Ano05o, GCS06, Ano06s]. Visualizing [CO01, KS07a, Kat03a, Zie01-53]. Vittinghoff [Gol06b]. Vladimir [Ano08k, Lor01, Ruk02]. Vlatko [Lew04]. Voakes [Ano06-38]. Vol [Ano07r, Ano07s, Ano08f, Liu09b, Mic05, Ord07a, Ord07b, Sca04, Ye03, Zie00d]. Volcanic [BCL^{+}09]. Volume [Ano00j, Ano01i, Ano02i, Ano03-56, Ano04-37, Ano05-63, Ano06-49, Ano07-29, Ano08y, Bur02, McC06c, Xu06, Zie00m, Zie00-42]. Volumes [Ano06m]. Vonta [Ano09l]. Vose [Zie02-37]. Voss [Ami00]. Vuchkov [Rob02].

W [Ahm04, Ahm08c, Ahm08h, Alt03, Ano03q, Ano03-28, Ano03-37, Ano04e, Ano04m, Ano05m, Ano05-40, Ano05-41, Ano05-54, Ano06-34, Ano06g, Ano06-39, Ano06-41, Ano07h, Ano08m, Bar08c, Bre01, Cab09, Car00, Cha03, Che03, Cob02, Con02a, Gar00, Gat02, Glo06, Gro00, Gup00, Hec01, Hec08, Hec09, Hly09a, Hof00, Jos04, Kat03a, Law01, Liu02, Qu07a, Sca02, Zie00-49, Zie00-42, Zie01-27, Zie01r, Zie01-52, Zie01-31, Zie01c, Zie02y, Zie02f, Zie04k, Zie05b]. Wackernagel [Ano04p, Zie00-36]. Wagner [Ano05n]. Wai [McC05a]. Walczak [Zie01-54]. Walden [Sca02, Che01c]. Walgren [Bar08b, Bar08b]. Walker [Ano03j, Mor05b]. Wallace [Ano03-50, Olw01b]. Wallenstein [Zie02-39]. Walls [She07]. Walter [Ano04f, Ano06g, Chi06, Moo03, Zie02y]. Walters [Kal03]. Wand [Wal05]. Wang [Zie00-28, Ano03-40, Ano03z, Bar07a, Bur07, Fot07b, Gan06]. Ware [Ano05e]. Warner [Zie04r]. Warranty [CLL09, FL07, WM02]. Warren [Glo01, Zie02g]. Washington [Cob04]. Wasserman [Fot07a, Sta04, Zie04b]. Watanabe [Sin06]. Water [Ano06-35]. Wavelet [Boo05a, HC00, JLH^{+}06, Nas02, Sen02, Whi04, Cha02b, Che01c]. Wavelet-Based [JLH^{+}06, Sen02, Whi04]. Wavelets [Gho01, Ket00, Zie01-54]. Way [FKQS01, Gri04, KM04, LLI05, Zie02-48, Boo05b]. Wayne [Ano05-52, Mee03, Zie04g]. Wearden [Bre05b]. Web [Ste09b, Mag09, Owe04]. Webb [Dav03]. Webster [Zie01u]. Weerahandi [Gho05, Lit01]. Wegman [Ano06s]. Weibull [BSG00, Kun08, NCB04, NM02, Dog04]. Weight [Mul08]. Weight-of [Mul08]. Weighted [AL03b, CM03, CE02, GB01, HMT07, HMT08, JCR01, RG07, RS05, RS06, TTL07, TJ09, Zie06, ZTW07]. Weisberg [Gra00, Mal05]. Weiss [Bur06]. Welch [Hec09]. Well [BSC09]. Well-Formulated [BSC09]. Wellek [Laz03a]. Wells [Zie02-45]. Wen [Nas02]. Wen-Liang [Nas02]. Wenclawiak [Ano05-40]. Wendell [Gol06a]. Wendy [Ano08o, Kus02, Kus06]. Weng [Ano06h]. Wenyu [Ano03-40]. Wermuth [Gho07]. Werner [Iye02]. Werwatz [Lor05]. West [Hec09, Zie00d]. Westfall [Zie00-34]. Wetterling [Lip09]. Wheeler [Bar01, Jar04, Jen09c, Nel01, Pfa04, Rum06, Vuk00]. Whitaker [Ano06-29]. Whitney [Liu09b, Ord07a, Ord07b]. Whittington [Ano05i]. Who [So01]. Whole [GV03]. Wickens [Mor05a]. Wilcox [Ano05y, Bab02b, Nem04]. Wild [Kok01]. Wilder [Kaf01b]. Wiley [Mai05, Zie01-52]. Wilkinson [Bai07b]. Willem [Ke09a]. Willett [Ano06-39]. William [Aba01a, Ano03-33, Ano03-41, Ano05-56, Ano08i, Bai04, Den04, Gan05, Glo06, Kar07b, Lip09, Mye01, O’N03, Pig06, Qu07b, Rut02, Sar00b, Zie00o, Zie00h, Zie01h, Zie04s, Zie04k, Zie04y]. Williams [Ano03-27, Den04, Kal06, Kim04]. Willmot [Ano06v]. Wilrich [Zie02-30]. Wilson [Ano08l, Hor09a, Aba01b, Sen06b, Zie04s]. WinBUGS [Ano05-56]. Wind [Sán06]. Window [JW04b, ZMW08, Cla04b]. Window-Observation [ZMW08]. Winker [Zie01-29]. Winston [Zie04g]. Wireless [BJL^{+}07, KLM08]. Within [Khu06]. Within-Block [Khu06]. without [CL05]. Wittes [Che07]. Wludyka [Esp06b]. Woitek [Jar09]. Wojtek [Rob00b]. Wold [Nem03, Nem03]. Wolf [Str01]. Wolfe [Ram00b]. Wolfgang [Ano07v, Hec05, Lor05]. Wolfinger [Zie00-34]. Wolstenholme [Hil00]. Wolter [Zie04j]. Wong [Ano06h, DS00, Qia07]. Wood [Zie01r, Chi07]. Woodward [Ano06o, Zie00n]. Woodworth [Ano05k]. Woolf [Gru08]. Woolson [Ano03-41]. Words [Das07b, Das07b]. Work [GM01]. Working [Ano06-38]. Workshop [HI00, HG02]. World [Ale01, Ano03-27, Liu09b, Ord07a, Ord07b, Sar06, Dra04, Zie01-55, Zie04z]. Wright [Ano05-54]. Wu [Alt01, Bar07c, DS00, Loe07]. Wynn [Zie00-32].

X [Hen01, Sar00a, Zie01-35, MN06]. X- [MN06]. X-11 [Zie01-35]. Xavier [Ano06r]. Xiao [Ano03-42, Ye05]. Xiao-Hua [Ano03-42]. Xiao-Li [Ye05]. Xibilia [Zie02-40]. Xie [Hor07]. Xin [Kuh03]. Xitao [Ano03-34]. xla [Zie00-27]. Xue [Boo05a]. Xue-Guang [Boo05a]. xv [Mai05]. xviii [Zie05b]. xx [Zie01-52].

Yadohisa [Mar06]. Yadolah [Ano04-30, Gho07]. Yakir [Qu08]. Yamaguchi [Sin06]. Yan [Sha04]. Yang [Kuh05b, Ng08]. Yao [Lu04a]. Years [Kaf00c, Mal06, Bri06, Buj06, Efr06, Hub06, Lan06, Zie02t]. Yee [Ano03h, Qia02]. Yefim [Tob03]. Yeong [KS07b]. Yi [Boo05a]. Yi-Zeng [Boo05a]. Yin [Ruk01]. Ying [Bar07a, Qia07]. Yingdong [Ano05-54]. York [Zie01-52]. Yosef [Zie00-34]. Young [Von06, Edg04, Sca02]. Youngjo [Gho08a]. Yudi [Gho08a, Sau02a]. Yuin [Bar07c]. Yum [Cha08b]. Yuri [Laz05]. Yvonne [Ano08l].

Z [Gho02, Zie00h]. Zakkula [Kat08b]. Zappe [Zie04g]. Zaven [Weh02]. Zdravko [Mag09]. Zeger [Ano03-49, Ano03h]. Zehua [Zha05]. Zeileis [Mye09a]. Zelterman [Ano05q, Ano07m, Kia00]. Zeng [Boo05a]. Zhang [Qiu08, Ruk01]. Zhao [Ano05-54]. Zhengxin [Ano03l]. Zhidong [Zha05]. Zhigljavsky [Ruk02]. Zhou [Ano03-42]. Zhuang [Bar07a]. Ziberna [Pic07]. Zimmerman [Ano03-43, Zie00-47]. Zio [Kar07f]. Zivot [Ano03z, Bur07, Fot07b]. Zotteri [Boo09c]. Zoubir [Bab05]. Zouwen [Ano03-37]. Zünd [Ano05-46]. Zuo [Gri04]. Zurada [Ano05-55]. Zwillinger [Zie01-38].

References

Allen:2003:SDM

[AB03] Theodore T. Allen and Mikhail Bernshteyn. Supersaturated designs that maximize the probability of identifying active factors. Technometrics, 45(1):90–97, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

           jstor.org/stable/25046978.

Arbogast:2004:MCT

[AB04] Patrick G. Arbogast and Edward J. Bedrick. Model-checking techniques for linear models with parametric variance functions. Technometrics, 46(4):404–410, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

           jstor.org/stable/25470888.

Abate:2001:BRBb

[Aba01a] Marcey Abate. Book review: Measuring the Software Process by William A. Florac; Anita D. Carleton. Technometrics, 43(4):495, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

           2723 (electronic). URL http://www.jstor.org/stable/1270835.

Abate:2001:BRBa

[Aba01b] Marcey Abate. Book review: Statistical Methods in Software Engineering by Nozer D. Singpurwalla; Simon P. Wilson. Technometrics, 43(1):108, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270878.

Anderson-Cook:2002:BRB

[AC02] Christine M. Anderson-Cook. Book review: Statistical Consulting: A Guide to Effective Communication by Janice Derr. Technometrics, 44(1):89–90, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270701.

Anderson-Cook:2009:GTA

[ACBJ09] Christine M. Anderson-Cook, Connie M. Borror, and Bradley Jones. Graphical tools for assessing the sensitivity of response surface designs to model misspecification. Technometrics, 51(1):75– 87, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586565.

Ahmed:2004:BRB

[Ahm04] S. Ejaz Ahmed. Book review: Applied Adaptive Statistical Methods: Tests of Significance and Confidence Intervals by Thomas W. O’Gorman. Technometrics, 46(4):484–485, November 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470898.

Ahmed:2006:BRB

[Ahm06] S. Ejaz Ahmed. Book review: Effect Sizes for Research: A Broad Application Approach by Robert J. Grissom; John J. Kim. Technometrics, 48(4):573, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471257.

Ahmed:2007:BRBb

[Ahm07a] S. E. Ahmed. Book review: Handbook of Statistical Distributions with Applications by K. Krishnamoorthy. Technometrics, 49(2): 231, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471334.

Ahmed:2007:BRBa

[Ahm07b] S. E. Ahmed. Book review: Life Time Data: Statistical Models and Methods by Jayant V. Deshpande; Sudha G. Purohit. Technometrics, 49(2):231, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471333.

Ahmed:2007:BRBc

[Ahm07c] S. E. Ahmed. Book review: Missing Data and Small-Area Estimation: Modern Analytical Equipment for the Survey Statistician by Nicholas T. Longford. Technometrics, 49(3):367, August 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471373.

Ahmed:2008:BRBa

[Ahm08a] S. E. Ahmed. Book review: A Handbook of Statistical Analysis Using Stata by S. Rabe-Hesketh; B. S. Everitt. Technometrics, 50(1):96–97, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471445.

Ahmed:2008:BRBc

[Ahm08b] S. E. Ahmed. Book review: Bayesian Networks and Decision Graphs by F. V. Jenson; T. D. Nielsen. Technometrics, 50(1): 97, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471447.

Ahmed:2008:BRBe

[Ahm08c] S. E. Ahmed. Book review: Life Distributions: Structure on Nonparametric, Semiparametric and Parametric Families by Albert W. Marshall; Ingram Olkin. Technometrics, 50(2):237–238, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471474.

Ahmed:2008:BRBb

[Ahm08d] S. E. Ahmed. Book review: Markov Chain Monte Carlo: Stochastic Simulation for Bayesian Inference by Dani Gamerman; Hedibert F. Lopes. Technometrics, 50(1):97, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471446.

Ahmed:2008:BRBg

[Ahm08e] S. E. Ahmed. Book review: Matrix Algebra: Theory, Computations, and Applications in Statistics by James E. Gentle. Technometrics, 50(2):238–239, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471476.

Ahmed:2008:BRBf

[Ahm08f] S. E. Ahmed. Book review: Reliable Reasoning: Induction and Statistical Learning Theory by Gilbert Harman; Sanjeev Kulkarni. Technometrics, 50(2):238, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471475.

Ahmed:2008:BRBd

[Ahm08g] S. E. Ahmed. Book review: Univariate and Multivariate General Linear Models: Theory and Applications with SAS by K. Kim; N. Timm. Technometrics, 50(1):97–98, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471448.

Ahmed:2008:BRBh

[Ahm08h] S. Ejaz Ahmed. Book review: Encyclopedia and Handbook of Process Capability Indices: A Comprehensive Exposition of Quality Control Measures by W. L. Pearn; Samuel Kotz. Technometrics, 50(4):548–549, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471546.

Ahmed:2009:BRB

[Ahm09] S. Ejaz Ahmed. Book review: The Science of Bradley Efron, Selected Papers by Carl N. Morris; Robert Tibshirani. Technometrics, 51(1):98–99, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586568.

Apley:2004:CCI

[AK04] Daniel W. Apley and Jeongbae Kim. Cautious control of industrial process variability with uncertain input and disturbance model parameters. Technometrics, 46(2):188–199, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470804.

Apley:2003:ISV

[AL03a] Daniel W. Apley and Ho Young Lee. Identifying spatial variation patterns in multivariate manufacturing processes: A blind separation approach. Technometrics, 45(3):220–234, August 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047049.

Apley:2003:DEW

[AL03b] Daniel W. Apley and Hyun Cheol Lee. Design of exponentially weighted moving average control charts for autocorrelated processes with model uncertainty. Technometrics, 45(3):187–198, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047046.

Alexander:2001:BRB

[Ale01] Melvin Alexander. Book review: Six Sigma: The Breakthrough Management Strategy Revolutionizing the World’s Top Corporations by Mikel Harry; Richard Schroeder. Technometrics, 43(3): 370, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271225.

Alexander:2002:BRB

[Ale02] Melvin Alexander. Book review: Six Sigma Simplified: Quantum Improvement Made Easy by Jay Arthur. Technometrics, 44(2):189– 190, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271267.

Ankenman:2002:CED

[ALKP02] Bruce E. Ankenman, Hui Liu, Alan F. Karr, and Jeffrey D. Picka. A class of experimental designs for estimating a response surface and variance components. Technometrics, 44(1):45–54, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270683.

Allen:2001:BRB

[All01] Theodore Allen. Book review: Quality Engineering Handbook by Thomas Pyzdek. Technometrics, 43(2):243, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271057.

Altland:2001:BRB

[Alt01] Henry W. Altland. Book review: Experiments: Planning, Analysis, and Parameter Design Optimization by C. F. Jeff Wu; Michael Hamada. Technometrics, 43(3):369, August 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271224.

Altland:2002:BRB

[Alt02] Henry W. Altland. Book review: Applied Regression Analysis for Business and Economics by Terry E. Dielman. Technometrics, 44(2):192–193, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271272.

Altland:2003:BRB

[Alt03] Henry W. Altland. Book review: Applied Functional Data Analysis by J. O. Ramsay; B. W. Silverman. Technometrics, 45(1): 101–102, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25046982.

Altland:2006:BRB

[Alt06] Henry W. Altland. Book review: Computer-Based Robust Engineering: Essentials for DFSS by Genichi Taguchi; Rajesh Jugulum; Shin Taguchi. Technometrics, 48(1):153, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471136.

Amidan:2000:BRB

[Ami00] Brett Amidan. Book review: Design and Analysis of Experiments by Angela Dean; Daniel Voss. Technometrics, 42(3):308, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271085.

Amin:2004:BRB

[Ami04] Raid W. Amin. Book review: Applied Statistics: Analysis of Variance and Regression by Ruth M. Mickey; Olive Jean Dunn; Virginia A. Clark. Technometrics, 46(4):491–492, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470909.

Andrews:2008:BRB

[And08] Chris Andrews. Book review: Competing Risks: A Practical Perspective by Melania Pintilie. Technometrics, 50(3):411, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471507.

Annis:2001:BRB

[Ann01] Charles Annis. Book review: Statistical Distributions in Engineering by Karl Bury. Technometrics, 43(1):101–102, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270868.

Annis:2003:BRB

[Ann03] Charles Annis. Book review: Mechanical Reliability Improvement: Probability and Statistics for Experimental Testing by R. E. Little. Technometrics, 45(4):367–368, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047102.

Annis:2006:BRB

[Ann06] Charles Annis. Book review: Reliability, Quality, and Safety for Engineers by B. S. Dhillon. Technometrics, 48(1):151–152, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471134.

Anonymous:2000:BMa

[Ano00a] Anonymous. Back matter. Technometrics, 42(1):108–110, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271441.

Anonymous:2000:BMb

[Ano00b] Anonymous. Back matter. Technometrics, 42(2):??, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271496.

Anonymous:2000:BMc

[Ano00c] Anonymous. Back matter. Technometrics, 42(3):??, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271127.

Anonymous:2000:BMd

[Ano00d] Anonymous. Back matter. Technometrics, 42(4):446–447, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270988.

Anonymous:2000:CCS

[Ano00e] Anonymous. Correction: “A Cepstral F Statistic for Detecting Delay-Fired Seismic Signals”. Technometrics, 42(2):226, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271495. See

[SBD98].

Anonymous:2000:FMa

[Ano00f] Anonymous. Front matter. Technometrics, 42(1):??, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271423.

Anonymous:2000:FMb

[Ano00g] Anonymous. Front matter. Technometrics, 42(2):203–225, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271442.

Anonymous:2000:FMc

[Ano00h] Anonymous. Front matter. Technometrics, 42(3):307, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271077.

Anonymous:2000:FMd

[Ano00i] Anonymous. Front matter. Technometrics, 42(4):425, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270940.

Anonymous:2000:VI

[Ano00j] Anonymous. Volume information. Technometrics, 42(4):448–452, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270939.

Anonymous:2001:BMa

[Ano01a] Anonymous. Back matter. Technometrics, 43(1):117–118, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270898.

Anonymous:2001:BMb

[Ano01b] Anonymous. Back matter. Technometrics, 43(2):??, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271076.

Anonymous:2001:BMc

[Ano01c] Anonymous. Back matter. Technometrics, 43(3):??, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271252.

Anonymous:2001:BMd

[Ano01d] Anonymous. Back matter. Technometrics, 43(4):502–503, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270850.

Anonymous:2001:FMa

[Ano01e] Anonymous. Front matter. Technometrics, 43(1):96, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270851.

Anonymous:2001:FMb

[Ano01f] Anonymous. Front matter. Technometrics, 43(2):230, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271024.

Anonymous:2001:FMc

[Ano01g] Anonymous. Front matter. Technometrics, 43(3):368, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271210.

Anonymous:2001:FMd

[Ano01h] Anonymous. Front matter. Technometrics, 43(4):485, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270808.

Anonymous:2001:VI

[Ano01i] Anonymous. Volume information. Technometrics, 43(4):504–509, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270807.

Anonymous:2002:BMa

[Ano02a] Anonymous. Back matter. Technometrics, 44(1):97, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270719.

Anonymous:2002:BMb

[Ano02b] Anonymous. Back matter. Technometrics, 44(2):203, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271294.

Anonymous:2002:BMc

[Ano02c] Anonymous. Back matter. Technometrics, 44(3):??, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270524.

Anonymous:2002:BMd

[Ano02d] Anonymous. Back matter. Technometrics, 44(4):412, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271577.

Anonymous:2002:FMa

[Ano02e] Anonymous. Front matter. Technometrics, 44(1):80, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270678.

Anonymous:2002:FMb

[Ano02f] Anonymous. Front matter. Technometrics, 44(2):185, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271253.

Anonymous:2002:FMc

[Ano02g] Anonymous. Front matter. Technometrics, 44(3):286, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270484.

Anonymous:2002:FMd

[Ano02h] Anonymous. Front matter. Technometrics, 44(4):396, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271536.

Anonymous:2002:VI

[Ano02i] Anonymous. Volume information. Technometrics, 44(4):413–417, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271535.

Anonymous:2003:ARE

[Ano03a] Anonymous. Addendum: Recurrent events data analysis for product repairs, disease recurrences, and other applications. Technometrics, 45(4):376, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047120.

Anonymous:2003:BMa

[Ano03b] Anonymous. Back matter. Technometrics, 45(1):??, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047006.

Anonymous:2003:BMb

[Ano03c] Anonymous. Back matter. Technometrics, 45(2):??, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047043.

Anonymous:2003:BMc

[Ano03d] Anonymous. Back matter. Technometrics, 45(3):??, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047079.

Anonymous:2003:BMd

[Ano03e] Anonymous. Back matter. Technometrics, 45(4):??, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047121.

Anonymous:2003:BRBj

[Ano03f] Anonymous. Book review: A Manager’s Guide to the Design and Conduct of Clinical Trials by Phillip I. Good. Technometrics, 45(1):112, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047003.

Anonymous:2003:BRBaq

[Ano03g] Anonymous. Book review: Advanced Calculus with Applications in Statistics by André I. Khuri. Technometrics, 45(4):375, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047118.

Anonymous:2003:BRBn

[Ano03h] Anonymous. Book review: Analysis of Longitudinal Data by Peter J. Diggle; Patrick Heagerty; Kung-Yee Liang; Scott L. Zeger. Technometrics, 45(2):181, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047033.

Anonymous:2003:BRBa

[Ano03i] Anonymous. Book review: Categorical Data Analysis by Alan Agresti. Technometrics, 45(1):109, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25046994.

Anonymous:2003:BRBt

[Ano03j] Anonymous. Book review: Common Statistical Methods for Clinical Research with SAS(R) Examples by Glenn A. Walker. Technometrics, 45(2):182–183, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047038.

Anonymous:2003:BRBp

[Ano03k] Anonymous. Book review: Cross-Cultural Survey Methods by Janet A. Harkness; Fons J. R. van de Vijver; Peter Ph. Mohler. Technometrics, 45(2):182, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047037.

Anonymous:2003:BRBk

[Ano03l] Anonymous. Book review: Data Mining and Uncertain Reasoning by Zhengxin Chen. Technometrics, 45(1):112–113, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047005.

Anonymous:2003:BRBaa

[Ano03m] Anonymous. Book review: Data Mining: Concepts, Models, Methods, and Algorithms by Mehmed Kantardzic. Technometrics, 45(3): 277, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047072.

Anonymous:2003:BRBm

[Ano03n] Anonymous. Book review: Factor Analysis in Chemistry by Edmund R. Malinowski. Technometrics, 45(2):180–181, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047031.

Anonymous:2003:BRBi

[Ano03o] Anonymous. Book review: Handbook of Statistical Genetics by D. J. Balding; M. Bishop; C. Cannings. Technometrics, 45(1):

112, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047004.

Anonymous:2003:BRBai

[Ano03p] Anonymous. Book review: Handbook of Statistics 20: Advances in Reliability by N. Balakrishnan; C. R. Rao. Technometrics, 45 (4):372, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047111.

Anonymous:2003:BRBaj

[Ano03q] Anonymous. Book review: Implementing Six Sigma by Forrest W. Breyfogle. Technometrics, 45(4):372, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047110.

Anonymous:2003:BRBab

[Ano03r] Anonymous. Book review: Introduction to Survey Quality by Paul P. Biemer; Lars E. Lyberg. Technometrics, 45(3):277, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047073.

Anonymous:2003:BRBd

[Ano03s] Anonymous. Book review: Introduction to Time Series and Forecasting by Peter J. Brockwell; Richard A. Davis. Technometrics, 45(1):110, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046997.

Anonymous:2003:BRBao

[Ano03t] Anonymous. Book review: Introductory Biostatistics by Chap T. Le. Technometrics, 45(4):374, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047117.

Anonymous:2003:BRBap

[Ano03u] Anonymous. Book review: Introductory Biostatistics for the Health Sciences by Michael R. Chernick; Robert H. Friis. Technometrics, 45(4):374, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047116.

Anonymous:2003:BRBc

[Ano03v] Anonymous. Book review: Logistic Regression: A Self-Learning Text by David G. Kleinbaum; Mitchel Klein. Technometrics, 45 (1):109–110, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046996.

Anonymous:2003:BRBah

[Ano03w] Anonymous. Book review: Methods and Applications of Linear Models by Ronald R. Hocking. Technometrics, 45(4):371, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047109.

Anonymous:2003:BRBb

[Ano03x] Anonymous. Book review: Methods of Multivariate Analysis by Alvin C. Rencher. Technometrics, 45(1):109, February 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25046995.

Anonymous:2003:BRBo

[Ano03y] Anonymous. Book review: Model Selection and Multimodel Inference by Kenneth P. Burnham; David R. Anderson. Technometrics, 45(2):181, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047032.

Anonymous:2003:BRBam

[Ano03z] Anonymous. Book review: Modeling Financial Time Series with S-PLUS(R) by Eric Zivot; Jiahui Wang. Technometrics, 45(4): 373–374, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047114.

Anonymous:2003:BRBac

[Ano03-27] Anonymous. Book review: Models for Investors in Real World Markets by James R. Thompson; Edward E. Williams; M. Chapman Findlay. Technometrics, 45(3):277–278, August 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047074.

Anonymous:2003:BRBe

[Ano03-28] Anonymous. Book review: Modern Applied Statistics with S by W. N. Venables; B. D. Ripley. Technometrics, 45(1):111, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25046999.

Anonymous:2003:BRBq

[Ano03-29] Anonymous. Book review: Multidimensional Scaling by Trevor F. Cox; Michael A. A. Cox. Technometrics, 45(2):182, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047034.

Anonymous:2003:BRBr

[Ano03-30] Anonymous. Book review: Observational Studies by Paul R. Rosenbaum. Technometrics, 45(2):182, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047035.

Anonymous:2003:BRBy

[Ano03-31] Anonymous. Book review: Principal Component Analysis by I. T. Jolliffe. Technometrics, 45(3):276, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047070.

Anonymous:2003:BRBz

[Ano03-32] Anonymous. Book review: Probability, Statistics, and Reliability for Engineers and Scientists by Bilal M. Ayyub; Richard H. Mc-Cuen. Technometrics, 45(3):276, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047069.

Anonymous:2003:BRBaf

[Ano03-33] Anonymous. Book review: Randomization in Clinical Trials by William F. Rosenberger; John M. Lachin. Technometrics, 45(3): 279, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047078.

Anonymous:2003:BRBal

[Ano03-34] Anonymous. Book review: SAS(R) for Monte Carlo Studies by Xitao Fan; Ákos Felsővályi; Stephen A. Sivo; Sean C. Keenan. Technometrics, 45(4):373, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047113.

Anonymous:2003:BRBag

[Ano03-35] Anonymous. Book review: Six Sigma and Beyond: Statistics and Probability by D. H. Stamatis. Technometrics, 45(3):279, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047077.

Anonymous:2003:BRBae

[Ano03-36] Anonymous. Book review: Six Sigma Quality for Business & Manufacture by M. Joseph Gordon. Technometrics, 45(3):278, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047075.

Anonymous:2003:BRBs

[Ano03-37] Anonymous. Book review: Standardization and Tacit Knowledge by Douglas W. Maynard; Hanneke Houtkoop-Steenstra; Nora Cate Schaeffer; Johannes van de Zouwen. Technometrics, 45(2):182, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047036.

Anonymous:2003:BRBx

[Ano03-38] Anonymous. Book review: Statistical Design and Analysis of Experiments by Robert L. Mason; Richard F. Gunst; James L. Hess. Technometrics, 45(3):275–276, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047068.

Anonymous:2003:BRBf

[Ano03-39] Anonymous. Book review: Statistical Inference by Paul H. Garthwaite; Ian T. Jolliffe; Byron Jones. Technometrics, 45(1):111, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047000.

Anonymous:2003:BRBak

[Ano03-40] Anonymous. Book review: Statistical Methods for Survival Data Analysis by Elisa T. Lee; John Wenyu Wang. Technometrics, 45 (4):372–373, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047112.

Anonymous:2003:BRBg

[Ano03-41] Anonymous. Book review: Statistical Methods for the Analysis of Biomedical Data by Robert F. Woolson; William R. Clarke.

Technometrics, 45(1):111, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047001.

Anonymous:2003:BRBh

[Ano03-42] Anonymous. Book review: Statistical Methods in Diagnostic Medicine by Xiao-Hua Zhou; Nancy A. Obuchowski; Donna K. McClish. Technometrics, 45(1):111–112, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047002.

Anonymous:2003:BRBv

[Ano03-43] Anonymous. Book review: Statistical Quality Control Using Excel by Steven M. Zimmerman; Marjorie L. Icenogle. Technometrics, 45(2):183, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047040.

Anonymous:2003:BRBl

[Ano03-44] Anonymous. Book review: Subset Selection in Regression by Alan Miller. Technometrics, 45(2):180, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047030.

Anonymous:2003:BRBan

[Ano03-45] Anonymous. Book review: The Analysis of Interpretation of Multivariate Data for Social Scientists by David J. Bartholomew; Fiona Steele; Irini Moustaki; Jane I. Galbraith. Technometrics, 45(4):374, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047115.

Anonymous:2003:BRBu

[Ano03-46] Anonymous. Book review: The Basics of S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 45(2):183, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047039.

Anonymous:2003:BRBw

[Ano03-47] Anonymous. Book review: The Cambridge Dictionary of Statistics by B. S. Everitt. Technometrics, 45(2):183–184, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047041.

Anonymous:2003:BRBad

[Ano03-48] Anonymous. Book review: The Power of Ultimate Six Sigma^{T M} by Keki R. Bhote. Technometrics, 45(3):278, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047076.

Anonymous:2003:BRBar

[Ano03-49] Anonymous. Book reviews: The Analysis of Gene Expression Data by Giovanni Parmigiani; Elizabeth S. Garrett; Rafael A. Irizzary; Scott L. Zeger; Statistical Analysis of Gene Expression Microarray Data by Terry Speed. Technometrics, 45(4):375, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047119.

Anonymous:2003:BSR

[Ano03-50] Anonymous. Case Studies in Reliability and Maintenance by Wallace R. Blischke; D. N. Prabhakar Murthy. Technometrics, 45 (3):276–277, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047071.

Anonymous:2003:BEE

[Ano03-51] Anonymous. Statistics for Environmental Engineers by Paul Mac Berthouex; Linfield C. Brown. Technometrics, 45(1):110–111, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25046998.

Anonymous:2003:FMa

[Ano03-52] Anonymous. Front matter. Technometrics, 45(1):??, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25046970.

Anonymous:2003:FMb

[Ano03-53] Anonymous. Front matter. Technometrics, 45(2):??, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047008.

Anonymous:2003:FMc

[Ano03-54] Anonymous. Front matter. Technometrics, 45(3):??, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047045.

Anonymous:2003:FMd

[Ano03-55] Anonymous. Front matter. Technometrics, 45(4):??, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047082.

Anonymous:2003:VI

[Ano03-56] Anonymous. Volume information. Technometrics, 45(4):??, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047081.

Anonymous:2004:BMa

[Ano04a] Anonymous. Back matter. Technometrics, 46(1):??, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470795.

Anonymous:2004:BMb

[Ano04b] Anonymous. Back matter. Technometrics, 46(2):??, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470839.

Anonymous:2004:BMc

[Ano04c] Anonymous. Back matter. Technometrics, 46(3):??, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470881.

Anonymous:2004:BMd

[Ano04d] Anonymous. Back matter. Technometrics, 46(4):??, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470926.

Anonymous:2004:BRBc

[Ano04e] Anonymous. Book review: An Introduction to Multivariate Statistical Analysis by T. W. Anderson. Technometrics, 46(1):119, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470783.

Anonymous:2004:BRBt

[Ano04f] Anonymous. Book review: Applied Econometric Time Series by Walter Enders. Technometrics, 46(2):264, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470830.

Anonymous:2004:BRBq

[Ano04g] Anonymous. Book review: Designing Experiments and Analyzing Data by Scott E. Maxwell; Harold D. Delaney. Technometrics, 46(2):263, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470827.

Anonymous:2004:BRBn

[Ano04h] Anonymous. Book review: Engineering Statistics by Douglas C. Montgomery. Technometrics, 46(2):261–262, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470824.

Anonymous:2004:BRBh

[Ano04i] Anonymous. Book review: Extreme Values in Finance, Telecommunications, and the Environment by Bärbel Finkenstädt; Holger Rootzén. Technometrics, 46(1):121, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470789.

Anonymous:2004:BRBi

[Ano04j] Anonymous. Book review: Handbook of Statistics 21: Stochastic Processes: Modelling and Simulation by D. N. Shanbhag; C. R. Rao. Technometrics, 46(1):121, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470790.

Anonymous:2004:BRBr

[Ano04k] Anonymous. Book review: Handbook of Statistics 22: Statistics in Industry by R. Khattree; C. R. Rao. Technometrics, 46(2):263, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470828.

Anonymous:2004:BRBa

[Ano04l] Anonymous. Book review: Introduction to the Practice of Statistics by David S. Moore; George P. McCabe. Technometrics, 46 (1):118–119, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470782.

Anonymous:2004:BRBl

[Ano04m] Anonymous. Book review: Leading Six Sigma by Ronald D. Snee; Roger W. Hoerl. Technometrics, 46(1):122, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470794.

Anonymous:2004:BRBd

[Ano04n] Anonymous. Book review: Modelling Binary Data by David Collett. Technometrics, 46(1):119–120, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470785.

Anonymous:2004:BRBw

[Ano04o] Anonymous. Book review: Modelling Survival Data in Medical Research by David Collett. Technometrics, 46(2):265, May 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470834.

Anonymous:2004:BRBv

[Ano04p] Anonymous. Book review: Multivariate Geostatistics by Hans Wackernagel. Technometrics, 46(2):264–265, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470832.

Anonymous:2004:BRBo

[Ano04q] Anonymous. Book review: Nonparametric Statistical Inference by Jean Dickinson Gibbons; Subhabrata Chakraborti. Technometrics, 46(2):262, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470826.

Anonymous:2004:BRBm

[Ano04r] Anonymous. Book review: Quantitative Methods in Population Health by Mari Palta. Technometrics, 46(1):122, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470793.

Anonymous:2004:BRBf

[Ano04s] Anonymous. Book review: Sample Survey: Principles & Methods by Vic Barnett. Technometrics, 46(1):120, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470786.

Anonymous:2004:BRBg

[Ano04t] Anonymous. Book review: SAS(R) for Forecasting Time Series by John C. Brocklebank; David A. Dickey. Technometrics, 46 (1):120–121, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470788.

Anonymous:2004:BRBu

[Ano04u] Anonymous. Book review: Six Sigma for Financial Professionals by D. H. Stamatis. Technometrics, 46(2):264, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470831.

Anonymous:2004:BRBe

[Ano04v] Anonymous. Book review: Small Area Estimation by J. N. K. Rao. Technometrics, 46(1):120, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470787.

Anonymous:2004:BRBj

[Ano04w] Anonymous. Book review: Statistical Data Mining and Knowledge Discovery by Hamparsum Bozdogan. Technometrics, 46(1): 121, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470791.

Anonymous:2004:BRBy

[Ano04x] Anonymous. Book review: Statistical Matching by Susanne Rässler. Technometrics, 46(2):265, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470833.

Anonymous:2004:BRBs

[Ano04y] Anonymous. Book review: Statistical Methods for Rates and Proportions by Joseph L. Fleiss; Bruce Levin; Myunghee Cho Paik. Technometrics, 46(2):263–264, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470829.

Anonymous:2004:BRBk

[Ano04z] Anonymous. Book review: Statistical Modeling and Analysis for Database Marketing by Bruce Ratner. Technometrics, 46(1):121– 122, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470792.

Anonymous:2004:BRBp

[Ano04-27] Anonymous. Book review: Statistical Tools for Nonlinear Regression by S. Huet; A. Bouvier; M.-A. Poursat; E. Jolivet. Technometrics, 46(2):262, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470825.

Anonymous:2004:BRBx

[Ano04-28] Anonymous. Book review: Statistics in Drug Research by Shein-Chung Chow; Jun Shao. Technometrics, 46(2):265, May 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470835.

Anonymous:2004:BRBb

[Ano04-29] Anonymous. Book review: The Analysis of Time Series by Chris Chatfield. Technometrics, 46(1):119, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470784.

Anonymous:2004:BRBaa

[Ano04-30] Anonymous. Book review: The Oxford Dictionary of Statistical Terms by Yadolah Dodge. Technometrics, 46(2):266–267, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470837.

Anonymous:2004:BRBz

[Ano04-31] Anonymous. Book reviews: Cross-Over Trials in Clinical Research by Stephen Senn; Design and Analysis of Cross-Over Trials by Byron Jones; Michael G. Kenward. Technometrics, 46(2):265–266, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470836.

Anonymous:2004:EC

[Ano04-32] Anonymous. Editorial correction. Technometrics, 46(2):268, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470838.

Anonymous:2004:FMa

[Ano04-33] Anonymous. Front matter. Technometrics, 46(1):??, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470750.

Anonymous:2004:FMb

[Ano04-34] Anonymous. Front matter. Technometrics, 46(2):??, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470797.

Anonymous:2004:FMc

[Ano04-35] Anonymous. Front matter. Technometrics, 46(3):??, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470841.

Anonymous:2004:FMd

[Ano04-36] Anonymous. Front matter. Technometrics, 46(4):??, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470884.

Anonymous:2004:VI

[Ano04-37] Anonymous. Volume information. Technometrics, 46(4):??, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470883.

Anonymous:2005:BMa

[Ano05a] Anonymous. Back matter. Technometrics, 47(1):??, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470970.

Anonymous:2005:BMb

[Ano05b] Anonymous. Back matter. Technometrics, 47(2):??, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471015.

Anonymous:2005:BMc

[Ano05c] Anonymous. Back matter. Technometrics, 47(3):??, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471059.

Anonymous:2005:BMd

[Ano05d] Anonymous. Back matter. Technometrics, 47(4):??, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471107.

Anonymous:2005:BRBu

[Ano05e] Anonymous. Book review: Applied Longitudinal Analysis by Garrett M. Fitzmaurice; Nan M. Laird; James H. Ware. Technometrics, 47(2):245, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471008.

Anonymous:2005:BRBab

[Ano05f] Anonymous. Book review: Applied Regression Analysis by Terry E. Dielman. Technometrics, 47(3):380, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471047.

Anonymous:2005:BRBap

[Ano05g] Anonymous. Book review: Applied Statistics for Engineers and Scientists by Jay Devore; Nicholas Farnum. Technometrics, 47 (4):526, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471094.

Anonymous:2005:BRBz

[Ano05h] Anonymous. Book review: Aspects of Multivariate Statistical Analysis in Geology by Richard A. Reyment; Enrico Savazzi. Technometrics, 47(2):247, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471013.

Anonymous:2005:BRBaaa

[Ano05i] Anonymous. Book review: Audit Sampling: An Introduction by Don M. Guy; Douglas R. Carmichael; Ray Whittington. Technometrics, 47(4):530, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471105.

Anonymous:2005:BRBaq

[Ano05j] Anonymous. Book review: Basic Statistical Ideas for Managers by David K. Hildebrand; R. Lyman Ott; J. Brian Gray. Technometrics, 47(4):526, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471096.

Anonymous:2005:BRBak

[Ano05k] Anonymous. Book review: Biostatistics: A Bayesian Introduction by George G. Woodworth. Technometrics, 47(3):382–383, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471056.

Anonymous:2005:BRBt

[Ano05l] Anonymous. Book review: Biostatistics by Gerald van Belle; Lloyd D. Fisher; Patrick J. Heagerty; Thomas Lumley. Technometrics, 47(2):244–245, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471007.

Anonymous:2005:BRBav

[Ano05m] Anonymous. Book review: Breakthrough Business Results with MVT(R) by Charles W. Holland; David Cochran. Technometrics, 47(4):528, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471100.

Anonymous:2005:BRBaz

[Ano05n] Anonymous. Book review: Chances Are: A Primer on Probability by Susan F. Wagner. Technometrics, 47(4):529–530, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471104.

Anonymous:2005:BRBai

[Ano05o] Anonymous. Book review: Data Visualization in the Geosciences by James R. Carr. Technometrics, 47(3):382, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471055.

Anonymous:2005:BRBd

[Ano05p] Anonymous. Book review: Design for Six Sigma by C. M. Creveling; J. L. Slutsky; D. Antis. Technometrics, 47(1):105, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470958.

Anonymous:2005:BRBah

[Ano05q] Anonymous. Book review: Discrete Distributions by Daniel Zelterman. Technometrics, 47(3):381–382, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471053.

Anonymous:2005:BRBat

[Ano05r] Anonymous. Book review: Geostatistical Reservoir Modeling by Clayton V. Deutsch. Technometrics, 47(4):527, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471098.

Anonymous:2005:BRBau

[Ano05s] Anonymous. Book review: Geostatistics Rio 2000 by M. Armstrong; C. Bettini; N. Champigny; A. Galli; A. Remacre. Technometrics, 47(4):527–528, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471099.

Anonymous:2005:BRBam

[Ano05t] Anonymous. Book review: Handbook of Computational Statistics by James E. Gentle. Technometrics, 47(3):383–384, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471058.

Anonymous:2005:BRBj

[Ano05u] Anonymous. Book review: Handbook of Statistics 23: Advances in Survival Analysis by N. Balakrishnan; C. R. Rao. Technometrics, 47(1):107, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470964.

Anonymous:2005:BRBa

[Ano05v] Anonymous. Book review: Intelligent Data Analysis: An Introduction by Michael Berthold; David J. Hand. Technometrics, 47(1):104, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470955.

Anonymous:2005:BRBf

[Ano05w] Anonymous. Book review: Internet Companion for Statistics by Michael D. Larsen. Technometrics, 47(1):106, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470963.

Anonymous:2005:BRBaw

[Ano05x] Anonymous. Book review: Introduction to Applied Econometrics by Kenneth G. Stewart. Technometrics, 47(4):528, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471101.

Anonymous:2005:BRBao

[Ano05y] Anonymous. Book review: Introduction to Robust Estimation and Hypothesis Testing by Rand R. Wilcox. Technometrics, 47 (4):525–526, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471093.

Anonymous:2005:BRBc

[Ano05z] Anonymous. Book review: Introduction to Spectral Analysis by Petre Stoica; Randolph Moses. Technometrics, 47(1):104–105, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470957.

Anonymous:2005:BRBaf

[Ano05-27] Anonymous. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 47(3):381, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471051.

Anonymous:2005:BRBq

[Ano05-28] Anonymous. Book review: JMP Start Statistics by John Sall; Lee Creighton; Ann Lehman. Technometrics, 47(2):244, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471006.

Anonymous:2005:BRBx

[Ano05-29] Anonymous. Book review: Methods for Testing and Evaluating Survey Questionnaires by Stanley Presser; Jennifer M. Rothgeb; Mick P. Couper; Judith T. Lessler; Elizabeth Martin; Jean Martin; Eleanor Singer. Technometrics, 47(2):246, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471012.

Anonymous:2005:BRBb

[Ano05-30] Anonymous. Book review: Mind on Statistics by Jessica M. Utts; Robert F. Heckard. Technometrics, 47(1):104, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470956.

Anonymous:2005:BRBp

[Ano05-31] Anonymous. Book review: Monte Carlo Statistical Methods by Christian P. Robert; George Casella. Technometrics, 47(2):243, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471003.

Anonymous:2005:BRBac

[Ano05-32] Anonymous. Book review: Multivariate Statistical Methods by Bryan F. J. Manly. Technometrics, 47(3):380, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471049.

Anonymous:2005:BRBg

[Ano05-33] Anonymous. Book review: Nature-Inspired Methods in Chemometrics: Genetic Algorithms and Artificial Neural Networks by Riccardo Leardi. Technometrics, 47(1):106, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470960.

Anonymous:2005:BRBag

[Ano05-34] Anonymous. Book review: Nonlinear Time Series Analysis by Holger Kantz; Thomas Schreiber. Technometrics, 47(3):381, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471050.

Anonymous:2005:BRBan

[Ano05-35] Anonymous. Book review: Permutation, Parametric, and Bootstrap Tests of Hypotheses by Phillip Good. Technometrics, 47 (4):525, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471092.

Anonymous:2005:BRBaj

[Ano05-36] Anonymous. Book review: Practical Business Forecasting by Michael K. Evans. Technometrics, 47(3):382, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471054.

Anonymous:2005:BRBr

[Ano05-37] Anonymous. Book review: Practical Genetic Algorithms by Randy L. Haupt; Sue Ellen Haupt. Technometrics, 47(2):244, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471005.

Anonymous:2005:BRBw

[Ano05-38] Anonymous. Book review: Practical Methods for Design and Analysis of Complex Surveys by Risto Lehtonen; Erkki Pahkinen. Technometrics, 47(2):245–246, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471010.

Anonymous:2005:BRBh

[Ano05-39] Anonymous. Book review: Practical Statistics for Environmental and Biological Scientists by John Townend. Technometrics, 47(1):106, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470962.

Anonymous:2005:BRBe

[Ano05-40] Anonymous. Book review: Quality Assurance in Analytical Chemistry by B. W. Wenclawiak; M. Koch; E. Hadjicostas. Technometrics, 47(1):105, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470959.

Anonymous:2005:BRBi

[Ano05-41] Anonymous. Book review: Quantitative Methods for Current Environmental Issues by Clive W. Anderson; Vic Barnett; Philip C. Chatwin; Abdel H. El-Shaarawi. Technometrics, 47(1):106, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470961.

Anonymous:2005:BRBad

[Ano05-42] Anonymous. Book review: Regression with Social Data: Modeling Continuous and Limited Response Variables by Alfred De-Maris. Technometrics, 47(3):380, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471048.

Anonymous:2005:BRBal

[Ano05-43] Anonymous. Book review: Risk Assessment and Decision Making in Business and Industry: A Practical Guide by Glenn Koller. Technometrics, 47(3):383, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471057.

Anonymous:2005:BRBar

[Ano05-44] Anonymous. Book review: Seeing through Statistics by Jessica M. Utts. Technometrics, 47(4):526, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471095.

Anonymous:2005:BRBl

[Ano05-45] Anonymous. Book review: Statistical Estimation of Epidemiological Risk by Kung-Jong Lui. Technometrics, 47(1):107–108, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470966.

Anonymous:2005:BRBaa

[Ano05-46] Anonymous. Book review: Statistical Methods in Analytical Chemistry by Peter C. Meier; Richard E. Zünd. Technometrics, 47 (2):247, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471014.

Anonymous:2005:BRBas

[Ano05-47] Anonymous. Book review: Statistics and Data Analysis in Geology by John C. Davis. Technometrics, 47(4):526–527, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471097.

Anonymous:2005:BRBs

[Ano05-48] Anonymous. Book review: Statistics and Finance: An Introduction by David Ruppert. Technometrics, 47(2):244, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471004.

Anonymous:2005:BRBay

[Ano05-49] Anonymous. Book review: Stereology for Statisticians by Adrian Baddeley; Eva B. Vedel Jensen. Technometrics, 47(4):529, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471103.

Anonymous:2005:BRBy

[Ano05-50] Anonymous. Book review: Survey Methodology by Robert M. Groves; Floyd J. Fowler,; Mick P. Couper; James M. Lepkowski. Technometrics, 47(2):246, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471011.

Anonymous:2005:BRBv

[Ano05-51] Anonymous. Book review: The Statistical Evaluation of Medical Tests for Classification and Prediction by Margaret Sullivan Pepe. Technometrics, 47(2):245, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471009.

Anonymous:2005:BRBae

[Ano05-52] Anonymous. Book reviews: Applied Life Data Analysis by Wayne B. Nelson; Accelerated Testing: Statistical Models, Test Plans, and Data Analysis by Wayne B. Nelson. Technometrics, 47(3):381, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471052.

Anonymous:2005:BRBo

[Ano05-53] Anonymous. Book reviews: Data Analysis Tools for DNA Microarrays by Sorin Drähici; Analyzing Microarray Gene Expression Data by Geoffrey J. McLachlan; Kim-Anh Do; Christophe Ambroise. Technometrics, 47(1):109, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470969.

Anonymous:2005:BRBn

[Ano05-54] Anonymous. Book reviews: Design and Analysis of Clinical Trials by Shein-Chung Chow; Jen-Pei Liu; Design and Analysis of DNA Microarray Investigations by Richard M. Simon; Edward L. Korn; Lisa M. McShane; Michael D. Radmacher; George W. Wright; Yingdong Zhao. Technometrics, 47(1):108–109, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470968.

Anonymous:2005:BRBax

[Ano05-55] Anonymous. Book reviews: Discovering Knowledge in Data by Daniel Larose; Next Generation of Data-Mining Applications by Mehmed M. Kantardzic; Jozef Zurada. Technometrics, 47(4):

528–529, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471102.

Anonymous:2005:BRBm

[Ano05-56] Anonymous. Book reviews: Disease Mapping with WinBUGS and MLwiN by Andrew B. Lawson; William J. Browne; Carmen L. Vidal Rodeiro; Bayesian Approaches to Clinical Trials and Health Care Evaluation by David J. Spiegelhalter; Keith R. Abrams; Jonathan P. Myles. Technometrics, 47(1):108, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470967.

Anonymous:2005:BRBk

[Ano05-57] Anonymous. Book reviews: Statistics for Epidemiology by Nicholas P. Jewell; Multivariate Methods in Epidemiology by Theodore R. Holford. Technometrics, 47(1):107, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470965.

Anonymous:2005:CLF

[Ano05-58] Anonymous. Correction: “Large Factorial Designs for Product Engineering and Marketing Research Applications”. Technometrics, 47(4):532, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471106. See [KT05].

Anonymous:2005:FMa

[Ano05-59] Anonymous. Front matter. Technometrics, 47(1):??, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470928.

Anonymous:2005:FMb

[Ano05-60] Anonymous. Front matter. Technometrics, 47(2):??, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470972.

Anonymous:2005:FMc

[Ano05-61] Anonymous. Front matter. Technometrics, 47(3):??, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471017.

Anonymous:2005:FMd

[Ano05-62] Anonymous. Front matter. Technometrics, 47(4):??, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471062.

Anonymous:2005:VI

[Ano05-63] Anonymous. Volume information. Technometrics, 47(4):??, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471061.

Anonymous:2006:BMa

[Ano06a] Anonymous. Back matter. Technometrics, 48(1):??, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471151.

Anonymous:2006:BMb

[Ano06b] Anonymous. Back matter. Technometrics, 48(2):??, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471197.

Anonymous:2006:BMc

[Ano06c] Anonymous. Back matter. Technometrics, 48(3):??, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471233.

Anonymous:2006:BMd

[Ano06d] Anonymous. Back matter. Technometrics, 48(4):??, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471267.

Anonymous:2006:BRBam

[Ano06e] Anonymous. Book review: A Handbook of Statistical Analysis Using R by Brian Everitt; Torster Northorn. Technometrics, 48 (4):577, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471266.

Anonymous:2006:BRBu

[Ano06f] Anonymous. Book review: Analysis of Financial Time Series by Ruey S. Tsay. Technometrics, 48(2):316, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471192.

Anonymous:2006:BRBae

[Ano06g] Anonymous. Book review: Analyzing Environmental Data by Walter W. Piegorsch; A. John Bailer. Technometrics, 48(3):450– 451, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471228.

Anonymous:2006:BRBq

[Ano06h] Anonymous. Book review: Applied Optimal Designs by Martijn P. F. Berger; Weng Kee Wong. Technometrics, 48(2):315, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471189.

Anonymous:2006:BRBal

[Ano06i] Anonymous. Book review: Bayesian Thinking: Modeling and Computation by D. K. Dey; C. R. Rao. Technometrics, 48(4): 576–577, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471265.

Anonymous:2006:BRBaf

[Ano06j] Anonymous. Book review: Computational Intelligence in Time Series Forecasting by Ajoy K. Palit; Dobrivoje Popovic. Technometrics, 48(3):451, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471229.

Anonymous:2006:BRBi

[Ano06k] Anonymous. Book review: Data Mining Techniques by Michael J. A. Berry; Gordon S. Linoff. Technometrics, 48(1):159, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471147.

Anonymous:2006:BRBh

[Ano06l] Anonymous. Book review: Design and Analysis of Experiments by Douglas C. Montgomery. Technometrics, 48(1):158–159, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471146.

Anonymous:2006:BRBk

[Ano06m] Anonymous. Book review: Encyclopedia of Biostatistics: Eight Volumes by Peter Armitage; Theodore Colton. Technometrics, 48(1):159–160, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471149.

Anonymous:2006:BRBd

[Ano06n] Anonymous. Book review: Engineering Design Reliability Handbook by Efstratios Nikolaidis; Dan M. Ghiocel; Suren Singhal. Technometrics, 48(1):156–157, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471142.

Anonymous:2006:BRBw

[Ano06o] Anonymous. Book review: Epidemiology: Study Design and Data Analysis by Mark Woodward. Technometrics, 48(2):316–317, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471194.

Anonymous:2006:BRBaj

[Ano06p] Anonymous. Book review: Evaluation of Mineral Reserves by Andre Journel; Phaedon Kyriakdis. Technometrics, 48(4):576, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471264.

Anonymous:2006:BRBf

[Ano06q] Anonymous. Book review: Fixed Effects Regression Methods for Longitudinal Data Using SAS(R) by Paul D. Allison. Technometrics, 48(1):157–158, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471144.

Anonymous:2006:BRBai

[Ano06r] Anonymous. Book review: geoENV IV: Geostatistics for Environmental Applications by Xavier Sanchez-Vila; Jesus Carrera; José Jaime Gómez-Hernández. Technometrics, 48(3):452, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471232.

Anonymous:2006:BRBp

[Ano06s] Anonymous. Book review: Handbook of Statistics 24: Data Mining and Data Visualization by C. R. Rao; Edward Wegman; Jeffrey Solka. Technometrics, 48(2):314–315, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471187.

Anonymous:2006:BRBab

[Ano06t] Anonymous. Book review: Introduction to Modern Portfolio Optimization by Bernd Scherer; Douglas Martin. Technometrics, 48(3):450, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471226.

Anonymous:2006:BRBn

[Ano06u] Anonymous. Book review: Introduction to Statistics through Resampling Methods and R/S-PLUS(R) by Phillip I. Good. Technometrics, 48(2):314, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471185.

Anonymous:2006:BRBx

[Ano06v] Anonymous. Book review: Loss Models: From Data to Decisions by Stuart A. Klugman; Harry H. Panjer; Gordon E. Willmot. Technometrics, 48(2):317, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471196.

Anonymous:2006:BRBaa

[Ano06w] Anonymous. Book review: Multilevel Statistical Models by Harvey Goldstein. Technometrics, 48(3):449, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471223.

Anonymous:2006:BRBag

[Ano06x] Anonymous. Book review: Pattern Recognition in Industry by Phiroz Bhagat. Technometrics, 48(3):451, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471230.

Anonymous:2006:BRBj

[Ano06y] Anonymous. Book review: Probability Models for Economic Decisions by Roger B. Myerson. Technometrics, 48(1):159, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471148.

Anonymous:2006:BRBak

[Ano06z] Anonymous. Book review: Resampling Methods by Phillip Good. Technometrics, 48(4):576, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471263.

Anonymous:2006:BRBb

[Ano06-27] Anonymous. Book review: State Space and Unobserved Component Models by Andrew Harvey; Siem Jan Koopman; Neil Shephard. Technometrics, 48(1):156, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471141.

Anonymous:2006:BRBs

[Ano06-28] Anonymous. Book review: Statistical Models: Theory and Practice by David A. Freedman. Technometrics, 48(2):315, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471190.

Anonymous:2006:BRBv

[Ano06-29] Anonymous. Book review: Statistical Thinking in Business by J. A. John; D. Whitaker; D. G. Johnson. Technometrics, 48(2):316, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471193.

Anonymous:2006:BRBr

[Ano06-30] Anonymous. Book review: Statistics for Fission Track Analysis by Rex F. Galbraith. Technometrics, 48(2):315, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471188.

Anonymous:2006:BRBe

[Ano06-31] Anonymous. Book review: Structural Equation Modeling by Bruce H. Pugesek; Adrian Tomer; Alexander von Eye. Technometrics, 48(1):157, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471143.

Anonymous:2006:BRBy

[Ano06-32] Anonymous. Book review: Survival Analysis: A Self-Learning Text by David G. Kleinbaum; Mitchel Klein. Technometrics, 48(2):317, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471195.

Anonymous:2006:BRBa

[Ano06-33] Anonymous. Book review: Time Series Analysis and Applications to Geophysical Systems by David R. Brillinger; Enders Anthony Robinson; Frederic Paik Schoenberg. Technometrics, 48 (1):155–156, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471139.

Anonymous:2006:BRBad

[Ano06-34] Anonymous. Book review: Univariate Discrete Distributions by Norman L. Johnson; Adrienne W. Kemp; Samuel Kotz. Technometrics, 48(3):450, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471227.

Anonymous:2006:BRBc

[Ano06-35] Anonymous. Book review: Using Statistical Methods for Water Quality Management by G. B. McBride. Technometrics, 48 (1):156, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471140.

Anonymous:2006:BRBt

[Ano06-36] Anonymous. Book review: What Is Six Sigma Process Management? by Rowland Hayler; Michael Nichols. Technometrics, 48 (2):315–316, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471191.

Anonymous:2006:BRBo

[Ano06-37] Anonymous. Book reviews: A Step-by-Step Approach to Using SAS(R) for Univariate & Multivariate Statistics by Norm O’Rourke; Larry Hatcher; Edward J. Stepanski; JMP(R) for Basic Univariate and Multivariate Statistics by Ann Lehman; Norm O’Rourke; Larry Hatcher; Edward J. Stepanski. Technometrics, 48(2):314, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471186.

Anonymous:2006:BRBl

[Ano06-38] Anonymous. Book reviews: Anthology of Statistics in Sports by Jim Albert; Jay Bennett; James J. Cochran; Working with Numbers and Statistics: A Handbook for Journalists by Charles Livingston; Paul Voakes. Technometrics, 48(1):160, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471150.

Anonymous:2006:BRBg

[Ano06-39] Anonymous. Book reviews: Applied Longitudinal Data Analysis for Epidemiology by Jos W. R. Twisk; Longitudinal and Panel Data by Edward W. Frees; Applied Longitudinal Data Analysis by Judith D. Singer; John B. Willett. Technometrics, 48(1):158, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471145.

Anonymous:2006:BRBz

[Ano06-40] Anonymous. Book reviews: Applied Statistics and the SAS(R) Programming Language by Ronald P. Cody; Jeffrey K. Smith; Statistical Analysis of Medical Data Using SAS by Geoff Der; Brian S. Everitt. Technometrics, 48(3):449, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471224.

Anonymous:2006:BRBm

[Ano06-41] Anonymous. Book reviews: Functional Data Analysis by J. O. Ramsay; B. W. Silverman; S+Functional Data Analysis User’s Guide by Douglas B. Clarkson; Chris Fraley; Charles C. Gu. Technometrics, 48(2):313, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471183.

Anonymous:2006:BRBac

[Ano06-42] Anonymous. Book reviews: Statistical Demography and Forecasting by Juha M. Alho; Bruce D. Spencer; Applied Mathematical Demography by Nathan Keyfitz; Hal Caswell. Technometrics, 48(3):450, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471225.

Anonymous:2006:BRBah

[Ano06-43] Anonymous. Book reviews: Statistical Methods in Molecular Evolution by Rasmus Nielsen; The Evaluation of Surrogate Endpoints by Tomasz Burzykowski; Geert Molenberghs; Marc Buyse. Technometrics, 48(3):451–452, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471231.

Anonymous:2006:BBP

[Ano06-44] Anonymous. The Basics of S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 48(2):314, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471184.

Anonymous:2006:FMa

[Ano06-45] Anonymous. Front matter. Technometrics, 48(1):??, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471109.

Anonymous:2006:FMb

[Ano06-46] Anonymous. Front matter. Technometrics, 48(2):??, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471153.

Anonymous:2006:FMc

[Ano06-47] Anonymous. Front matter. Technometrics, 48(3):??, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471199.

Anonymous:2006:FMd

[Ano06-48] Anonymous. Front matter. Technometrics, 48(4):??, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471236.

Anonymous:2006:VI

[Ano06-49] Anonymous. Volume information. Technometrics, 48(4):??, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471235.

Anonymous:2007:BMa

[Ano07a] Anonymous. Back matter. Technometrics, 49(1):??, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471302.

Anonymous:2007:BMb

[Ano07b] Anonymous. Back matter. Technometrics, 49(2):??, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471340.

Anonymous:2007:BMc

[Ano07c] Anonymous. Back matter. Technometrics, 49(3):??, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471378.

Anonymous:2007:BMd

[Ano07d] Anonymous. Back matter. Technometrics, 49(4):??, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471415.

Anonymous:2007:BRBm

[Ano07e] Anonymous. Book review: Advances in Statistical Methods for the Health Sciences: Applications to Cancer and AIDS Studies, Genome Sequence Analysis, and Survival Analysis by Jean-Louis Auget; N. Balakrishnan; Mounir Mesbah; Geert Molenberghs. Technometrics, 49(3):368, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471375.

Anonymous:2007:BRBr

[Ano07f] Anonymous. Book review: Chance and Choice: Memorabilia by Kai Lai Chung. Technometrics, 49(4):501, November 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471413.

Anonymous:2007:BRBl

[Ano07g] Anonymous. Book review: Direct and Large-Eddy Simulation VI by Eric Lamballais; Rainer Friedrich; Bernard J. Geurts; Olivier Métais. Technometrics, 49(3):367–368, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471374.

Anonymous:2007:BRBn

[Ano07h] Anonymous. Book review: Exploring Data Tables, Trends, and Shapes by David C. Hoaglin; Frederick Mosteller; John W. Tukey. Technometrics, 49(3):368, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471376.

Anonymous:2007:BRBq

[Ano07i] Anonymous. Book review: Handbook of Parametric and Nonparametric Statistical Procedures: Introduction to Linear Regression by David J. Sheskin. Technometrics, 49(4):500–501, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471411.

Anonymous:2007:BRBk

[Ano07j] Anonymous. Book review: Introduction to Linear Regression Analysis by Douglas C. Montgomery; Elizabeth A. Peck; G. Geoffrey Vining. Technometrics, 49(2):232–233, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471339.

Anonymous:2007:BRBa

[Ano07k] Anonymous. Book review: Math Refresher for Scientist and Engineers by J. R. Fanchi. Technometrics, 49(1):109–110, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471296.

Anonymous:2007:BRBs

[Ano07l] Anonymous. Book review: Methods and Models in Statistics, in Honour of Professor John Nelder, FRS by Niall M. Adams; Martin Crowder; D. J. Hand; Dave Stephens. Technometrics, 49 (4):501, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471414.

Anonymous:2007:BRBc

[Ano07m] Anonymous. Book review: Models for Discrete Data by D. Zelterman. Technometrics, 49(1):110, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471298.

Anonymous:2007:BRBh

[Ano07n] Anonymous. Book review: Regression Analysis by Example by Samprit Chatterjee; Ali S. Hadi. Technometrics, 49(2):232, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471338.

Anonymous:2007:BRBt

[Ano07o] Anonymous. Book review: Response Surface Methodology and Related Topics by André I. Khuri. Technometrics, 49(4):501, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471412.

Anonymous:2007:BRBp

[Ano07p] Anonymous. Book review: Response Surfaces, Mixtures and Ridge Analyses by George E. P. Box; Norman R. Draper. Technometrics, 49(4):500, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471410.

Anonymous:2007:BRBg

[Ano07q] Anonymous. Book review: Selected Papers of Frederick Mosteller by Stephen E. Fienberg; David C. Hoaglin. Technometrics, 49 (2):231, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471335.

Anonymous:2007:BRBd

[Ano07r] Anonymous. Book review: Selected Statistical Papers of Sir David Cox, Vol. 1: Design of Investigations, Statistical Methods and Applications by D. J. Hand; A. M. Hertzberg. Technometrics, 49(1):110–111, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471299.

Anonymous:2007:BRBf

[Ano07s] Anonymous. Book review: Selected Statistical Papers of Sir David Cox, Vol. 2: Foundations of Statistical Inference, Theoretical Statistics, Time Series and Stochastic Processes by D. J. Hand; A. M. Hertzberg. Technometrics, 49(1):111, February 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471300.

Anonymous:2007:BRBi

[Ano07t] Anonymous. Book review: Series Approximation Methods in Statistics by J. E. Kolassa. Technometrics, 49(2):232, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471337.

Anonymous:2007:BRBo

[Ano07u] Anonymous. Book review: Survival and Event History Analysis by Per Kragh Andersen; Niels Keiding. Technometrics, 49(3):368, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471377.

Anonymous:2007:BRBe

[Ano07v] Anonymous. Book review: The Art of Semiparametrics by S. Sperlich; Wolfgang Hardle; Gokhan Aydinli. Technometrics, 49 (1):111, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471301.

Anonymous:2007:BRBb

[Ano07w] Anonymous. Book review: The Cambridge Dictionary of Statistics by B. S. Everitt. Technometrics, 49(1):110, February 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471297.

Anonymous:2007:BRBj

[Ano07x] Anonymous. Book review: Tools for Construction Chronologies: Crossing Disciplinary Boundaries by Caitlin E. Buck; Andrew R. Millard. Technometrics, 49(2):232, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471336.

Anonymous:2007:FMa

[Ano07y] Anonymous. Front matter. Technometrics, 49(1):??, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471269.

Anonymous:2007:FMb

[Ano07z] Anonymous. Front matter. Technometrics, 49(2):??, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471304.

Anonymous:2007:FMc

[Ano07-27] Anonymous. Front matter. Technometrics, 49(3):??, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471342.

Anonymous:2007:FMd

[Ano07-28] Anonymous. Front matter. Technometrics, 49(4):??, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471381.

Anonymous:2007:VI

[Ano07-29] Anonymous. Volume information. Technometrics, 49(4):??, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471380.

Anonymous:2008:BMa

[Ano08a] Anonymous. Back matter. Technometrics, 50(1):??, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471450.

Anonymous:2008:BMb

[Ano08b] Anonymous. Back matter. Technometrics, 50(2):??, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471480.

Anonymous:2008:BMc

[Ano08c] Anonymous. Back matter. Technometrics, 50(3):??, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471515.

Anonymous:2008:BMd

[Ano08d] Anonymous. Back matter. Technometrics, 50(4):??, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471552.

Anonymous:2008:BRBm

[Ano08e] Anonymous. Book review: Bootstrap Methods: A Guide for Practitioners and Researchers by Michael R. Chernick. Technometrics, 50(4):550, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471551.

Anonymous:2008:BRBk

[Ano08f] Anonymous. Book review: Design and Analysis of Experiments, Vol. 1, Introduction to Experimental Design by Klaus Hinkelmann; Oscar Kempthorne. Technometrics, 50(4):549, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471548.

Anonymous:2008:BRBd

[Ano08g] Anonymous. Book review: Exact Analysis of Discrete Data by Karim F. Hirji. Technometrics, 50(3):412, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471509.

Anonymous:2008:BRBg

[Ano08h] Anonymous. Book review: Handbook of Parallel Computing and Statistics by Erricos John Kontoghiorghes. Technometrics, 50(3): 413, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471513.

Anonymous:2008:BRBn

[Ano08i] Anonymous. Book review: Introduction to Bayesian Statistics by William M. Bolstad. Technometrics, 50(4):550, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471550.

Anonymous:2008:BRBa

[Ano08j] Anonymous. Book review: JMP Start Statistics: A Guide to Statistics and Data Analysis Using JMP by John Sall; Lee Creighton; Ann Lehman. Technometrics, 50(2):239, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471479.

Anonymous:2008:BRBb

[Ano08k] Anonymous. Book review: Learning from Data: Concepts, Theory, and Methods by Vladimir Cherkassky; Filip M. Mulier. Technometrics, 50(2):239, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471478.

Anonymous:2008:BRBh

[Ano08l] Anonymous. Book review: Modern Statistical and Mathematical Methods in Reliability by Alyson Wilson; Nikolaos Limnios; Sallie Keller-McNulty; Yvonne Armijo. Technometrics, 50(3):413, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471512.

Anonymous:2008:BRBc

[Ano08m] Anonymous. Book review: Permutation Methods: A Distance Function Approach by Paul W. Mielke,; Kenneth J. Berry. Technometrics, 50(2):239, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471477.

Anonymous:2008:BRBi

[Ano08n] Anonymous. Book review: Predicting Structured Data by Gokhan Bakir; Thomas Hofmann; Bernard Scholkopf; Alexander J. Smola; Ben Taskar; S. V. N. Vishwanathan. Technometrics, 50(3):413– 414, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471514.

Anonymous:2008:BRBe

[Ano08o] Anonymous. Book review: SAS Graphics for Java: Examples Using SAS AppDev Studio and the Output Delivery System by Wendy Bohnenkamp; Jackie Iverson. Technometrics, 50(3):412, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471510.

Anonymous:2008:BRBl

[Ano08p] Anonymous. Book review: Statistical Analysis of Extreme Values with Applications to Insurance, Finance, Hydrology and Other Fields by R.-D. Reiss; M. Thomas. Technometrics, 50(4):549, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471549.

Anonymous:2008:BRBf

[Ano08q] Anonymous. Book review: Stochastic Orders by Moshe Shaked; J. George Shanthikumar. Technometrics, 50(3):412–413, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471511.

Anonymous:2008:BRBj

[Ano08r] Anonymous. Book review: The Analysis of Cross-Classified Categorical Data by Stephen E. Fienberg. Technometrics, 50(4):549, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471547.

Anonymous:2008:CCA

[Ano08s] Anonymous. Correction: “A Critical Assessment of Two-Stage Group Screening through Industrial Experimentation”. Technometrics, 50(2):230, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471463. See [VDLB08].

Anonymous:2008:CAF

[Ano08t] Anonymous. Correction: “Advanced Fuzzy Logic Technologies in Industrial Applications”. Technometrics, 50(1):99, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471449. See

[Bar07a].

Anonymous:2008:FMa

[Ano08u] Anonymous. Front matter. Technometrics, 50(1):??, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471417.

Anonymous:2008:FMb

[Ano08v] Anonymous. Front matter. Technometrics, 50(2):??, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471452.

Anonymous:2008:FMc

[Ano08w] Anonymous. Front matter. Technometrics, 50(3):??, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471482.

Anonymous:2008:FMd

[Ano08x] Anonymous. Front matter. Technometrics, 50(4):??, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471518.

Anonymous:2008:VI

[Ano08y] Anonymous. Volume information. Technometrics, 50(4):??, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471517.

Anonymous:2009:BRBi

[Ano09a] Anonymous. Book review: Acceptance Sampling in Quality Control (2nd ed.) by Edward G. Schilling; Dean V. Neubauer. Technometrics, 51(3):346–347, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586639.

Anonymous:2009:BRBh

[Ano09b] Anonymous. Book review: Applied Statistics for Engineers and Physical Scientists (3rd ed.) by Johannes Ledolter; Robert V. Hogg. Technometrics, 51(3):346, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586638.

Anonymous:2009:BRBb

[Ano09c] Anonymous. Book review: Bayesian Bounds for Parameter Estimation and Nonlinear Filtering/Tracking by Harry L. Van Trees; Kristine L. Bell. Technometrics, 51(1):108, February 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586585.

Anonymous:2009:BRBk

[Ano09d] Anonymous. Book review: Continuous Bivariate Distributions (2nd ed.) by N. Balakrishnan; Chin-Diew Lai. Technometrics, 51 (4):487, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586669.

Anonymous:2009:BRBl

[Ano09e] Anonymous. Book review: Design Theory (2nd ed.) by C. C. Lindner; C. A. Rodger. Technometrics, 51(4):487–488, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586670.

Anonymous:2009:BRBm

[Ano09f] Anonymous. Book review: Handbook of Statistical Analyses Using SAS (3rd ed.) by Geoff Der; Brian S. Everitt. Technometrics, 51(4):488, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586671.

Anonymous:2009:BRBc

[Ano09g] Anonymous. Book review: Introduction to Design of Experiments With JMP Examples (3rd ed.) by Jacques Goupy; Lee Creighton. Technometrics, 51(1):108, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586584.

Anonymous:2009:BRBa

[Ano09h] Anonymous. Book review: Introduction to Discrete Event Systems (2nd ed.) by Christos G. Cassandras; Stephane Lafortune. Technometrics, 51(1):107–108, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586583.

Anonymous:2009:BRBf

[Ano09i] Anonymous. Book review: Introductory Statistics With R (2nd ed.) by Peter Dalgaard. Technometrics, 51(2):225, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586615.

Anonymous:2009:BRBg

[Ano09j] Anonymous. Book review: Random Effect and Latent Variable Model Selection by David Dunson. Technometrics, 51(2):225, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586614.

Anonymous:2009:BRBe

[Ano09k] Anonymous. Book review: Recent Advances in Linear Models and Related Areas by Shalabh Heumann; Christian Heumann. Technometrics, 51(2):224–225, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586613.

Anonymous:2009:BRBd

[Ano09l] Anonymous. Book review: Statistical Models and Methods for Biomedicai and Technical Systems by Filia Vonta; Mikhail Nikulin; Nikolaos Limnios; Catherine Huber-Carol. Technometrics, 51(2): 224, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586612.

Anonymous:2009:BRBj

[Ano09m] Anonymous. Book review: The Elements of Statistical Learning: Data Mining, Inference, and Prediction (2nd ed.) by Trevor Hastie; Robert Tibshirani; Jerome Friedman. Technometrics, 51 (3):347, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586640.

Anonymous:2009:EC

[Ano09n] Anonymous. Editorial collaborators. Technometrics, 51(4):490– 491, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586672.

Anscombe:2003:LE

[Ans03] F. R. Anscombe. Letter to the Editor. Technometrics, 45(1):45– 46, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25046973.

Arslan:2005:BRB

[Ars05] Olcay Arslan. Book review: Measurement and Multivariate Analysis by S. Nishisato; Y. Baba; H. Bozdogan; K. Kanefuji. Technometrics, 47(1):95–96, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470941.

Apley:2001:FAM

[AS01] Daniel W. Apley and Jianjun Shi. A factor-analysis method for diagnosing variability in multivariate manufacturing processes. Technometrics, 43(1):84–95, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270860.

Anderson:2006:DLG

[ASW^{+}06] Dale N. Anderson, David C. Stromswold, Sharon C. Wunschel, Anthony J. Peurrung, and Randy R. Hansen. Detection and location of gamma-ray sources with a modulating coded mask. Technometrics, 48(2):252–261, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471161.

Atkinson:2005:ROD

[Atk05] Anthony C. Atkinson. Robust optimum designs for transformation of the responses in a multivariate chemical kinetic model. Technometrics, 47(4):478–487, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471072.

Aviles:2001:BRB

[Avi01] Ana Ivelisse Avilés. Book review: Linear Mixed Models for Longitudinal Data by Geert Verbeke; Geert Molenberghs. Technometrics, 43(3):375, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271233.

Anderson:2001:BRB

[AW01] Mark Anderson and Patrick Whitcomb. Book review: Design of Experiments: Statistical Principles of Research Design and Analysis by Robert O. Kuehl. Technometrics, 43(2):236–237, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271044.

Anderson:2005:BRB

[AW05] Mark Anderson and Patrick Whitcomb. Book review: Practical Guide to Designed Experiments: A Unified Modular Approach by Paul D. Funkenbusch. Technometrics, 47(4):515, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471076.

Babu:2002:BRBb

[Bab02a] G. Jogesh Babu. Book review: Multivariate Permutation Tests by Fortunato Pesarin. Technometrics, 44(3):290, August 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270498.

Babu:2002:BRBa

[Bab02b] Gutti J. Babu. Book review: Fundamentals of Modern Statistical Methods by Rand R. Wilcox. Technometrics, 44(1):83, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270691.

Babu:2005:BRB

[Bab05] Gutti J. Babu. Book review: Bootstrap Techniques for Signal Processing by Abdelhak M. Zoubir; D. Robert Iskander. Technometrics, 47(3):374–375, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471036.

Bagui:2005:BRB

[Bag05] Subhash C. Bagui. Book review: Combining Pattern Classifiers: Methods and Algorithms by Ludmila I. Kuncheva. Technometrics, 47(4):517–518, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471080.

Bailey:2004:BRB

[Bai04] Mark Bailey. Book review: Statistical Reasoning for Everyday Life by Jeffrey O. Bennett; William L. Briggs; Mario F. Triola. Technometrics, 46(4):490–491, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470907.

Bailey:2005:BRB

[Bai05] Mark Bailey. Book review: The Complete Idiot’s Guide to Statistics by Robert A. Donnelly. Technometrics, 47(4):524, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471090.

Bailey:2007:BRBb

[Bai07a] Mark Bailey. Book review: Six Sigma Quality Improvement with MINITAB by G. Robin Henderson. Technometrics, 49(1):105, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471289.

Bailey:2007:BRBa

[Bai07b] Mark Bailey. Book review: The Grammar of Graphics by Leland Wilkinson. Technometrics, 49(1):104, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471286.

Bajorski:2000:BRB

[Baj00] Peter Bajorski. Book review: Statistics in Civil Engineering by Andrew V. Metcalfe. Technometrics, 42(2):210–211, May 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271461.

Bakir:2000:BRB

[Bak00] Saad T. Bakir. Book review: Nonparametric Regression and Spline Smoothing by Randall L. Eubank. Technometrics, 42(3):318, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271102.

Bakir:2001:BRB

[Bak01] Saad T. Bakir. Book review: The Practice of Statistics: Putting the Pieces Together by John D. Spurrier. Technometrics, 43(3):373, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271230.

Barrett:2001:BRB

[Bar01] John D. Barrett. Book review: The Process Evaluation Handbook by Donald J. Wheeler. Technometrics, 43(4):494–495, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270834.

Barker:2002:BRB

[Bar02] Lawrence Barker. Book review: Pattern Classification by Richard O. Duda; Peter E. Hart; David G. Stork. Technometrics, 44(1):87, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270697.

Barrett:2004:BRB

[Bar04] John D. Barrett. Book review: Quality from Customer Needs to Customer Satisfaction by Bo Bergman; Bengt Klefsjö. Technometrics, 46(1):118, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470781.

Barker:2005:BRB

[Bar05] Lawrence E. Barker. Book review: Logit Models from Economics and Other Fields by J. S. Cramer. Technometrics, 47(1):97–98, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470945.

Barrett:2007:BRBc

[Bar07a] J. Douglas Barrett. Book review: Advanced Fuzzy Logic Technologies in Industrial Applications by Ying Bai; Hangi Zhuang; Dali Wang. Technometrics, 49(4):494–495, November 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471401. See correction

[Ano08t].

Barrett:2007:BRBb

[Bar07b] J. Douglas Barrett. Book review: Diagnosis and Fault-Tolerant Control by Mogens Blanke; Michael Kinnaert; Jan Lunze; Marcel Staroswiecki. Technometrics, 49(4):493–494, November 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471399.

Barrett:2007:BRBa

[Bar07c] J. Douglas Barrett. Book review: Taguchi’s Quality Engineering Handbook by Genichi Taguchi; Subir Chowdhury; Yuin Wu. Technometrics, 49(2):224–225, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471320.

Barker:2008:BRBa

[Bar08a] Lawrence Barker. Book review: Margins of Error: A Study of Reliability in Survey Measurements by Duane F. Alwin. Technometrics, 50(2):237, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471472.

Barker:2008:BRBb

[Bar08b] Lawrence Barker. Book review: Register-Based Statistics, Administrative Data for Statistical Purposes by Anders Walgren; Britt Walgren. Technometrics, 50(4):546, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471541.

Barrett:2008:BRB

[Bar08c] J. Douglas Barrett. Book review: Process Control Performance Assessment: From Theory to Implementation by Andrzej W. Ordys; Damien Uduehi; Michael Johnson. Technometrics, 50(1):92– 93, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471439.

Barrett:2009:BRB

[Bar09] J. Douglas Barrett. Book review: Applied Nonparametric Statistical Methods (4th ed.) by Peter Sprent; Nigel C. Smeeton. Technometrics, 51(2):217, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586600.

Basu:2000:BRB

[Bas00] Asit P. Basu. Book review: The Inverse Gaussian Distribution by V. Seshadri. Technometrics, 42(4):436, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270966.

Basu:2002:BRB

[Bas02] Asit P. Basu. Book review: Probabilistic Risk Analysis: Foundations and Methods by Tim Bedford; Roger Cooke. Technometrics, 44(4):402, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271558.

Bates:2008:FSC

[Bat08] Douglas Bates. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):439–440, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471523.

Bayne:2002:BRB

[Bay02] Charles K. Bayne. Book review: Multivariate Analysis of Quality: An Introduction by Harald Martens; Magni Martens. Technometrics, 44(2):186–187, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271263.

Badcock:2004:TPM

[BBJK04] Julie Badcock, Trevor C. Bailey, Philip Jonathan, and Wojtek J. Krzanowski. Two projection methods for use in the analysis of multivariate process data with an illustration in petrochemical production. Technometrics, 46(4):392–403, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470887.

Bayarri:2007:FVC

[BBP^{+}07] Maria J. Bayarri, James O. Berger, Rui Paulo, Jerry Sacks, John A. Cafeo, James Cavendish, Chin-Hsu Lin, and Jian Tu. A framework for validation of computer models. Technometrics, 49(2):138–154, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471307.

Bayarri:2009:SIC

[BBS09] M. J. Bayarri, Jim Berger, and David M. Steinberg. Special issue on computer modeling: [introduction]. Technometrics, 51 (4):353, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586645.

Bingham:2007:IPI

[BC07] Derek R. Bingham and Hugh A. Chipman. Incorporating prior information in optimal design for model selection. Technometrics, 49(2):155–163, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471308.

Brettschneider:2008:QASa

[BCBS08a] Julia Brettschneider, François Collin, Benjamin M. Bolstad, and Terence P. Speed. Quality assessment for short oligonucleotide microarray data. Technometrics, 50(3):241–264, August 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471483.

Brettschneider:2008:QASb

[BCBS08b] Julia Brettschneider, François Collin, Benjamin M. Bolstad, and Terence P. Speed. [Quality Assessment for Short Oligonucleotide Microarray Data]: Rejoinder. Technometrics, 50(3):279–283, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471489.

Bayarri:2009:USC

[BCL^{+}09] M. J. Bayarri, Eliza S. Calder, Simon Lunagomez, E. Bruce Pitman, James O. Berger, Keith Dalbey, Abani K. Patra, Elaine T. Spiller, and Robert L. Wolpert. Using statistical and computer models to quantify volcanic hazards. Technometrics, 51(4):402– 413, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586650.

Bhattacharjee:2004:UTG

[BDNN04] M. Bhattacharjee, J. V. Deshpande, and U. V. Naik-Nimbalkar. Unconditional tests of goodness of fit for the intensity of timetruncated nonhomogeneous Poisson processes. Technometrics, 46 (3):330–338, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470847.

Basu:2003:BSR

[BE03] Sanjib Basu and Nader Ebrahimi. Bayesian software reliability models based on martingale processes. Technometrics, 45(2):150– 158, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047012.

Bedrick:2000:CLF

[Bed00] Edward J. Bedrick. Checking for lack of fit in linear models with parametric variance functions. Technometrics, 42(3):227–236, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271078.

Belunis:2001:BRB

[Bel01] Nancy Belunis. Book review: Statistical Process Analysis by Layth C. Alwan. Technometrics, 43(2):231, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271036.

Burr:2005:BRC

[BF05] Tom L. Burr and Herbert A. Fry. Biased regression: The case for cautious application. Technometrics, 47(3):284–296, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471022.

Boente:2008:RMT

[BF08] Graciela Boente and Andrés Farall. Robust multivariate tolerance regions: influence function and Monte Carlo study. Technometrics, 50(4):487–500, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471532.

Beattie:2002:TSB

[BFL02] Scott D. Beattie, Duncan K. H. Fong, and Dennis K. J. Lin. A twostage Bayesian model selection strategy for supersaturated designs.

Technometrics, 44(1):55–63, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270684.

Bhaumik:2004:UPL

[BG04] Dulal Kumar Bhaumik and Robert David Gibbons. An upper prediction limit for the arithmetic mean of a lognormal random variable. Technometrics, 46(2):239–248, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470808.

Bhaumik:2005:CRR

[BG05] Dulal K. Bhaumik and Robert D. Gibbons. Confidence regions for random-effects calibration curves with heteroscedastic errors. Technometrics, 47(2):223–230, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470984.

Bhaumik:2006:OSA

[BG06] Dulal Kumar Bhaumik and Robert David Gibbons. One-sided approximate prediction intervals for at least p of m observations from a gamma population at each of r locations. Technometrics, 48(1):112–119, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471120.

Bayarri:2005:BSL

[BGD05] M. J. Bayarri and G. García-Donato. A Bayesian sequential look at μ-control charts. Technometrics, 47(2):142–151, May 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470976.

Ben-Gal:2003:CBS

[BGMS03] Irad Ben-Gal, Gail Morag, and Armin Shmilovici. Context-based statistical process control: A monitoring procedure for statedependent processes. Technometrics, 45(4):293–311, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047086.

Bates:2003:GSP

[BGW03] Ron A. Bates, Beatrice Giglio, and Henry P. Wynn. A global selection procedure for polynomial interpolators. Technometrics, 45(3):246–255, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047051.

Box:2000:FFD

[BH00] G. E. P. Box and J. S. Hunter. The 2^{k−p} fractional factorial designs Part I. Technometrics, 42(1):28–47, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271430.

Bao:2008:QAS

[BH08] Lei Bao and Ina Hoeschele. [Quality Assessment for Short Oligonucleotide Microarray Data]: Comment. Technometrics, 50(3):268– 271, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471485.

Bingham:2001:BRB

[Bin01] Derek R. Bingham. Book review: Graphical Methods for the Design of Experiments by Russell R. Barton. Technometrics, 43(1):97, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270861.

Buvaneswari:2007:SVT

[BJL^{+}07] A. Buvaneswari, David A. James, Chuanhai Liu, John M. Graybeal, Diane Lambert, and W. Michael MacDonald. A statistical view of the transient signals that support a wireless call. Technometrics, 49(3):305–317, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471350.

Bae:2004:NRC

[BK04] Suk Joo Bae and Paul H. Kvam. A nonlinear random-coefficients model for degradation testing. Technometrics, 46(4):460–469, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470894.

Bhaumik:2009:TPG

[BKG09] Dulal K. Bhaumik, Kush Kapur, and Robert D. Gibbons. Testing parameters of a gamma distribution for small samples. Technometrics, 51(3):326–334, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586626.

Blackwood:2001:BRB

[Bla01] Larry G. Blackwood. Book review: Risk Modeling for Determining Value and Decision Making by Glenn Koller. Technometrics, 43(3):377–378, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271236.

Baron:2001:MRF

[BLC01] Michael Baron, Choudur K. Lakshminarayan, and Zhenwu Chen. Markov random fields in pattern recognition for semiconductor manufacturing. Technometrics, 43(1):66–72, February 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270858.

Browne:2009:TSL

[BMS09] Ryan P. Browne, R. Jock MacKay, and Stefan H. Steiner. Twostage leveraged measurement system assessment. Technometrics, 51(3):239–249, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586619.

Brenneman:2001:MID

[BN01] William A. Brenneman and Vijayan N. Nair. Methods for identifying dispersion effects in unreplicated factorial experiments: A critical analysis and proposed strategies. Technometrics, 43(4): 388–405, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270810.

Bastos:2009:DGP

[BO09] Leonardo S. Bastos and Anthony O’Hagan. Diagnostics for Gaussian process emulators. Technometrics, 51(4):425–438, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586652.

Bodt:2002:BRBb

[Bod02a] Barry A. Bodt. Book review: Computer Intrusion Detection and Network Monitoring by David J. Marchette. Technometrics, 44 (3):294–295, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270506.

Bodt:2002:BRBa

[Bod02b] Barry A. Bodt. Book review: Modern Statistics for Engineering and Quality Improvement by John Lawson; John Erjavec. Technometrics, 44(2):186, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271262.

Booth:2000:BRB

[Boo00] David E. Booth. Book review: Analysis of Incomplete Multivariate Data by J. L. Schafer. Technometrics, 42(2):213–214, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271466.

Booth:2003:BRB

[Boo03] David E. Booth. Book review: Applied Multivariate Analysis by Neil H. Timm. Technometrics, 45(2):174, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047021.

Booth:2004:BRBa

[Boo04a] David E. Booth. Book review: Chemometrics: Data Analysis for the Laboratory and Chemical Plant by Richard G. Brereton. Technometrics, 46(1):110, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470767.

Booth:2004:BRBb

[Boo04b] David E. Booth. Book review: Teaching Statistics by Andrew Gelman; Deborah Nolan. Technometrics, 46(1):114, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470774.

Booth:2005:BRBa

[Boo05a] David E. Booth. Book review: Chemometrics: From Basics to Wavelet Transform by Foo-Tim Chau; Yi-Zeng Liang; Junbin Gao; Xue-Guang Shao. Technometrics, 47(2):240, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470998.

Booth:2005:BRBb

[Boo05b] David E. Booth. Book review: Multi-Way Analysis: Applications in the Chemical Sciences by Age Smilde; Rasmus Bro; Paul Geladi. Technometrics, 47(4):518–519, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471081.

Booth:2007:BRBa

[Boo07a] David E. Booth. Book review: Cancer Diagnostics with DNA Microarrays by Steen Knudsen. Technometrics, 49(4):492–493, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471398.

Booth:2007:BRBb

[Boo07b] David E. Booth. Book review: Data Mining Methods and Models by Daniel T. Larose. Technometrics, 49(4):500, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471409.

Booth:2008:BRB

[Boo08] David E. Booth. Book review: The Cross-Entropy Method by Reuven Y. Rubinstein; Dirk P. Kroese. Technometrics, 50(1): 92, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471438.

Booth:2009:BRBc

[Boo09a] David E. Booth. Book review: A History of Mathematics (3rd ed.) by Victor J. Katz. Technometrics, 51(4):487, November 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586668.

Booth:2009:BRBb

[Boo09b] David E. Booth. Book review: Advances in Mathematical and Statistical Modeling (Statistics for Industry and Technology) by Barry C. Arnold; N. Balakrishnan; Jose-Maria Sarabia; Robert Minquez. Technometrics, 51(3):346, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586637.

Booth:2009:BRBa

[Boo09c] David E. Booth. Book review: Introduction to Distribution Logistics by Paolo Brandimarte; Giulio Zotteri. Technometrics, 51 (1):105–106, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586579.

Borkowf:2000:BRB

[Bor00] Craig B. Borkowf. Book review: Random Number Generation and Monte Carlo Methods by James E. Gentle. Technometrics, 42(4):431–432, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270960.

Borowiak:2001:BRB

[Bor01] Dale Borowiak. Book review: Linear Models, Least Squares and Alternatives by C. Radhakrishna Rao; Helge Toutenburg. Technometrics, 43(1):99, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270864.

Borkowf:2002:BRB

[Bor02] Craig B. Borkowf. Book review: Time-Series Forecasting by Chris Chatfield. Technometrics, 44(2):194–195, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271275.

Boyles:2001:GCP

[Boy01] Russell A. Boyles. Gauge capability for pass-fail inspection. Technometrics, 43(2):223–229, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271035.

Bresky:2001:BRB

[Bre01] Norman Bresky. Book review: Achieving Quality through Continual Improvement by Claude W. Burrill; Johannes Ledolter. Technometrics, 43(1):108, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270877.

Breneman:2002:BRB

[Bre02] James E. Breneman. Book review: Probability: The Science of Uncertainty by Michael A. Bean. Technometrics, 44(3):294, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270504.

Bresky:2004:BRB

[Bre04] Norman P. Bresky. Book review: Risk Analysis in Engineering and Economics by Bilal M. Ayyub. Technometrics, 46(3):361, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470854.

Breneman:2005:BRB

[Bre05a] James Breneman. Book review: Kernel Methods for Pattern Analysis by John Shawe-Taylor; Nello Cristianini. Technometrics, 47 (2):237–238, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470995.

Brenneman:2005:BRB

[Bre05b] William A. Brenneman. Book review: Statistics for Research by Shirley Dowdy; Stanley Wearden; Daniel Chilko. Technometrics, 47(1):100, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470948.

Breneman:2006:BRB

[Bre06] James E. Breneman. Book review: Flowgraph Models for Multistate Time-to-Event Data by Aparna V. Huzurbazar. Technometrics, 48(1):154–155, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471138.

Bresky:2007:BRB

[Bre07] Norman Bresky. Book review: Root Cause Analysis: Simplified Tools and Techniques by Bjorn Andersen; Tom Fagerhaug. Technometrics, 49(3):364, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471367.

Brillinger:2001:DAK

[Bri01] David R. Brillinger. Does anyone know when the correlation coefficient is useful? A study of the times of extreme river flows. Technometrics, 43(3):266–273, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271213.

Brill:2003:BRB

[Bri03] Robert V. Brill. Book review: Statistics in Plain English by Timothy C. Urdan. Technometrics, 45(2):179–180, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047029.

Brill:2004:BRB

[Bri04] Robert V. Brill. Book review: Applied Statistics and Probability for Engineers by Douglas C. Montgomery; George C. Runger. Technometrics, 46(1):112–113, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470772.

Brill:2005:BRB

[Bri05] Robert V. Brill. Book review: Basic Statistics and Data Analysis by Larry J. Kitchens. Technometrics, 47(1):99–100, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470947.

Brillinger:2006:DTP

[Bri06] David R. Brillinger. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):325–327, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471201. See [Mal06].

Brugger:2000:BRB

[Bru00] Richard M. Brugger. Book review: Quality Sampling and Reliability: New Uses for the Poisson Distribution by John J. Heldt. Technometrics, 42(3):314, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271096.

Bingham:2003:FFS

[BS03] Derek Bingham and Randy R. Sitter. Fractional factorial splitplot designs for robust parameter experiments. Technometrics, 45 (1):80–89, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25046977.

Brusco:2009:EAH

[BSC09] Michael J. Brusco, Douglas Steinley, and J. Dennis Cradit. An exact algorithm for hierarchically well-formulated subsets in secondorder polynomial regression. Technometrics, 51(3):306–315, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586624.

Balasooriya:2000:PCR

[BSG00] Uditha Balasooriya, Sutaip L. C. Saw, and Veeresh Gadag. Progressively censored reliability sampling plans for the Weibull distribution. Technometrics, 42(2):160–167, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271448.

Buja:2006:DTP

[Buj06] Andreas Buja. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):327–330, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471202. See [Mal06].

Bukkapatnam:2009:BRB

[Buk09] Satish T. S. Bukkapatnam. Book review: Stream of Variation Modeling and Analysis for Multistage Manufacturing Processes by Jianjun Shi. Technometrics, 51(4):479–480, November 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586657.

Burr:2000:BRB

[Bur00] Tom Burr. Book review: Modelling and Estimation of Measurement Errors by Michèle Neuilly; Michèle Cetama. Technometrics, 42(4):427–428, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270955.

Burdick:2001:BRB

[Bur01a] Richard K. Burdick. Book review: Linear Models in Statistics by Alvin C. Rencher. Technometrics, 43(2):234–236, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271042.

Burr:2001:BRB

[Bur01b] Tom Burr. Book review: Introduction to Matrix Analytic Methods in Stochastic Modeling by G. Latouche; V. Ramaswami. Technometrics, 43(3):379–380, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271239.

Burdick:2002:BRB

[Bur02] Richard K. Burdick. Book review: Analysis of Messy Data Volume III: Analysis of Covariance by George A. Milliken; Dallas E. Johnson. Technometrics, 44(3):288, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270495.

Burr:2003:BRB

[Bur03] Tom Burr. Book review: Causation, Prediction, and Search by Peter Spirtes; Clark Glymour; Richard Scheines. Technometrics, 45(3):272–273, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047063.

Burr:2004:BRB

[Bur04] Thomas L. Burr. Book review: Bayesian Inference: Parameter Estimation and Decisions by H. L. Harney. Technometrics, 46(2):250–251, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470810.

Burr:2006:BRB

[Bur06] Tom L. Burr. Book review: Modeling Longitudinal Data by Robert E. Weiss. Technometrics, 48(3):447–448, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471220.

Burr:2007:BRB

[Bur07] Thomas L. Burr. Book review: Modeling Financial Time Series with S-Plus by Eric Zivot; Jiahui Wang. Technometrics, 49(1): 105–106, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471290.

Burr:2008:BRBa

[Bur08a] Thomas Burr. Book review: Multidimensional Nonlinear Descriptive Analysis (MUNDA) by Shizuhiko Nishisato. Technometrics, 50(1):88–89, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471432.

Burr:2008:BRBb

[Bur08b] Thomas Burr. Book review: Statistics for Fission Track Analysis by Rex Galbreith. Technometrics, 50(4):543, November 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471537.

Burr:2009:BRBa

[Bur09a] Thomas L. Burr. Book review: Simulation and Inference for Stochastic Differential Equations With R Examples by Stefano Iacus. Technometrics, 51(3):340–341, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586630.

Burr:2009:BRBb

[Bur09b] Tom Burr. Book review: Introduction to Machine Learning and Bioinformatics by S. Mitra; S. Datta; T. Perkins; G. Michalidis. Technometrics, 51(4):482–483, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586662.

Butler:2004:CTL

[But04] Neil A. Butler. Construction of two-level split-lot fractional factorial designs for multistage processes. Technometrics, 46(4):445– 451, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470892.

Borkowski:2001:CDO

[BV01] John J. Borkowski and Elsie S. Valeroso. Comparison of design optimality criteria of reduced models for response surface designs in the hypercube. Technometrics, 43(4):468–477, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270818.

Bzik:2004:BRB

[Bzi04] Thomas J. Bzik. Book review: Six Sigma and Beyond: Design of Experiments by D. H. Stamatis. Technometrics, 46(4): 489–490, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470905.

Chin:2006:ODS

[CA06] Chang-Ho Chin and Daniel W. Apley. Optimal design of secondorder linear filters for control charting. Technometrics, 48(3):337– 348, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471206.

Caby:2000:BRB

[Cab00] Errol Caby. Book review: Regression and Time Series Model Selection by Allan D. R. McQuarrie; Chih-Ling Tsai. Technometrics, 42(2):214–216, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271469.

Caby:2003:BRB

[Cab03] Errol Caby. Book review: Elements of Applied Stochastic Processes by U. Narayan Bhat; Gregory K. Miller. Technometrics, 45(3):270–271, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047061.

Caby:2004:BRB

[Cab04] Errol C. Caby. Book review: Data Mining Using SAS Applications by George Fernandez. Technometrics, 46(2):260–261, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470822.

Caby:2006:BRB

[Cab06] Errol C. Caby. Book review: An Introduction to Statistical Signal Processing by Robert M. Gray; Lee D. Davidson. Technometrics, 48(4):572–573, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471256.

Caby:2008:BRB

[Cab08] Errol Caby. Book review: Probability Theory in Finance: A Mathematical Guide to the Black-Scholes Formula by Sean Dineen. Technometrics, 50(2):234–235, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471467.

Caby:2009:BRB

[Cab09] Errol Caby. Book review: Vertical Density Representation and Its Applications by Marvin D. Troutt; W. K. Pang; S. H. Hou. Technometrics, 51(2):218, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586602.

Cahya:2006:BRB

[Cah06] Suntara Cahya. Book review: Applied Data Analysis for Process Improvement by James L. Lamprecht. Technometrics, 48 (2):306–307, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471172.

Campbell:2001:BRBa

[Cam01a] Katherine Campbell. Book review: Flood Frequency Analysis by A. Ramachandra Rao; Khaled H. Hamed. Technometrics, 43(2):238, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271047.

Campbell:2001:BRBb

[Cam01b] Katherine Campbell. Book review: Nonlinear Time Series Analysis by H. Kantz; T. Schreiber. Technometrics, 43(4):491, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270829.

Campbell:2002:BRB

[Cam02] Katherine Campbell. Book review: Statistical Analysis of Extreme Values by R.-D. Reiss; M. Thomas. Technometrics, 44(3):295–296, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270507.

Carpenter:2000:BRB

[Car00] Mark Carpenter. Book review: The New Statistical Analysis of Data by T. W. Anderson; Jeremy D. Finn. Technometrics, 42 (2):205–206, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271454.

Carbno:2007:BRBb

[Car07a] Collin Carbno. Book review: Actuarial Theory for Dependent Risks: Measures, Orders, and Models by M. Denuit; J. Dahaene; M. Goovaerts; R. Kaas. Technometrics, 49(4):495, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471402.

Carbno:2007:BRBc

[Car07b] Collin Carbno. Book review: Business Statistics: Contemporary Decision Making by Ken Black. Technometrics, 49(4):495–496, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471403.

Carbno:2007:BRBd

[Car07c] Collin Carbno. Book review: Thinking Statistically (Elephants Go to School) by Sarjinder Singh. Technometrics, 49(4):496, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471405.

Carbno:2007:BRBa

[Car07d] Collin Carbno. Book review: Uncertainty Analysis with High Dimensional Dependence Modelling by K. Kurowicka; R. Cooke. Technometrics, 49(1):108, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471294.

Castelli:2005:BRB

[Cas05] Vittorio Castelli. Book review: Bayesian Nonparametrics via Neural Networks by Herbert K. H. Lee. Technometrics, 47(2):235–236, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470991.

Cho:2009:CBP

[CCD09] Jung Jin Cho, Yong Chen, and Yu Ding. Calculating the breakdown point of sparse linear models. Technometrics, 51(1):34– 46, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586561.

Chen:2002:DPA

[CE02] Argon Chen and E. A. Elsayed. Design and performance analysis of the exponentially weighted moving average mean estimate for processes subject to random step changes. Technometrics, 44 (4):379–389, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271545.

Carreras:2008:RSD

[CFP08] Francesc Carreras, Josep Freixas, and María Albina Puente. On the reliability of the self-dual k-out-of-n systems. Technometrics, 50(1):79–85, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471428.

Cumming:2009:SSB

[CG09] Jonathan A. Cumming and Michael Goldstein. Small sample Bayesian designs for complex high-dimensional models based on information gained using fast approximations. Technometrics, 51 (4):377–388, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586648.

Castillo:2001:SAF

[CGHL01] Enrique Castillo, José Manuel Gutiérrez, Ali S. Hadi, and Beatriz Lacruz. Some applications of functional networks in statistics and engineering. Technometrics, 43(1):10–24, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270853.

Chaubey:2000:BRB

[Cha00] Yogendra P. Chaubey. Book review: Resampling Methods: A Practical Guide to Data Analysis by Phillip I. Good. Technometrics, 42(3):311, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271090.

Chakraborti:2001:BRB

[Cha01a] Subhabrata Chakraborti. Book review: Stat Labs: Mathematical Statistics through Applications by Deborah Nolan; Terry Speed. Technometrics, 43(3):378–379, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271238.

Chang:2001:BRB

[Cha01b] Ted Chang. Book review: Introduction to Geostatistics: Applications in Hydrogeology by P. K. Kitanidis. Technometrics, 43 (1):109–110, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270880.

Chatfield:2001:BRB

[Cha01c] Chris Chatfield. Book reviews: Fourier Analysis of Time Series by Peter Bloomfield; Random Data: Analysis and Measurement Procedures by Julian S. Bendat; Allan G. Piersol. Technometrics, 43(2):238, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271046.

Chan:2002:BRB

[Cha02a] Victor Chan. Book review: Calculated Bets by Steven Skiena. Technometrics, 44(4):404, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271561.

Chatfield:2002:BRB

[Cha02b] Chris Chatfield. Book review: Wavelet Transforms and Time-Frequency Signal Analysis by L. Debnath. Technometrics, 44(1):87, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270698.

Chao:2003:BRB

[Cha03] Edward C. Chao. Book review: Generalized Estimating Equations by James W. Hardin; Joseph M. Hilbe. Technometrics, 45 (4):363–364, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047096.

Chatterjee:2004:BRB

[Cha04] Snigdhansu Chatterjee. Book review: Order Statistics by H. A. David; H. N. Nagaraja. Technometrics, 46(3):364, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470859.

Chatterjee:2005:BRB

[Cha05] Snigdhansu Chatterjee. Book review: Procrustes Problems by John C. Gower; Garmt B. Dijksterhuis. Technometrics, 47(3):376, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471040.

Chambers:2008:FSC

[Cha08a] John M. Chambers. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):435–437, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471521.

Chatterjee:2008:BRB

[Cha08b] Snigdhansu Chatterjee. Book review: Structural Equation Modeling, a Bayesian Approach by Sik-Yum Lee. Technometrics, 50 (3):411–412, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471508.

Chatterjee:2009:BRB

[Cha09] Snigdhansu Chatterjee. Book review: Understanding Complex Datasets: Data Mining With Matrix Decompositions by David Skillicorn. Technometrics, 51(1):106, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586581.

Castillo:2004:GML

[CHCFC04] Enrique Castillo, Ali S. Hadi, Antonio Conejo, and Alfonso Fernández-Canteli. A general method for local sensitivity analysis with application to regression models and other optimization problems. Technometrics, 46(4):430–444, November 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470891.

Chen:2001:BRB

[Che01a] Dechang Chen. Book review: The Theory of the Design of Experiments by D. R. Cox; N. Reid. Technometrics, 43(4):497, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270839.

Chernick:2001:BRBa

[Che01b] Michael R. Chernick. Book review: Statistical Modelling Using GENSTAT(R) by K. J. McConway; M. C. Jones; P. C. Taylor. Technometrics, 43(2):241, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271053.

Chernick:2001:BRBb

[Che01c] Michael R. Chernick. Book review: Wavelet Methods for Time Series Analysis by Donald B. Percival; Andrew T. Walden. Technometrics, 43(4):491, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270828.

Chernick:2002:BRBb

[Che02a] Michael R. Chernick. Book review: Runs and Scans with Applications by N. Balakrishnan; Markos V. Koutras. Technometrics, 44(4):399, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271551.

Chernick:2002:BRBa

[Che02b] Michael R. Chernick. Book review: Variance Components Estimation: Mixed Models, Methodologies and Applications by Poduri S. R. S. Rao. Technometrics, 44(1):82–83, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270690.

Chen:2003:BRB

[Che03] Ming-Hui Chen. Book review: A Contingency Table Approach to Nonparametric Testing by J. C. W. Rayner; D. J. Best. Technometrics, 45(1):105–106, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25046988.

Chen:2006:AMT

[Che06] Yong Chen. Application of matroid theory for diagnosability study of coordinate sensing systems in discrete-part manufacturing processes. Technometrics, 48(3):386–398, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471210.

Chernick:2007:BRB

[Che07] Michael R. Chernick. Book review: Statistical Monitoring of Clinical Trials: A Unified Approach by Michael A. Proschan; K. K. Gordon Lan; Janet T. Wittes. Technometrics, 49(2):223–224, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471319.

Chernick:2008:BRBc

[Che08a] Michael Chernick. Book review: Global Sensitivity Analysis, the Primer by A. Saltelli; M. Ratto; T. Andres; F. Campolongo; J. Cariboni; D. Gatelli; M. Saisana; S. Tarantola. Technometrics, 50(4):548, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471545.

Chernick:2008:BRBb

[Che08b] Michael R. Chernick. Book review: An Introduction to Stein’s Method by A. D. Barbour; L. H. Y. Chen. Technometrics, 50 (2):236, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471470.

Chernick:2008:BRBa

[Che08c] Michael R. Chernick. Book review: Pharmaceutical Statistics Using SAS(R): A Practical Guide by Alex Dmitrienko; Christy Chuang-Stein; Ralph D’Agostino Sr. Technometrics, 50(1):96, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471444.

Chiang:2001:SGM

[Chi01] Andy K. L. Chiang. A simple general method for constructing confidence intervals for functions of variance components. Technometrics, 43(3):356–367, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1271223. See comments [IM02].

Chiang:2006:BRB

[Chi06] Alan Y. Chiang. Book review: Analysis of Clinical Trials Using SAS(R): A Practical Guide by Alex Dmitrienko; Geert Molenberghs; Christy Chuang-Stein; Walter Offen. Technometrics, 48 (2):311, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471180.

Chiang:2007:BRB

[Chi07] Alan Y. Chiang. Book review: Generalized Additive Models: An Introduction with R by Simon N. Wood. Technometrics, 49(3):360– 361, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471358.

Chow:2006:BRB

[Cho06] Mosuk Chow. Book review: The Essence of Multivariate Thinking by Lisa L. Harlow. Technometrics, 48(1):152, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471135.

Chakraborti:2006:NSS

[CHvdW06] S. Chakraborti, B. Hong, and M. A. van de Wiel. A note on sample size determination for a nonparametric test of location. Technometrics, 48(1):88–94, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471117.

Champ:2005:PCC

[CJFR05] Charles W. Champ, L. Allison Jones-Farmer, and Steven E. Rigdon. Properties of the T^{2} control chart when parameters are estimated. Technometrics, 47(4):437–445, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471068.

Conklin:2003:BRB

[CL03] W. Michael Conklin and Stan Lipovetsky. Book review: Chaos: A Statistical Perspective by Kung-Sik Chan; Howell Tong. Technometrics, 45(4):371, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047108.

Craiu:2005:MSC

[CL05] Radu V. Craiu and Thomas C. M. Lee. Model selection for the competing-risks model with and without masking. Technometrics, 47(4):457–467, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471070.

Cioppa:2007:ENO

[CL07] Thomas M. Cioppa and Thomas W. Lucas. Efficient nearly orthogonal and space-filling Latin hypercubes. Technometrics, 49 (1):45–55, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471274.

Chai:2009:DAT

[CL09] Feng-Shun Chai and Chen-Tuo Liao. Design and analysis of twolevel factorial experiments with partial replication. Technometrics, 51(1):66–74, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586564.

Clarkson:2004:BRB

[Cla04a] Douglas B. Clarkson. Book review: Measurement Error and Misclassification in Statistics and Epidemiology: Impact and Bayesian Adjustments by Paul Gustafson. Technometrics, 46(4):493–494, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470912.

Clausing:2004:OWE

[Cla04b] Don P. Clausing. Operating window: An engineering measure for robustness. Technometrics, 46(1):25–29, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470757.

Crowder:2009:ARW

[CLL09] M. J. Crowder, J. F. Lawless, and K.-A. Lee. Analysis of reliability and warranty claims in products with age and usage scales. Technometrics, 51(1):14–24, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586559.

Chan:2004:BBE

[CLM04] Victor Chan, Soumendra N. Lahiri, and William Q. Meeker. Block bootstrap estimation of the distribution of cumulative outdoor degradation. Technometrics, 46(2):215–224, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470806.

Cheng:2004:BNT

[CLY04] Shao-Wei Cheng, William Li, and Kenny Q. Ye. Blocked nonregular two-level factorial designs. Technometrics, 46(3):269–279, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470842.

Capizzi:2003:AEW

[CM03] Giovanna Capizzi and Guido Masarotto. An adaptive exponentially weighted moving average control chart. Technometrics, 45 (3):199–207, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047047.

Capizzi:2008:PDG

[CM08] Giovanna Capizzi and Guido Masarotto. Practical design of generalized likelihood ratio control charts for autocorrelated data. Technometrics, 50(3):357–370, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471496.

Cook:2001:NVR

[CO01] R. Dennis Cook and David J. Olive. A note on visualizing response transformations in regression. Technometrics, 43(4):443– 449, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270815.

Coburn:2000:BRB

[Cob00] Timothy C. Coburn. Book review: Geostatistics for Natural Resources Evaluation by Pierre Goovaerts. Technometrics, 42(4): 437–438, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270969.

Coburn:2002:BRB

[Cob02] Timothy C. Coburn. Book review: Practical Geostatistics: Modeling and Spatial Analysis by Simon W. Houlding. Technometrics, 44(1):83–84, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270692.

Coburn:2004:BRB

[Cob04] Timothy C. Coburn. Book review: Statistical and Econometric Methods for Transportation Data Analysis by Simon P. Washington; Matthew G. Karlaftis; Fred L. Mannering. Technometrics, 46(4):492–493, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470911.

Conklin:2000:BRB

[Con00] Joseph Conklin. Book review: Statistical Analysis Simplified by H. James Harrington; Glen D. Hoffherr; Robert P. Reid. Technometrics, 42(2):205, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271453.

Conklin:2001:BRB

[Con01] Michael Conklin. Book review: Monte Carlo Methods in Bayesian Computation by Ming-Hui Chen; Qi-Man Shao; Joseph G. Ibrahim. Technometrics, 43(2):240–241, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271052.

Conklin:2002:BRBa

[Con02a] Joseph D. Conklin. Book review: Applied Logistic Regression by David W. Hosmer; Stanley Lemeshow. Technometrics, 44(1):81– 82, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270688.

Conklin:2002:BRBb

[Con02b] Joseph D. Conklin. Book review: Classical Competing Risks by Martin Crowder. Technometrics, 44(3):298, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270511.

Conklin:2004:BRB

[Con04] Joseph D. Conklin. Book review: Sample Survey Theory: Some Pythagorean Perspectives by Paul Knottnerus. Technometrics, 46 (2):257–258, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470820.

Conklin:2005:BRB

[Con05] Michael Conklin. Book review: Multivariate Bayesian Statistics: Models for Source Separation and Signal Unmixing by Daniel B. Rowe. Technometrics, 47(1):95, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470940.

Cook:2000:DIO

[Coo00a] R. Dennis Cook. Detection of influential observation in linear regression. Technometrics, 42(1):65–68, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271434.

Cook:2000:RDD

[Coo00b] R. Dennis Cook. [Robust Diagnostic Data Analysis: Transformations in Regression]: Discussion. Technometrics, 42(4):395–398, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270949.

See [RA00].

Choi:2008:SCP

[COR08] Hyunyoung Choi, Hernando Ombao, and Bonnie Ray. Sequential change-point detection methods for nonstationary time series. Technometrics, 50(1):40–52, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471424.

Cheng:2006:ELR

[CR06] Chi-Lun Cheng and Jordi Riu. On estimating linear relationships when both variables are subject to heteroscedastic measurement errors. Technometrics, 48(4):511–519, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471243.

Choulakian:2001:GFT

[CS01] V. Choulakian and M. A. Stephens. Goodness-of-fit tests for the generalized Pareto distribution. Technometrics, 43(4):478–484, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270819.

Christensen:2002:ADF

[CS02] William F. Christensen and Stephan R. Sain. Accounting for dependence in a flexible multivariate receptor model. Technometrics, 44(4):328–337, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271540.

Cutler:2000:BRB

[Cut00] Adele Cutler. Book review: Neural Networks by Hervé Abdi; Dominique Valentin; Betty Edelman. Technometrics, 42(4):432, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270961.

Cheng:2002:COB

[CW02] Shao-Wei Cheng and C. F. J. Wu. Choice of optimal blocking schemes in two-level and three-level designs. Technometrics, 44 (3):269–277, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270491.

Cook:2008:FSC

[CW08] Dianne Cook and Hadley Wickham. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):442–443, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471525.

Dankovic:2001:BRB

[Dan01] Daniel D. Dankovic. Book review: Root Cause Analysis by Bjorn Andersen; Tom Fagerhaug. Technometrics, 43(3):370–371, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271226.

Dasgupta:2000:BRB

[Das00] Nairanjana Dasgupta. Book review: Statistics with STATA 5 by Lawrence C. Hamilton. Technometrics, 42(4):441, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270973.

Dasgupta:2004:BRB

[Das04] Nairanjana Dasgupta. Book review: Analyzing Categorical Data by Jeffrey S. Simonoff. Technometrics, 46(2):255–256, May 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470816.

Dasgupta:2007:BRBb

[Das07a] Nairanjana Dasgupta. Book review: Bayesian Models for Categorical Data by Peter Congdon. Technometrics, 49(2):230–231, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471332.

Dasgupta:2007:BRBa

[Das07b] Nairanjana Dasgupta. Book review: DNA, Words and Models, Statistics of Exceptional Words by S. Robin; F. Rodolphe; S. Schbath. Technometrics, 49(1):106–107, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471291.

Davis:2003:BRB

[Dav03] J. Wade Davis. Book review: Statistical Pattern Recognition by Andrew Webb. Technometrics, 45(4):368, November 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047103.

delCastillo:2002:CLD

[dC02] Enrique del Castillo. Closed-loop disturbance identification and controller tuning for discrete manufacturing processes. Technometrics, 44(2):134–141, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271257.

deCastro:2005:FSB

[dCGM05] B. Fernández de Castro, S. Guillas, and W. González Manteiga. Functional samples and bootstrap for predicting sulfur dioxide levels. Technometrics, 47(2):212–222, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470983.

deCastro:2008:LE

[dCGRB08] Mário de Castro, Manuel Galea-Rojas, and Heleno Bolfarine. Letter to the Editor. Technometrics, 50(1):86, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471429.

Deng:2000:BRB

[Den00] Lih-Yuan Deng. Book review: Orthogonal Arrays: Theory and Applications by A. S. Hedayat; N. J. A. Sloane; John Stufken. Technometrics, 42(4):440, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270971.

Deng:2004:BRB

[Den04] Lih-Yuan Deng. Book review: The Design and Analysis of Computer Experiments by Thomas J. Santner; Brian J. Williams; William I. Notz. Technometrics, 46(4):488–489, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470903.

Deng:2006:BRB

[Den06] Lih-Yuan Deng. Book review: The Cross-Entropy Method: A Unified Approach to Combinatorial Optimization, Monte-Carlo Simulation, and Machine Learning by Reuven Y. Rubinstein; Dirk P. Kroese. Technometrics, 48(1):147–148, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471127.

Denuit:2007:KST

[DGS07] Michel Denuit, Anne-Cécile Goderniaux, and Olivier Scaillet. A Kolmogorov-Smirnov-type test for shortfall dominance against parametric alternatives. Technometrics, 49(1):88–98, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471278.

Ding:2005:BRB

[Din05] Yu Ding. Book review: Design and Analysis of Accelerated Tests for Mission-Critical Reliability by Michael J. LuValle; Bruce G. Lefevre; SriRaman Kannan. Technometrics, 47(2):240–241, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470999.

Dass:2007:FBR

[DJ07] Sarat C. Dass and Anil K. Jain. Fingerprint-based recognition. Technometrics, 49(3):262–276, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471347.

Denby:2007:SAA

[DLM^{+}07] Lorraine Denby, James M. Landwehr, Colin L. Mallows, Jean Meloche, John Tuck, Bowei Xi, George Michailidis, and Vijayan N. Nair. Statistical aspects of the analysis of data networks. Technometrics, 49(3):318–334, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471351.

Dalal:2008:OSE

[DM08a] S. R. Dalal and C. L. Mallows. Optimal stopping with exact confidence on remaining defects. Technometrics, 50(3):397–406, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471499.

Dasgupta:2008:EPP

[DM08b] Tirthankar Dasgupta and Abhyuday Mandal. Estimation of process parameters to determine the optimum diagnosis interval for control of defective items. Technometrics, 50(2):167–181, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471457.

Dobrow:2002:BRB

[Dob02] Robert P. Dobrow. Book review: Applied Stochastic Modelling by Byron J. T. Morgan. Technometrics, 44(1):91, February 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270705.

Doganaksoy:2001:BRB

[Dog01] Necip Doganaksoy. Book review: HALT, HASS and HASA Explained: Accelerated Reliability Techniques by Harry McLean. Technometrics, 43(4):489–490, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270825.

Doganaksoy:2004:BRB

[Dog04] Necip Doganaksoy. Book review: Weibull Models by D. N. Prabhakar Murthy. Technometrics, 46(4):485–486, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470899.

Doksum:2008:BRB

[Dok08] Kjell A. Doksum. Book review: Principles of Statistical Inference by D. R. Cox. Technometrics, 50(2):235–236, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471468.

Doksum:2009:BRB

[Dok09] Kjell Doksum. Book review: Reminiscences of a Statistician: The Company I Kept (Kindle ed.) by Erich Lehmann. Technometrics, 51(4):480–481, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586658.

Draper:2002:GRA

[DP02] Norman R. Draper and Friedrich Pukelsheim. Generalized ridge analysis under linear restrictions, with particular applications to mixture experiments problems. Technometrics, 44(3):250–259, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270489.

Drain:2004:BRB

[Dra04] David Drain. Book review: Data Matters: Conceptual Statistics for a Random World by Nicholas Maxwell. Technometrics, 46(2):253–254, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470814.

Drignei:2006:EBA

[Dri06] Dorin Drignei. Empirical Bayesian analysis for high-dimensional computer output. Technometrics, 48(2):230–240, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471159.

Doganaksoy:2000:PAC

[DS00] Necip Doganaksoy and Josef Schmee. Practical aspects of corrected likelihood ratio confidence intervals: A discussion of Jeng–Meeker and Wong–Wu. Technometrics, 42(2):156–159, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271447.

Dror:2006:RED

[DS06] Hovav A. Dror and David M. Steinberg. Robust experimental design for multivariate generalized linear models. Technometrics, 48 (4):520–529, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471244.

Deng:2002:DSC

[DT02] Lih-Yuan Deng and Boxin Tang. Design selection and classification for Hadamard matrices using generalized minimum aberration criteria. Technometrics, 44(2):173–184, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271261.

Dupuis:2005:OCR

[Dup05] Debbie J. Dupuis. Ozone concentrations: A robust analysis of multivariate extremes. Technometrics, 47(2):191–201, May 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470981.

Dasgupta:2006:RPD

[DW06] Tirthankar Dasgupta and C. F. Jeff Wu. Robust parameter design with feedback control. Technometrics, 48(3):349–360, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471207.

DaVeiga:2009:LPE

[DWG09] Sebastien Da Veiga, François Wahl, and Fabrice Gamboa. Local polynomial estimation for sensitivity analysis on models with correlated inputs. Technometrics, 51(4):452–463, November 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586654.

Easterling:2007:BRB

[Eas07] Robert G. Easterling. Book review: Modern Experimental Design by Thomas P. Ryan. Technometrics, 49(4):497–499, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471407.

Edgeman:2004:BRB

[Edg04] Rick L. Edgeman. Book review: Multivariate Statistical Process Control with Industrial Applications by Robert L. Mason; John C. Young. Technometrics, 46(4):484, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470897.

Efron:2006:DTP

[Efr06] Bradley Efron. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):330–332, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471203. See [Mal06].

Epifanio:2008:SDC

[Epi08] Irene Epifanio. Shape descriptors for classification of functional data. Technometrics, 50(3):284–294, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471490.

Esmaili:2005:BRB

[Esm05] Ali Esmaili. Book review: Probability and Random Processes by Scott L. Miller; Donald G. Childers. Technometrics, 47(3):375, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471038.

Esmaili:2006:BRB

[Esm06] Ali Esmaili. Book review: Probability Models in Engineering and Science by Haym Benaroya; Sean Mi Han. Technometrics, 48 (4):575, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471261.

Esmaili:2009:BRBa

[Esm09a] Ali Esmaili. Book review: Applied Multiway Data Analysis by Pieter M. Kroonenberg. Technometrics, 51(2):219, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586604.

Esmaili:2009:BRBb

[Esm09b] Ali Esmaili. Book review: Applied Multiway Data Analysis by Pieter M. Kroonenberg. Technometrics, 51(4):481, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586660.

Espejo:2006:BRBa

[Esp06a] Mariano Ruiz Espejo. Book review: Design of Experiments for Engineers and Scientists by Jiju Antony. Technometrics, 48(2): 304–305, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471168.

Espejo:2006:BRBb

[Esp06b] Mariano Ruiz Espejo. Book review: The Analysis of Means: A Graphical Method for Comparing Means, Rates, and Proportions by Peter R. Nelson; Peter S. Wludyka; Karen A. F. Copeland. Technometrics, 48(2):307, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471173.

Espejo:2007:BRB

[Esp07] Mariano Ruiz Espejo. Book review: Quality Makes Money by Pat Townsend; Joan Gebhardt. Technometrics, 49(2):222–223, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471317.

Evans:2001:BRB

[Eva01] Brad Evans. Book review: Misused Statistics by Herbert F. Spirer; Louise Spirer; A. J. Jaffe. Technometrics, 43(1):106, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270874.

Fahmy:2007:BRB

[Fah07] Hesham M. Fahmy. Book review: Introduction to Statistics for Forensic Scientists by David Lucy. Technometrics, 49(3):362, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471361.

Famoye:2003:BRB

[Fam03] Felix Famoye. Book review: Plane Answers to Complex Questions: Theory of Linear Models by Ronald Christensen. Technometrics, 45(2):174–175, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047022.

Forbes:2006:CDD

[FFG06] Michael G. Forbes, J. Fraser Forbes, and Martin Guay. Controller design for discrete-time stochastic processes with nonquadratic loss. Technometrics, 48(3):361–372, August 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471208.

Fernholz:2001:CCT

[FG01] Luisa T. Fernholz and John A. Gillespie. Content-corrected tolerance limits based on the bootstrap. Technometrics, 43(2):147–155, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271028. Sta-

tistical methodology for industrial problems.

Field:2006:MD

[FG06] Christopher Field and Marc G. Genton. The multivariate g-and-h distribution. Technometrics, 48(1):104–111, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471119.

Fougeres:2006:PCC

[FHR06] A.-L. Fougères, S. Holm, and H. Rootzén. Pitting corrosion: Comparison of treatments with extreme-value: Distributed responses. Technometrics, 48(2):262–272, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471162.

Finkelstein:2002:BRB

[Fin02] Dianne M. Finkelstein. Book review: Analysis of Failure and Survival Data by Peter J. Smith. Technometrics, 44(4):397, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271548.

Finkelstein:2005:BRB

[Fin05] Dianne M. Finkelstein. Book review: A Beginner’s Guide to Structural Equation Modeling by Randall E. Schumacker; Richard G. Lomax. Technometrics, 47(4):522, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471087.

Fang:2008:RPM

[FJ08] Yi Fang and Myong K. Jeong. Robust probabilistic multivariate calibration model. Technometrics, 50(3):305–316, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471492.

Feiveson:2000:RSS

[FK00] Alan H. Feiveson and Pandurang M. Kulkarni. Reliability of space-Shuttle pressure vessels with random batch effects. Technometrics, 42(4):332–344, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270943.

Fokianos:2001:SAO

[FKQS01] Konstantinos Fokianos, Benjamin Kedem, Jing Qin, and David A. Short. A semiparametric approach to the one-way layout. Technometrics, 43(1):56–65, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270857.

Fredette:2007:FHP

[FL07] Marc Fredette and J. F. Lawless. Finite-horizon prediction of recurrent events, with application to forecasts of warranty claims. Technometrics, 49(1):66–80, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471276.

Flegal:2009:BRB

[Fle09] James M. Flegal. Book review: Statistical Thinking in Sports by Jim Albert; Ruud H. Koning. Technometrics, 51(1):106, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586580.

Fang:2000:UDT

[FLWZ00] Kai-Tai Fang, Dennis K. J. Lin, Peter Winker, and Yong Zhang. Uniform design: Theory and application. Technometrics, 42(3): 237–248, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271079.

Fang:2004:OSA

[FM04] Kai-Tai Fang and Rahul Mukerjee. Optimal selection of augmented pairs designs for response surface modeling. Technometrics, 46(2):147–152, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470800.

Fokianos:2007:BRB

[Fok07] Konstantinos Fokianos. Book review: Semiparametric Theory and Missing Data by Anastasios A. Tsiatis. Technometrics, 49(2):228– 229, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471328.

Forkman:2008:MDN

[For08] Johannes Forkman. A method for designing nonlinear univariate calibration. Technometrics, 50(4):479–486, November 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471531.

Fotopoulos:2001:BRB

[Fot01] Stergios B. Fotopoulos. Book review: Discrete-Time Dynamic Models by Ronald K. Pearson. Technometrics, 43(1):110–111, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270882.

Fotopoulos:2004:BRB

[Fot04] Stergios B. Fotopoulos. Book review: Introduction to Modern Nonparametric Statistics by James J. Higgins. Technometrics, 46 (4):488, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470902.

Fotopoulos:2007:BRBa

[Fot07a] Stergios B. Fotopoulos. Book review: All of Nonparametric Statistics by Larry Wasserman. Technometrics, 49(1):103, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471284.

Fotopoulos:2007:BRBb

[Fot07b] Stergios B. Fotopoulos. Book review: Modeling Financial Time Series with S-PLUS by Eric Zivot; Jiahui Wang. Technometrics, 49(3):360, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471357.

Fotopoulos:2007:BRBc

[Fot07c] Stergios B. Fotopoulos. Book review: Probability and Random Processes by Venkatarama Krishnan. Technometrics, 49(3):365, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471369.

Fotopoulos:2009:BRB

[Fot09] Stergios B. Fotopoulos. Book review: A Course in Applied Stochastic Processes by A. Goswami; B. V. Rao. Technometrics, 51(1): 100, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586570.

Faraway:2007:SDH

[FR07] Julian Faraway and Matthew P. Reed. Statistics for digital human motion modeling in ergonomics. Technometrics, 49(3):277– 290, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471348.

Frey:2000:BRB

[Fre00] Michael Frey. Book review: Stochastic Dynamic Programming and Control of Queueing Systems by Linn I. Sennott. Technometrics, 42(3):321–322, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271107.

Frey:2003:BRB

[Fre03] Michael Frey. Book review: Smoothing Spline ANOVA Models by Chong Gu. Technometrics, 45(3):269, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047059.

Fricker:2000:BRB

[Fri00] Ronald D. Fricker, Jr. Book review: Statistical Process Control and Quality Improvement by Gerald M. Smith. Technometrics, 42(2):209–210, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271460.

Fricker:2001:BRB

[Fri01] Ronald D. Fricker, Jr. Book review: A Handbook of Statistical Analyses Using Stata by Sophia Rabe-Hesketh; Brian Everitt. Technometrics, 43(4):496–497, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270838.

Fricker:2005:BRB

[Fri05] Ronald D. Fricker, Jr. Book review: Cognition and Chance: The Psychology of Probabilistic Reasoning by Raymond S. Nickerson. Technometrics, 47(3):376, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471039.

Fokianos:2008:CSS

[FS08] Konstantinos Fokianos and Alexios Savvides. On comparing several spectral densities. Technometrics, 50(3):317–331, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471493.

Fligner:2002:MJT

[FVB02] Michael A. Fligner, Joseph S. Verducci, and Paul E. Blower. A modification of the Jaccard-Tanimoto similarity index for diverse selection of chemical compounds using binary strings. Technometrics, 44(2):110–119, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271255.

Furnival:2000:RLB

[FW00] George M. Furnival and Robert W. Wilson, Jr. Regressions by leaps and bounds. Technometrics, 42(1):69–79, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271435.

Frey:2006:AOF

[FW06] Daniel D. Frey and Hungjen Wang. Adaptive one-factor-at-a-time experimentation and expected value of improvement. Technometrics, 48(3):418–431, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471213.

Ganocy:2002:BRB

[Gan02] Stephen J. Ganocy. Book review: Bayesian Statistical Modelling by Peter Congdon. Technometrics, 44(3):291–292, August 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270501.

Ganocy:2005:BRB

[Gan05] Stephen J. Ganocy. Book review: Introduction to Bayesian Statistics by William M. Bolstad. Technometrics, 47(2):236, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470992.

Ganocy:2006:BRB

[Gan06] Stephen J. Ganocy. Book review: Numerical Methods for Nonlinear Estimating Equations by Christopher G. Small; Jinfang Wang. Technometrics, 48(4):569, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471249.

Gardner:2000:BRB

[Gar00] Martha M. Gardner. Book review: Implementing Six Sigma: Smarter Solutions Using Statistical Methods by Forrest W. Breyfogle. Technometrics, 42(3):309–310, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271088.

Gatliffe:2002:BRB

[Gat02] Thomas R. Gatliffe. Book review: Calibration and Reliability in Groundwater Modelling: Coping with Uncertainty by F. Stauffer; W. Kinzelbach; K. Kovar; E. Hoehn. Technometrics, 44(1):84–85, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270693.

Gatliffe:2006:BRB

[Gat06] Thomas R. Gatliffe. Book review: Nondetects and Data Analysis by Dennis R. Helsel. Technometrics, 48(1):145, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471123.

Gibbons:2001:WRE

[GB01] Robert D. Gibbons and Dulal K. Bhaumik. Weighted randomeffects regression models with application to interlaboratory calibration. Technometrics, 43(2):192–198, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271032. Statistical methodol-

ogy for industrial problems.

Grimshaw:2001:EFI

[GCLH01] Scott D. Grimshaw, Bruce J. Collings, Wayne A. Larsen, and Carolyn R. Hurt. Eliciting factor importance in a designed experiment. Technometrics, 43(2):133–146, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271027. Statistical methodol-

ogy for industrial problems.

Ghosh:2006:CUK

[GCS06] Anil K. Ghosh, Probal Chaudhuri, and Debasis Sengupta. Classification using kernel density estimates: Multiscale analysis and visualization. Technometrics, 48(1):120–132, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471121.

Gilmour:2003:CRA

[GD03] Steven. G. Gilmour and Norman R. Draper. Confidence regions around the ridge of optimal response on fitted second-order response surfaces. Technometrics, 45(4):333–339, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047089.

Gilmour:2004:LER

[GD04] Steven G. Gilmour and Norman R. Draper. [Letter to the Editor]: Response. Technometrics, 46(3):358, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470851.

Gelman:2006:MHM

[Gel06] Andrew Gelman. Multilevel (hierarchical) modeling: What it can and cannot do. Technometrics, 48(3):432–435, August 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471214.

George:2000:BRB

[Geo00] Larry George. Book review: Mathematical Statistics, a Unified Introduction by George R. Terrell. Technometrics, 42(3):319–321, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271105.

George:2003:BRB

[Geo03] Laurence L. George. Book review: The Statistical Analysis of Failure Time Data by John D. Kalbfleisch; Ross L. Prentice. Technometrics, 45(3):265–266, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047055.

Gijbels:2004:BSC

[GG04] Irène Gijbels and Anne-Cécile Goderniaux. Bandwidth selection for changepoint estimation in nonparametric regression. Technometrics, 46(1):76–86, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470763.

Graves:2007:EPU

[GHB^{+}07] Todd Graves, Michael Hamada, Jane Booker, Michele Decroix, Kathy Chilcoat, and Clint Bowyer. Estimating a proportion using stratified data from both convenience and random samples. Technometrics, 49(2):164–171, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471309.

Ghosh:2001:BRB

[Gho01] Sujit K. Ghosh. Book review: Essential Wavelets for Statistical Applications and Data Analysis by R. Todd Ogden. Technometrics, 43(1):105, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270872.

Ghosh:2002:BRB

[Gho02] Subir Ghosh. Book review: Elements of Sampling Theory and Methods by Z. Govindarajulu. Technometrics, 44(2):191, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271270.

Ghosh:2003:BRB

[Gho03] Subir Ghosh. Book review: Nonparametric Analysis of Longitudinal Data in Factorial Experiments by Edgar Brunner; Sebastian Domhof; Frank Langer. Technometrics, 45(2):171–172, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047017.

Ghosh:2005:BRB

[Gho05] Subir Ghosh. Book review: Generalized Inference in Repeated Measures: Exact Methods in MANOVA and Mixed Models by Samaradasa Weerahandi. Technometrics, 47(2):233, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470987.

Ghosh:2007:BRB

[Gho07] Subir Ghosh. Book review: Celebrating Statistics: Papers in Honour of Sir David Cox on His 80^{th} Birthday by Anthony C. Davison; Yadolah Dodge; Nanny Wermuth. Technometrics, 49 (3):362–363, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471363.

Ghosh:2008:BRBb

[Gho08a] Subir Ghosh. Book review: Generalized Linear Models with Random Effects by Youngjo Lee; John A. Nelder; Yudi Pawitan. Technometrics, 50(4):542–543, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471536.

Ghosh:2008:BRBa

[Gho08b] Subir Ghosh. Book review: Reliability and Risk: A Bayesian Perspective by Nozer D. Singpurwalla. Technometrics, 50(1):88, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471431.

Ghosh:2009:BRB

[Gho09] Subir Ghosh. Book review: Bayesian Process Monitoring, Control and Optimization by Bianca M. Colosimo; Enrique Del Castillo. Technometrics, 51(2):216–217, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586599.

Gibbons:2001:BRB

[Gib01] Ned M. Gibbons. Book review: Practical Strategies for Experimenting by G. K. Robinson. Technometrics, 43(4):486, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270820.

Gotwalt:2009:FCD

[GJS09] Christopher M. Gotwalt, Bradley A. Jones, and David M. Steinberg. Fast computation of designs robust to parameter uncertainty for nonlinear settings. Technometrics, 51(1):88–95, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586566.

Goos:2009:LE

[GL09a] Peter Goos and James M. Lucas. Letter to the Editor. Technometrics, 51(1):96–97, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586567.

Gramacy:2009:ADA

[GL09b] Robert B. Gramacy and Herbert K. H. Lee. Adaptive design and analysis of supercomputer experiments. Technometrics, 51 (2):130–145, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586591.

Genkin:2007:LSB

[GLM07] Alexander Genkin, David D. Lewis, and David Madigan. Largescale Bayesian logistic regression for text categorization. Technometrics, 49(3):291–304, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471349.

Glosup:2001:BRB

[Glo01] Jeffrey Glosup. Book review: Statistical Modelling with Quantile Functions by Warren Gilchrist. Technometrics, 43(4):488–489, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270824.

Glosup:2005:BRB

[Glo05] Jeffrey Glosup. Book review: Generalized Linear Models: An Applied Approach by John P. Hoffmann. Technometrics, 47(2):232, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470985.

Glosup:2006:BRB

[Glo06] Jeffrey G. Glosup. Book review: Statistical Methods in Computer Security by William W. S. Chen. Technometrics, 48(4): 570–571, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471252.

Gluhovsky:2006:SIA

[Glu06] Ilya Gluhovsky. Smooth isotonic additive interaction models with application to computer system architecture design. Technometrics, 48(2):176–192, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471155.

Golowich:2002:IBP

[GLV02] Steven Golowich, James Landwehr, and Scott Vander Wiel. Interplay between physics and statistics for modeling optical fiber bandwidth. Technometrics, 44(3):215–229, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270486.

Graves:2001:IPD

[GM01] Todd L. Graves and Audris Mockus. Identifying productivity drivers by modeling work units using partial data. Technometrics, 43(2):168–179, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271030.

Goldfarb:2004:BRB

[Gol04] Heidi B. Goldfarb. Book review: Statistical Methods for Six Sigma in R&D and Manufacturing by Anand M. Joglekar. Technometrics, 46(3):362, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470856.

Goldfarb:2006:BRB

[Gol06a] Heidi B. Goldfarb. Book review: Experimental Design for Formulation by Wendell F. Smith. Technometrics, 48(2):304, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471167.

Goldstein:2006:BRB

[Gol06b] Richard Goldstein. Book review: Regression Methods in Biostatistics: Linear, Logistic, Survival and Repeated Measures Models by Eric Vittinghoff; David V. Glidden; Stephen C. Shiboski; Charles E. McCulloch. Technometrics, 48(1):149–150, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471131.

Goldstein:2007:BRB

[Gol07] Richard Goldstein. Book review: Statistical Evidence in Medical Trials: What Do the Data Really Tell Us? by Stephen D. Simon. Technometrics, 49(1):107–108, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471293.

Goldstein:2008:QAS

[Gol08] Darlene R. Goldstein. [Quality Assessment for Short Oligonucleotide Microarray Data]: Comment. Technometrics, 50(3):276– 279, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471488.

Goodman:2002:BRB

[Goo02] Jay H. Goodman. Book review: Applying and Interpreting Statistics: A Comprehensive Guide by Glen McPherson. Technometrics, 44(3):296, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270508.

Girard:2001:BAA

[GP01] Philippe Girard and Eric Parent. Bayesian analysis of autocorrelated ordered categorical data for industrial quality monitoring. Technometrics, 43(2):180–191, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1271031. Statistical methodology for industrial problems.

Gelman:2006:BME

[GP06] Andrew Gelman and Iain Pardoe. Bayesian measures of explained variance and pooling in multilevel (hierarchical) models. Technometrics, 48(2):241–251, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471160.

Guttman:2005:BAP

[GPR05] Irwin Guttman, Daniel Peña, and Dolores Redondas. A Bayesian approach for predicting with polynomial regression of unknown degree. Technometrics, 47(1):23–33, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470931.

Gray:2000:BRB

[Gra00] J. Brian Gray. Book review: Applied Regression Including Computing and Graphics by R. Dennis Cook; Sanford Weisberg. Technometrics, 42(4):426–427, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270953.

Gray:2002:BRB

[Gra02] J. Brian Gray. Book review: Introduction to Linear Regression Analysis by Douglas C. Montgomery; Elizabeth A. Peck; G. Geoffrey Vining. Technometrics, 44(2):191–192, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271271.

Grego:2004:BRB

[Gre04] John Grego. Book review: Bayesian Data Analysis by Andrew Gelman; John B. Carlin; Hal S. Stern; Donald B. Rubin. Technometrics, 46(3):363–364, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470858.

Grice:2000:BRB

[Gri00] John V. Grice. Book review: Design and Analysis of Experiments by Douglas C. Montgomery. Technometrics, 42(2):208–209, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271458.

Griffith:2004:BRB

[Gri04] William S. Griffith. Book review: Optimal Reliability Modeling: Principles and Applications by Way Kuo; Ming J. Zuo. Technometrics, 46(1):112, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470771.

Griffith:2007:BRB

[Gri07] William S. Griffith. Book review: Decisions under Uncertainty: Probabilistic Analysis for Engineering Decisions by Ian Jordaan. Technometrics, 49(1):102–103, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471283.

Groggel:2000:BRB

[Gro00] David J. Groggel. Book review: Practical Nonparametric Statistics by W. J. Conover. Technometrics, 42(3):317–318, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271101.

Gruber:2002:BRB

[Gru02] Marvin H. J. Gruber. Book review: The Subjectivity of Scientists and the Bayesian Approach by S. James Press; Judith M. Tanur. Technometrics, 44(3):292–293, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270502.

Gruber:2005:BRB

[Gru05] Marvin H. J. Gruber. Book review: Optimization by Kenneth Lange. Technometrics, 47(3):379, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471046.

Gruber:2008:BRB

[Gru08] Marvin H. J. Gruber. Book review: Bayes Linear Statistics Theory and Methods by Michael Goldstein; David Woolf. Technometrics, 50(4):547–548, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471544.

Gilbert:2007:ADS

[GS07] A. C. Gilbert and M. J. Strauss. Analysis of data streams: Computational and algorithmic challenges. Technometrics, 49(3):346– 356, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471353.

Grigg:2008:EAN

[GS08] O. A. Grigg and D. J. Spiegelhalter. An empirical approximation to the null unbounded steady-state distribution of the cumulative sum statistic. Technometrics, 50(4):501–511, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471533.

Guirguis:2008:NCP

[Gui08] Georges H. Guirguis. A note on computing the probability and critical values for the half-normal plot. Technometrics, 50(2):228–229, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471462.

Gunst:2000:CSR

[Gun00] Richard F. Gunst. Classical studies that revolutionized the practice of regression analysis. Technometrics, 42(1):62–64, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271433.

Gunst:2003:BRB

[Gun03] Richard F. Gunst. Book review: Regression and ANOVA: An Integrated Approach Using SAS Software by Keith E. Muller; Bethel A. Fetterman. Technometrics, 45(2):170–171, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047015.

Gunter:2004:BRB

[Gun04] Berton H. Gunter. Book review: Data Analysis and Graphics Using R: An Example-Based Approach by John Maindonald; John Braun. Technometrics, 46(2):258–260, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470821.

Gunter:2006:BRB

[Gun06] Berton H. Gunter. Book review: R Graphics by Paul Murrell. Technometrics, 48(2):308–309, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471176.

Gupta:2000:BRB

[Gup00] Sudhir Gupta. Book review: Statistical Regression with Measurement Error by Chi-Lun Cheng; John W. van Ness. Technometrics, 42(4):427, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270954.

Goos:2003:OSP

[GV03] Peter Goos and Martina Vandebroek. D-optimal split-plot designs with given numbers and sizes of whole plots. Technometrics, 45(3):235–245, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047050.

Gluhovsky:2007:CME

[GV07] Ilya Gluhovsky and David Vengerov. Constrained multivariate extrapolation models with application to computer cache rates. Technometrics, 49(2):129–137, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471306.

Hanrahan:2008:FSC

[Han08] Pat Hanrahan. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):440–441, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471524.

Hapeman:2000:BRB

[Hap00] David E. Hapeman. Book review: Statistical Analysis of Categorical Data by Chris J. Lloyd. Technometrics, 42(3):313, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271094.

He:2008:QAS

[HB08] Wenqing He and Shelley B. Bull. [Quality Assessment for Short Oligonucleotide Microarray Data]: Comment. Technometrics, 50 (3):271–274, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471486.

Hartless:2003:LIP

[HBL03] Glen Hartless, James G. Booth, and Ramon C. Littell. Local influence of predictors in multiple linear regression. Technometrics, 45(4):326–332, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047088.

Huang:2000:DSW

[HC00] H.-C. Huang and N. Cressie. Deterministic/stochastic wavelet decomposition for recovery of signal from noisy data. Technometrics, 42(3):262–276, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271081.

Hodges:2007:SBS

[HCSC07] James S. Hodges, Yue Cui, Daniel J. Sargent, and Bradley P. Carlin. Smoothing balanced single-error-term analysis of variance. Technometrics, 49(1):12–25, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471271.

Heavlin:2003:DEC

[Hea03] William D. Heavlin. Designing experiments for causal networks. Technometrics, 45(2):115–129, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047009.

Heckler:2001:BRB

[Hec01] Charles E. Heckler. Book review: Graphical Analysis of Multiresponse Data, Illustrated with a Plant Breeding Trial, Interdisciplinary Statistics, by K. E. Basford and J. W. Tukey. Technometrics, 43(1):97–98, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270862.

Heckler:2005:BRB

[Hec05] Charles E. Heckler. Book review: Applied Multivariate Statistical Analysis by Wolfgang Härdle; Léopold Simar. Technometrics, 47(4):517, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471079.

Heckler:2007:BRB

[Hec07] Charles E. Heckler. Book review: Correspondence Analysis and Data Coding with Java and R by Fionn Murtagh. Technometrics, 49(1):100, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471279.

Heckler:2008:BRB

[Hec08] Charles E. Heckler. Book review: Introduction to Mixed Modelling. Beyond Regression and Analysis of Variance by Nicholas W. Galwey. Technometrics, 50(3):410–411, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471506.

Heckler:2009:BRB

[Hec09] Charles Heckler. Book review: Linear Mixed Models: A Practical Guide Using Statistical Software by Brady T. West; Kathleen B. Welch; Andrzej T. Galecki; Brenda W. Gillespie. Technometrics, 51(2):216, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586598.

Hembree:2002:BRB

[Hem02] G. Barry Hembree. Book review: The Handbook of Applied Acceptance Sampling: Plans, Procedures, and Principles by Kenneth S. Stephens. Technometrics, 44(1):90–91, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270704.

Hendrickson:2001:BRB

[Hen01] John A. Hendrickson, Jr. Book review: Probability, Random Signals, and Statistics by X. Rong Li. Technometrics, 43(1):103–104, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270870.

Hesterberg:2002:BRB

[Hes02] Tim Hesterberg. Book review: Monte Carlo Strategies in Scientific Computing by Jun S. Liu. Technometrics, 44(4):403–404, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271559.

Hesterberg:2004:BRB

[Hes04] Tim Hesterberg. Book review: Introduction to Stochastic Search and Optimization: Estimation, Simulation, and Control by James C. Spall. Technometrics, 46(3):367–368, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470865.

Heydorn:2001:BRB

[Hey01] Richard P. Heydorn. Book review: Reliability Engineering Handbook by Brian Dodson; Dennis Nolan. Technometrics, 43(2):243– 244, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271058.

Heydorn:2002:BRB

[Hey02] Richard P. Heydorn. Book review: Design for Reliability by Dana Crowe; Alec Feinberg. Technometrics, 44(4):401, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271555.

Heydorn:2006:BRB

[Hey06] Richard P. Heydorn. Book review: Reliability Engineering: Probabilistic Models and Maintenance Methods by Joel A. Nachlas. Technometrics, 48(3):446, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471218.

Heydorn:2009:BRB

[Hey09] Richard P. Heydorn. Book review: Applied Regression Analysis and Generalized Linear Models (2nd ed.) by John Fox. Technometrics, 51(3):342–343, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586633.

Huzurbazar:2002:BRB

[HG02] Snehalata V. Huzurbazar and Mark C. Greenwood. Book review: Workshop Statistics: Discovery with Data, a Bayesian Approach by James H. Albert; Allan J. Rossman. Technometrics, 44(3):293, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270503.

Hagwood:2006:CDS

[HG06] Charles Hagwood and William Guthrie. Combining data in small multiple-method studies. Technometrics, 48(2):293–301, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471165.

Huzurbazar:2000:BRB

[HI00] S. Huzurbazar and Lynne Ipiña. Book review: Workshop Statistics: Discovery with Data and Minitab by Allan J. Rossman; Beth L. Chance. Technometrics, 42(4):440–441, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270972.

Hill:2000:BRB

[Hil00] Jeanne S. Hill. Book review: Reliability Modelling by Linda C. Wolstenholme. Technometrics, 42(4):434–435, November 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270964.

Hinich:2005:BRB

[Hin05] Melvin J. Hinich. Book review: Time Series Analysis by State Space Methods by J. Durbin; S. J. Koopman. Technometrics, 47(3):373, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471032.

Hinich:2006:BRB

[Hin06] Melvin J. Hinich. Book review: Developing Statistical Software in Fortran 95 by David R. Lemmon; Joseph L. Schafer. Technometrics, 48(4):571–572, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471254.

Hung:2009:DAC

[HJM09] Ying Hung, V. Roshan Joseph, and Shreyes N. Melkote. Design and analysis of computer experiments with branching and nested factors. Technometrics, 51(4):354–365, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586646.

Hamada:2004:BPI

[HJMW04] Michael Hamada, Valen Johnson, Leslie M. Moore, and Joanne Wendelberger. Bayesian prediction intervals and their relationship to tolerance intervals. Technometrics, 46(4):452–459, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470893.

Hoerl:2000:RRB

[HK00] Arthur E. Hoerl and Robert W. Kennard. Ridge regression: Biased estimation for nonorthogonal problems. Technometrics, 42 (1):80–86, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271436.

Heredia-Langner:2006:BRB

[HL06] Alejandro Heredia-Langner. Book review: Quality by Experimental Design by Thomas B. Barker. Technometrics, 48(4):574, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471259.

Hlynka:2007:BRB

[Hly07] Myron Hlynka. Book review: Queueing Networks and Markov Chains (Modeling and Performance Evaluation with Computer Science Applications) by Gunter Bolch; Stefan Greiner; Hermann De Meer; Kishor Trivedi. Technometrics, 49(1):104–105, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471287.

Hlynka:2009:BRBa

[Hly09a] Myron Hlynka. Book review: A First Course in Statistical Programming With R by W. John Braun; Duncan J. Murdoch. Technometrics, 51(3):340, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586629.

Hlynka:2009:BRBb

[Hly09b] Myron Hlynka. Book review: Level Crossing Methods in Stochastic Models (International Series in Operations Research & Management Science) by Percy H. Brill. Technometrics, 51(3):343, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586634.

Hlynka:2009:BRBc

[Hly09c] Myron Hlynka. Book review: Probability Models for DNA Sequences Evolution (2nd ed.) by Richard Durrett. Technometrics, 51(4):483, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586663.

Hong:2008:APN

[HME08] Yili Hong, William Q. Meeker, and Luis A. Escobar. Avoiding problems with normal approximation confidence intervals for probabilities. Technometrics, 50(1):64–68, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471426.

Hawkins:2007:SSM

[HMT07] Douglas M. Hawkins and Edgard M. Maboudou-Tchao. Selfstarting multivariate exponentially weighted moving average control charting. Technometrics, 49(2):199–209, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471313.

Hawkins:2008:MEW

[HMT08] Douglas M. Hawkins and Edgard M. Maboudou-Tchao. Multivariate exponentially weighted moving covariance matrix. Technometrics, 50(2):155–166, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471456.

Hwang:2003:PCR

[HN03] J. T. Gene Hwang and Dan Nettleton. Principal components regression with data-chosen components and related methods. Technometrics, 45(1):70–79, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25046976.

Hoffman:2000:BRB

[Hof00] Lorrie L. Hoffman. Book review: Statistical Quality Control, Strategies and Tools for Continual Improvement by Johannes Ledolter; Claude W. Burrill. Technometrics, 42(4):433–434, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270963.

Holt:2001:BRB

[Hol01] Melinda Miller Holt. Book review: Statistics and Data Analysis from Elementary to Intermediate by A. C. Tamhane; D. D. Dunlop. Technometrics, 43(2):237–238, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271045.

Holt:2008:BRB

[Hol08] Melinda Holt. Book review: Learning SAS by Example: A Programmer’s Guide by Ron Cody. Technometrics, 50(1):91–92, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471436.

Hornikova:2006:BRB

[Hor06] Adriana Hornikova. Book review: Statistics for the Quality Control Chemistry Laboratory by Eamonn Mullins. Technometrics, 48(1):151, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471133.

Hornikova:2007:BRB

[Hor07] Adriana Hornikova. Book review: Stochastic Ageing and Dependence for Reliability by Chin-Diew Lai; Min Xie. Technometrics, 49(2):222, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471316.

Hornikova:2009:BRBb

[Hor09a] Adriana Horníková. Book review: Bayesian Reliability by Michael S. Hamada; Alyson G. Wilson; C. Shane Reese; Harry F. Martz. Technometrics, 51(4):483–484, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586664.

Hornikova:2009:BRBa

[Hor09b] Adriana Horníková. Book review: Optimum Experimental Designs, With SAS by A. C. Atkinson; A. N. Donev; R. D. Tobias. Technometrics, 51(3):341–342, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586631.

Hornikova:2009:BRBc

[Hor09c] Adriana Horníková. Book review: The Little SAS Book: A Primer (4th ed.) by Lora D. Delwiche; Susan J. Slaughter. Technometrics, 51(4):485, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586666.

Hall:2001:FDA

[HPP01] Peter Hall, D. S. Poskitt, and Brett Presnell. A functional dataanalytic approach to signal discrimination. Technometrics, 43(1): 1–9, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270852.

Hoffman:2005:BRB

[HR05] Lorrie L. Hoffman and Elayne Reiss. Book review: Statistical Models by A. C. Davison. Technometrics, 47(1):97, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470944.

Hubert:2005:RNA

[HRV05] Mia Hubert, Peter J. Rousseeuw, and Karlien Vanden Branden. ROBPCA: A new approach to robust principal component analysis. Technometrics, 47(1):64–79, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470935.

Han:2009:PCE

[HSNB09] Gang Han, Thomas J. Santner, William I. Notz, and Donald L. Bartel. Prediction for computer experiments having quantitative and qualitative input variables. Technometrics, 51(3):278–288, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586622.

Han:2009:SDT

[HSR09] Gang Han, Thomas J. Santner, and Jeremy J. Rawlinson. Simultaneous determination of tuning and calibration parameters for computer experiments. Technometrics, 51(4):464–474, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586655.

Hengartner:2008:EEE

[HTMW08] Nicolas W. Hengartner, Bruce E. Takala, Sarah E. Michalak, and Stephen A. Wender. Evaluating experiments for estimating the bit failure cross-section of semiconductors using a colored spectrum neutron beam. Technometrics, 50(1):8–14, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471420.

Huang:2000:BRB

[Hua00] Peng Huang. Book review: Stochastic Models in Reliability by Terje Aven; Uwe Jensen. Technometrics, 42(3):314–315, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271097.

Hubert:2001:MOD

[Hub01] Mia Hubert. [Multivariate Outlier Detection and Robust Covariance Matrix Estimation]: Discussion. Technometrics, 43(3):303– 306, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

1271217. See [PP01a].

Huber:2006:DTP

[Hub06] Peter J. Huber. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):332–334, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471204. See [Mal06].

Hunter:2000:FET

[Hun00] J. Stuart Hunter. Foreword: The enduring tradition of “Technometrics”. Technometrics, 42(1):1, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271424.

Hussein:2008:BRB

[Hus08] Abdulkadir Hussein. Book review: Statistical Monitoring of Clinical Trials by Lemuel A. Moyé. Technometrics, 50(1):87–88, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471430.

Hutson:2004:BRB

[Hut04] Alan D. Hutson. Book review: Resampling Methods for Dependent Data by S. N. Lahiri. Technometrics, 46(2):252, May 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470812.

Huzurbazar:2000:MAE

[Huz00] Aparna V. Huzurbazar. Modeling and analysis of engineering systems data using flowgraph models. Technometrics, 42(3):300–306, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271084.

Huzurbazar:2001:BRB

[Huz01] Aparna V. Huzurbazar. Book review: Modeling, Analysis, Design, and Control of Stochastic Systems by V. G. Kulkarni. Technometrics, 43(3):378, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271237.

Huzurbazar:2003:BRB

[Huz03] Snehalata V. Huzurbazar. Book review: Statistics of the Galaxy Distribution by Vincent J. Martinez; Enn Saar. Technometrics, 45(1):108–109, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046993.

Heuchenne:2007:NRC

[HV07] Cédric Heuchenne and Ingrid Van Keilegom. Nonlinear regression with censored data. Technometrics, 49(1):34–44, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471273.

Huwang:2002:ICM

[HWY02] Longcheen Huwang, C. F. J. Wu, and C. H. Yen. The idle column method: Design construction, properties and comparisons. Technometrics, 44(4):347–355, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271542.

Hawkins:2005:SPC

[HZ05] Douglas M. Hawkins and K. D. Zamba. Statistical process control for shifts in mean or variance using a changepoint formulation. Technometrics, 47(2):164–173, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470978.

Iyer:2002:CSG

[IM02] Hari Iyer and Thomas Mathew. Comments on: “A simple general method for constructing confidence intervals for functions of variance components” [Technometrics 43 (2001), no. 3, 356–367; 1940069] by A. K. L. Chiang. Technometrics, 44(3):284–285, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270493. See

[Chi01].

Ingrassia:2005:NNM

[IM05] Salvatore Ingrassia and Isabella Morlini. Neural network modeling for small datasets. Technometrics, 47(3):297–311, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471023.

Iyer:2002:BRB

[Iye02] Hari Iyer. Book review: Collecting Spatial Data: Optimum Design of Experiments for Random Fields by Werner G. Müller. Technometrics, 44(3):296–297, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270509.

Jarrett:2004:BRB

[Jar04] Jeffrey E. Jarrett. Book review: Making Sense of Data by Donald J. Wheeler. Technometrics, 46(1):117–118, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470780.

Jarrett:2008:BRB

[Jar08] Jeffrey E. Jarrett. Book review: The Nature of Statistical Evidence by Bill Thomson. Technometrics, 50(3):409, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471503.

Jarrett:2009:BRB

[Jar09] Jeffrey E. Jarrett. Book review: Statistical Principles and Techniques in Scientific and Social Investigations by Woitek J. Krzanowski. Technometrics, 51(2):219–220, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586605.

Jeske:2006:EBB

[JC06] Daniel R. Jeske and Arunava Chakravartty. Effectiveness of bootstrap bias correction in the context of clock offset estimators. Technometrics, 48(4):530–538, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471245.

Jones:2001:PEW

[JCR01] L. Allison Jones, Charles W. Champ, and Steven E. Rigdon. The performance of exponentially weighted moving average charts with estimated parameters. Technometrics, 43(2):156–167, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271029. Statisti-

cal methodology for industrial problems.

Joseph:2007:FIP

[JD07] V. Roshan Joseph and James Dillon Delaney. Functionally induced priors for the analysis of experiments. Technometrics, 49 (1):1–11, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471270.

Jensen:2002:BRB

[Jen02] Donald R. Jensen. Book review: Binomial Distribution Handbook for Scientists and Engineers by Elart von Collani; Klaus Dräger. Technometrics, 44(3):290–291, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270500.

Jensen:2008:BRB

[Jen08] Willis A. Jensen. Book review: Decision Trees for Business Intelligence and Data Mining: Using SAS(R) Enterprise Miner^{T M} by Barry De Ville. Technometrics, 50(3):409–410, August 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471505.

Jensen:2009:BRBb

[Jen09a] Willis A. Jensen. Book review: Data Mining Using SAS(R) Enterprise Miner^{T M} by R. Matignon. Technometrics, 51(1):105, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586577.

Jensen:2009:BRBc

[Jen09b] Willis A. Jensen. Book review: Elementary Statistics Using JMP(R) by Sandra D. Schlotzhauer. Technometrics, 51(2):223– 224, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586610.

Jensen:2009:BRBa

[Jen09c] Willis A. Jensen. Book review: EMP III: Using Imperfect Data by D. J. Wheeler. Technometrics, 51(1):104–105, February 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586576.

Jiang:2003:BRB

[Jia03] Wei Jiang. Book review: The Mahalanobis–Taguchi Strategy by Genichi Taguchi; Rajesh Jugulum. Technometrics, 45(3):266– 267, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047056.

Jeske:2007:MTM

[JL07] Daniel R. Jeske and Regina Y. Liu. Mining and tracking massive text data: Classification, construction of tracking statistics, and inference under misclassification. Technometrics, 49(2):116–128, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471305.

Jeong:2006:WBD

[JLH^{+}06] Myong K. Jeong, Jye-Chyi Lu, Xiaoming Huo, Brani Vidakovic, and Di Chen. Wavelet-based data reduction techniques for process fault detection. Technometrics, 48(1):26–40, February 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471112.

Jeske:2008:CMT

[JLSZ08] D. R. Jeske, R. A. Lockhart, M. A. Stephens, and Q. Zhang. Cramér–von Mises tests for the compatibility of two software operating environments. Technometrics, 50(1):53–63, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471425.

Jeng:2000:CAC

[JM00] Shuen-Lin Jeng and William Q. Meeker. Comparisons of approximate confidence interval procedures for Type I censored data. Technometrics, 42(2):135–148, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271445.

Jeng:2001:PSC

[JM01] Shuen-Lin Jeng and William Q. Meeker. Parametric simultaneous confidence bands for cumulative distributions from censored data. Technometrics, 43(4):450–461, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270816.

Johnston:2000:BRB

[Joh00] Gordon Johnston. Book review: Reliability for Technology, Engineering, and Management by Paul Kales. Technometrics, 42(2):

207, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271456.

Johnston:2003:BRB

[Joh03] Gordon Johnston. Book review: Statistical Models and Methods for Lifetime Data by Jerald F. Lawless. Technometrics, 45 (3):264–265, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047054.

Jones:2004:MCM

[Jon04] Geoffrey Jones. Markov chain Monte Carlo estimation for the two-component model. Technometrics, 46(1):99–107, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470765.

Jones:2008:QAS

[Jon08] Wendell Jones. [Quality Assessment for Short Oligonucleotide Microarray Data]: Comment. Technometrics, 50(3):265–267, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471484.

Joseph:2003:RPD

[Jos03] V. Roshan Joseph. Robust parameter design with feed-forward control. Technometrics, 45(4):284–292, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047085.

Joseph:2004:BRB

[Jos04] V. Roshan Joseph. Book review: Statistics and Probability for Engineering Applications with Microsoft(R) Excel by W. J. De-Coursey. Technometrics, 46(3):361–362, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470855.

Joseph:2006:BAD

[Jos06a] V. Roshan Joseph. A Bayesian approach to the design and analysis of fractionated experiments. Technometrics, 48(2):219–229, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471158.

Joseph:2006:LK

[Jos06b] V. Roshan Joseph. Limit kriging. Technometrics, 48(4):458–466, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471238.

Juan:2001:UAI

[JP01] Jesus Juan and Francisco J. Prieto. Using angles to identify concentrated multivariate outliers. Technometrics, 43(3):311–322, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271219.

Joo:2009:JDR

[JQ09] Jong-Hoon Joo and Peihua Qiu. Jump detection in a regression curve and its derivative. Technometrics, 51(3):289–305, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586623.

Jeske:2003:ECO

[JS03] Daniel R. Jeske and Ashwin Sampath. Estimation of clock offset using bootstrap bias-correction techniques. Technometrics, 45 (3):256–261, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047052.

Jun:2007:APS

[JS07] Mikyoung Jun and Michael L. Stein. An approach to producing space: Time covariance functions on spheres. Technometrics, 49(4):468–479, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471392.

Jiang:2000:NSM

[JTW00] Wei Jiang, Kwok-Leung Tsui, and William H. Woodall. A new SPC monitoring method: The ARMA chart. Technometrics, 42 (4):399–410, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270950.

Joseph:2002:RPD

[JW02] V. Roshan Joseph and C. F. Jeff Wu. Robust parameter design of multiple-target systems. Technometrics, 44(4):338–346, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271541.

Joseph:2004:FAM

[JW04a] V. Roshan Joseph and C. F. Jeff Wu. Failure amplification method: An information maximization approach to categorical response optimization. Technometrics, 46(1):1–12, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470751. See discussion

[LM04, ME04, Miy04, NW04, Sud04].

Joseph:2004:OWE

[JW04b] V. Roshan Joseph and C. F. Jeff Wu. [Operating Window: An Engineering Measure for Robustness]: Response. Technometrics, 46 (1):30–31, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470758.

Jiang:2002:PID

[JWT^{+}02] Wei Jiang, Huaiqing Wu, Fugee Tsung, Vijayan N. Nair, and Kwok-Leung Tsui. Proportional integral derivative charts for process monitoring. Technometrics, 44(3):205–214, August 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270485.

Kafadar:2000:EA

[Kaf00a] Karen Kafadar. Editorial announcement. Technometrics, 42(4): 329, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270941.

Kafadar:2000:ER

[Kaf00b] Karen Kafadar. Editor’s report. Technometrics, 42(4):330–331, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270942.

Kafadar:2000:FYB

[Kaf00c] Karen Kafadar. Forty years of “Technometrics”: Past, present, and future. Technometrics, 42(1):2–4, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271425.

Kafadar:2001:ER

[Kaf01a] Karen Kafadar. Editor’s report. Technometrics, 43(4):385–387, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270809.

Kafadar:2001:MJW

[Kaf01b] Karen Kafadar. In memoriam: John Wilder Tukey, June 16, 1915–July 26, 2000. Technometrics, 43(3):251–255, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.allbusiness.com/technology/
808081-1.html;   http://www.jstor.org/stable/1271211. Spe-

cial Tukey Memorial issue.

Kafadar:2001:SMI

[Kaf01c] Karen Kafadar. Statistical methodology for industrial problems. Technometrics, 43(2):119, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271025.

Kafadar:2003:BRB

[Kaf03] Karen Kafadar. Book review: Statistical Process Control: The Deming Paradigm and Beyond by James R. Thompson; Jacek Koronacki. Technometrics, 45(1):103–104, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25046985.

Kalemkarian:2003:BRB

[Kal03] George S. Kalemkarian. Book review: The Six Sigma Journey from Art to Science: A Business Novel by Larry Walters. Technometrics, 45(3):273, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047064.

Kalemkarian:2006:BRB

[Kal06] George S. Kalemkarian. Book review: Six Sigma for Dummies(R) by Craig Gygi; Neil DeCarlo; Bruce Williams. Technometrics, 48(2):305–306, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471170.

Kapatou:2003:BRB

[Kap03] Alexandra Kapatou. Book review: Testing for Normality by Henry C. Thode. Technometrics, 45(2):179, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047028.

Karlsson:2007:BRBf

[Kar07a] Andreas Karlsson. Book review: Benchmarking, Temporal Distribution, and Reconciliation Methods for Time Series by E. B. Dagum; P. A. Cholette. Technometrics, 49(4):491, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471395.

Karlsson:2007:BRBa

[Kar07b] Andreas Karlsson. Book review: Elementary Survey Sampling by Richard L. Schaeffer; William Mendenhall,; R. Lyman Ott. Technometrics, 49(2):225, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471321.

Karlsson:2007:BRBc

[Kar07c] Andreas Karlsson. Book review: Estimation in Surveys with Nonresponse by Carl-Erik Särndal; Sixten Lundström. Technometrics, 49(2):227, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471326.

Karlsson:2007:BRBb

[Kar07d] Andreas Karlsson. Book review: Matrix Analysis for Statistics by James R. Schott. Technometrics, 49(2):226, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471324.

Karlsson:2007:BRBd

[Kar07e] Andreas Karlsson. Book review: Precedence-Types Tests and Applications by N. Balakrishnan; H. K. Tony Ng. Technometrics, 49(3):359–360, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471356.

Karlsson:2007:BRBe

[Kar07f] Andreas Karlsson. Book review: Statistical Matching: Theory and Practice by M. D’Orazio; M. Di Zio; M. Scanu. Technometrics, 49(3):361–362, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471360.

Katti:2001:BRB

[Kat01] S. K. Katti. Book review: Robust Bayesian Analysis by David Rios Insua; Fabrizio Ruggeri. Technometrics, 43(4):493, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270832.

Katsaounis:2003:BRBb

[Kat03a] Tena I. Katsaounis. Book review: Visualizing Statistical Models and Concepts by R. W. Farebrother. Technometrics, 45(2): 175–176, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047024.

Katsaounis:2003:BRBa

[Kat03b] Tena Ipsilantis Katsaounis. Book review: Methods of Multivariate Statistics by M. S. Srivastava. Technometrics, 45(1):100–101, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25046981.

Katsaounis:2004:BRBa

[Kat04a] Tena I. Katsaounis. Book review: Analyzing Multivariate Data by James M. Lattin; J. Douglas Carroll; Paul E. Green. Technometrics, 46(2):254–255, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470815.

Katsaounis:2004:BRBb

[Kat04b] Tena Ipsilantis Katsaounis. Book review: Exploring Multivariate Data with the Forward Search by Anthony C. Atkinson; Marco Riani; Andrea Cerioli. Technometrics, 46(4):483–484, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470896.

Katsaounis:2008:BRB

[Kat08a] T. I. Katsaounis. Book review: Introductory Statistical Inference by Nitis Mukhopadhyay. Technometrics, 50(1):89–90, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471433.

Katti:2008:BRB

[Kat08b] S. K. Katti. Book review: Sequential Analysis by Zakkula Govindarajulu. Technometrics, 50(4):545–546, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471540.

Katsaounis:2009:BRB

[Kat09] Tena I. Katsaounis. Book review: Statistical and Probabilistic Methods in Actuarial Science by Philip J. Boland. Technometrics, 51(1):101, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586572.

Kulahci:2005:UPB

[KB05] Murat Kulahci and Søren Bisgaard. The use of Plackett-Burman designs to construct split-plot designs. Technometrics, 47(4):495– 501, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471074.

Kimel:2008:FDR

[KBS08] Maria Tripolski Kimel, Yoav Benjamini, and David M. Steinberg. The false discovery rate for multiple testing in factorial experiments. Technometrics, 50(1):32–39, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471423.

Kuhn:2000:INF

[KCM00] Andrew M. Kuhn, Walter H. Carter, and Raymond H. Myers. Incorporating noise factors into experiments with censored data. Technometrics, 42(4):376–383, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270947.

Kowalski:2002:SPD

[KCV02] Scott M. Kowalski, John A. Cornell, and G. Geoffrey Vining. Splitplot designs and estimation methods for mixture experiments with process variables. Technometrics, 44(1):72–79, February 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270686.

Kim:2005:OES

[KD05] Pansoo Kim and Yu Ding. Optimal engineering system design guided by data-mining methods. Technometrics, 47(3):336– 348, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471026.

Ke:2005:BRB

[Ke05] Weiming Ke. Book review: The Optimal Design of Blocked and Split-Plot Experiments by Peter Goos. Technometrics, 47(1):94, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470938.

Ke:2006:BRB

[Ke06] Weiming Ke. Book review: Statistical Modelling in GLIM4 by Murray Aitkin; Brian Francis; John Hinde. Technometrics, 48 (3):446–447, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471219.

Ke:2007:BRB

[Ke07] Weiming Ke. Book review: Reliability and Risk Models: Setting Reliability Requirements by Michael Todinov. Technometrics, 49 (4):497, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471406.

Ke:2008:BRB

[Ke08] Weiming Ke. Book review: Probability, Statistics, and Reliability for Engineers and Scientists by Bilal M. Ayyub; Richard H. McCuen. Technometrics, 50(1):91, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471435.

Ke:2009:BRBa

[Ke09a] Weiming Ke. Book review: Design, Evaluation, and Analysis of Questionnaires for Survey Research by Willem E. Saris; Irmtraud N. Gallhofer. Technometrics, 51(1):105, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586578.

Ke:2009:BRBb

[Ke09b] Weiming Ke. Book review: Introduction to Stochastic Calculus Applied to Finance (2nd ed.) by Damien Lamberton; Bernard Lapeyre. Technometrics, 51(2):215–216, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586597.

Kenyon:2002:BRB

[Ken02] James R. Kenyon. Book review: Analysis of Multivariate Survival Data by Philip Hougaard. Technometrics, 44(1):86–87, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270696.

Kenyon:2003:BRB

[Ken03] James R. Kenyon. Book review: Statistical Methods for the Analysis of Repeated Measurements by Charles S. Davis. Technometrics, 45(1):99–100, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046980.

Kerkering:2002:BRB

[Ker02] J. Charles Kerkering. Book review: Eliciting and Analyzing Expert Judgment. A Practical Guide by Mary A. Meyer; Jane M. Booker. Technometrics, 44(2):193–194, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271273.

Kerkering:2003:BRB

[Ker03] J. Charles Kerkering. Book review: Subjective and Objective Bayesian Statistics: Principles, Models, and Applications by S. James Press. Technometrics, 45(4):369–370, November 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047106.

Kettaneh:2000:BRB

[Ket00] Nouna Kettaneh. Book review: Statistical Modeling by Wavelets by Brani Vidakovic. Technometrics, 42(3):316–317, August 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271099.

Keyes:2001:BRB

[Key01] Tim K. Keyes. Book review: Applied Regression Analysis and Multivariable Methods by D. Kleinbaum; L. Kupper; K. Muller; A. Nizam. Technometrics, 43(1):101, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270867.

Karr:2007:SPP

[KFV^{+}07] Alan F. Karr, William J. Fulp, Francisco Vera, S. Stanley Young, Xiaodong Lin, and Jerome P. Reiter. Secure, privacy-preserving analysis of distributed databases. Technometrics, 49(3):335–345, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471352.

Khuri:2006:MRS

[Khu06] André I. Khuri. Mixed response surface models with heterogeneous within-block error variances. Technometrics, 48(2):206–218, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471157.

Kianifard:2000:BRB

[Kia00] Farid Kianifard. Book review: Models for Discrete Data by D. Zelterman. Technometrics, 42(3):313–314, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271095.

Kianifard:2001:BRB

[Kia01] Farid Kianifard. Book review: Models for Repeated Measurements by J. K. Lindsey. Technometrics, 43(2):234, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271041.

Kibria:2007:BRB

[Kib07] B. M. Golam Kibria. Book review: Bayesian Statistics and Marketing by P. E. Rossi; G. M. Allenby; R. McCulloch. Technometrics, 49(2):230, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471331.

Kim:2000:BRB

[Kim00] Hoon Kim. Book review: Monte Carlo Statistical Methods by Christian P. Robert; George Casella. Technometrics, 42(4):430– 431, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270959.

Kim:2004:BRB

[Kim04] Hoon Kim. Book review: Probability, Statistics, and Random Processes for Engineers by Richard H. Williams. Technometrics, 46(3):360, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470852.

Kim:2005:BRB

[Kim05] Hoon Kim. Book review: A First Course in Statistical Methods by R. Lyman Ott; Michael T. Longnecker. Technometrics, 47 (1):98–99, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470946.

Kang:2009:BOS

[KJ09] Lulu Kang and V. Roshan Joseph. Bayesian optimal single arrays for robust parameter design. Technometrics, 51(3):250–261, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586620.

Kabaila:2002:IDS

[KL02] Paul Kabaila and Chris J. Lloyd. The importance of the designated statistic on Buehler upper limits on a system failure probability. Technometrics, 44(4):390–395, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271546.

Kendall:2008:QAS

[KL08] Jude Kendall and B. Lakshmi. [Quality Assessment for Short Oligonucleotide Microarray Data]: Comment. Technometrics, 50 (3):274–276, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471487.

Klaassen:2001:CAS

[Kla01] Chris A. J. Klaassen. Credit in acceptance sampling on attributes. Technometrics, 43(2):212–222, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www. jstor.org/stable/1271034. Statistical methodology for industrial problems.

Katenka:2008:RTL

[KLM08] Natallia Katenka, Elizaveta Levina, and George Michailidis. Robust target localization from binary decisions in wireless sensor networks. Technometrics, 50(4):448–461, November 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471528.

Krishnamoorthy:2004:OST

[KM04] K. Krishnamoorthy and Thomas Mathew. One-sided tolerance limits in balanced and unbalanced one-way random models based on generalized confidence intervals. Technometrics, 46(1):44–52, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470760.

Kaistha:2004:SPC

[KML04] Nitin Kaistha, Charles F. Moore, and Mary G. Leitnaker. A statistical process control framework for the characterization of variation in batch profiles. Technometrics, 46(1):53–68, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470761.

Krishnamoorthy:2008:NBM

[KMM08] K. Krishnamoorthy, Thomas Mathew, and Shubhabrata Mukherjee. Normal-based methods for a gamma distribution: Prediction and tolerance intervals and stress-strength reliability. Technometrics, 50(1):69–78, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471427.

Kokoska:2001:BRB

[Kok01] Stephen Kokoska. Book review: Chance Encounters: A First Course in Data Analysis and Inference by Christopher J. Wild; George A. F. Seber. Technometrics, 43(2):242–243, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271056.

Koru-Sengul:2007:BRBb

[KS07a] Tulay Koru-Sengul. Book review: Graphics of Large Datasets: Visualizing a Million by Antony Unwin; Martin Theus; Heike Hofmann. Technometrics, 49(3):364–365, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471368.

Koru-Sengul:2007:BRBa

[KS07b] Tulay Koru-Sengul. Book review: Statistics in the Pharmaceutical Industry by C. Ralph Buncher; Jia-Yeong Tsay. Technometrics, 49(2):228, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471327.

Ke:2003:SDU

[KT03] Weiming Ke and Boxin Tang. Selecting 2^{m−p} designs using a minimum aberration criterion when some two-factor interactions are important. Technometrics, 45(4):352–360, November 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047092.

Kuhfeld:2005:LFD

[KT05] Warren F. Kuhfeld and Randall D. Tobias. Large factorial designs for product engineering and marketing research applications. Technometrics, 47(2):132–141, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470975. See correction [Ano05-58].

Kuhn:2003:BRB

[Kuh03] Andrew M. Kuhn. Book review: Growth Curve Models and Statistical Diagnostics by Jian-Xin Pan; Kai-Tai Fang. Technometrics, 45(3):270, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047060.

Kuhn:2005:BRBb

[Kuh05a] Andrew M. Kuhn. Book review: Code Complete by Steve Mc-Connell. Technometrics, 47(4):524–525, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471091.

Kuhn:2005:BRBa

[Kuh05b] Andrew M. Kuhn. Book review: Multivariate Statistical Methods in Quality Management by Kai Yang; Jayant Trewn. Technometrics, 47(1):94–95, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470939.

Kundu:2008:BIL

[Kun08] Debasis Kundu. Bayesian inference and life testing plan for the Weibull distribution in presence of progressive censoring. Technometrics, 50(2):144–154, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471455.

Kuo:2000:BRB

[Kuo00] Lynn Kuo. Book review: Markov Chain Monte Carlo by Dani Gamerman. Technometrics, 42(2):216, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271470.

Kushary:2000:BRB

[Kus00] Debashis Kushary. Book review: Bootstrap Methods and Their Application by A. C. Davison; D. V. Hinkley. Technometrics, 42(2):216–217, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271471.

Kushler:2002:BRB

[Kus02] Robert H. Kushler. Book review: Computational Statistics Handbook with MATLAB(R) by Wendy L. Martinez; Angel R. Martinez. Technometrics, 44(4):405–406, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271563.

Kushler:2003:BRB

[Kus03] Robert Kushler. Book review: Statistical Computing: An Introduction to Data Analysis Using S-PLUS by Michael J. Crawley. Technometrics, 45(4):369, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047104.

Kushler:2006:BRB

[Kus06] Robert H. Kushler. Book review: Exploratory Data Analysis with MATLAB(R) by Wendy L. Martinez; Angel R. Martinez. Technometrics, 48(2):311, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471179.

Kvam:2008:LBM

[Kva08] Paul Kvam. Length bias in the measurements of carbon nanotubes. Technometrics, 50(4):462–467, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471529.

LaLonde:2001:BRB

[LaL01] Steven M. LaLonde. Book review: Multivariate Statistical Methods: A First Course by George A. Marcoulides; Scott L. Hershberger. Technometrics, 43(4):491–492, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270830.

LaLonde:2004:BRBa

[LaL04a] Steven M. LaLonde. Book review: Combined Survey Sampling Inference by Ken Brewer. Technometrics, 46(1):117, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470779.

LaLonde:2004:BRBb

[LaL04b] Steven M. LaLonde. Book review: Probability and Statistics for Computer Science by James L. Johnson. Technometrics, 46(4): 491, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470908.

LaLonde:2009:BRB

[LaL09] Steven M. LaLonde. Book review: Statistics Using SAS(R) Enterprise Guide by James B. Davis. Technometrics, 51(1):103, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586574.

Langsrud:2001:ISE

[Lan01] Øyvind Langsrud. Identifying significant effects in fractional factorial multiresponse experiments. Technometrics, 43(4):415–424, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270812.

Landwehr:2006:DTP

[Lan06] James M. Landwehr. Discussion: “Tukey’s paper after 40 years” [Technometrics 48 (2006), no. 3, 319–325; MR2248365] by C. Mallows. Technometrics, 48(3):334–336, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471205. See [Mal06].

Larsen:2008:BRB

[Lar08] Greg Larsen. Book review: Excel for Scientists and Engineers: Numerical Methods by E. Joseph Billo. Technometrics, 50(2):236–237, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471471.

Lawless:2000:ITC

[Law00] Jerry Lawless. Introduction to two classics in reliability theory. Technometrics, 42(1):5–6, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271426.

Lawson:2001:BRB

[Law01] John S. Lawson. Book review: Improving Performance through Statistical Thinking by Galen C. Britz; Donald W. Emerling; Lynne B. Hare; Roger W. Hoerl; Stuart J. Janis; Janice E. Shade. Technometrics, 43(2):231–232, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271037.

Lazar:2003:BRBa

[Laz03a] Nicole Lazar. Book review: Testing Statistical Hypotheses of Equivalence by Stefan Wellek. Technometrics, 45(3):271–272, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047062.

Lazar:2003:BRBb

[Laz03b] Nicole A. Lazar. Book review: Statistical Analysis with Missing Data by R. J. A. Little; D. B. Rubin. Technometrics, 45 (4):364–365, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047098.

Lazar:2005:BRB

[Laz05] Nicole A. Lazar. Book review: Statistics of Extremes: Theory and Applications by Jan Beirlant; Yuri Goegebeur; Johan Segers; Jozef Teugels. Technometrics, 47(3):376–377, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471041.

Linkletter:2006:VSG

[LBH^{+}06] Crystal Linkletter, Derek Bingham, Nicholas Hengartner, David Higdon, and Kenny Q. Ye. Variable selection for Gaussian process models in computer experiments. Technometrics, 48(4):478– 490, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471240.

Lin:2004:AMD

[LBHB04] Nan Lin, Barbara A. Bailey, Xuming He, and William G. Buttlar. Adjustment of measuring devices with linear models. Technometrics, 46(2):127–134, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470798.

Lucas:2000:FIR

[LC00] James M. Lucas and Ronald B. Crosier. Fast initial response for CUSUM quality-control schemes: Give your CUSUM a head start. Technometrics, 42(1):102–107, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271440.

Lian:2006:SAM

[LCdC06] Zilong Lian, Bianca M. Colosimo, and Enrique del Castillo. Setup adjustment of multiple lots using a sequential Monte Carlo method. Technometrics, 48(3):373–385, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471209.

Liu:2004:CSD

[LD04] Yufeng Liu and Angela Dean. k-circulant supersaturated designs. Technometrics, 46(1):32–43, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470759.

Lindqvist:2003:TRP

[LEH03] B. H. Lindqvist, G. Elvebakk, and K. Heggland. The trend-renewal process for statistical analysis of repairable systems. Technometrics, 45(1):31–44, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046972.

Lewis:2004:BRB

[Lew04] Richard A. Lewis. Book review: Applied Simulation Modeling by Andrew F. Seila; Vlatko Ceric; Pandu Tadikamalla. Technometrics, 46(3):360–361, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470853.

Li:2002:FBA

[LH02] Ta-Hsin Li and Melvin J. Hinich. A filter bank approach for modeling and forecasting seasonal patterns. Technometrics, 44(1):1–14, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270679.

Liang:2006:BRB

[LH06] Faming Liang and Jianhua Huang. Book review: Statistical and Computational Inverse Problems by Jari Kaipio; Erkki Somersalo. Technometrics, 48(1):146, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471125.

Lee:2002:MRF

[LHB^{+}02] Herbert K. H. Lee, Dave M. Higdon, Zhuoxin Bi, Marco A. R. Ferreira, and Mike West. Markov random field models for highdimensional parameters in simulations of fluid flow in porous media. Technometrics, 44(3):230–241, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270487.

Li:2005:BIO

[Li05] Baibing Li. Bayesian inference for origin-destination matrices of transport networks using the EM algorithm. Technometrics, 47 (4):399–408, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471065.

Li:2006:BRB

[Li06] Runze Li. Book review: Introduction to Nonparametric Regression by Kunio Takezawa. Technometrics, 48(4):575–576, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471262.

Lang:2008:FSC

[LI08] Duncan Temple Lang and Ross Ihaka. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):443–446, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471526.

Li:2009:NGK

[Li09] Baibing Li. A non-Gaussian Kalman filter with application to the estimation of vehicular speed. Technometrics, 51(2):162–172, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586593.

Lin:2002:BRBb

[Lin02] Yachen Lin. Book review: Geometric Data Analysis: An Empirical Approach to Dimensionality Reduction and the Study of Patterns by Michael Kirby. Technometrics, 44(2):196–197, May 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271278.

Lipovetsky:2001:BRB

[Lip01] Stan Lipovetsky. Book review: Latent Variable Models and Factor Analysis by David J. Bartholomew; Martin Knott. Technometrics, 43(1):111, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270883.

Lipovetsky:2005:BRB

[Lip05] Stan Lipovetsky. Book review: Generalized Latent Variable Modeling: Multilevel, Longitudinal, and Structural Equation Models by Anders Skrondal; Sophia Rabe-Hesketh. Technometrics, 47(2):237, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470994.

Lipovetsky:2006:BRBa

[Lip06a] Stan Lipovetsky. Book review: Probability, Statistics, and Stochastic Processes by Peter Olofsson. Technometrics, 48(1):150–151, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471132.

Lipovetsky:2006:BRBb

[Lip06b] Stan Lipovetsky. Book review: Quantile Regression by Roger Koenker. Technometrics, 48(3):445–446, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471216.

Lipovetsky:2007:BRBc

[Lip07a] Stan Lipovetsky. Book review: Designing Economic Mechanisms by Leonid Hurwicz; Stanley Reiter. Technometrics, 49(2):229, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471330.

Lipovetsky:2007:BRBb

[Lip07b] Stan Lipovetsky. Book review: Prediction, Learning, and Games by Nicolo Cesa-Bianchi; Gabor Lugosi. Technometrics, 49(2):225, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471322.

Lipovetsky:2007:BRBa

[Lip07c] Stan Lipovetsky. Book review: The Structural Representation of Proximity Matrices with MATLAB by Lawrence Hubert; Phipps Arabie; Jacqueline Meulman. Technometrics, 49(1):107, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471292.

Lipovetsky:2009:BRB

[Lip09] Stan Lipovetsky. Book review: Numerical Recipes: The Art of Scientific Computing (3rd ed.) by William H. Press; Saul A. Teukolsky; William T. Wetterling; Brian P. Flannery. Technometrics, 51(4):481–482, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586661.

Littell:2001:BRB

[Lit01] Ramon C. Littell. Book review: Exact Statistical Methods for Data Analysis by Samaradasa Weerahandi. Technometrics, 43 (1):106–107, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270875.

Liu:2002:BRB

[Liu02] Shin Ta Liu. Book review: Permutation Methods: A Distance Function Approach by Paul W. Mielke,; Kenneth J. Berry. Technometrics, 44(3):289–290, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270497.

Liu:2004:BRB

[Liu04] Shin Ta Liu. Book review: SAS(R) Survival Analysis Techniques for Medical Research by Alan B. Cantor. Technometrics, 46(2):261, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470823.

Liu:2006:BRB

[Liu06] Shin Ta Liu. Book review: Nonlinear Signal Processing: A Statistical Approach by Gonzalo R. Arce. Technometrics, 48(1):148– 149, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471129.

Liu:2007:BRBa

[Liu07a] Shin Ta Liu. Book review: Counting Processes and Survival Analysis by Thomas R. Fleming; David P. Harrington. Technometrics, 49(3):362, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471362.

Liu:2007:BRBb

[Liu07b] Shin Ta Liu. Book review: Springer Handbook of Engineering Statistics by Hoang Pham. Technometrics, 49(4):494, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471400.

Liu:2009:BRBa

[Liu09a] Shin Ta Liu. Book review: A Statistical Approach to Neural Networks for Pattern Recognition by Robert A. Dunne. Technometrics, 51(1):103–104, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586575.

Liu:2009:BRBb

[Liu09b] Shuangzhe Liu. Book review: Advances in Economics and Econometrics: Theory and Applications, Ninth World Congress, Vol. III by Richard Blundell; Whitney K. Newey; Torsten Persson. Technometrics, 51(1):107, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586582.

Li:2003:OFPa

[LL03] William Li and Dennis K. J. Lin. Optimal foldover plans for twolevel fractional factorial designs. Technometrics, 45(2):142–149, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047011.

Lawrence:2008:BIM

[LLBN08] Earl Lawrence, Chuanhai Liu, Derek Bingham, and Vijayan N. Nair. Bayesian inference for multivariate ordinal data using parameter expansion. Technometrics, 50(2):182–191, May 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471458.

Liao:2005:OTS

[LLI05] C. T. Liao, T. Y. Lin, and H. K. Iyer. One- and two-sided tolerance intervals for general balanced mixed models and unbalanced one-way random models. Technometrics, 47(3):323–335, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471025.

Li:2003:OFPb

[LLY03] William Li, Dennis K. J. Lin, and Kenny Q. Ye. Optimal foldover plans for two-level nonregular orthogonal designs. Technometrics, 45(4):347–351, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047091.

Li:2002:BFF

[LM02] Huo Li and Robert W. Mee. Better foldover fractions for resolution III 2^{k−p} designs. Technometrics, 44(3):278–283, August 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270492.

Leitnaker:2004:FAM

[LM04] Mary G. Leitnaker and Robert W. Mee. [Failure Amplification Method: An Information Maximization Approach to Categorical Response Optimization]: Discussion. Technometrics, 46(1):12– 14, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

25470752. See [JW04a].

Li:2000:MRF

[LN00] William Li and Christopher J. Nachtsheim. Model-robust factorial designs. Technometrics, 42(4):345–352, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270944.

Li:2006:SSI

[LN06] Lexin Li and Christopher J. Nachtsheim. Sparse sliced inverse regression. Technometrics, 48(4):503–510, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471242.

Loeppky:2007:BRB

[Loe07] Jason Loeppky. Book review: A Modern Theory of Factorial Design by Rahul Mukerjee; C. F. Jeff Wu. Technometrics, 49 (3):365–366, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471370.

Lombard:2005:NCB

[Lom05] F. Lombard. Nonparametric confidence bands for a quantile comparison function. Technometrics, 47(3):364–369, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471028.

Lordo:2001:BRB

[Lor01] Robert A. Lordo. Book review: Learning from Data: Concepts, Theory, and Methods by Vladimir Cherkassky; Filip Mulier. Technometrics, 43(1):105–106, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270873.

Lordo:2005:BRB

[Lor05] Robert Lordo. Book review: Nonparametric and Semiparametric Models by Wolfgang Härdle; Marlene Müller; Stefan Sperlich; Axel Werwatz. Technometrics, 47(2):234–235, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470990.

Lordo:2006:BRB

[Lor06] Robert A. Lordo. Book review: Image Processing and Jump Regression Analysis by Peihua Qiu. Technometrics, 48(2):312–313, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471182.

Loughin:2001:BRB

[Lou01] Thomas M. Loughin. Book review: Data Analysis by Resampling: Concepts and Applications by Clifford E. Lunneborg. Technometrics, 43(3):373–374, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271231.

Lockhart:2006:TND

[LP06] Richard A. Lockhart and Chandanie W. Perera. Testing normality in designs with many parameters. Technometrics, 48(3):436– 444, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471215.

Luceno:2000:ERL

[LPP00] Alberto Luceno and Jaime Puig-Pey. Evaluation of the run-length probability distribution for CUSUM charts: Assessing chart performance. Technometrics, 42(4):411–416, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270951.

Liu:2000:AIC

[LS00] Chuanhai Liu and Don X. Sun. Analysis of interval-censored data from fractionated experiments using covariance adjustment. Technometrics, 42(4):353–365, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270945.

Li:2005:ACE

[LS05] Runze Li and Agus Sudjianto. Analysis of computer experiments using penalized likelihood in Gaussian kriging models. Technometrics, 47(2):111–120, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470973.

Luo:2006:TVS

[LSB06] Xiaohui Luo, Leonard A. Stefanski, and Dennis D. Boos. Tuning variable selection procedures by adding noise. Technometrics, 48(2):165–175, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471154.

Loeppky:2007:NDD

[LST07] J. L. Loeppky, R. R. Sitter, and B. Tang. Nonregular designs with desirable projection properties. Technometrics, 49(4):454– 467, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471391.

Loeppky:2009:CSS

[LSW09] Jason L. Loeppky, Jerome Sacks, and William J. Welch. Choosing the sample size of a computer experiment: A practical guide. Technometrics, 51(4):366–376, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586647.

Li:2009:FDR

[LT09] Yanting Li and Fugee Tsung. False discovery rate-adjusted charting schemes for multistage process monitoring and fault identification. Technometrics, 51(2):186–205, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586595.

Lu:2004:BRBa

[Lu04a] Z.-Q. John Lu. Book review: Nonlinear Time Series: Nonparametric and Parametric Methods by Jianqing Fan; Qiwei Yao. Technometrics, 46(1):114–115, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470775.

Lu:2004:BRBb

[Lu04b] Z.-Q. Jonh Lu. Book review: Modelling and Forecasting Financial Data: Techniques of Nonlinear Dynamics by Abdol S. Soofi; Liangyue Cao. Technometrics, 46(1):116–117, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470778.

Lu:2007:BRB

[Lu07] Z. Q. John Lu. Book review: Nonparametric Functional Data Analysis: Theory and Practice by Frédéric Ferraty; Philippe Vieu. Technometrics, 49(2):226, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471323.

Lu:2008:BRB

[Lu08] Z. Q. John Lu. Book review: Nonparametric Statistics with Applications to Science and Engineering by Paul H. Kvam; Brani Vidakovic. Technometrics, 50(2):237, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471473.

Lumley:2001:BRB

[Lum01] Thomas Lumley. Book review: Environmental Statistics with S-PLUS by Steven P. Millard; Nagaraj K. Neerchal. Technometrics, 43(4):495, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270836.

Logsdon:2003:PET

[LWS03] John Logsdon, Granville Tunnicliffe Wilson, and Carl Scarrott. Prediction of extreme temperatures in a reactor using measurements affected by control action. Technometrics, 45(2):159–168, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047013.

Lam:2002:UCD

[LWY02] Raymond L. H. Lam, William J. Welch, and S. Stanley Young. Uniform coverage designs for molecule selection. Technometrics, 44(2):99–109, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271254.

Lin:2002:BRBa

[LZ02] Haiqun Lin and Daniel Zelterman. Book review: Modeling Survival Data: Extending the Cox Model by T. M. Therneau; P. M. Grambsch. Technometrics, 44(1):85–86, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270694.

Magoun:2000:BRB

[Mag00] A. Dale Magoun. Book review: Data Driven Statistical Methods by Peter Sprent. Technometrics, 42(2):218, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271474.

Magnan:2009:BRB

[Mag09] Shon Magnan. Book review: Data Mining the Web by Zdravko Markov; Daniel T. Larose. Technometrics, 51(4):481, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586659.

Maitra:2001:CMD

[Mai01] Ranjan Maitra. Clustering massive datasets with applications in software metrics and tomography. Technometrics, 43(3):336–346, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271221.

Maitra:2005:BRN

[Mai05] Ranjan Maitra. Book review: Numerical Issues in Statistical Computing for the Social Scientist by Micah Altman; Jeff Gill; Michael P. McDonald. Hoboken, NJ: Wiley, 2004, ISBN 0-471-23633-0, xv + 323 pp. Technometrics, 47(2):241–242, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471001.

Maitra:2007:BRB

[Mai07] Ranjan Maitra. Book review: Independent Component Analysis: A Tutorial Introduction by James V. Stone. Technometrics, 49 (3):358–359, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471354.

Maitra:2008:BRB

[Mai08] Ranjan Maitra. Book review: Directional Statistics by Kanti V. Mardia; Peter E. Jupp. Technometrics, 50(4):546–547, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471543.

Mallows:2000:SC

[Mal00] C. L. Mallows. Some comments on C_{P}. Technometrics, 42(1):87– 94, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271437.

Malik:2005:BRB

[Mal05] Malik Beshir Malik. Book review: Applied Linear Regression by Sanford Weisberg. Technometrics, 47(3):371–372, August 2005.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471030.

Mallows:2006:TPA

[Mal06] Colin Mallows. Tukey’s paper after 40 years. Technometrics, 48(3): 319–325, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/

25471200. See discussion [Bri06, Buj06, Efr06, Hub06, Lan06].

Mamon:2008:BRB

[Mam08] Rogemar S. Mamon. Book review: Parameter Estimation for Scientists and Engineers by A. van den Bos. Technometrics, 50 (4):546, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471542.

Mamon:2009:BRB

[Mam09] Rogemar S. Mamon. Book review: Elements of Applied Probability for Engineering, Mathematics and Systems Science by D. McDonald. Technometrics, 51(1):100, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586571.

Marchette:2003:BRB

[Mar03] David J. Marchette. Book review: Bayesian Networks and Decision Graphs by Finn V. Jensen. Technometrics, 45(2):178–179, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047027.

Maronna:2005:PCO

[Mar05] Ricardo Maronna. Principal components and orthogonal regression based on robust scales. Technometrics, 47(3):264–273, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471020.

Marchette:2006:BRB

[Mar06] David J. Marchette. Book review: Data Analysis of Asymmetric Structures: Advanced Approaches in Computational Statistics by Takayuki Saito; Hiroshi Yadohisa. Technometrics, 48(2):310–311, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471178.

Mathew:2001:BRB

[Mat01] Thomas Mathew. Book review: The Analysis of Variance by Hardeo Sahai; Mohammed I. Ageel. Technometrics, 43(2):233–234, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271040.

Mazumdar:2000:BRB

[Maz00] Mainak Mazumdar. Book review: The Art of Causal Conjecture by Glenn Shafer. Technometrics, 42(3):321, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271106.

Mazu:2002:BRB

[Maz02] Michael J. Mazu. Book review: Sampling Methodologies with Applications by Poduri S. R. S. Rao. Technometrics, 44(4):398–399, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271550.

Mazu:2006:BRBb

[Maz06a] Michael J. Mazu. Book review: Design and Analysis of Gauge R&R Studies by Richard K. Burdick; Connie M. Borror; Douglas C. Montgomery. Technometrics, 48(2):305, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471169.

Mazu:2006:BRBa

[Maz06b] Michael J. Mazu. Book review: Elementary Statistical Quality Control by John T. Burr. Technometrics, 48(1):153–154, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471137.

McLeod:2004:DBF

[MB04] Robert G. McLeod and John F. Brewster. The design of blocked fractional factorial split-plot experiments. Technometrics, 46(2): 135–146, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470799.

Mease:2006:LHS

[MB06] David Mease and Derek Bingham. Latin hyperrectangle sampling for computer experiments. Technometrics, 48(4):467–477, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471239.

Molina:2005:SIA

[MBB05] German Molina, M. J. Bayarri, and James O. Berger. Statistical inverse analysis for a network microsimulator. Technometrics, 47(4):388–398, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471064.

Mckay:2000:CTM

[MBC00] M. D. Mckay, R. J. Beckman, and W. J. Conover. A comparison of three methods for selecting values of input variables in the analysis of output from a computer code. Technometrics, 42(1):55–61, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271432.

Morris:2009:CRS

[MBDD09] Max D. Morris, Stuart J. Birrell, Brad Dilts, and Philip M. Dixon. Composite response surface designs for factors with jointly symmetric effects. Technometrics, 51(2):206–214, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586596.

Malloy:2007:DSF

[MBG07] Elizabeth J. Malloy, Edward J. Bedrick, and Tim Goldsmith. Diagnostics for the scale of functional predictors in generalized linear models. Technometrics, 49(4):480–489, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471393.

McComb:2000:BRB

[McC00] Mark A. McComb. Book review: A Practical Guide to Heavy Tails by R. Adler; R. Feldman; M. Taqqu. Technometrics, 42(2):218, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271475.

McCormack:2002:BRB

[McC02] D. W. McCormack, Jr. Book review: Basic Statistical Methods and Models for the Sciences by Judah Rosenblatt. Technometrics, 44(4):401–402, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271556.

McComb:2003:BRB

[McC03a] Mark A. McComb. Book review: Comparison Methods for Stochastic Models and Risks by A. Müller; D. Stoyan. Technometrics, 45(4):370–371, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047107.

McCool:2003:BRB

[McC03b] John I. McCool. Book review: Probability and Statistics with Reliability, Queuing and Computer Science Applications by Kishor S. Trivedi. Technometrics, 45(1):107, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25046991.

McCullough:2003:BRB

[McC03c] B. D. McCullough. Book review: Predictions in Time Series Using Regression Models by Frantisek Stulajter. Technometrics, 45(1):102, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25046983.

McCullough:2004:BRB

[McC04] B. D. McCullough. Book review: Random Number Generation and Monte Carlo Methods by James E. Gentle. Technometrics, 46(2):252–253, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470813.

McCullough:2005:BRBa

[McC05a] B. D. McCullough. Book review: Diagnostic Checks in Time Series by Wai Kung Li. Technometrics, 47(1):100, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470949.

McCullough:2005:BRBb

[McC05b] B. D. McCullough. Book review: Statistical Analysis of Stochastic Processes in Time by J. K. Lindsey. Technometrics, 47(3):373– 374, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471034.

McCool:2006:BRB

[McC06a] John I. McCool. Book review: Statistical Engineering: An Algorithm for Reducing Variation in Manufacturing Processes by Stefan H. Steiner; R. Jock MacKay. Technometrics, 48(2):312, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471181.

McCool:2006:TDF

[McC06b] John I. McCool. Testing for dependency of failure times in life testing. Technometrics, 48(1):41–48, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471113.

McCormack:2006:BRB

[McC06c] D. W. McCormack, Jr. Book review: Analysis of Variance for Random Models, Volume II: Unbalanced Data, Theory, Methods, Applications, and Data Analysis by Hardeo Sahai; Mario Miguel Ojeda. Technometrics, 48(1):145–146, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471124.

McComb:2007:BRB

[McC07] Mark A. McComb. Book review: SPC for Right-Brain Thinkers by Lon Roberts. Technometrics, 49(1):101–102, February 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471281.

McComb:2008:BRB

[McC08] Mark A. McComb. Book review: The Quality Toolbox by Nancy R. Tague. Technometrics, 50(3):408, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471501.

McGuire:2002:BRB

[McG02] Stephen A. McGuire. Book review: Interpreting ISO 9001:2000 with Statistical Methodology by James L. Lamprecht. Technometrics, 44(4):406, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271564.

McGuire:2007:BRB

[McG07] Stephen A. McGuire. Book review: Introductory Statistics by Sheldon M. Ross. Technometrics, 49(1):100–101, February 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471280.

Meeker:2004:FAM

[ME04] William Q. Meeker and Luis A. Escobar. [Failure Amplification Method: An Information Maximization Approach to Categorical Response Optimization]: Discussion. Technometrics, 46 (1):15–16, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470753. See [JW04a].

Marx:2005:MPS

[ME05] Brian D. Marx and Paul H. C. Eilers. Multidimensional penalized signal regression. Technometrics, 47(1):13–22, February 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470930.

Mee:2001:NCD

[Mee01] Robert Mee. Noncentral composite designs. Technometrics, 43 (1):34–43, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270855.

Meeker:2003:BRB

[Mee03] William Q. Meeker. Book review: Recurrent Events Data Analysis for Product Repairs, Disease Recurrences, and Other Applications by Wayne B. Nelson. Technometrics, 45(3):263–264, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047053.

Meeker:2009:UAL

[MEH09] William Q. Meeker, Luis A. Escobar, and Yili Hong. Using accelerated life tests results to predict product field reliability. Technometrics, 51(2):146–161, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586592.

McKane:2005:SSN

[MEM05] Scott W. McKane, Luis A. Escobar, and William Q. Meeker. Sample size and number of failure requirements for demonstration tests with log-location-scale distributions and failure censoring. Technometrics, 47(2):182–190, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470980.

Michelson:2001:BRB

[Mic01] Diane K. Michelson. Book review: Applied Statistics for Engineers and Scientists by Jay Devore; Nicholas Farnum. Technometrics, 43(4):487, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270822.

Michelson:2003:BRB

[Mic03] Diane K. Michelson. Book review: Components of Variance by D. R. Cox; P. J. Solomon. Technometrics, 45(4):363, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047095.

Michelson:2005:BRB

[Mic05] Diane K. Michelson. Book review: Analysis of Variance for Random Models, Vol. 1: Balanced Data, Theory, Methods, Applications and Data Analysis by Hardeo Sahai; Mario Miguel Ojeda. Technometrics, 47(2):233, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470986.

Miller:2005:AUF

[Mil05] Arden Miller. The analysis of unreplicated factorial experiments using all possible comparisons. Technometrics, 47(1):51–63, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470934.

Minton:2002:BRB

[Min02] Paul D. Minton. Book review: Statistics: The Exploration and Analysis of Data by Jay Devore; Roxy Peck. Technometrics, 44 (3):297–298, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270510.

Mitchell:2000:ACO

[Mit00] Toby J. Mitchell. An algorithm for the construction of “D-optimal” experimental designs. Technometrics, 42(1):48–54, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271431.

Mitra:2002:BRB

[Mit02] Amit Mitra. Book review: Statistical Quality Control by M. Jeya Chandra. Technometrics, 44(4):397–398, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271549.

Miyakawa:2004:FAM

[Miy04] Masami Miyakawa. [Failure Amplification Method: An Information Maximization Approach to Categorical Response Optimization]: Discussion. Technometrics, 46(1):16–19, February 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470754. See [JW04a].

Morton-Jones:2000:GLS

[MJH00] Tony Morton-Jones and Robin Henderson. Generalized least squares with ignored errors in variables. Technometrics, 42(4): 366–375, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270946.

McGrath:2001:TMD

[ML01] Richard N. McGrath and Dennis K. J. Lin. Testing multiple dispersion effects in unreplicated fractional factorial designs. Technometrics, 43(4):406–414, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270811.

Mee:2009:LER

[MLC09] Robert Mee, Chen-Tuo Liao, and Feng-Shun Chai. Letter to the Editor [with response]. Technometrics, 51(4):475–478, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586656.

Maitra:2000:BRB

[MM00] Ranjan Maitra and Thomas Mathew. Book review: Analysis of Variance in Statistical Image Processing by Ludwik Kurz; M. Hafed Benteftifa. Technometrics, 42(2):212–213, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271465.

Morris:2008:UOA

[MMM08] Max D. Morris, Leslie M. Moore, and Michael D. McKay. Using orthogonal arrays in the sensitivity analysis of computer models. Technometrics, 50(2):205–215, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471460.

Mease:2006:EXT

[MN06] David Mease and Vijayan N. Nair. Extreme (X-) testing with binary data and applications to reliability demonstration. Technometrics, 48(3):399–410, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471211.

Mease:2004:SAM

[MNS04] David Mease, Vijayan N. Nair, and Agus Sudjianto. Selective assembly in manufacturing: Statistical issues and optimal binning strategies. Technometrics, 46(2):165–175, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470802.

Moore:2001:BRB

[Moo01] Terri Moore. Book review: Regression Analysis by Example by Samprit Chatterjee; Ali S. Hadi; Bertram Price. Technometrics, 43(2):236, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271043.

Moore:2003:BRB

[Moo03] Terri L. Moore. Book review: SAS(R) for Linear Models by Ramon Littell; Walter Stroup; Rudolf Freund. Technometrics, 45 (2):172–173, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047018.

Morris:2000:CTL

[Mor00a] Max D. Morris. A class of three-level experimental designs for response surface modeling. Technometrics, 42(2):111–121, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271443.

Morris:2000:TBE

[Mor00b] Max D. Morris. Three “Technometrics” experimental design classics. Technometrics, 42(1):26–27, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271429.

Morgan:2004:BRB

[Mor04] Walter T. Morgan. Book review: Six Sigma and Beyond: Statistical Process Control by D. H. Stamatis. Technometrics, 46 (3):362–363, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470857.

Morgan:2005:BRBa

[Mor05a] J. P. Morgan. Book review: Design and Analysis: A Researcher’s Handbook by Geoffrey Keppel; Thomas D. Wickens. Technometrics, 47(4):522–523, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471088.

Morgan:2005:BRBb

[Mor05b] Walter T. Morgan. Book review: Applied Statistics for the Six Sigma Green Belt by Bhisham C. Gupta; H. Fred Walker. Technometrics, 47(4):523–524, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471089.

Mee:2000:SD

[MP00] Robert W. Mee and Marta Peralta. Semifolding 2^{k−p} designs. Technometrics, 42(2):122–134, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271444.

Mudivarthy:2000:BRB

[MR00] Surekha Mudivarthy and M. Bhaskara Rao. Book review: Model Selection and Inference by Kenneth P. Burnham; David R. Anderson. Technometrics, 42(3):319, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271104.

Mignani:2001:MCM

[MR01] Stefania Mignani and Rodolfo Rosa. Markov chain Monte Carlo in statistical mechanics: the problem of accuracy. Technometrics, 43(3):347–355, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271222.

McKean:2000:PRP

[MS00] Joseph W. McKean and Simon J. Sheather. Partial residual plots based on robust fits. Technometrics, 42(3):249–261, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271080.

Miller:2001:UFR

[MS01] A. Miller and R. R. Sitter. Using the folded-over 12-run Plackett– Burman design to consider interactions. Technometrics, 43(1):44– 55, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1270856.

Miller:2005:UFN

[MS05] A. Miller and R. R. Sitter. Using folded-over nonorthogonal designs. Technometrics, 47(4):502–513, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471075.

Michailidis:2007:BRB

[MS07] George Michailidis and Stilian Stoev. Book review: Extreme Value Theory: An Introduction by Laurens De Haan; Ana Ferreira. Technometrics, 49(4):491–492, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471396.

Merrick:2003:BSA

[MSM03] Jason R. W. Merrick, Refik Soyer, and Thomas A. Mazzuchi. A Bayesian semiparametric analysis of the reliability and maintenance of machine tools. Technometrics, 45(1):58–69, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25046975.

Mulekar:2001:BRB

[Mul01] Madhuri Mulekar. Book review: Statistical Inference in Science by D. A. Sprott. Technometrics, 43(3):376–377, August 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271235.

Mulrow:2002:BRB

[Mul02] Edward J. Mulrow. Book review: The Visual Display of Quantitative Information by Edward R. Tufte. Technometrics, 44(4):400, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271552.

Mulekar:2004:BRB

[Mul04] Madhuri S. Mulekar. Book review: Data Mining: Multimedia, Soft Computing, and Bioinformatics by Sushmita Mitra; Tinku Acharya. Technometrics, 46(3):368–369, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470866.

Mulekar:2008:BRB

[Mul08] Madhuri S. Mulekar. Book review: Weight-of Evidence for Forensic DNA Profiles by David J. Balding. Technometrics, 50(3):407– 408, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471500.

Murty:2007:BRB

[Mur07] Katta G. Murty. Book review: Nonlinear Programming Theory and Algorithms by M. S. Bazaraa; H. D. Sherali; C. M. Shetty. Technometrics, 49(1):105, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471288.

McLeod:2005:PCI

[MV05] A. I. McLeod and E. R. Vingilis. Power computations for intervention analysis. Technometrics, 47(2):174–181, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470979.

Mahmoud:2004:PAL

[MW04] Mahmoud A. Mahmoud and William H. Woodall. Phase I analysis of linear profiles with calibration applications. Technometrics, 46(4):380–391, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470886.

Mathew:2005:GVC

[MW05] Thomas Mathew and David W. Webb. Generalized p values and confidence intervals for variance components: applications to Army test and evaluation. Technometrics, 47(3):312–322, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471024.

Mandal:2006:SSE

[MWJ06] Abhyuday Mandal, C. F. Jeff Wu, and Kjell Johnson. SELC: Sequential elimination of level combinations by means of modified genetic algorithms. Technometrics, 48(2):273–283, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471163.

Mee:2008:SAM

[MX08] Robert W. Mee and Jihua (Regina) Xiao. Steepest ascent for multiple-response applications. Technometrics, 50(3):371–382, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471497.

Maronna:2008:RLR

[MY08] Ricardo A. Maronna and Victor J. Yohai. Robust low-rank approximation of data matrices with elementwise contamination. Technometrics, 50(3):295–304, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471491.

Myers:2000:BRB

[Mye00] Donald E. Myers. Book review: Basic Linear Geostatistics by Margaret Armstrong. Technometrics, 42(4):437, November 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270968.

Myers:2001:BRB

[Mye01] Donald E. Myers. Book review: Practical Geostatistics 2000 by Isobel Clark; William V. Harper. Technometrics, 43(4):492, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270831.

Myers:2005:BRBa

[Mye05a] Donald E. Myers. Book review: Reliability and Statistics in Geotechnical Engineering by Gregory B. Baecher; John T. Christian. Technometrics, 47(1):103–104, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470954.

Myers:2005:BRBb

[Mye05b] Donald E. Myers. Book review: Time Series Analysis and Inverse Theory for Geophysicists by David Gubbins. Technometrics, 47(3):374, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471035.

Myers:2008:BRB

[Mye08] Donald E. Myers. Book review: Linear and Generalized Linear Mixed Models and Their Applications by Jiming Jiang. Technometrics, 50(1):93–94, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471440.

Myers:2009:BRBc

[Mye09a] Donald E. Myers. Book review: Applied Econometrics With R by Christian Klieber; Achim Zeileis. Technometrics, 51(4):484– 485, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586665.

Myers:2009:BRBa

[Mye09b] Donald E. Myers. Book review: Sampling Techniques for Forest Inventories by Daniel Mandallaz. Technometrics, 51(2):222–223, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586609.

Myers:2009:BRBb

[Mye09c] Donald E. Myers. Book review: Statistical Methods for Environmental Epidemiology With R: A Case Study in Air Pollution and Health by Roger D. Peng; Francesca Dominici. Technometrics, 51(3):344–346, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586636.

Maronna:2002:REL

[MZ02] Ricardo A. Maronna and Ruben H. Zamar. Robust estimates of location and dispersion for high-dimensional datasets. Technometrics, 44(4):307–317, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271538.

Nagaraja:2004:BRB

[Nag04] H. N. Nagaraja. Book review: Measures of Interobserver Agreement by Mohamed M. Shoukri. Technometrics, 46(3):366–367, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470864.

Nagaraja:2006:BRB

[Nag06] Haikady N. Nagaraja. Book review: Inference in Hidden Markov Models by Olivier Cappé; Eric Moulines; Tobias Rydén. Technometrics, 48(4):574–575, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471260.

Nair:2007:SIS

[Nai07] Vijay Nair. Special issue on statistics in information technology. Technometrics, 49(3):236, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471344.

Nash:2001:BRB

[Nas01] Maliha S. Nash. Book review: Handbook of Parametric and Nonparametric Statistical Procedures by David J. Sheskin. Technometrics, 43(3):374, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271232.

Nash:2002:BRB

[Nas02] Maliha S. Nash. Book review: Practical Time-Frequency Analysis, Gabor and Wavelet Transforms with an Implementation in S by René Carmona; Wen-Liang Hwang; Bruno Torresani. Technometrics, 44(2):196, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271277.

Nash:2004:BRB

[Nas04] Maliha S. Nash. Book review: Spatial Statistics and Computational Methods by Jesper Møller. Technometrics, 46(1):115–116, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470776.

Nash:2008:BRB

[Nas08] Maliha S. Nash. Book review: Data Preparation for Analytics Using SAS by Gerhard Svolba. Technometrics, 50(1):92, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471437.

Nguyen:2008:NOS

[NC08] Nam-Ky Nguyen and Ching-Shui Cheng. New E(s^{2})-optimal supersaturated designs constructed from incomplete block designs. Technometrics, 50(1):26–31, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471422.

Ng:2004:OPC

[NCB04] H. K. T. Ng, P. S. Chan, and N. Balakrishnan. Optimal progressive censoring plans for the Weibull distribution. Technometrics, 46(4):470–481, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470895.

Nelson:2002:EAM

[ND02] Peter R. Nelson and Edward J. Dudewicz. Exact analysis of means with unequal variances. Technometrics, 44(2):152–160, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271259.

Neal:2007:BRB

[Nea07] Radford M. Neal. Book review: Pattern Recognition and Machine Learning by Christopher M. Bishop. Technometrics, 49(3):366, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471371.

Nelson:2000:BRB

[Nel00a] Peter R. Nelson. Book review: Relating Statistics and Experimental Design by Irwin P. Levin. Technometrics, 42(3):308–309, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271086.

Nelson:2000:TAH

[Nel00b] Wayne Nelson. Theory and applications of hazard plotting for censored failure data. Technometrics, 42(1):12–25, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271428.

Nelson:2001:BRB

[Nel01] Peter R. Nelson. Book review: Normality and the Process Behavior Chart by Donald J. Wheeler. Technometrics, 43(3):371, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271227.

Nemeth:2000:BRB

[Nem00] Margaret A. Nemeth. Book review: Applied Multivariate Methods for Data Analysis by Dallas E. Johnson. Technometrics, 42(2):211, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271463.

Nemeth:2002:BRBb

[Nem02a] Margaret A. Nemeth. Book review: A Primer for Sampling Solids, Liquids, and Gases by Patricia L. Smith. Technometrics, 44(2):190, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271269.

Nemeth:2002:BRBa

[Nem02b] Margaret A. Nemeth. Book review: Design and Analysis in Chemical Research by Roy L. Tranter. Technometrics, 44(2):190, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271268.

Nemeth:2003:BRB

[Nem03] Margaret A. Nemeth. Book review: Multi- and Megavariate Data Analysis by L. Eriksson; E. Johansson; N. Kettaneh-Wold; S. Wold. Technometrics, 45(4):362, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047093.

Nembhard:2004:BRB

[Nem04] Harriet Black Nembhard. Book review: Applying Contemporary Statistical Techniques by Rand R. Wilcox. Technometrics, 46 (4):490, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470906.

Nembhard:2008:BRB

[Nem08] Harriet Black Nembhard. Book review: Statistical Methods for Dose-Finding Experiments by Sylvie Chevret. Technometrics, 50 (3):409, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471504.

Neubauer:2001:BRB

[Neu01] Dean V. Neubauer. Book review: Statistical Procedures for Machine and Process Qualification by Edgar Dietrich; Alfred Schulze. Technometrics, 43(3):372–373, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271229.

Neubauer:2003:BRB

[Neu03] Dean V. Neubauer. Book review: Statistical Process Adjustment for Quality Control by Enrique del Castillo. Technometrics, 45 (1):102–103, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046984.

Neubauer:2005:BRB

[Neu05] Dean V. Neubauer. Book review: Statistics for the Sciences by Martin Buntinas; Gerald M. Funk. Technometrics, 47(3):377– 378, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471042.

Neubauer:2008:BRBb

[Neu08a] Dean V. Neubauer. Book review: Chemical Process Performance Evaluation by Ali Cinar; Ahmet Palazoglu; Ferhan Kayihan. Technometrics, 50(1):95–96, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471443.

Neubauer:2008:BRBc

[Neu08b] Dean V. Neubauer. Book review: Practical PID Control by Antonio Visioli. Technometrics, 50(2):232–233, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471465.

Neubauer:2008:BRBa

[Neu08c] Dean V. Neubauer. Book review: Statistical Design of Experiments with Engineering Applications by Kamel Rekab; Muzaffar Shaikh. Technometrics, 50(1):90–91, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471434.

Neubauer:2009:BRBb

[Neu09a] Dean V. Neubauer. Book review: Chance Rules: An Informal Guide to Probablity, Risk and Statistics by Brian Everitt. Technometrics, 51(4):485–487, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586667.

Neubauer:2009:BRBa

[Neu09b] Dean V. Neubauer. Book review: Missing Data: A Gentle Introduction by Patrick E. McKnight; Katherine M. McKnight; Souraya Sidani; Aurelio José Flgueredo. Technometrics, 51(1):101–103, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586573.

Ng:2005:BRB

[Ng05] Hon Keung Tony Ng. Book review: Computational Methods in Statistics and Econometrics by Hisashi Tanizaki. Technometrics, 47(4):521–522, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471086.

Ng:2006:BRB

[Ng06] H. K. Tony Ng. Book review: Statistics: An Introduction Using R by Michael J. Crawley. Technometrics, 48(2):308, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471175.

Ng:2008:BRB

[Ng08] H. K. Tony Ng. Book review: Life Cycle Reliability Engineering by Guangbin Yang. Technometrics, 50(1):94–95, February 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471442.

Ng:2009:BRB

[Ng09] Hon Keung Tony Ng. Book review: Probability and Statistics With R by Maria Dolores Ugarte; Ana F. Militino; Alan T. Arnholt. Technometrics, 51(2):220, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586606.

Nembhard:2003:AFB

[NK03] Harriet Black Nembhard and Ming Shu Kao. Adaptive forecastbased monitoring for dynamic systems. Technometrics, 45(3):208– 219, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047048.

Nkurunziza:2007:BRB

[Nku07] Sévérien Nkurunziza. Book review: Fundamentals of Actuarial Mathematics by S. David Promislow. Technometrics, 49(3): 363, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471365.

Nordman:2002:WPI

[NM02] Daniel J. Nordman and William Q. Meeker. Weibull prediction intervals for a future number of failures. Technometrics, 44(1):15–23, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270680.

Natarajan:2004:MBG

[NM04] Rajeshwari Natarajan and Govind S. Mudholkar. Moment-based goodness-of-fit tests for the inverse Gaussian distribution. Technometrics, 46(3):339–347, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470848.

Notz:2002:ER

[Not02] William Notz. Editor’s report. Technometrics, 44(4):305–306, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271537.

Notz:2003:EA

[Not03a] William Notz. Editorial announcement. Technometrics, 45(4):283, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047084.

Notz:2003:ER

[Not03b] William Notz. Editor’s report. Technometrics, 45(4):281–282, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047083.

Notz:2004:ER

[Not04] William Notz. Editor’s report. Technometrics, 46(4):377–379, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470885.

Nair:2004:FAM

[NW04] Vijayan N. Nair and Xiao Wang. [Failure Amplification Method: An Information Maximization Approach to Categorical Response Optimization]: Discussion. Technometrics, 46(1):19–23, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470755. See

[JW04a].

Oakley:2009:DTS

[Oak09] Jeremy E. Oakley. Decision-theoretic sensitivity analysis for complex computer models. Technometrics, 51(2):121–129, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586590.

Ojelund:2002:PBM

[ÖBMT02] H. Öjelund, P. J. Brown, H. Madsen, and P. Thyregod. Prediction based on mean subset. Technometrics, 44(4):369–378, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271544.

Odencrantz:2000:BRBb

[Ode00a] John Odencrantz. Book review: Markov Chains: Gibbs Fields, Monte Carlo Simulation, and Queues by P. Brémaud. Technometrics, 42(4):438–439, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270970.

Odencrantz:2000:BRBa

[Ode00b] John Odencrantz. Book review: Nonparametric Statistics for Stochastic Processes: Estimation and Prediction by D. Bosq. Technometrics, 42(4):429–430, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270958.

Olive:2005:VSR

[OH05] David J. Olive and Douglas M. Hawkins. Variable selection for 1D regression models. Technometrics, 47(1):43–50, February 2005.

CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470933.

Oldford:2008:FSC

[Old08] R. Wayne Oldford. [The Future of Statistical Computing]: Comment. Technometrics, 50(4):437–439, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471522.

Olive:2002:ARD

[Oli02] David J. Olive. Applications of robust distances for regression. Technometrics, 44(1):64–71, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270685.

Olive:2003:BRB

[Oli03] David J. Olive. Book review: Linear Regression Analysis by George A. F. Seber; Alan J. Lee. Technometrics, 45(4):362–363, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047094.

Olive:2004:BRB

[Oli04] David J. Olive. Book review: Introduction to Applied Statistics: A Modelling Approach by J. K. Lindsey. Technometrics, 46 (4):492, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470910.

Olive:2005:BRB

[Oli05] David J. Olive. Book review: Introduction to Regression Analysis by M. A. Golberg; H. A. Cho. Technometrics, 47(2):233–234, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470988.

Olive:2007:BRB

[Oli07] David J. Olive. Book review: Robust Statistical Methods with R by J. Jureä ová; J. Picek. Technometrics, 49(4):496, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471404.

Olive:2008:BRB

[Oli08] David J. Olive. Book review: Beyond Beta: Other Continuous Families of Distributions with Bounded Support and Applications by S. Kotz; J. R. van Dorp. Technometrics, 50(2):236, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471469.

Olive:2009:BRBa

[Oli09a] David J. Olive. Book review: Asymptotic Theory of Statistics and Probability by A. DasGupta. Technometrics, 51(2):218–219, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586603.

Olive:2009:BRBc

[Oli09b] David J. Olive. Book review: Model Selection and Model Averaging by G. Claeskens; N. L. Hjort. Technometrics, 51(3):343–344, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586635.

Olive:2009:BRBb

[Oli09c] David J. Olive. Book review: Observed Confidence Levels by A. M. Polansey. Technometrics, 51(2):224, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586611.

Olive:2009:BED

[Oli09d] David J. Olive. Testing 1–2–3: Experimental Design With Applications in Marketing and Service Operations by Johannes Ledolter; Arthur J. Swersey. Technometrics, 51(2):221–222, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586608.

Olwell:2001:BRBa

[Olw01a] David Olwell. Book review: Reliability Engineering and Risk Analysis by Mohammad Modarres; Mark Kaminskiy; Vasiliy Krivtsov. Technometrics, 43(1):104–105, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270871.

Olwell:2001:BRBb

[Olw01b] David Olwell. Book review: Reliability Modeling, Prediction, and Optimization by Wallace R. Blischke; D. N. Prabhakar Murthy. Technometrics, 43(4):486–487, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270821.

ONeill:2001:BRB

[O’N01] Julia O’Neill. Book review: Introduction to the Design and Analysis of Experiments by Geoffrey M. Clarke; Robert E. Kempson. Technometrics, 43(1):98–99, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270863.

ONeill:2003:BRB

[O’N03] Julia O’Neill. Book review: Box on Quality and Discovery with Design, Control, and Robustness by George C. Tiao; Søren Bisgaard; William J. Hill; Daniel Peña; Stephen M. Stigler. Technometrics, 45(1):105, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046986.

ONeill:2005:BRB

[O’N05] Julia C. O’Neill. Book review: Statistical Techniques for Data Analysis by John K. Taylor; Cheryl Cihon. Technometrics, 47(3): 379, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471045.

Ordaz:2007:BRBa

[Ord07a] José A. Ordaz. Book review: Advances in Economics and Econometrics: Theory and Applications, Ninth World Congress, Vol. I by Richard Blundell; Whitney K. Newey; Torsten Persson. Technometrics, 49(3):366–367, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471372.

Ordaz:2007:BRBb

[Ord07b] José A. Ordaz. Book review: Advances in Economics and Econometrics: Theory and Applications, Ninth World Congress, Vol. II by Richard Blundell; Whitney K. Newey; Torsten Persson. Technometrics, 49(4):499–500, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471408.

Owen:2000:BRB

[Owe00] W. Jason Owen. Book review: Statistical Data Analysis by Glen Cowan. Technometrics, 42(3):311–312, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271091.

Owen:2003:BRB

[Owe03] William J. Owen. Book review: Elements of Computational Statistics by James E. Gentle. Technometrics, 45(3):268–269, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047058.

Owen:2004:BRB

[Owe04] William J. Owen. Book review: Modeling the Internet and the Web by Pierre Baldi; Paolo Fransconi; Padhraic Smyth. Technometrics, 46(3):369, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470867.

Owen:2006:BRB

[Owe06] William Jason Owen. Book review: Computational Statistics by Geof H. Givens; Jennifer A. Hoeting. Technometrics, 48(2):309– 310, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471177.

Pan:2000:BRB

[Pan00] Guohua Pan. Book review: Local Regression and Likelihood by Clive Loader. Technometrics, 42(4):428–429, November 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270956.

Pascual:2003:TOT

[Pas03] Francis G. Pascual. Theory for optimal test plans for the random fatigue-limit model. Technometrics, 45(2):130–141, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047010.

Pascual:2006:ALT

[Pas06] Francis G. Pascual. Accelerated life test plans robust to misspecification of the stress-life relationship. Technometrics, 48(1):11– 25, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471111.

Priebe:2001:SSD

[PC01] Carey E. Priebe and Dalei Chen. Spatial scan density estimates. Technometrics, 43(1):73–83, February 2001. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270859.

Peterson:2004:LE

[PCdC04] John J. Peterson, Suntara Cahya, and Enrique del Castillo. Letter to the Editor. Technometrics, 46(3):355–357, August 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470850.

Prescott:2002:MEI

[PDDL02] P. Prescott, A. M. Dean, N. R. Draper, and S. M. Lewis. Mixture experiments: ILL-conditioning and quadratic model specification. Technometrics, 44(3):260–268, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270490.

Pederson:2001:BRB

[Ped01] Shane P. Pederson. Book review: Probabilistic Networks and Expert Systems by Robert G. Cowell; A. Philip Dawid; Steffen L. Lauritzen; David J. Spiegelhalter. Technometrics, 43(1):108–109, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270879.

Pena:2005:NSI

[Peñ05] Daniel Peña. A new statistic for influence in linear regression. Technometrics, 47(1):1–12, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470929.

Percival:2004:BRB

[Per04] Donald B. Percival. Book review: Probability, Statistical Optics, and Data Testing: A Problem Solving Approach by B. R. Frieden. Technometrics, 46(1):113–114, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470773.

Peterson:2003:BRB

[Pet03a] John J. Peterson. Book review: Statistical Group Comparison by Tim Futing Liao. Technometrics, 45(2):175, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047023.

Peterson:2003:LE

[Pet03b] John J. Peterson. Letter to the Editor. Technometrics, 45(2):185, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25047042.

Pfahler:2004:BRB

[Pfa04] Lori B. Pfahler. Book review: Range Based Analysis of Means by Donald J. Wheeler. Technometrics, 46(3):364, August 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470860.

Piepel:2002:MAC

[PHSL02] Greg F. Piepel, Ruel D. Hicks, Jeff M. Szychowski, and Jason L. Loeppky. Methods for assessing curvature and interaction in mixture experiments. Technometrics, 44(2):161–172, May 2002. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271260.

Picard:2005:ISS

[Pic05] Richard R. Picard. Importance sampling for simulation of Markovian physical processes. Technometrics, 47(2):202–211, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470982.

Picka:2006:BRB

[Pic06] Jeffrey D. Picka. Book review: Gaussian Markov Random Fields: Theory and Applications by Havard Rue; Leonhard Held. Technometrics, 48(1):146–147, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471126.

Picka:2007:BRB

[Pic07] Jeffrey D. Picka. Book review: Data Science and Classification by V. Batagelj; H.-H. Bock; A. Ferligoj; A. Žiberna. Technometrics, 49(3):363–364, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471366.

Picka:2009:BRB

[Pic09] Jeffrey Picka. Book review: Introduction to Probability With R by K. Baclawski. Technometrics, 51(2):221, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586607.

Pigeon:2006:BRB

[Pig06] Joseph G. Pigeon. Book review: Statistics for Experimenters: Design, Innovation and Discovery by George E. P. Box; J. Stuart Hunter; William G. Hunter. Technometrics, 48(2):303–304, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471166.

Peterson:2005:RAN

[PK05] John J. Peterson and Andrew M. Kuhn. Ridge analysis with noise variables. Technometrics, 47(3):274–283, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471021.

Parker:2007:CBE

[PKV07] Peter A. Parker, Scott M. Kowalski, and G. Geoffrey Vining. Construction of balanced equivalent estimation second-order split-plot designs. Technometrics, 49(1):56–65, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471275.

Placzkowski:2001:BRB

[Pla01] Gene Placzkowski. Book review: Statistical Process Control in Industry by R. J. M. M. Does; K. C. B. Roes; A. Trip. Technometrics, 43(2):232, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271038.

Placzkowski:2002:BRB

[Pla02] Gene Placzkowski. Book review: Running Your Machines with SPC by James C. Abbott. Technometrics, 44(1):90, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270703.

Pascual:2003:MRT

[PM03] Francis G. Pascual and Grace Montepiedra. Model-robust test plans with applications in accelerated life testing. Technometrics, 45(1):47–57, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046974.

Polansky:2001:SNA

[Pol01] Alan M. Polansky. A smooth nonparametric approach to multivariate process capability. Technometrics, 43(2):199–211, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271033. Statisti-

cal methodology for industrial problems.

Portier:2002:BRB

[Por02] Kenneth M. Portier. Book review: Robust Diagnostic Regression Analysis by Anthony Atkinson; Marco Riani. Technometrics, 44(1):81, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270687.

Pena:2001:MODa

[PP01a] Daniel Peña and Francisco J. Prieto. Multivariate outlier detection and robust covariance matrix estimation. Technometrics, 43 (3):286–300, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1271215. See discussion [RW01, Hub01] and response [PP01b].

Pena:2001:MODb

[PP01b] Daniel Peña and Francisco J. Prieto. [Multivariate Outlier Detection and Robust Covariance Matrix Estimation]: Response. Technometrics, 43(3):306–310, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271218. See [PP01a].

Proschan:2000:TEO

[Pro00] Frank Proschan. Theoretical explanation of observed decreasing failure rate. Technometrics, 42(1):7–11, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271427.

Prybutok:2003:BRB

[Pry03] Victor R. Prybutok. Book review: Statistical Rules of Thumb by Gerald van Belle. Technometrics, 45(1):107, February 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25046990.

Puig:2000:TFL

[PS00] Pedro Puig and Michael A. Stephens. Tests of fit for the Laplace distribution, with applications. Technometrics, 42(4):417–424, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270952.

Pronzato:2001:ESL

[PW01] Luc Pronzato and Éric Walter. Eliminating suboptimal local minimizers in nonlinear parameter estimation. Technometrics, 43(4): 434–442, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270814.

Pardoe:2007:GTQ

[PYC07] Iain Pardoe, Xiangrong Yin, and R. Dennis Cook. Graphical tools for quadratic discriminant analysis. Technometrics, 49(2):172–183, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471310.

Quesada:2004:DRA

[QD04] Guillermo Miró Quesada and Enrique Del Castillo. A dual-response approach to the multivariate robust parameter design problem. Technometrics, 46(2):176–187, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470803.

Qiu:2001:RBM

[QH01] Peihua Qiu and Douglas Hawkins. A rank-based multivariate CUSUM procedure. Technometrics, 43(2):120–132, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271026. Statis-

tical methodology for industrial problems.

Qian:2000:BRB

[Qia00] Lianfen Qian. Book review: Nonparametric Curve Estimation: Methods, Theory, and Applications by Sam Efromovich. Technometrics, 42(4):429, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270957.

Qian:2001:BRB

[Qia01] Lianfen Qian. Book review: Probability Applications in Mechanical Design by Franklin E. Fisher; Joy R. Fisher. Technometrics, 43(4):490, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270827.

Qian:2002:BRB

[Qia02] Lianfen Qian. Book review: Regularized Radial Basis Function Networks: Theory and Applications by Paul V. Yee; Simon Haykin. Technometrics, 44(3):294, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270505.

Qian:2005:BRB

[Qia05] Lianfen Qian. Book review: Statistical Curves and Parameters: Choosing an Appropriate Approach by Michael E. Tarter. Technometrics, 47(1):96–97, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470943.

Qian:2007:BRB

[Qia07] Lianfen Qian. Book review: Simulation Techniques in Financial Risk Management by Ngai Hang Chan; Hoi Ying Wong. Technometrics, 49(2):222, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471315.

Qiu:2000:BRB

[Qiu00] Peihua Qiu. Book review: Multivariate Image Analysis by Paul Geladi; Hans Grahn. Technometrics, 42(2):211–212, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271464.

Qiu:2002:BRB

[Qiu02] Peihua Qiu. Book review: Statistical Analysis of Microstructures in Materials Science by Joachin Ohser; Frank Mücklich. Technometrics, 44(1):86, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270695.

Qiu:2004:LPL

[Qiu04] Peihua Qiu. The local piecewisely linear kernel smoothing procedure for fitting jump regression surfaces. Technometrics, 46 (1):87–98, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470764.

Qiu:2005:BRB

[Qiu05] Peihua Qiu. Book review: Generalized Least Squares by Takeaki Kariya; Hiroshi Kurata. Technometrics, 47(4):519, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471082.

Qiu:2008:BRB

[Qiu08] Peihua Qiu. Book review: Fuzzy Modeling and Fuzzy Control by Huaguang Zhang; Derong Liu. Technometrics, 50(3):408–409, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471502.

Qu:2007:BRBb

[Qu07a] Xianggui Qu. Book review: Linear Model Theory: Univariate, Multivariate, and Mixed Models by Keith E. Muller; Paul W. Stewart. Technometrics, 49(2):223, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471318.

Qu:2007:BRBa

[Qu07b] Xianggui Qu. Book review: Multivariate Data Analysis by Joseph F. Hair; William C. Black; Barry J. Babin; Rolph E. Anderson; Ronald L. Tatham. Technometrics, 49(1):103–104, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471285.

Qu:2008:BRB

[Qu08] Xianggui Qu. Book review: The Statistics of Gene Mapping by David Siegmund; Benjamin Yakir. Technometrics, 50(1):94, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471441.

Qian:2008:BHM

[QW08] Peter Z. G. Qian and C. F. Jeff Wu. Bayesian hierarchical modeling for integrating low-accuracy and high-accuracy experiments.

Technometrics, 50(2):192–204, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471459.

Qian:2008:GPM

[QWW08] Peter Z. G. Qian, Huaiqing Wu, and C. F. Jeff Wu. Gaussian process models for computer experiments with qualitative and quantitative factors. Technometrics, 50(3):383–396, August 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471498.

Riani:2000:RDD

[RA00] Marco Riani and Anthony C. Atkinson. Robust diagnostic data analysis: Transformations in regression. Technometrics, 42(4): 384–394, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/1270948. See discussion [Coo00b].

Robinson:2009:BAS

[RACH09] Timothy J. Robinson, Christine M. Anderson-Cook, and Michael S. Hamada. Bayesian analysis of split-plot experiments with nonnormal responses for evaluating nonstandard performance criteria. Technometrics, 51(1):56–65, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586563.

Raghavachari:2001:BRB

[Rag01] Madabhushi Raghavachari. Book review: Applied Multivariate Statistics in Geohydrology and Related Sciences by Charles E. Brown. Technometrics, 43(1):110, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270881.

Ramirez:2000:BRB

[Ram00a] José G. Ramírez. Book review: Data Analysis: Statistical and Computational Methods for Scientists and Engineers by Siegmund Brandt; Glen Cowan. Technometrics, 42(3):312, August 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271092.

Ramsey:2000:BRB

[Ram00b] Philip H. Ramsey. Book review: Nonparametric Statistical Methods by Myles Hollander; Douglas A. Wolfe. Technometrics, 42(2):217– 218, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271473.

Randles:2003:BRB

[Ran03] Ronald H. Randles. Book review: Applied Nonparametric Statistical Methods by P. Sprent; N. C. Smeeton. Technometrics, 45 (4):369, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25047105.

Rao:2000:BRB

[Rao00] M. B. Rao. Book review: Feedforward Neural Network Methodology by Terrence L. Fine. Technometrics, 42(4):432–433, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270962.

Rao:2001:BRB

[Rao01] M. Bhaskara Rao. Book review: Conditional Specification of Statistical Models by Barry C. Arnold; Enrique Castillo; Jose Maria Sarabia. Technometrics, 43(4):490, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270826.

Ravishanker:2004:BRB

[Rav04] Nalini Ravishanker. Book review: Survival Analysis by John P. Klein; L. Moeschberger. Technometrics, 46(1):111–112, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470769.

Rayens:2000:BRB

[Ray00] Bill Rayens. Book review: An Introduction to Copulas by Roger B. Nelsen. Technometrics, 42(3):317, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271100.

Ray:2003:BRB

[Ray03a] Bonnie K. Ray. Book review: Regression Models for Time Series Analysis by Benjamin Kedem; Konstantinos Fokianos. Technometrics, 45(4):364, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047097.

Rayens:2003:BRB

[Ray03b] William S. Rayens. Book review: Independent Component Analysis: Principles and Practice by Stephen Roberts; Richard Everson. Technometrics, 45(1):107–108, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25046992.

Ranjan:2008:SED

[RBM08] Pritam Ranjan, Derek Bingham, and George Michailidis. Sequential experiment design for contour estimation from complex computer codes. Technometrics, 50(4):527–541, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471535. See er-

rata [RBM11].

Ranjan:2011:ESE

[RBM11] P. Ranjan, D. Bingham, and G. Michailidis. Errata: “Sequential experiment design for contour estimation from complex computer codes” [MR2655651]. Technometrics, 53(1):109–110, February 2011. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40997315. See

[RBM08].

Ramdani:2006:SAD

[RCGD06] Nacim Ramdani, Yves Candau, Gilles Guyon, and Christophe Dalibart. Sensitivity analysis of dynamic models to uncertainties in inputs data with time-varying variances. Technometrics, 48(1):74–87, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471116.

Rajagopal:2005:MRP

[RdC05] Ramkumar Rajagopal and Enrique del Castillo. Model-robust process optimization using Bayesian model averaging. Technometrics, 47(2):152–163, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470977.

Ringrose:2005:SMS

[RF05] Trevor J. Ringrose and Shaun A. Forth. Simplifying multivariate second-order response surfaces by fitting constrained models using automatic differentiation. Technometrics, 47(3):249–259, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471018.

Reilly:2007:WCV

[RG07] Cavan Reilly and Andrew Gelman. Weighted classical variogram estimation for data with clustering. Technometrics, 49(2):184–194, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471311.

Remlinger:2006:SDP

[RHOYL06] Katja S. Remlinger, Jacqueline M. Hughes-Oliver, S. Stanley Young, and Raymond L. Lam. Statistical design of pools using optimal coverage and minimal collision. Technometrics, 48 (1):133–143, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471122.

Ray:2006:DRM

[RLR06] Bonnie K. Ray, Zhaohui Liu, and Nalini Ravishanker. Dynamic reliability models for software using time-dependent covariates. Technometrics, 48(1):1–10, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471110.

Rougier:2009:EKM

[RMGR09] Jonathan Rougier, Astrid Maute, Serge Guillas, and Arthur D. Richmond. Expert knowledge and multivariate emulation: The Thermosphere-lonosphere Electrodynamics General Circulation Model (TIE-GCM). Technometrics, 51(4):414–424, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/40586651.

Roberts:2000:CCT

[Rob00a] S. W. Roberts. Control chart tests based on geometric moving averages. Technometrics, 42(1):97–101, February 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271439.

Robertson:2000:BRB

[Rob00b] Roy Robertson. Book review: An Introduction to Statistical Modelling by Wojtek J. Krzanowski. Technometrics, 42(4):435–436, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270965.

Robinson:2002:BRB

[Rob02] Timothy Robinson. Book review: Quality Improvement with Design of Experiments by Ivan N. Vuchkov; Lidia N. Boyadjieva. Technometrics, 44(2):188–189, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271266.

Robinson:2005:BRB

[Rob05] Timothy J. Robinson. Book review: Linear Models with R by Julian J. Faraway. Technometrics, 47(3):372–373, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471031.

Robinson:2006:BRB

[Rob06] Timothy Robinson. Book review: Introduction to Linear Models and Statistical Inference by Steven J. Janke; Frederick C. Tinsley. Technometrics, 48(4):573–574, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471258.

Romeu:2003:BRB

[Rom03] Jorge L. Romeu. Book review: Practical Reliability Engineering by Patrick D. T. O’Connor. Technometrics, 45(2):173, May 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047019.

Romeu:2006:BRB

[Rom06] Jorge L. Romeu. Book review: Statistical Methods for Communication Science by Andrew F. Hayes. Technometrics, 48(3):448, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471221.

Rose:2004:BRB

[Ros04] Nicholas Rose. Book review: Spatial Cluster Modelling by Andrew B. Lawson; David G. T. Denison. Technometrics, 46(1):116, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470777.

Rose:2005:BRB

[Ros05] Nicholas Rose. Book review: Highly Structured Stochastic Systems by Peter J. Green; Nils Lid Hjort; Sylvia Richardson. Technometrics, 47(1):101, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470951.

Reynolds:2004:CCE

[RS04] Marion R. Reynolds, Jr. and Zachary G. Stoumbos. Control charts and the efficient allocation of sampling resources. Technometrics, 46(2):200–214, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470805.

Reynolds:2005:SEW

[RS05] Marion R. Reynolds, Jr. and Zachary G. Stoumbos. Should exponentially weighted moving average and cumulative sum charts be used with Shewhart limits? Technometrics, 47(4):409–424, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471066.

Reynolds:2006:CSE

[RS06] Marion R. Reynolds, Jr. and Zachary G. Stoumbos. Comparisons of some exponentially weighted moving average control charts for monitoring the process mean and variance. Technometrics, 48 (4):550–567, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471247.

Reich:2009:VSB

[RSB09] Brian J. Reich, Curtis B. Storlie, and Howard D. Bondell. Variable selection in Bayesian smoothing spline ANOVA models: Application to deterministic computer codes. Technometrics, 51(2): 110–120, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586589.

Reichert:2002:EST

[RSS02] Peter Reichert, Mark Schervish, and Mitchell J. Small. An efficient sampling technique for Bayesian inference with computationally demanding models. Technometrics, 44(4):318–327, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271539.

Rukhin:2001:BRB

[Ruk01] Andrew L. Rukhin. Book review: Continuous-Time Markov Chains and Applications: A Singular Perturbation Approach by G. George Yin; Qing Zhang. Technometrics, 43(2):240, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271050.

Rukhin:2002:BRB

[Ruk02] Andrew L. Rukhin. Book review: Analysis of Time Series Structure SSA and Related Techniques by Nina E. Golyandina; Vladimir V. Nekrutkin; Anantoly A. Zhigljavsky. Technometrics, 44(3):290, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270499.

Rumpf:2000:BRB

[Rum00] David L. Rumpf. Book review: Your Statistical Consultant by Rae R. Newton; Kjell Erik Rudestam. Technometrics, 42(3):312–313, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271093.

Rumpf:2004:BRB

[Rum04] David L. Rumpf. Book review: Statistics for Dummies by Deborah Rumsey. Technometrics, 46(3):366, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470863.

Rumpf:2006:BRB

[Rum06] David Rumpf. Book review: The Six Sigma Practitioner’s Guide to Data Analysis by Donald J. Wheeler. Technometrics, 48(2):306, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471171.

Rutledge:2000:BRB

[Rut00] Jim Rutledge. Book review: An Introduction to Design of Experiments: A Simplified Approach by Larry B. Barrentine. Technometrics, 42(3):309, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271087.

Rutherford:2002:BRB

[Rut02] Jim Rutherford. Book review: Practical Experiment Designs for Engineers and Scientists by William J. Diamond. Technometrics, 44(4):400–401, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271553.

Rutherford:2005:BRB

[Rut05] Jim Rutherford. Book review: Planning, Construction, and Statistical Analysis of Comparative Experiments by Francis G. Giesbrecht; Marcia L. Gumpertz. Technometrics, 47(1):93, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470937.

Rutherford:2009:BRB

[Rut09] Jim Rutherford. Book review: Linear Models in Statistics (2nd ed.) by Alvin C. Rencher; G. Bruce Schaalje. Technometrics, 51(2):217–218, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586601.

Rousseeuw:2004:RMR

[RVVA04] Peter J. Rousseeuw, Stefan Van Aelst, Katrien Van Driessen, and Jose Agulló. Robust multivariate regression. Technometrics, 46 (3):293–305, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470844.

Rocke:2001:MOD

[RW01] David M. Rocke and David L. Woodruff. [Multivariate Outlier Detection and Robust Covariance Matrix Estimation]: Discussion. Technometrics, 43(3):300–303, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271216. See [PP01a].

Reese:2004:IAC

[RWH^{+}04] C. Shane Reese, Alyson G. Wilson, Michael Hamada, Harry F. Martz, and Kenneth J. Ryan. Integrated analysis of computer and physical experiments. Technometrics, 46(2):153–164, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470801.

Ruan:2004:SMM

[RWZF04] Shiling Ruan, Douglas A. Wolfe, Weilong Zhang, and Gerald S. Frankel. Statistical modeling of minimum intergranular corrosion path length in high-strength aluminum alloy. Technometrics, 46 (1):69–75, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470762.

Ryan:2008:LER

[Rya08] Thomas P. Ryan. Letter to the Editor response to Easterling’s review of “Modern Experimental Design”. Technometrics, 50(2):231, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471464.

Shan:2008:BIM

[SA08] Xuemei Shan and Daniel W. Apley. Blind identification of manufacturing variation patterns by combining source separation criteria. Technometrics, 50(3):332–343, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471494.

Sandry:2003:BRBb

[San03a] Thomas D. Sandry. Book review: Introductory Statistics with R by Peter Dalgaard. Technometrics, 45(3):274–275, August 2003. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25047066.

Sandry:2003:BRBa

[San03b] Thomas D. Sandry. Book review: Longitudinal Data and SAS(R): A Programmer’s Guide by Ron Cody. Technometrics, 45(2):173– 174, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25047020.

Sandry:2005:BRB

[San05] Thomas D. Sandry. Book review: Applied Data Mining by Paolo Giudici. Technometrics, 47(1):102–103, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470953.

Sanchez:2006:RED

[Sán06] Ismael Sánchez. Recursive estimation of dynamic models using Cook’s distance, with application to wind energy forecast. Technometrics, 48(1):61–73, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471115.

Sandry:2007:BRB

[San07] Thomas D. Sandry. Book review: Probabilistic and Randomized Methods for Design under Uncertainty by Giuseppe Calafiore; Fabrizio Dabbene. Technometrics, 49(2):227, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471325.

Sandry:2008:BRBb

[San08a] Thomas D. Sandry. Book review: Probability and Statistics Applications for Environmental Science by Stacey J. Shaefer; Louis Theodore. Technometrics, 50(4):543–545, November 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471538.

Sandry:2008:BRBa

[San08b] Thomas D. Sandry. Book review: Uncertainty Modeling and Analysis in Engineering and the Sciences by Bilal M. Ayyub; George J. Klir. Technometrics, 50(2):233–234, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471466.

Sarkar:2000:BRB

[Sar00a] Pradipta Sarkar. Book review: Elements of Forecasting by Francis X. Diebold. Technometrics, 42(3):318–319, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271103.

Sartori:2000:BRB

[Sar00b] David Sartori. Book review: Experimental Design Techniques in Statistical Practice: A Practical Software-Based Approach by William P. Gardiner; George Gettinby. Technometrics, 42(2):209, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271459.

Sarkar:2002:BRB

[Sar02] Pradipta Sarkar. Book review: Practical Time Series by Gareth Janacek. Technometrics, 44(2):195–196, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271276.

Sarkar:2003:BRB

[Sar03] Pradipta Sarkar. Book review: Sequential Monte Carlo Methods in Practice by Arnaud Doucet; Nando de Freitas; Neil Gordon. Technometrics, 45(1):106, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25046989.

Sarkar:2005:BRB

[Sar05] Pradipta Sarkar. Book review: An Introduction to Statistical Analysis for Business and Industry: A Problem Solving Approach by Michael Stuart. Technometrics, 47(2):238–239, May 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470996.

Sarkar:2006:BRB

[Sar06] Pradipta Sarkar. Book review: Applied Sequential Methodologies: Real-World Examples with Data Analysis by Nitis Mukhopadhyay; Sujay Datta; Saibal Chattopadhyay. Technometrics, 48(4):570, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471251.

Sauter:2002:BRBb

[Sau02a] Roger M. Sauter. Book review: In All Likelihood: Statistical Modeling and Inference Using Likelihood by Yudi Pawitan. Technometrics, 44(4):404, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271560.

Sauter:2002:BRBa

[Sau02b] Roger M. Sauter. Book review: Introduction to Statistics and Data Analysis by Roxy L. Peck; Chris Olsen; Jay L. Devore. Technometrics, 44(1):90, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270702.

Sauter:2005:BRB

[Sau05] Roger M. Sauter. Book review: Introduction to Probability and Statistics for Engineers and Scientists by Sheldon M. Ross. Technometrics, 47(3):378, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471043.

Sauter:2006:BRB

[Sau06] Roger M. Sauter. Book review: Elementary Probability with Applications by Larry Rabinowitz. Technometrics, 48(3):446, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471217.

Sauter:2007:BRB

[Sau07] Roger Sauter. Book review: Introduction to Engineering Statistics and Six Sigma by Theodore T. Allen. Technometrics, 49(2):229, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471329.

Steinberg:2004:DAT

[SB04] David M. Steinberg and Dizza Bursztyn. Data analytic tools for understanding random field regression models. Technometrics, 46 (4):411–420, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470889.

Steinberg:2008:BHI

[SB08] David M. Steinberg and Søren Bisgaard. “Technometrics”: How it all started. Technometrics, 50(1):2–7, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471419.

Shumway:1998:CSD

[SBD98] R. H. Shumway, D. R. Baumgardt, and Z. A. Der. A cepstral F statistic for detecting delay-fired seismic signals. Technometrics, 40(2):100–110, May 1998. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor. org/stable/1270644. See correction [Ano00e].

Steinberg:2008:FIS

[SBF^{+}08] David M. Steinberg, Søren Bisgaard, Nicholas Fisher, Gerald Hahn, Jon Kettenring, Douglas C. Montgomery, Necip Doganaksoy, Bert Gunter, Sallie Keller-McNulty, William Q. Meeker, and C. F. Jeff Wu. The future of industrial statistics: A panel discussion. Technometrics, 50(2):103–127, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471453.

Shmueli:2000:RRP

[SC00] Galit Shmueli and Ayala Cohen. Run-related probability functions applied to sampling inspection. Technometrics, 42(2):188– 202, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271451.

Setodji:2004:MIR

[SC04] C. Messan Setodji and R. Dennis Cook. K-means inverse regression. Technometrics, 46(4):421–429, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470890.

Scargle:2002:BRB

[Sca02] Jeffrey D. Scargle. Book review: Nonlinear and Nonstationary Signal Processing by W. J. Fitzgerald; Richard L. Smith; A. T. Walden; Peter Young. Technometrics, 44(1):88–89, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270700.

Scargle:2004:BRB

[Sca04] Jeffrey D. Scargle. Book review: An Introduction to the Theory of Point Processes, Vol. I: Elementary Theory and Methods by Daryl Daley; David Vere-Jones. Technometrics, 46(2):257, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470819.

Scargle:2005:BRB

[Sca05] Jeffrey D. Scargle. Book review: Chaos and Time-Series Analysis by Julien Clinton Sprott. Technometrics, 47(3):373, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471033.

Scargle:2007:BRB

[Sca07] Jeffrey Scargle. Book review: Fractal-Based Point Processes by Steven Bradley Lowen; Malvin Carl Teich. Technometrics, 49 (1):102, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471282.

Schaefer:2004:BRB

[Sch04] Robert L. Schaefer. Book review: Intro Stats by Richard D. De Veaux; Paul F. Velleman; David E. Bock. Technometrics, 46(2):256–257, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470818.

Scott:2001:PSM

[Sco01] David W. Scott. Parametric statistical modeling by minimum integrated square error. Technometrics, 43(3):274–285, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271214.

Stinstra:2003:CMD

[SdHSV03] Erwin Stinstra, Dick den Hertog, Peter Stehouwer, and Arjen Vestjens. Constrained maximin designs for computer experiments. Technometrics, 45(4):340–346, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047090.

Seaman:2001:BRB

[Sea01] John W. Seaman, Jr. Book review: Bayesian Methods by Thomas Leonard; John S. J. Hsu. Technometrics, 43(1):100, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270866.

Seaver:2002:BRB

[Sea02] William L. Seaver. Book review: Mind on Statistics by Jessica M. Utts; Robert F. Heckard. Technometrics, 44(4):404–405, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271562.

Seaver:2005:BRB

[Sea05] William L. Seaver. Book review: Random Graphs for Statistical Pattern Recognition by David J. Marchette. Technometrics, 47(1):96, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470942.

Seaver:2009:BRB

[Sea09] William Seaver. Book review: A First Course in Order Statistics by B. C. Arnold; N. Balakrishnan; H. N. Nagaraja. Technometrics, 51(3):342, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586632.

Shi:2009:ADD

[SEM09] Ying Shi, Luis A. Escobar, and William Q. Meeker. Accelerated destructive degradation test planning. Technometrics, 51 (1):1–13, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586558.

Sen:2000:BRB

[Sen00] Ananda Sen. Book review: Statistical Reliability Engineering by Boris Gnedenko; Igor Pavlov; Igor Ushakov; Sumantra Chakravarty. Technometrics, 42(3):315–316, August 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271098.

Sengupta:2002:BRB

[Sen02] Sailes K. Sengupta. Book review: Bayesian Inference in Wavelet-Based Models by Peter Muller; Brani Vidakovic. Technometrics, 44(1):87–88, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270699.

Sen:2005:BRB

[Sen05] Ananda Sen. Book review: Skew-Elliptical Distributions and Their Applications: A Journey beyond Normality by Marc G. Genton. Technometrics, 47(4):519–521, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471084.

Sengupta:2006:BRB

[Sen06a] Sailes K. Sengupta. Book review: Applied Research in Uncertainty Modeling and Analysis by Nil O. Attoh-Okine; Bilal M. Ayyub. Technometrics, 48(4):569–570, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471250.

Senturk:2006:BRB

[Sen06b] Deniz Senturk. Book review: Explanatory Item Response Models: A Generalized Linear and Nonlinear Approach by Paul De Boeck; Mark Wilson. Technometrics, 48(4):568–569, November 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471248.

Sendova:2007:BRB

[Sen07] Kristina Sendova. Book review: Operational Risk: Modeling Analytics by Harry Panjer. Technometrics, 49(4):492, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471397.

Shanmugam:2002:BRB

[Sha02] Ramalingam Shanmugam. Book review: Finite Mixture Models by Geoffrey McLachlan; David Peel. Technometrics, 44(1):82, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270689.

Shanmugam:2004:BRB

[Sha04] Ramalingam Shanmugam. Book review: The Stress-Strength Model and Its Generalizations: Theory and Applications by Samuel Kotz; Yan Lumelskii; Marianna Pensky. Technometrics, 46(1): 112, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470770.

Shea:2000:BRB

[She00] Gerald Shea. Book review: Risk Assessment and Decision Making in Business and Industry: A Practical Guide by Glenn Koller. Technometrics, 42(2):217, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271472.

Sheng:2007:BRB

[She07] Xiaoming Sheng. Book review: Models for Intensive Longitudinal Data by Theodore A. Walls; Joseph L. Schafer. Technometrics, 49(3):361, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471359.

Shen:2009:MFT

[She09] Haipeng Shen. On modeling and forecasting time series of smooth curves. Technometrics, 51(3):227–238, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586618.

Short:2003:BRB

[Sho03] Thomas H. Short. Book review: Applied Statistics with Microsoft Excel(R) by Gerald Keller. Technometrics, 45(3):275, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047067.

Short:2004:BRB

[Sho04] Thomas H. Short. Book review: Applied Bayesian Modelling by Peter Congdon. Technometrics, 46(2):250, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470809.

Short:2005:BRB

[Sho05] Thomas H. Short. Book review: Stats: Data and Models by Richard D. De Veaux; Paul F. Velleman; David E. Bock. Technometrics, 47(4):515–516, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471077.

Shen:2004:AMS

[SHY04] Xiaotong Shen, Hsin-Cheng Huang, and Jimmy Ye. Adaptive model selection and assessment for exponential family distributions. Technometrics, 46(3):306–317, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470845.

Singh:2001:BRB

[Sin01] Ashok K. Singh. Book review: Analysis of Customer Satisfaction Data by Derek R. Allen; Tanniru R. Rao. Technometrics, 43(4):493–494, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270833.

Singpurwalla:2005:DTH

[Sin05] Nozer D. Singpurwalla. Decelerated testing: A hierarchical Bayes approach. Technometrics, 47(4):468–477, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471071.

Singh:2006:BRB

[Sin06] Ashok K. Singh. Book review: The EM Algorithm and Related Statistical Models by Michiko Watanabe; Kazunori Yamaguchi. Technometrics, 48(1):148, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471128.

Singh:2007:BRB

[Sin07] Ashok Singh. Book review: A Manager’s Guide to the Design and Conduct of Clinical Trials by Phillip I. Good. Technometrics, 49(4):491, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471394.

Singh:2008:BRB

[Sin08] Ashok K. Singh. Book review: Introduction to Data Mining Using SAS Enterprise Miner by Patricia B. Cerrito. Technometrics, 50(4):545, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471539.

Singh:2009:BRB

[Sin09] Ashok K. Singh. Book review: Model-Based Geostatistics by Peter J. Diggle; Paulo J. Ribeiro. Technometrics, 51(3):339–340, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586628.

Sitter:2005:ER

[Sit05] Randy R. Sitter. Editor’s report. Technometrics, 47(4):385–387, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471063.

Sitter:2006:ER

[Sit06] Randy R. Sitter. Editor’s report. Technometrics, 48(4):455–457, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471237.

Sitter:2007:ER

[Sit07] Randy R. Sitter. Editor’s report. Technometrics, 49(4):370–372, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471383.

Sullivan:2002:SSC

[SJ02] Joe H. Sullivan and L. Allison Jones. A self-starting control chart for multivariate individual observations. Technometrics, 44 (1):24–33, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270681.

Slud:2003:BRB

[Slu03] Eric V. Slud. Book review: Graphical Models: Methods for Data Analysis and Mining by Christian Borgelt; Rudolf Kruse. Technometrics, 45(2):177–178, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047026.

Steiner:2004:SC

[SM04] Stefan H. Steiner and R. Jock MacKay. Scale counting. Technometrics, 46(3):348–354, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470849.

Sulieman:2001:PBA

[SMB01] H. Sulieman, P. J. McLellan, and D. W. Bacon. A profile-based approach to parametric sensitivity analysis of nonlinear regression models. Technometrics, 43(4):425–433, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270813.

Smith:2001:BRB

[Smi01] Patricia L. Smith. Book review: Geostatistical Error Management: Quantifying Uncertainty for Environmental Sampling and Mapping by Jeffrey C. Myers. Technometrics, 43(2):238–239, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271048.

Smith:2002:BRB

[Smi02] Eric P. Smith. Book review: An Introduction to Statistical Modeling of Extreme Values by Stuart Coles. Technometrics, 44(4):397, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271547.

Smith:2006:BRB

[Smi06] Eric P. Smith. Book review: Using R for Introductory Statistics by John Verzani. Technometrics, 48(2):307–308, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471174.

Smyth:2000:ESC

[Smy00] Gordon K. Smyth. Employing symmetry constraints for improved frequency estimation by eigenanalysis methods. Technometrics, 42(3):277–289, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271082.

Somboonsavatdee:2007:GEP

[SNS07] Anupap Somboonsavatdee, Vijayan N. Nair, and Ananda Sen. Graphical estimators from probability plots with right-censored data. Technometrics, 49(4):420–429, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471388.

So:2001:BRB

[So01] Ernest M. T. So. Book review: Who Counts? The Politics of Census-Taking in Contemporary America by Margo J. Anderson; Stephen E. Fienberg. Technometrics, 43(2):242, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271055.

Soyer:2004:BRB

[Soy04] Refik Soyer. Book review: Bayesian Methods for Nonlinear Classification and Regression by D. G. T. Denison; C. C. Holmes; B. K. Mallick; A. F. M. Smith. Technometrics, 46(2):251–252, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470811.

Soyer:2009:BRB

[Soy09] Refik Soyer. Book review: Bayesian Networks and Probabilistic Inference in Forensic Science by F. Tarconi; C. Aitken; P. Garbolino; A. Biedermann. Technometrics, 51(1):99–100, February 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586569.

Stocker:2007:GCP

[SP07] Russell S. Stocker IV and Edsel A. Peña. A general class of parametric models for recurrent event data. Technometrics, 49 (2):210–220, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471314.

Spiring:2007:BRB

[Spi07] Fred Spiring. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 49(1):108– 109, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471295.

Singh:2004:CLS

[SR04] Rita Singh and Bhiksha Raj. Classification in likelihood spaces. Technometrics, 46(3):318–329, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470846.

Scott:2001:KM

[SS01] David W. Scott and William F. Szewczyk. From kernels to mixtures. Technometrics, 43(3):323–335, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271220.

Sitter:2007:EEE

[SS07] Randy R. Sitter and David M. Steinberg. The Editor and Editor-Elect: On papers in information science and technology. Technometrics, 49(3):235, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471343.

Simonoff:2002:STS

[ST02] Jeffrey S. Simonoff and Chih-Ling Tsai. Score tests for the single index model. Technometrics, 44(2):142–151, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271258.

Stefanescu:2006:MFM

[ST06] Catalina Stefanescu and Bruce W. Turnbull. Multivariate frailty models for exchangeable survival data with covariates. Technometrics, 48(3):411–417, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471212.

Stanard:2000:BRB

[Sta00] Christopher Stanard. Book review: SPC Simplified: Practical Steps to Quality by Robert T. Amsden; Howard E. Butler; Davida M. Amsden. Technometrics, 42(3):310–311, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271089.

Stanard:2004:BRB

[Sta04] Christopher L. Stanard. Book review: Reliability Verification, Testing, and Analysis in Engineering Design by Gary S. Wasserman. Technometrics, 46(4):486–487, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470900.

Steiner:2000:SPC

[Ste00a] Stefan H. Steiner. Statistical process control using two measurement systems. Technometrics, 42(2):178–187, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271450.

Stephenson:2000:BRB

[Ste00b] Hal W. Stephenson. Book review: Quality Problem Solving by Gerald F. Smith. Technometrics, 42(2):211, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271462.

Stephenson:2004:BRB

[Ste04] W. Robert Stephenson. Book review: Nonparametric Statistical Methods for Complete and Censored Data by M. M. Desu; D. Raghavarao. Technometrics, 46(4):487–488, November 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470901.

Steinberg:2007:TNA

[Ste07] David M. Steinberg. Tragic news about Randy Sitter. Technometrics, 49(4):369, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471382.

Steinberg:2008:BF

[Ste08a] David M. Steinberg. “Technometrics” at fifty. Technometrics, 50(1):1, February 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25471418.

Steinberg:2008:ER

[Ste08b] David M. Steinberg. Editor’s report. Technometrics, 50(4):415– 417, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471519.

Steinberg:2009:EA

[Ste09a] David M. Steinberg. Editorial announcement. Technometrics, 51 (4):349, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/40586643.

Steinberg:2009:ENI

[Ste09b] David M. Steinberg. Editor’s note: Improved Web site with supplemental materials. Technometrics, 51(2):109, May 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/40586588.

Steinberg:2009:ER

[Ste09c] David M. Steinberg. Editor’s report. Technometrics, 51(4):350– 352, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586644.

Stromberg:2001:BRB

[Str01] Arnold J. Stromberg. Book review: Subsampling by Dimitris N. Politis; Joseph P. Romano; Michael Wolf. Technometrics, 43(2): 241–242, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
1271054.

Sudjianto:2004:FAM

[Sud04] Agus Sudjianto. [Failure Amplification Method: An Information Maximization Approach to Categorical Response Optimization]: Discussion. Technometrics, 46(1):23–24, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470756. See [JW04a].

Sullivan:2001:BRB

[Sul01] Joe H. Sullivan. Book review: Classification by A. D. Gordon. Technometrics, 43(1):111–112, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270884.

Shen:2007:DLM

[SX07] Qing Shen and Hongquan Xu. Diagnostics for linear models with functional responses. Technometrics, 49(1):26–33, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471272.

Symanzik:2001:BRB

[Sym01] Jürgen Symanzik. Book review: A Casebook for Spatial Statistical Data Analysis by Daniel A. Griffith; Larry J. Layne. Technometrics, 43(3):375–376, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271234.

Symanzik:2005:BRB

[Sym05] Jürgen Symanzik. Book review: Statistical Analysis of Spatial Point Patterns by Peter J. Diggle. Technometrics, 47(4):516–517, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471078.

Taam:2002:BRB

[Taa02] Winson Taam. Book review: Probability with Statistical Applications by Rinaldo B. Schinazi. Technometrics, 44(4):402, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271557.

Thomas:2000:DRM

[TG00] Edward V. Thomas and Nanxiang Ge. Development of robust multivariate calibration models. Technometrics, 42(2):168–177, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271449.

Trinca:2001:MRS

[TG01] Luzia A. Trinca and Steven G. Gilmour. Multistratum response surface designs. Technometrics, 43(1):25–33, February 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270854.

Taddy:2009:BGP

[TGLG09] Matthew A. Taddy, Genetha A. Gray, Herbert K. H. Lee, and Joshua D. Griffin. Bayesian guided pattern search for robust local optimization. Technometrics, 51(4):389–401, November 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586649.

Tsiamyrtzis:2005:BSD

[TH05] Panagiotis Tsiamyrtzis and Douglas M. Hawkins. A Bayesian scheme to detect changes in the mean of a short-run process. Technometrics, 47(4):446–456, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471069.

Thompson:2001:AT

[Tho01] James R. Thompson. The age of Tukey. Technometrics, 43(3):256– 265, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.allbusiness.com/
technology/808082-1.html;   http://www.jstor.org/stable/

1271212. Special Tukey Memorial issue.

Tseng:2009:TNS

[TJ09] Sheng-Tsaing Tseng and Bo-Yan Jou. A technical note on “Sample Size Determination for Achieving Stability of Double Multivariate Exponentially Weighted Moving Average Controller” [MR2414514]. Technometrics, 51(3):335–338, August 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586627.

Tobias:2000:BRB

[Tob00] Paul A. Tobias. Book review: Mechanical Reliability and Design by A. D. S. Carter. Technometrics, 42(2):207–208, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271457.

Tobias:2003:BRB

[Tob03] Paul A. Tobias. Book review: Process Improvement in the Electronics Industry by Yefim Fasser; Donald Brettner. Technometrics, 45(4):366, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047100.

Toman:2007:BAC

[Tom07] Blaza Toman. Bayesian approaches to calculating a reference value in key comparison experiments. Technometrics, 49(1):81– 87, February 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471277.

Tortorella:2001:BRB

[Tor01] M. Tortorella. Book review: Reliability Theory: with Applications to Preventive Maintenance by Ilya Gertsbakh. Technometrics, 43 (4):487–488, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270823.

Tseng:2007:SSD

[TTL07] Sheng-Tsaing Tseng, Jen Tang, and Chien-Hua Lin. Sample size determination for achieving stability of double multivariate exponentially weighted moving average controller. Technometrics, 49 (4):409–419, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471387.

Turlach:2005:SVS

[TVW05] Berwin A. Turlach, William N. Venables, and Stephen J. Wright. Simultaneous variable selection. Technometrics, 47(3):349–363, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471027.

Uys:2007:NTS

[UL07] Nadia Uys and Fred Lombard. A note on two-sided cusums for a normal mean. Technometrics, 49(2):195–198, May 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471312.

Umble:2001:BRB

[Umb01] Elisabeth J. Umble. Book review: Cumulative Sum Charts and Charting for Quality Improvement by Douglas M. Hawkins; David H. Olwell. Technometrics, 43(1):107, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270876.

Utlaut:2005:BRB

[Utl05] Theresa L. Utlaut. Book review: Statistical Analysis and Data Display: An Intermediate Course with Examples in S-PLUS. R, and SAS by Richard M. Heiberger; Burt Holland. Technometrics, 47(3):371, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471029.

Vangel:2002:LAC

[Van02] Mark G. Vangel. Lot acceptance and compliance testing using the sample mean and an extremum. Technometrics, 44(3):242–249, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270488.

VanBrackle:2004:BRB

[Van04] Lewis VanBrackle. Book review: Step-by-Step Basic Statistics Using SAS(R): Student Guide by Larry Hatcher. Technometrics, 46 (4):494–495, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470913.

Vardeman:2000:ITC

[Var00] Stephen B. Vardeman. Introduction to two classics in statistical process control. Technometrics, 42(1):95–96, February 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271438.

Bowen:2002:BRB

[vB02] Jacob van Bowen, Jr. Book review: Single-Case and Small-n Experimental Designs by John B. Todman; Pat Dugard. Technometrics, 44(4):401, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271554.

Vivacqua:2009:PFS

[VB09] Carla A. Vivacqua and Søren Bisgaard. Post-fractionated stripblock designs. Technometrics, 51(1):47–55, February 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586562.

Vanli:2007:CLS

[VD07] O. Arda Vanli and Enrique Del Castillo. Closed-loop system identification for small samples with constraints. Technometrics, 49 (4):382–394, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471385.

VanWieringen:2008:MSA

[VD08] Wessel N. Van Wieringen and Jeroen De Mast. Measurement system analysis for binary data. Technometrics, 50(4):468–478, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471530.

Vine:2008:CAT

[VDLB08] A. E. Vine, A. M. Dean, S. M. Lewis, and D. Brunson. A critical assessment of two-stage group screening through industrial experimentation. Technometrics, 50(1):15–25, February 2008. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471421. See correction

[Ano08s].

vanNostrand:2002:BRB

[vN02] Richard Craig van Nostrand. Book review: Design of Experiments Using the Taguchi Approach: 16 Steps to Product and Process Improvement by Ranjit K. Roy. Technometrics, 44(3):289, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270496.

Voelkel:2005:EFF

[Voe05] Joseph G. Voelkel. The efficiencies of fractional factorial designs. Technometrics, 47(4):488–494, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471073.

Volino:2005:BRB

[Vol05] Casey A. Volino. Book review: A First Course in Stochastic Models by Henk C. Tijms. Technometrics, 47(3):375, August 2005. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471037.

VonTress:2003:BRB

[Von03] Mark Von Tress. Book review: Generalized, Linear, and Mixed Models by Charles E. McCulloch; Shayle R. Searle. Technometrics, 45(1):99, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25046979.

VonTress:2006:BRB

[Von06] Mark Von Tress. Book review: Analyzing Rater Agreement by Alexander von Eye; Eun Young Mun. Technometrics, 48(3):448– 449, August 2006. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471222.

vonTress:2000:BRB

[vT00] Mark von Tress. Book review: Applied Categorical Analysis by Chap T. Le. Technometrics, 42(2):214, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271468.

Vukovinsky:2000:BRB

[Vuk00] Kimberly Vukovinsky. Book review: Beyond Capability Confusion by Donald J. Wheeler. Technometrics, 42(2):204, May 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271452.

Walker:2003:BRB

[Wal03] Esteban Walker. Book review: Regression Modeling Strategies by Frank E. Harrell. Technometrics, 45(2):170, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047014.

Walker:2005:BRB

[Wal05] Esteban Walker. Book review: Semiparametric Regression by David Ruppert; M. P. Wand; R. J. Carroll. Technometrics, 47 (2):234, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470989.

Wassell:2001:BRB

[Was01] James T. Wassell. Book review: Bootstrap Methods: A Practitioner’s Guide by Michael R. Chernick. Technometrics, 43(1): 99–100, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270865.

Wassell:2002:BRB

[Was02] James T. Wassell. Book review: Numerical Methods of Statistics by John F. Monahan. Technometrics, 44(1):91–92, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270706.

Wassell:2004:BRB

[Was04] James T. Wassell. Book review: Statistics with Applications in Biology and Geology by Preben Blæsild; Jørgen Granfeldt. Technometrics, 46(1):111, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470768.

Wassell:2005:BRB

[Was05] James T. Wassell. Book review: Sensitivity Analysis in Practice by Andrea Saltelli; Stefano Tarantola; Francesca Campolongo; Marco Ratto. Technometrics, 47(2):236–237, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470993.

Wikle:2005:CIA

[WB05] Christopher K. Wikle and L. Mark Berliner. Combining information across spatial scales. Technometrics, 47(1):80–91, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470936.

Wang:2009:BVC

[WCT09] Shuchun Wang, Wei Chen, and Kwok-Leung Tsui. Bayesian validation of computer models. Technometrics, 51(4):439–451, November 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/40586653.

Wehrly:2002:BRB

[Weh02] Thomas E. Wehrly. Book review: Fitting Statistical Distributions: The Generalized Lambda Distribution and Generalized Bootstrap Methods by Zaven A. Karian; Edward J. Dudewicz. Technometrics, 44(2):194, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271274.

Wehrly:2004:BRB

[Weh04] Thomas E. Wehrly. Book review: An Invariant Approach to Statistical Analysis of Shapes by Subhash R. Lele; Joan T. Richtsmeier. Technometrics, 46(3):364–365, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470861.

Weisman:2003:BRB

[Wei03] Doris A. Weisman. Book review: Experimental Design with Applications in Management, Engineering and the Sciences by Paul D. Berger; Robert E. Maurer. Technometrics, 45(1):105, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25046987.

Whitcher:2004:WBE

[Whi04] Brandon Whitcher. Wavelet-based estimation for seasonal longmemory processes. Technometrics, 46(2):225–238, May 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470807.

Wikle:2003:BRB

[Wik03] Christopher K. Wikle. Book review: Modeling Hydrologic Change: Statistical Methods by Richard H. McCuen. Technometrics, 45 (4):366–367, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25047101.

Wikle:2005:BRB

[Wik05] Christopher K. Wikle. Book review: Atmospheric Modeling, Data Assimilation, and Predictability by Eugenia Kalnay. Technometrics, 47(4):521, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471085.

Wills:2002:BRB

[Wil02] Graham J. Wills. Book review: Introduction to Graphical Modelling by David Edwards. Technometrics, 44(2):197, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271279.

Wilson:2005:BRB

[Wil05] William J. Wilson. Book review: Multivariate Statistical Methods by Donald F. Morrison. Technometrics, 47(3):378, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471044.

Wilkinson:2008:FSCa

[Wil08a] Leland Wilkinson. The future of statistical computing. Technometrics, 50(4):418–435, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471520.

Wilkinson:2008:FSCb

[Wil08b] Leland Wilkinson. [The Future of Statistical Computing]: Rejoinder. Technometrics, 50(4):446–447, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471527.

Woodall:2003:RAM

[WKT^{+}03] William H. Woodall, Rachelle Koudelik, Kwok-Leung Tsui, Seoung Bum Kim, Zachary G. Stoumbos, Christos P. Carvounis, Rajesh Jugulum, Genichi Taguchi, Shin Taguchi, James O.

Wilkins, Bovas Abraham, Asokan Mulayath Variyath, and Douglas M. Hawkins. A review and analysis of the Mahalanobis– Taguchi system [with discussion and response]. Technometrics, 45(1):1–30, February 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/ stable/25046971. With discussion and a response by the authors.

Woods:2006:DGL

[WLER06] D. C. Woods, S. M. Lewis, J. A. Eccleston, and K. G. Russell. Designs for generalized linear models with several variables and model uncertainty. Technometrics, 48(2):284–292, May 2006. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471164.

Wu:2007:SPI

[WLLZ07] Ying Nian Wu, Jinhui Li, Ziqiang Liu, and Song-Chun Zhu. Statistical principles in image modeling. Technometrics, 49(3):249– 261, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471346.

Wludyka:2003:BRB

[Wlu03] Peter Wludyka. Book review: Statistical Analysis of Designed Experiments by Helge Toutenburg. Technometrics, 45(2):171, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047016.

Wludyka:2006:BRB

[Wlu06] Peter Wludyka. Book review: Process Quality Control: Troubleshooting and Interpretation of Data by Ellis R. Ott; Edward G. Schilling; Dean V. Neubauer. Technometrics, 48(4):572, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471255.

Wu:2002:EDR

[WM02] Huaiqing Wu and William Q. Meeker. Early detection of reliability problems using information from warranty databases. Technometrics, 44(2):120–133, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271256.

Woodall:2005:IPQ

[WM05] William H. Woodall and Mahmoud A. Mahmoud. The inertial properties of quality control charts. Technometrics, 47(4):425– 436, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471067.

Wong:2004:BRB

[Won04] Aldous G. Wong. Book review: Design of Experiments and Other Six Sigma Related Methods Explained and Derived by Theodore T. Allen. Technometrics, 46(4):489, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470904.

Woodard:2000:BRB

[Woo00] Roger Woodard. Book review: Interpolation of Spatial Data: Some Theory for Kriging by Michael L. Stein. Technometrics, 42(4): 436–437, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270967.

Walsh:2002:DMM

[WR02] Daniel C. I. Walsh and Adrian E. Raftery. Detecting Mines in minefields with linear characteristics. Technometrics, 44(1):34–44, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270682.

Wang:2009:TII

[WT09] Hsiuying Wang and Fugee Tsung. Tolerance intervals with improved coverage probabilities for binomial and Poisson variables. Technometrics, 51(1):25–33, February 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/40586560.

Wong:2000:PSS

[WW00] A. C. M. Wong and J. Wu. Practical small-sample asymptotics for distributions used in life-data analysis. Technometrics, 42(2):149– 155, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271446.

Xu:2004:OPT

[XCW04] Hongquan Xu, Shao-Wei Cheng, and C. F. J. Wu. Optimal projective three-level designs for factor screening and interaction detection. Technometrics, 46(3):280–292, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470843.

Xu:2005:MAP

[XD05] Hongquan Xu and Lih-Yuan Deng. Moment aberration projection for nonregular fractional factorial designs. Technometrics, 47(2):121–131, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470974.

Xu:2002:ACO

[Xu02] Hongquan Xu. An algorithm for constructing orthogonal and nearly-orthogonal arrays with mixed levels and small runs. Technometrics, 44(4):356–368, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271543.

Xu:2006:BRB

[Xu06] Hongquan Xu. Book review: Design and Analysis of Experiments, Volume 2: Advanced Experimental Design by Klaus Hinkelmann; Oscar Kempthorne. Technometrics, 48(4):571, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471253.

Xu:2009:ACE

[Xu09] Hongquan Xu. Algorithmic construction of efficient fractional factorial designs with large run sizes. Technometrics, 51(3):262– 277, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
40586621.

Yates:2001:BRB

[Yat01] Phillip Yates. Book review: The Desk Reference of Statistical Quality Methods by Mark L. Crossley. Technometrics, 43(2):232–233, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271039.

Yates:2002:BRB

[Yat02] Phillip Yates. Book review: SPC: Practical Understanding of Capability by Implementing Statistical Process Control by James C. Abbott. Technometrics, 44(2):187–188, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271264.

Ye:2000:BRB

[Ye00] Kenny Q. Ye. Book review: Statistical Tests for Mixed Linear Models by André I. Khuri; Thomas Mathew; Bimal K. Sinha. Technometrics, 42(2):214, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271467.

Ye:2003:BRB

[Ye03] Kenny Q. Ye. Book review: Case Studies in Bayesian Statistics, Vol. VI by Constantine Gatsonis; Robert E. Kass; Alicia Carriquiry; Andrew Gelman; David Higdon; Donna K. Pauler; Isabella Verdinelli. Technometrics, 45(3):273–274, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047065.

Ye:2005:BRB

[Ye05] Keying Ye. Book review: Applied Bayesian Modeling and Causal Inference from Incomplete-Data Perspectives by Andrew Gelman; Xiao-Li Meng. Technometrics, 47(4):519, November 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471083.

Yeh:2005:BRB

[Yeh05] Arthur B. Yeh. Book review: Fundamentals of Probability and Statistics for Engineers by T. T. Soong. Technometrics, 47(2):239, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470997.

Yeh:2007:BRB

[Yeh07] Arthur B. Yeh. Book review: A Modern Introduction to Probability and Statistics by F. M. Dekking; C. Kraaikamp; H. P. Lopuhaa; L. E. Meester. Technometrics, 49(3):359, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471355.

Yuan:2007:EVS

[YJL07] Ming Yuan, V. Roshan Joseph, and Yi Lin. An efficient variable selection approach for analyzing designed experiments. Technometrics, 49(4):430–439, November 2007. CODEN TCMTA2.

ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471389.

Yang:2000:DRR

[YM00] Jiangbin Yang and Viliam Makis. Dynamic response of residuals to external deviations in a controlled production process. Technometrics, 42(3):290–299, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271083.

Yin:2005:SEM

[YS05] Xiangrong Yin and Lynne Seymour. Standard errors for the multiple roots in quadratic response surface models. Technometrics, 47(3):260–263, August 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471019.

Yu:2007:ESC

[Yu07a] Bin Yu. Embracing statistical challenges in the information technology age. Technometrics, 49(3):237–248, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471345.

Yu:2007:BRB

[Yu07b] Jihnhee Yu. Book review: Measures, Integrals and Martingales by René L. Schilling. Technometrics, 49(3):363, August 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25471364.

Yu:2005:MSM

[YW05] Qiqing Yu and George Y. C. Wong. Modified semiparametric maximum likelihood estimator in linear regression analysis with complete data or right-censored data. Technometrics, 47(1):34– 42, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470932.

Zhu:2006:DEP

[ZC06] Mu Zhu and Hugh A. Chipman. Darwinian evolution in parallel universes: A parallel genetic algorithm for variable selection. Technometrics, 48(4):491–502, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471241.

Zhang:2008:LFA

[ZCC08] Jian Zhang, Peter F. Craigmile, and Noel Cressie. Loss function approaches to predict a spatial quantile and its exceedance region. Technometrics, 50(2):216–227, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471461.

Zhou:2003:DSM

[ZDCS03] Shiyu Zhou, Yu Ding, Yong Chen, and Jianjun Shi. Diagnosability study of multistage manufacturing processes based on linear mixed-effects models. Technometrics, 45(4):312–325, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047087.

Zelterman:2000:BRB

[Zel00] Daniel Zelterman. Book review: Numerical Analysis for Statisticians by Kenneth Lange. Technometrics, 42(3):322, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271108.

Zelterman:2001:BRB

[Zel01] Daniel Zelterman. Book review: Causality: Models, Reasoning, and Inference by Judea Pearl. Technometrics, 43(2):239–240, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271049.

Zelterman:2005:BRB

[Zel05] Daniel Zelterman. Book review: Bayesian Artificial Intelligence by Kevin B. Korb; Ann E. Nicholson. Technometrics, 47(1): 101–102, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470952.

Zamba:2006:MCP

[ZH06] K. D. Zamba and Douglas M. Hawkins. A multivariate changepoint model for statistical process control. Technometrics, 48 (4):539–549, November 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471246.

Zukovic:2009:MNC

[ŽH09] Milan Žukovič and Dionissios T. Hristopulos. The method of normalized correlations: A fast parameter estimation method for random processes and isotropic random fields that focuses on shortrange dependence. Technometrics, 51(2):173–185, May 2009. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/40586594.

Zhang:2005:BRB

[Zha05] Chunming Zhang. Book review: Ranked Set Sampling: Theory and Applications by Zehua Chen; Zhidong Bai; Bimal K. Sinha. Technometrics, 47(1):100–101, February 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470950.

Zhang:2006:PTS

[Zha06] Jin Zhang. Powerful two-sample tests based on the likelihood ratio. Technometrics, 48(1):95–103, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471118.

Zhou:2001:RCE

[Zho01] Julie Zhou. A robust criterion for experimental designs for serially correlated observations. Technometrics, 43(4):462–467, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270817.

Ziegel:2000:BRBap

[Zie00a] Eric R. Ziegel. Book review: Applied Mixed Models in Medicine by Helen Brown; Robin Prescott. Technometrics, 42(4):442–443, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270978.

Ziegel:2000:BRBx

[Zie00b] Eric R. Ziegel. Book review: Applied Multivariate Statistics with SAS(R) Software by Ravindra Khattree; Dayanand N. Naik. Technometrics, 42(3):324, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271113.

Ziegel:2000:BRBab

[Zie00c] Eric R. Ziegel. Book review: Asymptotics, Nonparametrics, Time Series by Subir Ghosh. Technometrics, 42(3):325, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271118.

Ziegel:2000:BRBak

[Zie00d] Eric R. Ziegel. Book review: Case Studies in Bayesian Statistics, Vol. IV by Constantine Gatsonis; Robert E. Kass; Bradley Carlin; Alicia Carriquiry; Andrew Gelman; Isabella Verdinelli; Mike West. Technometrics, 42(3):327–328, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271126.

Ziegel:2000:BRBaf

[Zie00e] Eric R. Ziegel. Book review: Cognition and Survey Research by Monroe G. Sirken; Douglas J. Herrmann; Susan Schechter; Norbert Schwarz; Judith M. Tanur; Roger Tourangeau. Technometrics, 42(3):326, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271121.

Ziegel:2000:BRBao

[Zie00f] Eric R. Ziegel. Book review: Comparative Statistical Inference by Vic Barnett. Technometrics, 42(4):442, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270977.

Ziegel:2000:BRBan

[Zie00g] Eric R. Ziegel. Book review: Computer-Assisted Analysis of Mixtures and Applications by Dankmar Böhning. Technometrics, 42 (4):442, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270976.

Ziegel:2000:BRBq

[Zie00h] Eric R. Ziegel. Book review: Computer Assisted Survey Information Collection by Mick P. Couper; Reginald P. Baker; Jelke Bethlehem; Cynthia Z. F. Clark; Jean Martin; William L. Nicholls; James M. O’Reilly. Technometrics, 42(2):224, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271492.

Ziegel:2000:BRBai

[Zie00i] Eric R. Ziegel. Book review: Data Mining by Bhavani Thuraisingham. Technometrics, 42(3):327, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271124.

Ziegel:2000:BRBn

[Zie00j] Eric R. Ziegel. Book review: Design and Analysis of Animal Studies in Pharmaceutical Development by Shein-Chung Chow; Jen-Pei Liu. Technometrics, 42(2):223, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271489.

Ziegel:2000:BRBr

[Zie00k] Eric R. Ziegel. Book review: Elements of Large-Sample Theory by E. L. Lehmann. Technometrics, 42(2):224, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271493.

Ziegel:2000:BRBm

[Zie00l] Eric R. Ziegel. Book review: Encyclopedia of Biostatistics by P. Armitage; T. Colton. Technometrics, 42(2):222–223, May 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271488.

Ziegel:2000:BRBad

[Zie00m] Eric R. Ziegel. Book review: Encyclopedia of Statistical Sciences, Update Volume 3 by Samuel Kotz; Campbell B. Read; David L. Banks. Technometrics, 42(3):325–326, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271119.

Ziegel:2000:BRBah

[Zie00n] Eric R. Ziegel. Book review: Epidemiology: Study Design and Data Analysis by Mark Woodward. Technometrics, 42(3):327, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271123.

Ziegel:2000:BRBaj

[Zie00o] Eric R. Ziegel. Book review: Fitting Linear Relationships: A History of the Calculus of Observations 1750–1900 by Richard William Farebrother. Technometrics, 42(3):327, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271125.

Ziegel:2000:BRBac

[Zie00p] Eric R. Ziegel. Book review: Fundamental Concepts in the Design of Experiments by Charles R. Hicks; Kenneth V. Turner. Technometrics, 42(3):325, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271115.

Ziegel:2000:BRBau

[Zie00q] Eric R. Ziegel. Book review: Geostatistics and Petroleum Geology by Michael Edward Hohn. Technometrics, 42(4):444, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270983.

Ziegel:2000:BRBat

[Zie00r] Eric R. Ziegel. Book review: Geostatistics by Jean-Paul Chilès; Pierre Delfiner. Technometrics, 42(4):444, November 2000. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270982.

Ziegel:2000:BRBav

[Zie00s] Eric R. Ziegel. Book review: Geostatistics for Engineers and Earth Scientists by Ricardo A. Olea. Technometrics, 42(4):444–445, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270984.

Ziegel:2000:BRBo

[Zie00t] Eric R. Ziegel. Book review: Handbook of Applied Economic Statistics by Aman Ullah; David E. A. Giles. Technometrics, 42(2):223, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271490.

Ziegel:2000:BRBa

[Zie00u] Eric R. Ziegel. Book review: Handbook of Chemometrics and Qualimetrics, Part B by D. L. Massart; B. G. M. Vandeginste; L. M. C. Buydens; S. de Jong; P. J. Lewi; J. Smeyers Verbeke. Technometrics, 42(2):218–219, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271476.

Ziegel:2000:BRBc

[Zie00v] Eric R. Ziegel. Book review: Handbook of Simulation by Jerry Banks. Technometrics, 42(2):219–220, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271478.

Ziegel:2000:BRBb

[Zie00w] Eric R. Ziegel. Book review: Handbook of Statistics 13: Design and Analysis of Experiments by S. Ghosh; C. R. Rao. Technometrics, 42(2):219, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271477.

Ziegel:2000:BRBaw

[Zie00x] Eric R. Ziegel. Book review: Handbook of Statistics 15: Robust Inference by G. S. Maddala; C. R. Rao. Technometrics, 42(4):445, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270985.

Ziegel:2000:BRBax

[Zie00y] Eric R. Ziegel. Book review: Handbook of Statistics 16: Order Statistics: Theory and Methods by N. Balakrishnan; C. R. Rao. Technometrics, 42(4):445, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270986.

Ziegel:2000:BRBay

[Zie00z] Eric R. Ziegel. Book review: Handbook of Statistics 17: Order Statistics: Applications by N. Balakrishnan; C. R. Rao. Technometrics, 42(4):445, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270987.

Ziegel:2000:BRBh

[Zie00-27] Eric R. Ziegel. Book review: Insight. xla by Sam L. Savage. Technometrics, 42(2):221, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271483.

Ziegel:2000:BRBy

[Zie00-28] Eric R. Ziegel. Book review: Integrating Results by Morgan C. Wang; Brad J. Bushman. Technometrics, 42(3):324–325, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271114.

Ziegel:2000:BRBam

[Zie00-29] Eric R. Ziegel. Book review: Intelligent Data Analysis by Michael Berthold; David J. Hand. Technometrics, 42(4):442, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270975.

Ziegel:2000:BRBf

[Zie00-30] Eric R. Ziegel. Book review: Introduction to Simulation and Risk Analysis by James R. Evans; David L. Olson. Technometrics, 42(2):220–221, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271481.

Ziegel:2000:BRBw

[Zie00-31] Eric R. Ziegel. Book review: Logistic Regression Using the SAS(R) System by Paul D. Allison. Technometrics, 42(3):323–324, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271112.

Ziegel:2000:BRBk

[Zie00-32] Eric R. Ziegel. Book review: MODA 5: Advances in Model-Oriented Data Analysis and Experimental Design by A. C. Atkinson; L. Pronzato; H. P. Wynn. Technometrics, 42(2):222, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271486.

Ziegel:2000:BRBi

[Zie00-33] Eric R. Ziegel. Book review: Modern Applied Biostatistical Methods Using S-Plus by Steve Selvin. Technometrics, 42(2):221, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271484.

Ziegel:2000:BRBv

[Zie00-34] Eric R. Ziegel. Book review: Multiple Comparisons and Multiple Tests by Peter H. Westfall; Randall D. Tobias; Dror Rom; Russell D. Wolfinger; Yosef Hochberg. Technometrics, 42(3):323, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271111.

Ziegel:2000:BRBaa

[Zie00-35] Eric R. Ziegel. Book review: Multivariate Analysis, Design of Experiments, and Survey Sampling by Subir Ghosh. Technometrics, 42(3):325, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271117.

Ziegel:2000:BRBd

[Zie00-36] Eric R. Ziegel. Book review: Multivariate Geostatistics by Hans Wackernagel. Technometrics, 42(2):220, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271479.

Ziegel:2000:BRBl

[Zie00-37] Eric R. Ziegel. Book review: Pharmaceutical Experimental Design by Gareth A. Lewis; Didier Mathieu; Roger Phan-Tan-Luu. Technometrics, 42(2):222, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271487.

Ziegel:2000:BRBt

[Zie00-38] Eric R. Ziegel. Book review: Practical Statistics by Example Using Microsoft(R) Excel by Terry Sincich; David M. Levine; David Stephan. Technometrics, 42(3):322–323, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271109.

Ziegel:2000:BRBp

[Zie00-39] Eric R. Ziegel. Book review: Sampling: Design and Analysis by Sharon L. Lohr. Technometrics, 42(2):223–224, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271491.

Ziegel:2000:BRBae

[Zie00-40] Eric R. Ziegel. Book review: Sampling of Populations by P. Levy; S. Lemeshow. Technometrics, 42(3):326, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271120.

Ziegel:2000:BRBj

[Zie00-41] Eric R. Ziegel. Book review: SAS(R) Macro Programming Made Easy by Michele M. Burlew. Technometrics, 42(2):221–222, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271485.

Ziegel:2000:BRBu

[Zie00-42] Eric R. Ziegel. Book review: Statistical Analysis Using STAT-GRAPHICS Plus, Volume 2: Quality Control and Experimental Design by Neil W. Polhemus. Technometrics, 42(3):323, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271110.

Ziegel:2000:BRBz

[Zie00-43] Eric R. Ziegel. Book review: Statistical and Probabilistic Models in Reliability by D. C. Ionescu; N. Limnios. Technometrics, 42(3):325, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271116.

Ziegel:2000:BRBar

[Zie00-44] Eric R. Ziegel. Book review: Statistical Aspects of Health and the Environment, 4 by Vic Barnett; Alfred Stein; K. Feridun Turkman. Technometrics, 42(4):443, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270980.

Ziegel:2000:BRBag

[Zie00-45] Eric R. Ziegel. Book review: Statistical Methods for Health Sciences by M. M. Shoukri; C. A. Pause. Technometrics, 42(3):326–327, August 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271122.

Ziegel:2000:BRBe

[Zie00-46] Eric R. Ziegel. Book review: Statistical Procedures for Analysis of Environmental Monitoring Data and Risk Assessment by Edward A. McBean; Frank A. Rovers. Technometrics, 42(2):220, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271480.

Ziegel:2000:BRBg

[Zie00-47] Eric R. Ziegel. Book review: Statistical Quality Control Using Excel by Steven M. Zimmerman; Marjorie L. Icenogle. Technometrics, 42(2):221, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271482.

Ziegel:2000:BRBaq

[Zie00-48] Eric R. Ziegel. Book review: Statistics and Experimental Design for Toxicologists by Shayne C. Gad. Technometrics, 42(4):443, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270979.

Ziegel:2000:BRBas

[Zie00-49] Eric R. Ziegel. Book review: Statistics and Neural Networks by J. W. Kay; D. M. Titterington. Technometrics, 42(4):443–444, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270981.

Ziegel:2000:BRBs

[Zie00-50] Eric R. Ziegel. Book review: Statistics in Sports by Jay Bennett. Technometrics, 42(2):224, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271494.

Ziegel:2000:BRBal

[Zie00-51] Eric R. Ziegel. Book review: The Complete Guide to Six Sigma by Thomas Pyzdek. Technometrics, 42(4):441–442, November 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270974.

Ziegel:2001:BRBbb

[Zie01a] Eric R. Ziegel. Book review: A Course in Time Series Analysis by Daniel Peña; George C. Tiao; Ruey S. Tsay. Technometrics, 43(4):500, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270846.

Ziegel:2001:BRBw

[Zie01b] Eric R. Ziegel. Book review: Analysis of Health Surveys by Edward L. Korn; Barry I. Graubard. Technometrics, 43(2):247, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271067.

Ziegel:2001:BRBf

[Zie01c] Eric R. Ziegel. Book review: Applied Survival Analysis by David W. Hosmer,; Stanley Lemeshow. Technometrics, 43(1):113–114, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270890.

Ziegel:2001:BRBag

[Zie01d] Eric R. Ziegel. Book review: Basic Engineering Data Collection and Analysis by Stephen B. Vardeman; J. Marcus Jobe. Technometrics, 43(3):380, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271240.

Ziegel:2001:BRBv

[Zie01e] Eric R. Ziegel. Book review: Basic Statistics and Pharmaceutical Statistical Applications by James E. de Muth. Technometrics, 43(2):247, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271066.

Ziegel:2001:BRBs

[Zie01f] Eric R. Ziegel. Book review: Bayes and Empirical Bayes Methods for Data Analysis by Bradley P. Carlin; Thomas A. Louis. Technometrics, 43(2):246, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271063.

Ziegel:2001:BRBm

[Zie01g] Eric R. Ziegel. Book review: Bioenvironmental and Public Health Statistics by P. K. Sen; C. R. Rao. Technometrics, 43(1):115–116, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270897.

Ziegel:2001:BRBo

[Zie01h] Eric R. Ziegel. Book review: Box on Quality and Discovery by George C. Tiao; Søren Bisgaard; William J. Hill; Daniel Peña; Stephen M. Stigler. Technometrics, 43(2):244–245, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271059.

Ziegel:2001:BRBah

[Zie01i] Eric R. Ziegel. Book review: Business Statistics in Practice by Bruce Bowerman; Richard T. O’Connell; Michael L. Hand. Technometrics, 43(3):380, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271241.

Ziegel:2001:BRBab

[Zie01j] Eric R. Ziegel. Book review: Categorical Data Analysis Using the SAS(R) System by Maura E. Stokes; Charles S. Davis; Gary G. Koch. Technometrics, 43(2):249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271072.

Ziegel:2001:BRBk

[Zie01k] Eric R. Ziegel. Book review: Chance Rules by Brian S. Everitt. Technometrics, 43(1):115, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270895.

Ziegel:2001:BRBn

[Zie01l] Eric R. Ziegel. Book review: Chemometrics: Statistics and Computer Application in Analytical Chemistry by Matthias Otto. Technometrics, 43(2):240, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271051.

Ziegel:2001:BRBbc

[Zie01m] Eric R. Ziegel. Book review: Customer Centered Six Sigma by Earl Naumann; Steven H. Hoisington. Technometrics, 43(4):500–501, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270849.

Ziegel:2001:BRBj

[Zie01n] Eric R. Ziegel. Book review: Data Analysis and Signal Processing in Chromatography by Attila Felinger. Technometrics, 43(1):115, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270894.

Ziegel:2001:BRBh

[Zie01o] Eric R. Ziegel. Book review: Data Mining Techniques by Michael J. A. Berry; Gordon Linoff. Technometrics, 43(1):114, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270892.

Ziegel:2001:BRBp

[Zie01p] Eric R. Ziegel. Book review: Design and Analysis of Experiments by Douglas C. Montgomery. Technometrics, 43(2):245, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271060.

Ziegel:2001:BRBl

[Zie01q] Eric R. Ziegel. Book review: Encyclopedia of Biopharmaceutical Statistics by Shein-Chung Chow. Technometrics, 43(1):115, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270896.

Ziegel:2001:BRBaq

[Zie01r] Eric R. Ziegel. Book review: Fitting Equations to Data by Cuthbert Daniel; Fred S. Wood; John W. Gorman. Technometrics, 43(3):384, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271250.

Ziegel:2001:BRBak

[Zie01s] Eric R. Ziegel. Book review: Generalized Linear Models by Dipak K. Dey; Sujit K. Ghosh; Bani K. Mallick. Technometrics, 43(3):381–382, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271244.

Ziegel:2001:BRBi

[Zie01t] Eric R. Ziegel. Book review: geoENV II: Geostatistics for Environmental Applications by Jaime Gómez-Hernández; Amilcar Soares; Roland Froidevaux. Technometrics, 43(1):114–115, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270893.

Ziegel:2001:BRBax

[Zie01u] Eric R. Ziegel. Book review: Geostatistics for Environmental Scientists by Richard Webster; Margaret A. Oliver. Technometrics, 43(4):499, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270844.

Ziegel:2001:BRBat

[Zie01v] Eric R. Ziegel. Book review: Introduction to Statistical Quality Control by Douglas C. Montgomery. Technometrics, 43(4): 497–498, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270840.

Ziegel:2001:BRBq

[Zie01w] Eric R. Ziegel. Book review: Introduction to the Practice of Statistics by David S. Moore; George P. McCabe. Technometrics, 43(2): 245, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271061.

Ziegel:2001:BRBal

[Zie01x] Eric R. Ziegel. Book review: Managing Six Sigma by Forrest Breyfogle; James Cupello; Becki Meadows. Technometrics, 43(3):382, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271245.

Ziegel:2001:BRBas

[Zie01y] Eric R. Ziegel. Book review: Mastering Data Mining by Michael J. A. Berry; Gordon S. Linoff. Technometrics, 43(4):496, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270837.

Ziegel:2001:BRBe

[Zie01z] Eric R. Ziegel. Book review: Mixed-Effects Models in S and S-PLUS by José C. Pinheiro; Douglas M. Bates. Technometrics, 43(1):113, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270889.

Ziegel:2001:BRBad

[Zie01-27] Eric R. Ziegel. Book review: Modern Applied Statistics with S-PLUS by W. N. Venables; B. D. Ripley. Technometrics, 43(2):249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271074.

Ziegel:2001:BRBaa

[Zie01-28] Eric R. Ziegel. Book review: Multivariate Data Reduction and Discrimination with SAS(R) Software by Ravindra Khattree; Dayanand N. Naik. Technometrics, 43(2):248–249, May 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1271071.

Ziegel:2001:BRBba

[Zie01-29] Eric R. Ziegel. Book review: Optimization Heuristics in Econometrics by Peter Winker. Technometrics, 43(4):500, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270848.

Ziegel:2001:BRBg

[Zie01-30] Eric R. Ziegel. Book review: Permutation Tests by Phillip Good. Technometrics, 43(1):114, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270891.

Ziegel:2001:BRBav

[Zie01-31] Eric R. Ziegel. Book review: Principles of Multivariate Analysis by W. J. Krzanowski. Technometrics, 43(4):498, November 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270842.

Ziegel:2001:BRBx

[Zie01-32] Eric R. Ziegel. Book review: Public Policy and Statistics by Sally C. Morton; John E. Rolph. Technometrics, 43(2):247–248, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271068.

Ziegel:2001:BRBau

[Zie01-33] Eric R. Ziegel. Book review: SAS(R) System for Regression by Rudolf J. Freund; Ramon C. Littell. Technometrics, 43(4):498, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270841.

Ziegel:2001:BRBu

[Zie01-34] Eric R. Ziegel. Book review: Scan Statistics and Applications by Joseph Glaz; N. Balakrishnan. Technometrics, 43(2):246–247, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271065.

Ziegel:2001:BRBaz

[Zie01-35] Eric R. Ziegel. Book review: Seasonal Adjustment with the X-11 Method by Dominique Ladiray; Benoit Quenneville. Technometrics, 43(4):500, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270847.

Ziegel:2001:BRBaj

[Zie01-36] Eric R. Ziegel. Book review: Sensitivity Analysis by Andrea Saltelli; Karen Chan; E. Marian Scott. Technometrics, 43(3):381, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271243.

Ziegel:2001:BRBr

[Zie01-37] Eric R. Ziegel. Book review: Smoothing and Regression by Michael G. Schimek. Technometrics, 43(2):245–246, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271062.

Ziegel:2001:BRBae

[Zie01-38] Eric R. Ziegel. Book review: Standard Probability and Statistics Tables and Formulae by Daniel Zwillinger; Stephen Kokoska. Technometrics, 43(2):249–250, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271075.

Ziegel:2001:BRBap

[Zie01-39] Eric R. Ziegel. Book review: Statistical Aspects of BSE and vCJD by Christl A. Donnelly; Neil M. Ferguson. Technometrics, 43(3):383, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271249.

Ziegel:2001:BRBb

[Zie01-40] Eric R. Ziegel. Book review: Statistical Methods for Quality Improvement by Thomas P. Ryan. Technometrics, 43(1):112, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270886.

Ziegel:2001:BRBaf

[Zie01-41] Eric R. Ziegel. Book review: Statistical Methods for the Reliability of Repairable Systems by Steven E. Rigdon; Asit P. Basu. Technometrics, 43(3):371–372, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271228.

Ziegel:2001:BRBay

[Zie01-42] Eric R. Ziegel. Book review: Statistical Methods in Spatial Epidemiology by Andrew B. Lawson. Technometrics, 43(4):499, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270845.

Ziegel:2001:BRBa

[Zie01-43] Eric R. Ziegel. Book review: Statistical Process Monitoring and Optimization by Sung H. Park; G. Geoffrey Vining. Technometrics, 43(1):112, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270885.

Ziegel:2001:BRBam

[Zie01-44] Eric R. Ziegel. Book review: Statistical Science in the Courtroom by Joseph L. Gastwirth. Technometrics, 43(3):382–383, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271246.

Ziegel:2001:BRBz

[Zie01-45] Eric R. Ziegel. Book review: Statistics for Managers Using Microsoft(R) Excel by David M. Levine; Mark L. Berenson; David Stephan. Technometrics, 43(2):248, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271070.

Ziegel:2001:BRBd

[Zie01-46] Eric R. Ziegel. Book review: Statistics for the 21st Century by C. R. Rao; Gábor J. Székely. Technometrics, 43(1):113, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270888.

Ziegel:2001:BRBan

[Zie01-47] Eric R. Ziegel. Book review: Statistics in Ecotoxicology by Tim Sparks. Technometrics, 43(3):383, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271247.

Ziegel:2001:BRBy

[Zie01-48] Eric R. Ziegel. Book review: Statistics on the Table by Stephen M. Stigler. Technometrics, 43(2):248, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271069.

Ziegel:2001:BRBao

[Zie01-49] Eric R. Ziegel. Book review: Studies in the Atmospheric Sciences by L. Mark Berliner; Douglas Nychka; Timothy Hoar. Technometrics, 43(3):383, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271248.

Ziegel:2001:BRBai

[Zie01-50] Eric R. Ziegel. Book review: The Basic Practice of Statistics by David S. Moore. Technometrics, 43(3):381, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271242.

Ziegel:2001:BRBac

[Zie01-51] Eric R. Ziegel. Book review: The Basics of S and S-PLUS by Andreas Krause; Melvin Olson. Technometrics, 43(2):249, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271073.

Ziegel:2001:BRBar

[Zie01-52] Eric R. Ziegel. Book review: Understanding Robust and Exploratory Data Analysis, edited by David C. Hoaglin, Frederick Mosteller, and John W. Tukey, New York: Wiley, 2000, ISBN 0-471-38491-7, xx + 445 pp., $69.95. Technometrics, 43(3):384, August 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.allbusiness.com/technology/
808121-1.html;   http://www.jstor.org/stable/1271251. Spe-

cial Tukey Memorial issue.

Ziegel:2001:BRBaw

[Zie01-53] Eric R. Ziegel. Book review: Visualizing Categorical Data by Michael Friendly. Technometrics, 43(4):498–499, November 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270843.

Ziegel:2001:BRBt

[Zie01-54] Eric R. Ziegel. Book review: Wavelets in Chemistry by Beata Walczak. Technometrics, 43(2):246, May 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271064.

Ziegel:2001:BRBc

[Zie01-55] Eric R. Ziegel. Book review: World Class Quality by Keki R. Bhote; Adi K. Bhote. Technometrics, 43(1):112–113, February 2001. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1270887.

Ziegel:2002:BRBi

[Zie02a] Eric Ziegel. Book review: Biostatistical Methods by John M. Lachin. Technometrics, 44(1):95–96, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270715.

Ziegel:2002:BRBl

[Zie02b] Eric Ziegel. Book review: COMPSTAT: Proceedings in Computational Statistics, 2000 by Jelke G. Bethlehem; Peter G. M. van der Heijden. Technometrics, 44(1):96, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270718.

Ziegel:2002:BRBf

[Zie02c] Eric Ziegel. Book review: Data Mining Cookbook by Olivia Parr Rud. Technometrics, 44(1):94–95, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270712.

Ziegel:2002:BRBg

[Zie02d] Eric Ziegel. Book review: Genetic Algorithms and Engineering Optimization by Mitsuo Gen; Runwei Cheng. Technometrics, 44(1):95, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270713.

Ziegel:2002:BRBe

[Zie02e] Eric Ziegel. Book review: Multivariate Statistical Modelling Based on Generalized Linear Models by Ludwig Fahrmeir; Gerhard Tutz. Technometrics, 44(1):94, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270711.

Ziegel:2002:BRBd

[Zie02f] Eric Ziegel. Book review: Reliability Improvement with Design of Experiments by Lloyd W. Condra. Technometrics, 44(1):93–94, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270710.

Ziegel:2002:BRBj

[Zie02g] Eric Ziegel. Book review: Statistical Methods in Bioinformatics by Warren J. Ewens; Gregory R. Grant. Technometrics, 44(1):96, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270716.

Ziegel:2002:BRBb

[Zie02h] Eric Ziegel. Book review: Trusting Measurement Results in the Chemical and Process Industries by ASQ Chemical and Process Industries Division Chemical Interest Committee. Technometrics, 44(1):93, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270708.

Ziegel:2002:BRBag

[Zie02i] Eric R. Ziegel. Book review: A Handbook of Statistical Analyses Using S-PLUS by Brian S. Everitt. Technometrics, 44(3):300, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270517.

Ziegel:2002:BRBah

[Zie02j] Eric R. Ziegel. Book review: A Handbook of Statistical Analyses Using SAS by Geoff Der; Brian S. Everitt. Technometrics, 44(3):300–301, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270518.

Ziegel:2002:BRBp

[Zie02k] Eric R. Ziegel. Book review: Advanced Linear Modeling by Ronald Christensen. Technometrics, 44(2):198–199, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271283.

Ziegel:2002:BRBan

[Zie02l] Eric R. Ziegel. Book review: An Introduction to Generalized Linear Models by Annette J. Dobson. Technometrics, 44(4):406–407, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271565.

Ziegel:2002:BRBas

[Zie02m] Eric R. Ziegel. Book review: Analysis of Financial Time Series by Ruey Tsay. Technometrics, 44(4):408, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271570.

Ziegel:2002:BRBx

[Zie02n] Eric R. Ziegel. Book review: Analyzing Medical Data Using S-PLUS by B. S. Everitt; Sophia Rabe-Hesketh. Technometrics, 44(2):201–202, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1271291.

Ziegel:2002:BRBao

[Zie02o] Eric R. Ziegel. Book review: Applied Multivariate Data Analysis by Brian S. Everitt; Graham Dunn. Technometrics, 44(4):407, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271566.

Ziegel:2002:BRBc

[Zie02p] Eric R. Ziegel. Book review: Applied Statistics and Probability for Engineers by Douglas C. Montgomery; George C. Runger. Technometrics, 44(1):93, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270709.

Ziegel:2002:BRBax

[Zie02q] Eric R. Ziegel. Book review: Applied Statistics in the Pharmaceutical Industry by Steven Millard; Andreas Krause. Technometrics, 44(4):410, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271575.

Ziegel:2002:BRBt

[Zie02r] Eric R. Ziegel. Book review: Applying Statistics in the Courtroom by Phillip I. Good. Technometrics, 44(2):200, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271287.

Ziegel:2002:BRBw

[Zie02s] Eric R. Ziegel. Book review: Bayesian Survival Analysis by Joseph G. Ibrahim; Ming-Hui Chen; Debajyoti Sinha. Technometrics, 44(2):201, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271290.

Ziegel:2002:BRBam

[Zie02t] Eric R. Ziegel. Book review: Biometrika: One Hundred Years by D. M. Titterington; D. R. Cox. Technometrics, 44(3):302–303, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1270523.

Ziegel:2002:BRBau

[Zie02u] Eric R. Ziegel. Book review: Biostatistical Genetics and Genetic Epidemiology by Robert Elston; Jane Olson; Lyle Palmer. Technometrics, 44(4):409, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271572.

Ziegel:2002:BRBaj

[Zie02v] Eric R. Ziegel. Book review: Biostatistical Methods in Epidemiology by Stephen C. Newman. Technometrics, 44(3):301–302, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270520.

Ziegel:2002:BRBy

[Zie02w] Eric R. Ziegel. Book review: Biostatistics in Clinical Trials by Carol Redmond; Theodore Colton. Technometrics, 44(2):202, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271292.

Ziegel:2002:BRBk

[Zie02x] Eric R. Ziegel. Book review: Curve Ball by Jim Albert; Jay Bennett. Technometrics, 44(1):96, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270717.

Ziegel:2002:BRBat

[Zie02y] Eric R. Ziegel. Book review: Encyclopedia of Environmetrics by Abdel H. El-Shaarawi; Walter W. Piegorsch. Technometrics, 44 (4):408–409, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271571.

Ziegel:2002:BRBz

[Zie02z] Eric R. Ziegel. Book review: Encyclopedia of Epidemiologic Methods by Mitchell H. Gail; Jacques Benichou. Technometrics, 44(2): 202, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271293.

Ziegel:2002:BRBm

[Zie02-27] Eric R. Ziegel. Book review: Engineering Statistics by Douglas C. Montgomery; George C. Runger; Norma F. Hubele. Technometrics, 44(2):197, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271280.

Ziegel:2002:BRBad

[Zie02-28] Eric R. Ziegel. Book review: Experiments with Mixtures by John Cornell. Technometrics, 44(3):299, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270514.

Ziegel:2002:BRBn

[Zie02-29] Eric R. Ziegel. Book review: Fault Detection and Diagnosis in Industrial Systems by L. Chiang; E. Russell; R. Braatz. Technometrics, 44(2):197–198, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271281.

Ziegel:2002:BRBal

[Zie02-30] Eric R. Ziegel. Book review: Frontiers in Statistical Quality Control 6 by H.-J. Lenz; P.-T. Wilrich. Technometrics, 44(3):302, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1270522.

Ziegel:2002:BRBaa

[Zie02-31] Eric R. Ziegel. Book review: Generalized Linear Models by Raymond H. Myers; Douglas C. Montgomery; G. Geoffrey Vining. Technometrics, 44(3):287–288, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270494.

Ziegel:2002:BRBs

[Zie02-32] Eric R. Ziegel. Book review: Making Hard Decisions by Robert T. Clemen; Terence Reilly. Technometrics, 44(2):199–200, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271286.

Ziegel:2002:BRBav

[Zie02-33] Eric R. Ziegel. Book review: Modeling in Medical Decision Making by Giovanni Parmigiani. Technometrics, 44(4):409–410, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271573.

Ziegel:2002:BRBaw

[Zie02-34] Eric R. Ziegel. Book review: Multilevel Modelling of Health Statistics by A. H. Leyland; H. Goldstein. Technometrics, 44(4):410, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-

2723 (electronic). URL http://www.jstor.org/stable/1271574.

Ziegel:2002:BRBac

[Zie02-35] Eric R. Ziegel. Book review: Resampling Methods by Phillip I. Good. Technometrics, 44(3):299, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270513.

Ziegel:2002:BRBab

[Zie02-36] Eric R. Ziegel. Book review: Response Surface Methodology by Raymond H. Myers; Douglas C. Montgomery. Technometrics, 44 (3):298–299, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/1270512.

Ziegel:2002:BRBh

[Zie02-37] Eric R. Ziegel. Book review: Risk Analysis by David Vose. Technometrics, 44(1):95, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270714.

Ziegel:2002:BRBap

[Zie02-38] Eric R. Ziegel. Book review: Sampling by Steven K. Thompson. Technometrics, 44(4):407, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271567.

Ziegel:2002:BRBo

[Zie02-39] Eric R. Ziegel. Book review: Scan Statistics by Joseph Glaz; Joseph Naus; Sylvan Wallenstein. Technometrics, 44(2):198, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271282.

Ziegel:2002:BRBr

[Zie02-40] Eric R. Ziegel. Book review: Soft Computing by L. Fortuna; G. Rizzotto; M. Lavorgna; G. Nunnari; M. G. Xibilia; R. Caponetto.

Technometrics, 44(2):199, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271285.

Ziegel:2002:BRBaq

[Zie02-41] Eric R. Ziegel. Book review: Statistical Inference by George Casella; Roger L. Berger. Technometrics, 44(4):407–408, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271568.

Ziegel:2002:BRBa

[Zie02-42] Eric R. Ziegel. Book review: Statistical Thinking by Roger Hoerl; Ronald Snee. Technometrics, 44(1):92–93, February 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1270707.

Ziegel:2002:BRBv

[Zie02-43] Eric R. Ziegel. Book review: Statistics for Environmental Science and Management by Bryan F. J. Manly. Technometrics, 44(2):201, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271289.

Ziegel:2002:BRBu

[Zie02-44] Eric R. Ziegel. Book review: Statistics for Lawyers by Michael O. Finkelstein; Bruce Levin. Technometrics, 44(2):200, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/1271288.

Ziegel:2002:BRBak

[Zie02-45] Eric R. Ziegel. Book review: Statistics in the 21st Century by Adrian E. Raftery; Martin A. Tanner; Martin T. Wells. Technometrics, 44(3):302, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270521.

Ziegel:2002:BRBai

[Zie02-46] Eric R. Ziegel. Book review: Survey Nonresponse by Robert M. Groves; Don A. Dillman; John L. Eltinge; Roderick J. A. Little. Technometrics, 44(3):301, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270519.

Ziegel:2002:BRBq

[Zie02-47] Eric R. Ziegel. Book review: The Six Sigma Revolution by George Eckes. Technometrics, 44(2):199, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271284.

Ziegel:2002:BRBae

[Zie02-48] Eric R. Ziegel. Book review: The Six Sigma Way by Peter S. Pande; Robert P. Neuman; Roland R. Cavanagh. Technometrics, 44(3):299–300, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1270515.

Ziegel:2002:BRBay

[Zie02-49] Eric R. Ziegel. Book review: The Ultimate Six Sigma by Keki R. Bhote. Technometrics, 44(4):410–411, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/1271576.

Ziegel:2002:BRBar

[Zie02-50] Eric R. Ziegel. Book review: Time Series by Ngai Chan. Technometrics, 44(4):408, November 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1271569.

Ziegel:2002:BRBaf

[Zie02-51] Eric R. Ziegel. Book reviews: Six Sigma and beyond: Foundations of Excellent Performance by D. H. Stamatis; Six Sigma and beyond: Problem Solving and Basic Mathematics by D. H. Stamatis. Technometrics, 44(3):300, August 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/1270516.

Ziegel:2003:BRBc

[Zie03a] Eric R. Ziegel. Book review: Experimental Design for Combinatorial and High Throughput Materials Development by James N. Cawse. Technometrics, 45(4):365, November 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25047099.

Ziegel:2003:BRBa

[Zie03b] Eric R. Ziegel. Book review: Statistical Methods for Detection and Quantification of Environmental Contamination by Robert D.

Gibbons; David E. Coleman. Technometrics, 45(2):176–177, May 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25047025.

Ziegel:2003:BRBb

[Zie03c] Eric R. Ziegel. Book review: The Elements of Statistical Learning by Trevor Hastie; Robert Tibshirani; Jerome Friedman. Technometrics, 45(3):267–268, August 2003. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25047057.

Ziegel:2004:BRBa

[Zie04a] Eric R. Ziegel. Book review: A User-Friendly Guide to Multivariate Calibration and Classification by Tormod Næs; Tomas Isaksson; Tom Fearn; Tony Davies. Technometrics, 46(1):109– 110, February 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470766.

Ziegel:2004:BRBh

[Zie04b] Eric R. Ziegel. Book review: All of Statistics by Larry Wasserman. Technometrics, 46(3):371–372, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470873.

Ziegel:2004:BRBg

[Zie04c] Eric R. Ziegel. Book review: An R and S-PLUS Companion to Applied Regression by John Fox. Technometrics, 46(3):371, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25470872.

Ziegel:2004:BRBu

[Zie04d] Eric R. Ziegel. Book review: Analysis of Survey Data by R. L. Chambers; C. J. Skinner. Technometrics, 46(4):497, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470919.

Ziegel:2004:BRBl

[Zie04e] Eric R. Ziegel. Book review: Chemoinformatics by Johann Gasteiger; Thomas Engel. Technometrics, 46(3):373, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470877.

Ziegel:2004:BRBp

[Zie04f] Eric R. Ziegel. Book review: Computer-Aided Multivariate Analysis by Abdelmonem Afifi; Virginia A. Clark; Susanne May. Technometrics, 46(4):495, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470914.

Ziegel:2004:BRBy

[Zie04g] Eric R. Ziegel. Book review: Data Analysis for Managers with Microsoft(R) Excel by S. Christian Albright; Wayne L. Winston; Christopher Zappe. Technometrics, 46(4):499, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470923.

Ziegel:2004:BRBx

[Zie04h] Eric R. Ziegel. Book review: Environmental Statistics: Methods and Applications by Vic Barnett. Technometrics, 46(4):498–499, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470922.

Ziegel:2004:BRBc

[Zie04i] Eric R. Ziegel. Book review: Handbook of Parametric and Nonparametric Statistical Procedures by David J. Sheskin. Technometrics, 46(3):369–370, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470868.

Ziegel:2004:BRBo

[Zie04j] Eric R. Ziegel. Book review: Introduction to Variance Estimation by Kirk M. Wolter. Technometrics, 46(3):374–375, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470880.

Ziegel:2004:BRBi

[Zie04k] Eric R. Ziegel. Book review: Juran Institute’s Six Sigma Breakthrough and Beyond by Joseph A. De Feo; William W. Barnard. Technometrics, 46(3):372, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470874.

Ziegel:2004:BRBm

[Zie04l] Eric R. Ziegel. Book review: Multiple Analyses in Clinical Trials by Lemuel A. Moyé. Technometrics, 46(3):373–374, August 2004. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25470878.

Ziegel:2004:BRBt

[Zie04m] Eric R. Ziegel. Book review: Probability and Statistics for Engineering and the Sciences by Jay L. Devore. Technometrics, 46 (4):497, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470918.

Ziegel:2004:BRBf

[Zie04n] Eric R. Ziegel. Book review: Regression Using JMP by Rudolf Freund; Ramon Littell; Lee Creighton. Technometrics, 46(3):370– 371, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25470871.

Ziegel:2004:BRBj

[Zie04o] Eric R. Ziegel. Book review: Six Sigma and Beyond: Design for Six Sigma by D. H. Stamatis. Technometrics, 46(3):372, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470875.

Ziegel:2004:BRBr

[Zie04p] Eric R. Ziegel. Book review: Six Sigma and Beyond: The Implementation Process by D. H. Stamatis. Technometrics, 46(4): 495–496, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.org/

stable/25470916.

Ziegel:2004:BRBe

[Zie04q] Eric R. Ziegel. Book review: Spatial Statistics through Applications by J. Mateu; F. Montes. Technometrics, 46(3):370, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470870.

Ziegel:2004:BRBv

[Zie04r] Eric R. Ziegel. Book review: Spectral Analysis of Time-Series Data by Rebecca M. Warner. Technometrics, 46(4):497–498, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470920.

Ziegel:2004:BRBd

[Zie04s] Eric R. Ziegel. Book review: Statistical Methods by Rudolf J. Freund; William J. Wilson. Technometrics, 46(3):370, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470869.

Ziegel:2004:BRBz

[Zie04t] Eric R. Ziegel. Book review: Statistical Size Distributions in Economics and Actuarial Sciences by Christian Kleiber; Samuel Kotz. Technometrics, 46(4):499–500, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470924.

Ziegel:2004:BRBb

[Zie04u] Eric R. Ziegel. Book review: Statistical Tools for Environmental Quality Measurement by Michael E. Ginevan; Douglas E. Splitstone. Technometrics, 46(3):365–366, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470862.

Ziegel:2004:BRBw

[Zie04v] Eric R. Ziegel. Book review: Statistics and Chemometrics for Analytical Chemistry by James Miller; Jane Miller. Technometrics, 46(4):498, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25470921.

Ziegel:2004:BRBn

[Zie04w] Eric R. Ziegel. Book review: Statistics in Musicology by Jan Beran. Technometrics, 46(3):374, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470879.

Ziegel:2004:BRBq

[Zie04x] Eric R. Ziegel. Book review: System Reliability Theory: Models, Statistical Methods, and Applications by Marvin Rausand; Arnljot Høyland. Technometrics, 46(4):495, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470915.

Ziegel:2004:BRBk

[Zie04y] Eric R. Ziegel. Book review: The Practice of Business Statistics by David S. Moore; George P. McCabe; William M. Duckworth; Stanley L. Sclove. Technometrics, 46(3):372–373, August 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470876.

Ziegel:2004:BRBs

[Zie04z] Eric R. Ziegel. Book review: World Class Reliability by Keki R. Bhote; Adi K. Bhote. Technometrics, 46(4):496, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25470917.

Ziegel:2004:BRBaa

[Zie04-27] Eric R. Ziegel. Book reviews: Clinical Trials in Oncology by Stephanie Green; Jacqueline Benedetti; John Crowley; Design and Analysis of Quality of Life Studies in Clinical Trials by Diane L. Fairclough. Technometrics, 46(4):500, November 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL

http://www.jstor.org/stable/25470925.

Ziegel:2005:BRBb

[Zie05a] Eric R. Ziegel. Book review: Statistical Analysis of Financial Data in S-PLUS by René A. Carmona. Technometrics, 47(2):242–243, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471002.

Ziegel:2005:BRBa

[Zie05b] Eric R. Ziegel. Book review: Statistics for Petroleum Engineers and Geoscientists by Jerry L. Jensen; Larry W. Lake; Patrick W. M. Corbett; David J. Goggin. Amsterdam: Elsevier, 2000, ISBN 0-444-50552-0, xviii + 338 pp. Technometrics, 47(2):241, May 2005. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471000.

Ziegel:2006:BRB

[Zie06] Eric R. Ziegel. Book review: Geographically Weighted Regression by A. Stewart Fotheringham; Chris Brunsdon; Martin Charlton. Technometrics, 48(1):149, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471130.

Zimmer:2000:BRB

[Zim00] Lora Zimmer. Book review: Process Capability Indices in Theory and Practice by Samuel Kotz; Cynthia R. Lovelace. Technometrics, 42(2):206–207, May 2000. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/1271455.

Zimmer:2001:BRB

[Zim01] Lora Zimmer. Book review: Applied Statistics for Engineers and Scientists by Joseph D. Petruccelli; Balgobin Nandram; Minghui Chen. Technometrics, 43(1):103, February 2001. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/1270869.

Zimmer:2002:BRB

[Zim02] Lora Zimmer. Book review: Statistical Process Control and Quality Improvement by Gerald M. Smith. Technometrics, 44(2):188, May 2002. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/1271265.

Zimmer:2004:BRB

[Zim04] Lora S. Zimmer. Book review: Applied Probability by Kenneth Lange. Technometrics, 46(2):256, May 2004. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25470817.

Zhang:2006:BMP

[ZM06] Yao Zhang and William Q. Meeker. Bayesian methods for planning accelerated life tests. Technometrics, 48(1):49–60, February 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (elec-

tronic). URL http://www.jstor.org/stable/25471114.

Zuo:2008:AWO

[ZMW08] Jianying Zuo, William Q. Meeker, and Huaiqing Wu. Analysis of window-observation recurrence data. Technometrics, 50(2):128– 143, May 2008. CODEN TCMTA2. ISSN 0040-1706 (print),

1537-2723 (electronic). URL http://www.jstor.org/stable/
25471454.

Zhang:2009:NEE

[ZS09] Jin Zhang and Michael A. Stephens. A new and efficient estimation method for the generalized Pareto distribution. Technometrics, 51(3):316–325, August 2009. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/40586625.

Zhu:2006:LCE

[ZSC06] Mu Zhu, Wanhua Su, and Hugh A. Chipman. LAGO: A computationally efficient approach for statistical detection. Technometrics, 48(2):193–205, May 2006. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471156.

Zou:2008:CPA

[ZTL08] Changliang Zou, Fugee Tsung, and Yukun Liu. A change point approach for Phase I analysis in multistage processes. Technometrics, 50(3):344–356, August 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471495.

Zou:2007:MGL

[ZTW07] Changliang Zou, Fugee Tsung, and Zhaojun Wang. Monitoring general linear profiles using multivariate exponentially weighted moving average schemes. Technometrics, 49(4):395–408, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723

(electronic). URL http://www.jstor.org/stable/25471386.

Zou:2008:MPB

[ZTW08] Changliang Zou, Fugee Tsung, and Zhaojun Wang. Monitoring profiles based on nonparametric regression methods. Technometrics, 50(4):512–526, November 2008. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.

jstor.org/stable/25471534.

Zeng:2007:IIC

[ZZ07] Li Zeng and Shiyu Zhou. Inferring the interactions in complex manufacturing processes using graphical models. Technometrics, 49(4):373–381, November 2007. CODEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic). URL http://www.jstor.

org/stable/25471384.

Zhu:2007:OCO

[ZZJ07] Yu Zhu, Peng Zeng, and Kristofer Jennings. Optimal compound orthogonal arrays and single arrays for robust parameter design experiments. Technometrics, 49(4):440–453, November 2007. CO-DEN TCMTA2. ISSN 0040-1706 (print), 1537-2723 (electronic).

URL http://www.jstor.org/stable/25471390.